A Complete Bibliography of Publications in Mathematics and Computers in Simulation: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

22 August 2025 Version 1.00

Title word cross-reference

$107.25 [116]. $139 [138]. $14.95 [129]. $173 [7]. 2 + 1 [1015]. $24.95 [144]. $26.95 [28]. $27.95 [146]. $29.75 [186]. $29.95 [141, 89]. $34.50 [8]. $36.95 [125]. $37.75 [23]. $39.75 [33]. $44.50 [139, 140]. $45 [148, 200]. $49.95 [143, 214]. $53.75 [25]. $59.95 [147]. $64.00 [20]. $69.95 [29, 92]. $70 [150]. $79 [9]. $84 [36]. $89.75 [16]. $95 [126]. $99 [136, 127]. ^{m} [1027]. _{2} [1038]. E [527]. a [1277]. α [730, 683]. ARIMA(p, d, q) [1309]. β [1257]. N [158, 873, 854, 1156]. O(N) [913]. x [528]. y′′ = f(x, y) [482].

-adic [1257]. -dimensional [873, 1015, 158]. -link [1156]. -stable [730, 683]. -tuple [854].

0 [254, 267, 272, 257, 260, 273, 264, 269, 252, 262, 268, 261, 253, 255, 270, 266, 263, 271, 27, 25, 31, 34, 5, 6, 9, 23, 30, 24, 29, 4, 8, 26, 33, 7, 36, 381, 384, 383, 396, 391, 393, 388, 392, 380, 394, 390, 385, 382, 397, 398, 387, 487, 489, 499, 500, 501, 490, 485, 494, 498, 495, 496, 486, 492, 491, 520, 488, 502, 493, 88, 92, 126, 136, 152, 91, 144, 147, 150, 129, 139, 116, 118, 138, 87, 145, 89, 123, 131, 132, 134, 140, 143, 149, 125, 86, 120, 117, 151, 209, 217, 187]. 0 [196, 198, 208, 210, 225, 289, 203, 229, 286, 193, 211, 218, 224, 230, 282, 183, 219, 281, 214, 216, 204, 184, 185, 194, 201, 205, 221, 189, 202, 207, 212, 283, 226, 186, 190, 223, 197, 284, 182, 192, 199, 285, 429, 430, 446, 456, 356, 359, 423, 450, 354, 357, 360, 432, 418, 421, 427, 415, 353, 451, 352, 413, 424, 452, 358, 422, 366, 371, 420, 447, 453, 364, 349, 433, 434, 454, 367, 368, 369, 370, 426, 455, 457, 414]. 0

2

[417, 425, 428, 431, 448, 449, 529, 545, 558, 546, 530, 531, 550, 560, 522, 533, 532, 523, 561, 524, 521, 548, 552, 557, 527, 547, 525, 562, 526, 549, 551, 535, 555, 556, 559, 15, 16, 17, 11, 12, 14, 13, 534]. 0-02-403891-1 [257]. 0-0702-9213-2 [380]. 0-0804-0734-X [387]. 0-12-037101-4 [152]. 0-12-058470-0 [86]. 0-12-067570-6 [29]. 0-12-079064-5 [203]. 0-12-089040-2 [354]. 0-12-128470-0 [87]. 0-12-174583-X [267]. 0-12-218100-X [147]. 0-12-236430-9 [189]. 0-12-249051-7 [196]. 0-12-257960-7 [89]. 0-12-283955-2 [366]. 0-12-289255-0 [367]. 0-12-304752-8 [197]. 0-12-307350-2 [139]. 0-12-329650-1 [255]. 0-12-400465-2 [22]. 0-12-400466-0 [22]. 0-12-407475-8 [145, 214]. 0-12-460515-X [143]. 0-12-518400-X [144]. 0-12-545030-3 [266]. 0-12-670230-6 [194]. 0-12-675610-4 [91]. 0-12-762535-6 [92]. 0-120-41538-0 [533]. 0-120-44210-8 [502]. 0-120-56740-7 [534]. 0-120-78145-X [500]. 0-120-79061-0 [561]. 0-120-79065-3 [497]. 0-120-79066-1 [497]. 0-1205-6752-0 [394]. 0-121-17645-2 [527]. 0-121-74589-9 [499]. 0-1211-7650-9 [425]. 0-1219-2350-9 [393]. 0-122-08630-9 [562]. 0-122-15508-4 [556]. 0-122-82620-5 [485]. 0-122-89253-4 [535]. 0-123-29070-8 [532]. 0-123-52380-X [524]. 0-124-21850-4 [501]. 0-124-27610-5 [522]. 0-124-87650-1 [547]. 0-1244-0740-4 [415]. 0-125-15440-2 [552]. 0-125-87755-2 [525]. 0-125-98455-3 [549]. 0-1252-3270-5 [392]. 0-126-39360-5 [551]. 0-126-54040-3 [494]. 0-1270-1225-7 [453]. 0-1270-8240-9 [414]. 0-1273-2961-7 [454]. 0-1278-4391-4 [391]. 0-1303-3663-7 [446]. 0-131-76173-0 [495]. 0-1321-3489-6 [432]. 0-1381-2819-7 [385]. 0-19-853394-2 [273]. 0-19-853772-7 [224]. 0-19-856332-9 [221]. 0-201-05866-9 [26]. 0-201-06035-3 [31]. 0-201-06673-4 [27]. 0-201-08319-1 [30]. 0-201-10645-0 [25]. 0-201-12910-8 [34]. 0-201-15674-1 [13]. 0-201-17564-9 [286]. 0-201-19245-4 [33]. 0-201-19248-9 [28]. 0-201-19249-7 [23]. 0-201-50814-1 [183]. 0-201-51002-2 [223]. 0-201-52350-7 [184]. 0-201-54409-1 [526]. 0-201-55580-8 [361]. 0-201-55592-1 [361]. 0-201-56514-5 [358]. 0-2015-5703-7 [398]. 0-262-13235-4 [125]. 0-387-53714-7 [350]. 0-387-97180-7 [216]. 0-387-97502-0 [289]. 0-444-70365-9 [230]. 0-444-70366-7 [116]. 0-444-87473-9 [182]. 0-444-88110-7 [126]. 0-444-88720-2 [208]. 0-444-88721-0 [208]. 0-444-88722-9 [208]. 0-444-88818-7 [353]. 0-444-88861-6 [229]. 0-444-89761-5 [530]. 0-444-89762-3 [531]. 0-4448-8869-1 [382]. 0-470-21050-8 [135]. 0-470-21498-8 [137]. 0-471- [218]. 0-471-50023-2 [104]. 0-471-50357-6 [118]. 0-471-50916-7 [269]. 0-471-61163-8 [11]. 0-471-61294-4 [268]. 0-471-61298-7 [186]. 0-471-61842-X [88]. 0-471-62074-2 [202]. 0-471-62305-9 [132]. 0-471-62896-4 [210]. 0-471-9173-1 [6]. 0-471-91825-3 [219]. 0-471-92064-9 [12]. 0-471-92129-7 [204]. 0-471-92580-2 [205]. 0-471-92581-0 [205]. 0-4715-7127-X [426]. 0-4715-7278-0 [420]. 0-486-67085-6 [486]. 0-486-67102-X [488]. 0-486-67121-6 [487]. 0-4866-6893-2 [457]. 0-4866-7129-1 [449]. 0-521- [207]. 0-521-34340-2 [140]. 0-521-36734-4 [95]. 0-521-36760-3 [93]. 0-521-36791-3 [103]. 0-521-36830-8 [105]. 0-521-37181-3 [8]. 0-521-43064-X [520]. 0-534-92134-5 [545]. 0-534-92880-3 [546]. 0-534-93273-8 [550]. 0-582-05666-7 [211]. 0-7167-2231-3 [263]. 0-7167-8232-4 [252]. 0-74358-0805-0 [131]. 0-7458-0423-3 [101]. 0-7458-0477-2 [137]. 0-7458-0714-3 [134]. 0-7458-0715-1 [102]. 0-792- [521]. 0-792-31448-4 [557]. 0-792-31507-3 [489]. 0-792-31605-3 [493]. 0-792-31677-0 [490]. 0-792-31683-5 [492]. 0-792-31714-9 [496]. 0-792-31772-6 [523]. 0-792-31891-9 [559]. 0-792-31944-3 [555]. 0-792-32059-X [558]. 0-792-32061-1 [529]. 0-792-32064-6 [560]. 0-7923-0049-1 [138]. 0-7923-0051-3 [24]. 0-7923-0113-7 [36]. 0-7923-0127-7 [7]. 0-7923-0210-9 [4]. 0-7923-0216-8 [383].

0-7923-0219-2 [5]. 0-7923-0229-X [450]. 0-7923-0231-1 [136]. 0-7923-0233-8 [431]. 0-7923-0307-5 [9]. 0-7923-0340-7 [284]. 0-7923-0442-X [149]. 0-7923-0471-3 [151]. 0-7923-0472-1 [151]. 0-7923-0473-X [151]. 0-7923-0474-8 [151]. 0-7923-0494-2 [264]. 0-7923-0521-3 [272, 217]. 0-7923-0539-6 [254]. 0-7923-0543-4 [261]. 0-7923-0585-X [356]. 0-7923-0594-9 [199]. 0-7923-0595-7 [201]. 0-7923-0623-6 [260]. 0-7923-0675-9 [357]. 0-7923-0689-9 [187]. 0-7923-0721-6 [193]. 0-7923-0724-0 [198]. 0-7923-0728-3 [185]. 0-7923-0730-5 [190]. 0-7923-0822-0 [359]. 0-7923-0875-1 [253]. 0-7923-0901-4 [281]. 0-7923-0910-3 [396]. 0-7923-0992-8 [368]. 0-7923-1029-2 [270]. 0-7923-1036-5 [416]. 0-7923-1038-1 [369]. 0-7923-1042-X [451]. 0-7923-1113-2 [364]. 0-7923-1122-1 [417]. 0-7923-1152-3 [271]. 0-7923-1153-1 [424]. 0-7923-1155-8 [447]. 0-7923-1157-4 [371]. 0-7923-1218-X [370]. 0-7923-1236-8 [285]. 0-7923-1248-1 [360]. 0-7923-1265-1 [418]. 0-7923-1296-1 [429]. 0-7923-1317-8 [362]. 0-7923-1330-5 [352]. 0-7923-1336-4 [413]. 0-7923-1339-9 [390]. 0-7923-1409-3 [430]. 0-7923-1438-7 [421]. 0-7923-1441-7 [384]. 0-7923-1447-6 [397]. 0-7923-1483-2 [423]. 0-7923-1488-3 [452]. 0-7923-1517-0 [388]. 0-7923-1531-6 [448]. 0-7923-1546-4 [422]. 0-7923-1550-2 [428]. 0-7923-1574-X [381]. 0-7923-1593-6 [427]. 0-7923-1594-4 [455]. 0-7923-1680-0 [433]. 0-7923-1700-9 [456]. 0-7923-9055-5 [150]. 0-7923-9095-4 [282]. 0-805-80267-3 [536]. 0-805-80268-1 [536]. 0-8176-2641-7 [287]. 0-8176-3530-0 [288]. 0-8176-3555-6 [355]. 0-8176-3585-8 [351]. 0-8176-3586-6 [386]. 0-8247-7333-0 [117]. 0-8247-7556-2 [120]. 0-8247-7661-5 [98]. 0-8247-8049-3 [16]. 0-8247-8120-1 [17]. 0-8247-8125-2 [99]. 0-8247-8161-9 [14]. 0-8247-8347-6 [209]. 0-8247-8411-1 [226]. 0-8306-3137-2 [90]. 0-8306-9137-5 [90]. 0-8493-8608-X [389]. 0-8493-8610-1 [283]. 0-8493-8614-4 [389]. 0-85312-226-1 [15]. 0-867-20203-3 [548]. 0-867-20293-9 [498]. 0-87835-337-2 [129]. 0-89871-250-5 [212]. 0-938974-03-3 [349]. 07923-0781-X [191]. 07923-0782-8 [191].

1 [257, 262, 253, 255, 6, 30, 382, 115, 136, 138, 130, 124, 183, 222, 283, 226, 181, 429, 360, 418, 424, 434, 369, 419, 417, 449, 529, 554, 547, 526, 1120, 1181, 1267, 1294, 1160, 1125, 1226, 1232, 1173, 1206, 1287, 1275, 1409, 15, 17, 13, 877, 395]. 1-5555-8075-0 [434]. 1-5555-8090-4 [419]. 1-55558-047-5 [181]. 1-568-81000-8 [554]. 1-85296-354-9 [94]. 1-85312-062-6 [206]. 1-85312-065-0 [262]. 1-D [1401]. 1/4in [361]. 10 [556]. 10-VIII [556]. 10th [743]. 11th [1034]. 13th [235, 237]. 14.95 [94]. 15 [95]. 16.50 [104]. 1989/90 [38, 39]. 1990 [82, 174, 40, 107, 155]. 1991 [275, 303, 243, 291, 363]. 1992 [436, 435, 344, 400, 469, 373, 459, 438]. 1993 [504, 566, 478, 538, 516, 511]. 1994 [573, 592, 662, 671, 582, 626, 644, 616]. 1995 [727, 825, 703, 742, 784, 812, 1034]. 1996 [849, 883, 935]. 1997 [1120, 1033, 1020, 1086, 1094]. 1998 [1181, 1185, 1267, 1294, 1160, 1125, 1141, 1226, 1232, 1173, 1206, 1287, 1275]. 1999 [1347, 1409, 1418, 1301, 1328, 1338, 1394, 1363, 1357].

2 [273, 264, 270, 5, 380, 395, 129, 139, 130, 140, 120, 213, 205, 202, 223, 423, 354, 366, 364, 457, 428, 552, 527, 525, 180]. 2-225-81899-1 [130]. 2-D [114, 180]. 21 [626]. 22 [1418]. 24 [784]. 25 [102]. 28119 [127]. 2D [1372, 1028, 1229]. 2N [653]. 2N-wave [653]. 2nd [27, 391, 520, 122, 196, 210, 186, 434, 525].

3 [272, 265, 266, 263, 271, 31, 32, 24, 8, 386, 396, 489, 494, 491, 493, 115, 133, 122, 121, 119, 123, 124, 134, 142, 128, 213, 217, 225, 219, 215, 185, 222, 228, 192, 220, 227, 430, 351, 355, 365, 452, 349, 546, 531, 548, 553, 549, 555, 10, 16, 19, 18, 21]. 3-211-82110-4 [142]. 3-211-82157-0 [192]. 3-211-82222-4 [365]. 3-519-02089-0 [491]. 3-519-02176-5 [362]. 3-5190-2175-7 [416]. 3-540-12760-7 [228]. 3-540-17008-1 [115]. 3-540-18179-2 [213]. 3-540-50368-4 [19]. 3-540-50443-5 [10]. 3-540-50739-6 [96]. 3-540-51106-7 [97]. 3-540-51159-8 [100]. 3-540-51485-6 [119]. 3-540-51830-4 [133]. 3-540-52645-5 [206]. 3-540-52768-0 [225]. 3-540-53714-7 [350]. 3-540-96689-7 [128]. 3-540-96794-X [21]. 3-540-96919-5 [18]. 3-540-96981-0 [123]. 3-540-97003-7 [215]. 3-540-97019-3 [121]. 3-540-97042-8 [122]. 3-540-97195-5 [220]. 3-540-97202-1 [222]. 3-540-97342-7 [227]. 3-764-32872-X [553]. 3-7643-2384-1 [124]. 3-7643-2464-3 [265]. 3-7643-2641-7 [287]. 3-7643-3530-0 [288]. 3-7643-3555-6 [355]. 3-7643-3585-8 [351]. 3-7643-3586-6 [386]. 3-D [906, 843, 114]. 3-dimensional [656]. 30 [644, 812, 1185]. 31 [671, 1141]. 312in [361]. 3192 [521]. 3192-0 [521]. 39 [97]. 39.95 [101]. 395pp [188]. 39877 [207]. 39877-0 [207]. 3D [1372, 1378, 1385, 1190, 1350]. 3rd [205].

4 [252, 268, 261, 27, 33, 391, 501, 133, 152, 91, 142, 125, 210, 188, 282, 281, 365, 415, 413, 422, 371, 419, 455, 557, 535, 556, 19]. 4-nodes [1137]. 42 [1117]. 45 [1293]. 47 [1300]. 473 [1117]. 4D [1166]. 4in [361].

5 [259, 256, 9, 392, 485, 492, 150, 203, 212, 190, 220, 181, 352, 358, 545, 530, 522, 551, 662, 825, 10, 18]. 5-256-00304-6 [259]. 5th [549].

6 [254, 260, 259, 29, 386, 397, 487, 486, 92, 141, 118, 119, 209, 229, 193, 194, 355, 432, 427, 448, 560, 523]. 68 [96].

7 [269, 23, 7, 36, 384, 385, 398, 126, 116, 89, 128, 196, 211, 224, 215, 216, 204, 184, 201, 228, 186, 284, 227, 446, 421, 353, 453, 454, 20, 534]. 7.50 [103]. 7.95 [93].

8 [34, 383, 502, 122, 145, 195, 214, 197, 285, 351, 447, 368, 431, 550, 532, 554, 742, 11]. 8.95 [105]. 80th [234]. 81-224-0062-0 [135]. 85 [532].

9 [258, 4, 26, 35, 393, 390, 499, 498, 496, 132, 187, 208, 286, 230, 221, 189, 182, 199, 456, 357, 414, 425, 562, 559, 12, 14]. 90 [258, 256, 32, 35, 188, 195, 20]. 90-277-2565-9 [35]. 90-277-2722-8 [195]. 90-277-2733-3 [32]. 90-277-2758-9 [258]. 90-277-2772-4 [188]. 90-277-2776-7 [20]. 90-277-2830-5 [256]. 90277 [127]. 90277-28119 [127]. 92287 [218]. 92287-0 [218]. ’96 [1102].

Abaffy [101]. Abell [533]. ABLAK [247]. Abrahams [398]. ABS [101]. absorbing [1399]. Abstract [487]. Academic [267, 255, 266, 22, 5, 9, 32, 24, 29, 4, 35, 7, 36, 391, 393, 392, 394, 499, 500, 497, 501, 485, 494, 502, 92, 152, 91, 144, 147, 148, 139, 141, 87, 145, 146, 89, 143, 86, 196, 203, 214, 194, 189, 197, 354, 415, 366, 453, 454, 367, 414, 425, 522, 533, 532, 561, 524, 552, 527, 547, 525, 562, 549, 551, 535, 556, 20, 534]. according [1345]. accuracy [1278, 1282, 3, 165]. Accurate [648, 1230]. Achieser [449]. acid [1099]. acidification [45, 191]. Acoustic [208, 7, 531, 1396]. acoustics [208, 530, 531, 208]. activated [807]. Active [1200, 815, 720, 1194]. activity [1065]. actuator [1193]. actuators [718]. Ada [251]. Adams [502]. Adaptation [317]. adapted [1271]. Adaptive [714, 1373, 869, 1029, 1417, 638, 475, 603, 933, 1243, 542, 870, 1390]. Addison [27, 25, 31, 34, 23, 30, 28, 26, 33, 398, 286, 183, 184, 223, 358, 361, 526, 13]. Addison-Wesley [27, 34, 28, 398, 286, 183, 184, 223, 358, 361, 526, 13]. addition [909]. adhesive [905]. adic [1257]. adjustment [1391]. Admiral [514]. Adv [529, 555]. Advances [787, 253, 35, 521, 152]. Advancing [981]. advection [1121, 476, 111, 63]. Aeroacoustics [208]. aerosol [1234]. Affiliated [242, 242]. agency [1101]. agent [1082]. aggregates [839]. Aggregation [756]. ago [1413]. agricultural [1281, 748]. agriculture [952]. AGV [893]. AI [788, 1175]. aid [312]. aided [143, 196, 292, 296, 300, 250]. AIDS [1041]. air [1040]. AKNS [1018]. Akritas [11]. al [94, 100, 225, 229, 222, 10]. Alabama [1041]. Alcamo [191]. Alexandrov [127]. algae [1051]. Algebra [384, 553, 487, 450, 974, 11, 1132, 1129, 1002, 845, 960, 967, 1128, 979, 923, 1139, 1133]. Algebraic [1004, 1137, 983, 485, 195, 988, 1122, 296, 984, 976, 980, 962, 1133, 22]. Algebras [136]. algorithm [642, 974, 1154, 158, 587, 604, 476, 374, 1242, 971, 1122, 1014, 483, 159, 1304, 657, 1075, 977, 718, 1390, 1230, 1047]. Algorithms [31, 998, 126, 128, 1378, 1251, 866, 1252, 618, 1028, 1411, 484, 1255, 1081, 1355, 1072, 169, 980, 518, 878, 101, 199, 27]. allocation [1038, 62]. Allogower [228]. alloys [1377]. almost [356, 326]. along [921]. Alpha [446]. Alternate [88]. alternating [807, 930]. alternative [1321, 766]. Altman [390]. America [45]. américaines [690]. Ames [534]. amino [1099]. among [1320, 777]. amorphous [1356]. Amsterdam [382, 126, 116, 208, 229, 230, 182, 353, 530, 531]. analog [1202, 1263]. analyses [1041]. analysing [907]. Analysis [230, 1046, 822, 707, 1387, 377, 892, 719, 259, 253, 271, 331, 116, 97, 135, 189, 212, 418, 556, 676, 1291, 1412, 1080, 724, 16, 640, 1245, 1208, 1044, 444, 78, 852, 914, 1113, 1415, 1070, 603, 1030, 442, 480, 1243, 164, 901, 407, 297, 1383, 298, 1100, 853, 580, 815, 1326, 248, 60, 1053, 329, 1224, 1388, 1217, 1081, 479, 718, 176, 539, 800, 1212, 1171, 165, 1109, 994, 1397, 309, 1264, 780, 1325, 1392, 868, 64, 856, 809, 1393, 1332, 1140, 282, 267, 22, 31, 104, 209, 193, 560, 523, 525, 15, 19, 12, 500]. Analysis [115, 213]. Analytic [1245, 761, 3, 133, 429]. Analytical [1021, 113]. analyzing [988]. and/or [878]. Anderson [86]. anharmonic [1345]. animal [311, 951, 306, 309]. anisotropy [834]. annealing [1092, 62, 618]. Announcement [274, 233, 232]. annually [326]. annular [1190]. antibiotic [1355]. anticipative [678]. antithetic [1410, 1249]. Antoni [151]. Antoniou [358]. ANU [54]. ANU-CLOUD [54]. Apanasov [383]. apparent [624]. applicability [1037]. applicable [219, 1138]. Application [242, 241, 893, 905, 444, 815, 931, 512, 928, 331, 365, 898, 439, 1042, 1282, 1052, 60, 1128, 1116, 1397, 1024, 1110, 624, 870]. Applications [4, 136, 138, 149, 151, 187, 188, 195, 284, 260, 261, 258, 256, 9, 32, 24, 35, 322, 199, 371, 447, 522, 521, 559, 787, 534, 54, 845, 782, 1251, 866, 295, 1083, 1209, 951, 1106, 1041, 1040, 1135, 114, 472, 253, 381, 496, 493, 135, 211, 429, 420, 17, 11, 14, 126]. Applied [209, 235, 237, 1299, 6, 394, 215, 216, 186, 433, 13, 912, 1335, 272, 489, 217, 545, 15, 16]. apply [242]. Applying [1307]. approach [1269, 1031, 34, 562, 541, 277, 1043, 475, 1242, 635, 1245, 957, 774, 791, 1370, 77, 918, 597, 56, 1405, 473, 48, 603, 305, 857, 720, 542, 767, 1053, 329, 110, 955, 403, 772, 907, 1389, 1124, 1305, 165, 1264, 1110, 1271, 961, 1295, 140]. Approaches [969, 761, 1406, 940, 47]. Approximation [635, 124, 211, 158, 674, 1227, 1300, 1164, 874, 655, 462, 84, 685, 696, 449, 267, 381, 499, 456, 10]. approximations [92, 1213, 678, 1292, 1298, 281]. April [1181, 1185, 82, 344, 504, 573, 849]. APROS [792]. APSIM [748]. arbitrary [584]. archaeological [1297]. Architectural [1068]. architecture [446, 594, 933, 410, 423]. architectures [871]. area [1304]. areal [68]. areas [745]. arid [747]. ARIMA [1053]. arising [790]. arithmetic [507]. Arkhangel’skii [448]. arm [915]. ARMA [1060]. arms [922]. army [51]. Arnold [369]. array [872]. arrivals [1326]. art [520]. articulated [894]. artificial [1167, 50, 1132, 891, 880, 1051, 859, 871, 803]. Ashtekar [1003]. ASI [253, 456, 418, 364, 428]. aspects [1280, 620, 952, 576, 364]. assembly [923]. Assessing [1036, 306]. assessment [322, 878]. asset [691, 692]. assignment [1154, 1072]. assistant [606]. assisted [471, 979]. associate [242]. associated [463]. association [1024]. associative [822]. Asymmetric [1012]. asymptotic [647, 254, 92, 209, 529, 558]. ATESAT [1167]. Athreya [86]. Atkinson [104]. atmospheric [1290]. attraction [415, 1149]. attractors [730, 105]. Aubert [29]. auditory [819]. August [1267, 275, 400, 592, 727, 1347]. Austin [181]. Australia [342, 758, 1326, 76, 1057, 323, 1036]. Australian [339, 51, 1323, 60]. Author [38, 172, 467, 564, 615, 670, 726, 811, 863, 925, 1009, 1077, 700]. automated [905, 430]. Automatic [1127, 1354, 657, 502, 899, 519]. autonomous [483, 877, 906]. autoregression [1062, 754]. autoregressive [1061]. average [1351, 46, 751]. averaged [508]. averaging [138]. avoidance [893, 483, 737, 922]. Azizov [204].

B [5, 30, 383, 385, 398, 492, 520, 136, 148, 137, 118, 97, 132, 86, 227, 456, 418, 349, 368, 546, 553, 21]. Baase [31]. Babbage [514]. Bachmair [355]. back [1041, 818]. back-calculation [1041]. backpropagation [851]. backscattering [1258, 1265]. Backward [1266]. Bainov [496].

Bakhvalov [138]. balancing [894]. Banach [396]. band [1256]. Bandermer [523]. bandwidth [1342]. bar [53]. Barbeau [18]. Barbu [500]. Barnsley [497, 203, 554, 561]. baroclinic [348]. Bartlett [498, 548]. base [51]. based [1417, 765, 285, 541, 893, 180, 604, 598, 891, 1242, 1405, 1177, 603, 296, 789, 1085, 1252, 1371, 596, 904, 509, 823, 1217, 76, 906, 1389, 1258, 1193, 1124, 405, 595, 1212, 1022, 249, 600, 941, 1352, 856, 639, 568, 599, 288]. Basel [265, 386, 124, 288, 287, 351, 355, 553]. bases [1169, 1170, 982]. Basin [1059, 1036]. basis [673, 544, 1298]. batch [1099, 1100, 794]. bay [1331]. Bayesian [1106]. Bayesian-validated [1106]. be [774]. beam [1248, 1352]. Beams [208]. bearing [1116]. Beaufort [514]. bed [635]. beer [1283]. beginner [425]. Behavior [1088, 647, 1069, 334]. behaviour [1204, 781, 918, 1070, 1315, 1308]. beispielen [350]. Bell [221]. BEM [1393]. benchmark [1244]. bending [918]. benefits [319]. Benjamin [150]. Bennewitz [354]. Benutzerhandbuch [553]. Benzon [263]. Berezin [370]. Berger [185]. Berlin [262, 115, 133, 122, 121, 96, 97, 119, 123, 128, 100, 213, 206, 225, 289, 215, 216, 222, 228, 220, 227, 351, 355, 350, 10, 19, 18, 21]. Berman [202]. Berndt [21]. betas [1322]. between [1222, 544, 827, 576]. bias [150]. Biennial [1034, 743]. Bifurcation [832, 1332, 695]. Bifurcations [833]. bilinear [1148]. bilinearization [1080]. binary [818]. Biological [842, 5, 802, 1296]. biologically [815]. biophysical [749, 818]. bioprocess [941]. biopsies [843]. bioreactor [1100]. bioreactors [943, 512, 636]. biotechnical [942]. Birkhäuser [265, 386, 124, 288, 287, 351, 355, 553]. Birkhoff [234]. birthday [234]. bistable [841]. bivariate [1319]. Bl [281]. Bldg [242]. bleomycin [804]. Block [947, 583, 1367, 1271]. block-diagonalization [1271]. Blue [90, 1051]. blue-green [1051]. Board [41, 175, 345, 470, 567, 617, 702, 728, 813, 864, 926, 1010, 1078]. Boca [389, 283]. body [341]. Bollinger [25]. Boltzmann [688, 696]. Bombieri [29]. bond [787, 790]. bonds [1092, 1273]. bone [843]. Book [254, 267, 272, 257, 260, 273, 264, 269, 252, 265, 262, 268, 259, 261, 253, 258, 255, 256, 270, 266, 263, 271, 22, 27, 25, 31, 34, 5, 6, 9, 23, 30, 32, 24, 28, 29, 4, 8, 26, 35, 33, 7, 381, 384, 386, 383, 396, 391, 393, 388, 392, 380, 394, 389, 390, 395, 379, 385, 382, 397, 398, 487, 489, 499, 500, 497, 501, 490, 485, 494, 498, 495, 496, 486, 492, 491, 520, 488, 502, 493, 88, 92, 101, 115, 126, 133, 136, 152, 122, 91, 94, 95, 105, 121, 144, 147, 148, 150, 96, 90, 102, 129, 137, 139]. Book [127, 141, 93, 116, 118, 138, 87, 97, 103, 104, 119, 145, 98, 130, 146, 89, 123, 124, 131, 132, 134, 140, 142, 143, 99, 149, 125, 86, 120, 128, 117, 151, 135, 100, 209, 213, 217, 206, 187, 196, 198, 208, 210, 225, 289, 203, 229, 286, 200, 188, 193, 211, 218, 224, 230, 282, 288, 183, 219, 281, 195, 214, 215, 216, 204, 184, 185, 194, 201, 205, 221, 189, 202, 207, 212, 222, 228, 283, 287, 226, 186, 190, 223, 191]. Book [197, 284, 182, 192, 199, 285, 220, 227, 181, 429, 430, 446, 456, 356, 351, 355, 359, 423, 436, 450, 354, 357, 360, 365, 432, 418, 421, 427, 415, 353, 451, 352, 413, 424, 452, 358, 422, 361, 366, 371, 420, 447, 453, 364, 349, 350, 362, 416, 433, 434, 454, 367, 368, 369, 370, 419, 426, 455, 457, 414, 417, 425, 428, 431, 448, 449, 435, 529, 545, 558, 546, 530, 531, 550, 560, 522, 533, 532, 523, 554, 561, 524, 521, 548]. Book [552, 557, 527, 547, 553, 525, 562, 526, 549, 551, 535, 555, 556, 559, 10, 15, 16, 17, 19, 11, 20, 12, 14, 13, 18, 21, 534, 536]. Books [90]. Bootstrap [333, 761]. Borde [425, 527]. bordering [544]. Boston [267, 255, 266, 391, 393, 392, 394, 500, 494, 498, 502, 152, 144, 147, 148, 129, 145, 146, 143, 196, 288, 194, 189, 287, 197, 351, 355, 354, 366, 453, 454, 414, 545, 546, 550, 522, 533, 532, 561, 524, 548, 552, 527, 547, 525, 562, 549, 551, 535, 556]. boundaries [1, 555]. boundary [1269, 287, 371, 1375, 1353, 1368, 649, 1297, 1399, 544, 930, 1252, 681, 685, 666, 1110, 568, 206, 546]. Bounded [170, 1358, 480, 375, 460, 168, 165, 157, 356]. Bounded-error [170, 480, 375, 460, 168, 165, 157]. boundedness [1149]. bounding [1358, 163, 162]. bounds [19, 1316, 159, 1315, 973, 157]. Bowman [120, 17]. box [949]. Boyer [118, 430]. BR [1006]. brachistochrone [589]. Brackx [384]. brain [846]. brake [1199]. braking [797]. Branching [688]. Braselton [533]. Brawer [87]. Brayshaw [397]. break [1062, 1310]. breathing [78]. Brebbia [126, 206]. Bret [452]. bridges [1379]. broad [745]. broadband [1179]. broadening [736]. Broel [1043]. Brough [422]. Bruno [996]. brushless [1199]. Brusselator [736]. Bu [91]. Bu-quing [91]. bubble [1110]. bubble-structure [1110]. budget [1039]. Building [949, 205, 249]. buildings [1383, 1068, 146]. buoyancy [586]. Burgers [655]. Burlington [495, 446, 432, 434, 419]. Buschman [559]. Bush [389, 137]. business [757]. Butkovskiy [187]. Byrnes [253].

C

[267, 266, 28, 394, 390, 385, 499, 491, 493, 126, 148, 118, 124, 142, 117, 206, 198, 184, 192, 351, 423, 354, 447, 532, 555, 21, 242, 1270]. C.-c [493]. C.-G [493]. c/o [242]. CA [499, 497, 501, 485, 415, 425]. cables [67]. Cabri [1007]. Cabri-graph [1007]. CAD [494]. Cakmac [208]. calculating [768]. Calculation [1192, 583, 463, 1006, 1041, 586]. calculations [180, 1246, 1249, 695]. Calculus [805, 676, 260]. calendar [81, 154, 239, 240, 1076, 1093]. calibrating [323]. calibration [58]. calibrations [255]. Calif [22]. California [29]. call [274, 232]. Cambridge [8, 520, 94, 95, 105, 93, 103, 140, 125, 207]. canonical [1184, 355]. canyon [1039]. capacitance [1]. capillary [481]. capita [1057]. capital [330]. CAPM [1322]. car [1191]. carbon [339]. carcass [917]. card [524]. cardinals [95]. Carlo [142, 1031, 1236, 692, 73, 1238, 1239, 1410, 1242, 1241, 1244, 1245, 755, 1066, 1251, 1243, 1252, 687, 1254, 1253, 310, 1255, 1258, 539, 1260, 1261, 1262, 1263, 1235, 1264, 1266, 1265, 338]. Carroll [382]. cartesian [1198, 922]. cascade [633]. case [252, 140, 45, 603, 248, 160, 736]. cases [1041]. Casti [88]. cat [819]. catchment [1043, 305]. catchments [1050, 1036]. caustics [369]. cautionary [779]. cavities [1381]. cavity [1381]. CBF [1385]. cell [311, 1237, 905, 1252, 634, 887, 844]. cell-to-boundary [1252]. Cellier [289]. Center [436, 435]. centers [1298]. certain [1382]. Ch [350]. Chambers [94]. Chan [7]. Chandy [26]. Chang [91]. Change [150]. changes [335]. Chanturia [558]. chaos [646, 95, 1318, 827, 215]. chaotic [1012, 128, 1121, 1091, 650, 1317, 835, 105]. Characterisation [1196]. characteristic [1130, 963]. characteristics [327, 894]. Characterization [166, 160]. charge [590, 1286]. charges [1228]. Chaudouet [206]. Chaudouet-Miranda [206]. Chebyshev [210]. chemical [1046, 1043]. chemically [944]. chemistry [1290]. Chen [562]. Cheney [499]. Cheshire [200]. Chichester [269, 268, 6, 101, 102, 137, 134, 218, 219, 204, 205, 12]. Chidaglia [364]. Chief [363]. Chikan [264]. China [1320, 1070, 312, 1319]. Chinese [91]. Chiyokura [33]. chloride [1049]. chlorophyll [1277]. choice [765, 769]. CHRUs [1043]. Chua [128]. Chui [267, 499, 124]. Ciarlet [116, 230]. Cioranescu [396]. cipher [514]. circuit [1089, 801]. circuits [927, 447]. circulation [1121]. City [13]. Clarence [1036]. Clarendon [273, 221]. clarifiers [806]. class [1150, 464, 791, 901, 684, 1153, 1257]. classes [279]. classical [333, 628, 560]. Classics [420]. Classification [1114, 875, 816]. classifier [893]. clauses [288]. Clements [140]. Cleveland [436, 435]. Clifford [1109, 450]. Cliffs [385]. climate [305, 1037, 1285]. climatic [313]. climex [317]. clinical [859]. Clock [1304, 473]. Closed [912, 1090]. closure [319, 986, 966, 985]. closures [1105]. cloth [205]. Cloud [46, 508, 54]. Cloud-average [46]. cloud-averaged [508]. cluster [193]. Co [27, 25, 23, 28, 26, 149, 754, 1038]. co-integration [754]. co-published [149]. coagulation [1333, 1234]. coastal [80, 341, 1]. coastal-sand-body [341]. Code [351]. codes [1127, 506, 20]. coefficient [1128]. coefficients [1278, 571]. cognitive [536, 579]. Cohen [415]. coherence [846]. coil [1021]. Cointegration [758, 752, 767]. collaborative [608]. collapse [1091]. Collision [737]. Collocation [443, 544, 589, 512]. colour [1244]. Combes [100]. combination [991, 1085]. Combining [1042, 632]. combustion [412]. Comment [836]. Comments [276, 508]. commitees [238]. Common [1058]. communication [840, 1179]. community [59]. Comp [1293, 1300, 1117]. compact [158]. compactons [1268]. companies [247]. companion [361]. comparative [322, 197]. Comparing [1351]. Comparison [1299, 1198, 325, 463, 618, 460, 66, 755, 1072]. Comparisons [571]. compartmental [83]. compensators [713]. competition [775, 1329]. complementarity [393]. complete [148]. completely [490]. complex [577, 220, 578, 1108, 569, 1105, 45, 1082, 802, 178, 943, 1174, 659, 607]. Complexity [1211, 973, 975, 1222, 1242, 1282, 972, 1073, 386]. compliance [319]. component [838, 164, 1100]. components [1089, 305, 753, 776]. Composite [1318, 1373]. compression [554]. compressor [1219]. Computation [235, 237, 1273, 1107, 642, 1080, 1341, 991, 1183, 1239, 605, 1091, 992, 978, 624, 1122, 716, 985, 1117, 1212, 376, 507, 983, 1134, 198]. Computational [206, 262, 209, 1242, 612, 1252, 800, 817, 819, 828, 91, 208, 225, 436, 530, 531]. computations [1236, 987, 1241, 965, 296, 113, 1179, 98, 363]. compute [461, 159, 948]. Computer [28, 471, 11, 1132, 1129, 960, 894, 730, 683, 296, 967, 505, 1128, 1217, 300, 250, 263, 494, 126, 143, 196, 286, 181, 413, 358, 434, 414, 550, 292, 1002, 845, 1250, 1100, 1025, 321, 76, 1106, 279, 979, 176, 1298, 341, 1139, 1133, 273, 25, 31, 384]. Computer-aided [296, 300, 250, 196, 292]. Computer-assisted [471, 979]. computer-based [76]. computer-simulation [1106]. Computerized [843]. computers [575, 259, 121]. Computing [1092, 593, 654, 1130, 178, 1342, 269, 1348, 520, 502, 367, 535, 1340, 1004, 1413, 1082, 823, 977, 1390, 1346, 79, 518, 607, 183, 535, 423]. con [764]. concentration [46, 508]. Concept [631, 1254]. Concepts [665, 880]. conceptual [408, 948]. concise [94]. concrete [34]. condition [1353, 1399]. conditional [1313, 1319]. conditions [970, 894, 873, 1192, 685, 445, 1283]. conduction [179, 1230]. conduction-type [1230]. conductors [1400]. Conference [1365, 1034, 593, 743, 1118, 436, 362, 416, 435, 903, 704]. confidence [1282]. configurations [901]. confined [586].

Conformal [147]. congested [637]. Congestion [773]. Congress [235, 237]. connected [348, 166]. connection [872]. connectionist [857, 861]. connectivity [843]. consciousness [846, 847]. Conservation [342, 265, 199, 1370, 308, 1114, 1115]. consideration [719]. considerations [874]. consistency [144]. consistent [348]. Constales [384]. Constant [1319, 1021]. constrained [62, 347, 1171]. constraints [1231]. construct [1013]. constructible [966]. Constructing [1257]. Construction [44, 604, 1184]. constructions [1359]. constructive [732]. consumer [1070]. Consumption [840]. contact [1368, 1336, 1380]. contactless [1188]. contaminant [1253]. contamination [932]. Contemporary [550]. Contents [39, 173, 468, 565]. context [460]. contextual [871]. continuation [228, 735]. continued [133, 17]. continuization [472]. continuous [413, 1288, 852, 873, 1025, 580, 793, 338, 928, 809, 289]. continuous-time [413, 852]. contracts [1321]. Contribution [1265, 878]. Control [890, 877, 1417, 577, 331, 25, 5, 24, 500, 50, 360, 562, 714, 1080, 475, 1198, 732, 1359, 640, 1222, 465, 1360, 342, 723, 869, 915, 296, 629, 297, 299, 904, 1155, 278, 720, 634, 889, 637, 1220, 347, 1296, 906, 931, 718, 1156, 3, 870, 1096, 312, 641, 636, 1285, 1178, 632, 885, 472, 250, 868, 1191, 301, 639, 1195, 357, 187]. controllability [1143, 359]. controlled [1031, 1148, 406, 698, 707]. controller [711, 881, 876, 916, 1202, 879, 1218]. controllers [1330, 633, 631, 628, 861]. convection [1292]. conventional [712]. Convergence [1358, 1087, 1057, 1375, 991]. Convergency [1255]. convergent [1241]. converter [999, 1202, 712, 1187]. converter/induction [1187]. converters [1205, 706]. convex [542, 1182]. convolution [559]. Convolutional [260]. cooling [1021]. Cooper [95]. Cooperated [899]. cooperative [602]. Coordinating [242, 144]. copies [200]. Coron [364]. Corporation [349]. correction [441, 758, 777]. corrections [762]. corrector [680, 482]. correlated [439, 1246, 755, 1079]. Correlation [946, 1128, 1319]. cortical [818]. cost [1351, 463, 1360, 1155, 1098]. cost-optimization [1360]. costs [691]. Cottle [393]. cotton [316]. counterexamples [227]. counting [987, 1289]. coupled [1387, 833, 64]. Coupling [1039, 1016]. Cournot [1414]. course [257, 252, 212]. cover [389]. crack [1393]. CRC [389, 283]. crest [843]. crime [777]. Criteria [1052, 1148]. critical [529, 1123]. Criticality [1246]. critique [771, 770, 1065]. CRONE [886]. crop [312, 316]. cropping [315, 321]. cross [462]. cross-validation [462]. crossings [4]. cryptography [514, 207]. CT [200]. cube [481]. cubic [1017]. cultivated [314]. cultivations [634]. cultural [313]. cultures [945]. curiosity [229]. currency [1056, 1321]. current [712, 1224]. currents [1203, 834]. curve [91]. curved [1399, 1382]. curves [198, 1123, 1006, 1323, 988, 1005, 71, 485, 127, 196]. cyanobacteria [1051]. cycle [1324, 736, 757]. cyclical [1054, 776]. cycloconverter [1194]. cycloconverter-active [1194]. cylinder [1336]. cylindrical [1376].

D [273, 29, 35, 384, 391, 380, 385, 485, 498, 496, 94, 105, 150, 129, 137, 139, 127, 145, 125, 86, 208, 214, 221, 222, 283, 182, 423, 415, 352, 422, 366, 350, 434, 419, 414, 545, 530, 531, 550, 560, 552, 526, 15, 20, 536, 180, 906, 1401, 843, 114]. d.c [1199]. DAE [1231, 801]. DAE-index [801]. DAEs [506]. Dahmen [198]. daily [335, 1048, 323, 1305]. Dalecky [388]. dans [690]. Darling [321]. Darnton [495]. Data [553, 764, 7, 523, 1278, 324, 1307, 332, 325, 949, 69, 1316, 1202, 1416, 939, 1054, 480, 1313, 170, 298, 1028, 1324, 321, 509, 929, 375, 715, 309, 1024, 910, 1392, 157]. database [80, 246, 895]. databases [300]. Davey [660]. Davydov [270, 836]. Dawson [204]. Day [93]. dc [870]. deal [308]. December [662, 671, 825, 1120, 1294, 1409, 1418, 174, 303, 469, 566]. Decision [576, 264, 385, 521, 1097, 580, 749, 307, 320]. decision-making [749]. decomposition [1104, 1215, 609, 610, 1115, 453]. deconvolution [1390]. Decoupled [1191, 927]. Decoupled-control [1191]. Decoupling [1144, 1201]. decouplings [276]. definition [722]. deflections [1374]. degradation [1297]. degraded [747]. degree [904]. Dekker [98, 99, 120, 117, 209, 226, 16, 17, 14]. Delanghe [450]. Delay [1145, 455, 1088, 1148, 1142, 1152, 1155, 1153, 1157, 1146, 1144, 522]. Delay-dependent [1145]. Delay-differential [1144]. delayed [1113]. delays [668, 1151, 1147, 1154, 623]. Delfour [555]. delineation [1043]. delta [328]. Demand [1064, 758, 1071, 764, 770]. DEMANIR [1035]. Demon [406]. Denmark [242, 242]. densities [693]. density [634, 843]. dependence [1044]. dependent [1145]. Derivation [328, 1159]. derivative [1273]. derivatives [1006, 996, 352]. DERIVE [1140]. describe [1282]. describing [277]. description [167]. descriptor [790]. descriptors [611, 305]. Design [919, 1177, 177, 299, 31, 26, 497, 196, 203, 415, 434, 292, 760, 795, 1222, 874, 296, 407, 620, 170, 297, 1220, 1296, 879, 1106, 168, 405, 1335, 719, 300, 1223, 1392, 250, 409, 1218, 143]. DESIGNBASE [33]. Designing [50, 80, 861, 245]. desktop [497, 203]. detecting [1318]. detection [782, 629]. detection-isolation [629]. determinant [990]. determinants [431]. determination [738, 843]. determine [1165]. Deterministic [518, 19]. deterrence [777]. deterrent [330]. Developing [1360, 1081]. Development [634, 1035, 1248, 294, 767, 786, 745, 312, 602]. Developments [866, 201, 1343, 76, 162]. développement [130]. device [142, 1194]. devices [50, 1356, 989, 296, 705, 818]. DF [277]. Dfl [5, 9, 32, 24, 4, 35, 7, 36, 492, 136, 150, 138, 149, 151, 370, 431]. Dfl.100 [557]. Dfl.110 [397]. Dfl.116 [282]. Dfl.120 [271, 493, 447]. Dfl.124 [353]. Dfl.130 [427]. Dfl.135 [456]. Dfl.140 [448]. Dfl.150 [260, 190, 285, 422]. Dfl.159 [193]. Dfl.160 [384, 430, 423, 369]. Dfl.165 [559]. Dfl.170 [356]. Dfl.175 [396, 187, 429]. Dfl.180 [452, 523]. Dfl.185 [390, 496]. Dfl.190 [284]. Dfl.195 [382, 201, 359, 433]. Dfl.200 [191]. Dfl.210 [357]. Dfl.215 [364]. Dfl.220 [254, 272, 388, 185, 360, 424, 417]. Dfl.230 [281, 558]. Dfl.235 [451]. Dfl.240 [381, 188, 416, 368]. Dfl.245 [529]. Dfl.250 [270]. Dfl.255 [555]. Dfl.260 [261, 256, 195, 418, 421, 362, 428, 521]. Dfl.265 [198]. Dfl.270 [371]. Dfl.275 [413]. Dfl.280 [352]. Dfl.295 [258, 383, 182, 450, 455]. Dfl.300 [253]. Dfl.320 [379]. Dfl.325 [264]. Dfl.360 [489]. Dfl.425 [560]. Dfl.460 [395]. diagnosis [859]. diagnostic [577]. diagonalization [1271]. Diagrammatic [408]. diatomic [1345]. dichotomy [1352]. dictionary [94, 346]. Diego [22, 29, 499, 497, 501, 485, 92, 91, 139, 141, 87, 89, 86, 203, 415, 367, 425]. diesel [624]. diet [951]. Dieudonné [556]. difference [111, 543, 1017, 1018, 96]. Differences [377]. different [1205, 894, 1217]. differential [254, 981, 1127, 391, 388, 492, 96, 186, 199, 429, 456, 356, 427, 367, 455, 529, 545, 558, 546, 522, 526, 1150, 676, 974, 968, 463, 674, 1144, 571, 1168, 1013, 679, 1066, 1306, 441, 1122, 1158, 789, 1272, 1337, 686, 967, 1087, 443, 84, 993, 1171, 543, 1146, 501, 354, 357, 533, 534]. differential-algebraic [1122]. differential-difference [543]. differentialgleichungen [491]. differentiation [997]. différentielle [678]. diffusion [365, 464, 1278, 1240, 1353, 440, 1292, 684, 621, 697, 63, 586, 690]. diffusions [1410, 683, 675]. Diffusive [736]. Digital [495, 181, 446, 432, 434, 419, 709, 1202, 708, 879, 432]. dimension [1052, 837, 1311]. Dimensional [1140, 1092, 388, 500, 987, 1108, 677, 158, 325, 1403, 1353, 1244, 1405, 1052, 873, 687, 1331, 1411, 1015, 110, 656, 621, 830, 1192, 1335, 1262, 585, 340, 917, 584, 428]. dimensions [1312, 831, 665, 426]. Dimovski [260]. Ding [91]. Ding-Yuan [91]. dioxide [339]. Dipolar [830]. dipoles [844]. Dirac [450, 368, 328]. Direct [1407, 870, 1417, 1112, 624, 840]. direct-injection [624]. directory [181]. disciplinary [1124]. discipline [229]. discontinuities [717]. discontinuous [571, 1310, 1390]. discrepancy [1257]. Discrete [653, 29, 408, 1084, 791, 651, 1023, 920, 1329, 1315, 694, 472, 383, 286, 360]. discrete-event [791]. discrete-time [1315]. Discretization [684, 484, 993]. disease [956]. Disk [415, 252, 361, 1157, 361]. disks [550]. dispensing [905]. dispersion [1039, 1259]. dispersive [647]. displacement [1366]. Dissipation [319]. dissipative [647, 1401]. distance [689, 800]. distinguishability [937]. Distributed [1178, 402, 608, 1340, 1341, 1122, 540, 512, 636, 300]. distribution [254, 729, 1307, 335, 755, 956, 1128, 337, 622, 740, 539, 309, 47]. distributions [70, 65, 324]. District [342]. disturbance [888]. Dita [147]. divisions [954]. Dixon [1344]. DM [115, 133, 122, 121, 96, 97, 119, 123, 142, 128, 10, 19, 21]. DM.164 [365]. DM.78 [355]. DM118 [220]. DM128 [228]. DM128.00 [213]. DM228 [216]. DM37 [225]. DM84 [227]. DM94 [386]. DM98 [215, 222, 192]. DNA [837]. docking [738]. Does [1139]. doing [1291]. dollar [1351]. dollar-cost [1351]. Domain [609, 1026, 1104, 1405, 610, 1406, 1315, 717, 1393]. domains [774, 1145, 544, 1382, 348, 1149, 584, 1107]. DOMINO [596]. Doppler [1029]. Doran [16]. Dordrecht [254, 272, 260, 264, 261, 253, 258, 256, 270, 271, 9, 32, 24, 4, 35, 7, 36, 381, 384, 383, 396, 388, 390, 395, 379, 397, 489, 490, 496, 492, 493, 136, 150, 127, 138, 149, 151, 217, 187, 198, 188, 193, 282, 281, 195, 185, 201, 190, 191, 284, 199, 285, 429, 430, 456, 356, 359, 423, 450, 357, 360, 418, 421, 427, 451, 352, 413, 424, 452, 422, 371, 447, 364, 362, 416, 433, 368, 369, 370, 455, 417, 428, 431, 448, 529, 558, 560, 523, 521, 557, 555, 559]. Dordrecht [20]. double [1200]. Dover [487, 486, 488, 457, 449]. Downs [321]. drawing [1005]. DREAM [877]. DREAM-1 [877]. dressing [658]. drift [365, 1000]. drilling [250]. drills [1204]. drive [1201, 1187, 1195]. driven [650, 683]. drives [1199, 711, 1198, 723, 706]. dromions [660]. droughts [1037]. drying [947, 948]. DSMC [1289]. dual [1297]. Duality [1089, 396, 1389]. Dubinskii [429]. Dudewicz [252]. Duffie [25]. dune [52]. duopoly [1123]. dust [1221]. Dutre [379, 285]. Dynamic [1144, 986, 59, 889, 269, 1414, 202, 328, 332, 908, 1227, 1300, 1026, 876, 897, 966, 67, 327, 996, 894, 305, 915, 299, 1252, 1155, 1254, 777, 320, 786, 1217, 1317, 870, 1214, 938, 769, 1352, 1393]. Dynamical [995, 36, 215, 359, 1091, 635, 842, 1209, 1315, 607]. dynamics [262, 642, 911, 439, 838, 805, 1215, 1050, 775, 807, 792, 798, 747, 317, 1116, 1136, 1266, 844, 828, 436, 455, 522].

each [22]. earth [735, 1387]. earthing [1392]. Earthquake [1383]. East [260, 32, 24, 35, 151, 251]. Eastern [135, 321, 323]. easy [1210]. Ebbinghaus [222]. EC [1027]. ecological [75, 79]. Econometric [1056, 769, 772]. econometrics [330, 764, 1303, 16]. Economic [339, 795, 774, 773, 342, 756, 1055, 749, 767, 777, 60, 1053, 1068, 1065]. economic-biophysical [749]. economy [773]. Ed [7, 36, 14, 267, 272, 264, 262, 381, 391, 520, 115, 133, 122, 95, 144, 150, 209, 213, 217, 196, 210, 183, 219, 201, 205, 207, 287, 186, 199, 430, 446, 436, 354, 432, 418, 422, 362, 434, 414, 428, 561, 521, 525, 549, 555, 556]. Edition [91, 270, 27, 31, 30, 457]. Editor [152, 363, 699]. Editor-in-Chief [363]. Editorial [1126, 1161, 1186, 1207, 1233, 1276, 1302, 1339, 1364, 1395, 865, 936, 850, 41, 175, 345, 470, 567, 617, 702, 728, 813, 864, 926, 1010, 1078]. Eds [22, 5, 86, 10, 253, 392, 394, 499, 126, 121, 147, 137, 116, 145, 124, 143, 151, 100, 206, 198, 208, 225, 229, 230, 214, 190, 191, 423, 413, 424, 364, 416, 433, 435, 529, 531, 532, 534]. Effect [838, 713, 1269, 894, 1030, 1381, 330, 1016]. effective [1154, 1270, 698]. effector [915]. Effects [650, 70, 752, 1201, 314, 51, 1253, 1224, 871]. efficiency [773, 1249]. Efficient [1236, 674, 988, 1122, 1230]. efficiently [583]. effort [1222]. eigenproblem [1018]. eigenvalue [598]. eigenvalues [254]. Eisenstadt [397]. elastic [1092, 916, 278, 1258, 1261, 641, 1265]. elastic-joint [641]. elasticity [1378, 1385]. Elden [189]. electric [801]. electrical [206, 447, 724, 109, 927, 1137]. electrochemistry [374]. Electrohydrodynamic [588]. Electromagnetic [1199, 1189, 1405, 1192, 418]. electromagnetics [390]. electromechanical [1204]. electron [1248, 1256, 1258, 1265]. electrotechnical [710]. electrotherapy [804]. element [1348, 139, 211, 230, 1372, 1375, 1376, 1368, 1369, 664, 1297, 67, 603, 739, 920, 606, 596, 1279, 1192, 666]. elementary [178]. Elements [377, 206, 505, 421, 11]. Elimination [972, 583, 974, 973, 994, 878]. ellipsoidal [162]. ellipsometry [165]. elliptic [1369, 1104, 930, 6]. elliptiques [682]. Ellis [101, 102, 137, 134]. Elsevier [126, 229, 182]. Embedded [517]. Embedding [307]. emergence [847]. EMGP3 [379]. emission [1038]. emissions [339]. Empirical [729, 1355, 771, 1325, 1065]. EMTP [712]. encode [871]. end [561, 760, 915, 1192, 595]. end-effector [915]. end-use [760]. end-windings [1192]. Endogenous [771]. endplate [1224]. Energy [278, 509, 1285, 949, 66, 650, 1228, 918, 1039, 799, 1188, 839, 827, 800, 844]. Energy-based [509]. engineer [1217]. engineering [24, 126, 433, 435, 524, 604, 9, 206, 371]. Engineers [380, 389, 129, 132, 134, 283]. engines [624]. Englewood [385]. Enhanced [786]. Enterprise [495, 293]. environment [269, 608, 897, 53, 540, 601, 1098, 1346, 376, 665, 895]. environmental [273, 324, 51, 782, 72]. environments [739, 786, 910, 602]. Envisioning [200]. enzyme [1282]. Epidemic [43, 1147]. Equation [678, 108, 646, 1012, 365, 1398, 991, 1403, 688, 667, 649, 822, 652, 655, 1333, 1015, 657, 762, 993, 1401, 63, 1391, 1261, 1262, 585, 766, 1016, 85, 543, 659, 1017, 696, 1402, 370]. equational [355]. Equations [377, 981, 1127, 9, 388, 394, 96, 186, 199, 354, 367, 455, 529, 522, 533, 526, 1150, 676, 328, 647, 674, 534, 571, 109, 832, 569, 970, 1023, 1399, 1013, 679, 1066, 1306, 763, 1292, 971, 1158, 789, 1272, 1337, 686, 1014, 967, 1087, 443, 681, 1387, 84, 931, 1114, 1171, 963, 584, 962, 928, 474, 682, 391, 492, 96, 192, 456, 356, 457, 545, 558, 546, 559]. equilibrium [1183, 835, 1136]. equivalence [427, 544]. eradication [318]. Erlbaum [536]. eroded [1297]. eroding [749]. erosion [315]. Erratum [1293, 1300, 1117]. Error [1243, 471, 1358, 676, 167, 19, 161, 1177, 480, 170, 1371, 758, 777, 375, 1081, 460, 168, 1389, 1391, 165, 157, 1106]. error-correction [777]. errors [729, 755, 1079]. Establishing [820]. estimate [1307, 314]. estimated [1278]. Estimates [1242, 1322, 333, 1106, 1391]. Estimating [1309, 1066, 1306, 754, 70, 310]. Estimation [282, 332, 693, 1079, 715, 759, 1067, 1296, 740, 1321, 157, 471, 65, 1358, 1237, 73, 328, 161, 782, 689, 68, 480, 164, 71, 1381, 110, 540, 622, 162, 1389, 938, 769, 1024, 1311, 1295, 99]. estimator [337]. estimators [120, 197, 755, 1030, 1371, 1060, 909, 766]. Euclidean [548]. Euler [642, 676, 686, 1087, 276]. Eulerian [1259]. Europe [191, 45]. European [260, 32, 24, 35, 151, 362, 416, 903]. EURSOL [285]. evaluate [277, 321]. evaluating [729]. Evaluation [1003, 1282, 310, 3, 1166, 1349, 986, 1001, 176, 434]. evaporation [64]. event [1237, 408, 1084, 791]. events [81, 154, 239, 240, 1076, 1093, 732, 1037]. everywhere [561]. Evidence [1024, 1318, 777, 1314, 1319]. evisceration [917]. evolution [105, 970, 5]. evolutionary [188, 201]. Exact [1396, 1081, 160, 108, 1013, 755, 1015, 159]. example [425, 1080, 876, 527]. Examples [966, 108, 350, 1159]. exchange [335, 1021]. excitation [803]. Excited [1238, 1256]. exciter [1204]. exclusive [774]. existence [1317]. expectations [766]. expectative [1101]. Experience [792, 905, 1384, 126]. experiences [349]. Experiment [168]. experimental [1278, 949, 1050, 1223, 1195]. experiments [760, 404, 407]. Expert [593, 597, 606, 772, 312, 585, 114, 910]. experts [939]. explained [224].

Explanation [943]. explanatory [575, 766]. Explicit [441, 621, 111]. explicites [690]. Explorations [532]. Exploring [90]. Exponential [1152, 674]. exposure [1248]. expression [1131]. Extending [731, 723]. extension [75]. Exterior [427]. external [1278]. externality [773]. extinctions [306]. extrapolation [648]. Extreme [1037, 324]. extremes [729].

F [255, 384, 394, 497, 450, 364, 370, 554, 561, 524, 548, 534]. Faà [996]. fabrics [918]. Fabrikant [371]. factor [1194, 719]. factorisation [968]. factorizations [1367]. factors [945]. FAEST [1081]. Falconer [218]. families [970, 1382]. Farebrother [98]. Farin [196]. farm [314]. farmed [957]. Farmer [273]. Fast [159, 867, 1081]. Fault [629, 1192]. feasible [160]. features [632]. February [1160, 40, 478, 1033, 1301]. fed [999, 1099, 1100]. fed-batch [1099, 1100]. feed [867, 909]. feed-forward [867]. Feedback [876, 765, 562, 1148, 1152, 1155, 1220, 909, 1144]. feedback-controlled [1148]. fermentation [1099, 1355, 946]. Fermi [485]. fetch [53]. Feyman [568]. FF [130]. Fibonacci [102]. Field [23, 258, 1269, 195, 1129, 68, 1381, 985, 310, 55, 718, 1193, 1335, 1016, 1195]. fields [1398, 1199, 439, 965, 1105]. fifth [362]. Figueredo [562]. film [1349]. films [1248]. filter [720, 734, 1194, 928]. filtering [1200, 910]. filters [751, 1221]. final [806]. finance [778]. financial [779, 776, 338]. find [1210]. Finite [1348, 110, 377, 1018, 1372, 1154, 1376, 1369, 664, 67, 603, 739, 1384, 606, 596, 505, 1279, 111, 1081, 1192, 1017, 139, 211, 230]. finite-difference [111, 1017]. Finite-dimensional [110]. Fink [352, 560]. First [233, 232, 252, 432]. fish [957]. fitted [482]. fitting [729]. Fiume [89]. five [432]. Fixed [1098, 635, 776]. fixed-term [776]. Fixed-viewpoint [1098]. FL [389]. FLAC [296]. flame [476, 346, 586]. Flamig [148]. Flannery [520]. flawed [1316]. flaws [1400]. flexibility [630]. flexible [911, 664, 915, 920, 887, 1352, 301, 665]. flight [842, 410]. Floating [982]. floods [1037]. flow [180, 1021, 52, 611, 1105, 1050, 1137, 1096, 1229, 998, 1107, 137]. flow-based [180]. flows [1412, 1108, 830, 1234, 139]. fluctuations [953]. fluid [262, 436, 1021, 949, 1109, 828]. fluids [1108]. fluorescence [950, 1277]. flux [714]. fly [318, 317]. FOC [1193]. focusing [1016]. Fokker [569]. folds [1224]. following [804, 906]. Fomin [388, 360]. food [955]. foods [954]. forces [1192]. forcing [443]. forecast [757]. forecasts [1323]. forest [60]. Foreword [1103, 826, 645, 1011, 156]. form [1164, 1006, 912, 295, 1090]. formal [971, 1171]. formant [858]. formation [841, 53]. forming [654]. forms [1162, 1184]. formula [996]. formulae [441]. formulas [29]. formulation [67, 913, 951]. formulations [544]. Formules [690]. Fortran [520, 129, 224]. forty [1413]. forward [867, 909]. foundation [358, 26]. four [1299, 1085]. Fourier [253]. fourth [416, 704]. Fractal [1312, 497, 203, 415, 57, 1313, 1311, 218, 415, 554]. fractal-dimension [1311]. Fractals [561]. Fractional [374, 1309, 1159]. fractions [133, 17]. frame [707, 1252]. framework [1340, 293, 75]. Francois [289]. free [1269, 287, 555, 329, 1110]. freedom [904]. Freeman [252, 263]. frequency [421, 1404, 999, 170, 715, 1406, 1016]. frequency-domain [1406]. friction [1368]. Friedhoff [263]. front [595, 948]. front-end [595]. frontiers [422]. fronts [369]. Fuchsian [492]. Fuchssteiner [553]. Fuglede [456]. full [337]. Fumigation [48]. function [225, 226, 450, 426, 448, 277, 328, 1242, 874, 177, 815, 1315, 112, 1298, 1402, 1090, 120]. function-generating [177]. functional [267, 135, 1150, 15, 978, 675, 1158, 1071, 750, 110, 1117, 23]. Functionals [1087]. Functions [1165, 271, 402, 381, 395, 493, 450, 352, 463, 443, 178, 84, 835, 386, 131, 220, 490]. fundamental [1002, 54]. fundamentals [350, 188]. Further [508]. futures [1056, 1321]. Fuzzy [1193, 1417, 190, 1330, 890, 888, 633, 513, 628, 1211, 887, 855, 878, 9, 193, 523]. fuzzy-logic [628]. fuzzy-neural [888]. fuzzy-process-model [1417].

G [273, 256, 25, 495, 491, 493, 95, 129, 116, 138, 187, 196, 230, 183, 194, 353, 413, 358, 367, 529, 553, 562, 551, 535, 559, 11]. game [1414, 1123]. Gamkrelidze [115, 213]. gamma [120, 65]. Gap [834]. Garbey [529]. GARCH [1319]. Garrett [234]. Gasca [198]. gases [1289, 47]. Gaussian [208, 73]. GB [88]. GDP [1057]. Gelbaum [227]. General [1202, 706, 881, 1370, 336, 1081, 1182, 887, 1107, 928, 127]. generalised [1067]. generalization [426]. generalized [1112, 640, 1058, 1023, 655, 462, 734, 721, 474]. generated [766]. generating [177]. Generatingfunctionology [141]. generation [1112, 1224, 1081, 279, 405, 1174, 1182, 900, 1063, 1090]. Generative [494]. generator [109, 1085, 1192, 1335]. generators [1210]. Generic [985, 77, 76]. Genetic [1330, 878, 1411, 1355, 1075, 619]. genetics [188]. Geng [91]. Geng-Zhe [91]. geodynamic [1387]. geographic [78]. Geometric [1298, 35, 143, 196, 895]. Geometry [976, 485, 91, 90, 983, 1133, 218, 396, 20]. Georg [228]. Georgescu [147]. geoscientific [428]. Geostatistical [1313]. Germany [1043, 1042]. Giacoletto [495]. Gibbons [34]. Gieseker [485]. Gillespie [366]. Gini [768]. Girard [8]. Girko [431]. GIS [1043, 1042]. given [1130, 200].

Glasserman [282]. global

[417, 1037, 1387, 1290, 340]. GLS [755]. gobal [47]. golden [102]. Goldfeld [29]. Gol’dshte [261]. Goldstein [456]. Golub [367, 535]. Gomez [10]. Gonin [99]. Good [1210]. goods [59, 1064]. Gopalsamy [455]. Goppa [20]. Gordon [1386]. Gottheil [553]. governed [931]. GPSS [252]. GPSS/PC [252]. gradient [463, 1030, 620, 282]. grain [313]. Gram [1292]. graph [1007, 787, 790, 1096, 288]. Graph-based [288]. graphical [1135]. Graphics [200, 269, 494, 89]. Graphs [1005, 1342, 998]. grasping [920]. gravitation [1004]. gravitational [1384]. Gray [841]. grazing [1151, 746]. Greater [1320]. green [1051, 1242]. greenhouse [1285]. grey [949]. grid [180, 1373, 411, 1112, 1391, 1174, 1182, 1392]. grids [1104, 611]. Griffel [15]. Griffiths [283]. grinding [24]. Gröbner [982]. Groen [229]. Grossetie [423]. Grossman [491]. Groundwater [342, 1046, 1253]. grouping [1067]. groups [383, 395, 454, 977, 29]. growth [771, 753, 71, 767, 804, 943, 1223, 1332, 844, 490]. Gruber [197]. GTO [720]. Guaranteed [480, 1155]. guidance [891]. guide [196, 425]. guidelines [799]. gum [944]. Gunzburger [139].

H [260, 6, 30, 33, 36, 485, 491, 520, 488, 141, 99, 205, 207, 222, 190, 197, 227, 353, 416, 433, 367, 426, 435, 529, 523, 553, 559, 15, 16, 12]. H. [528]. H.-D [222]. H.-G [491]. H.-O [36]. Haar [1113, 1415]. Habitat [308]. half [1403]. half-space [1403]. Hall [385]. Halsted [137]. Hamiltonian [1184, 650, 827, 1171]. Hamiltonians [658]. Hammesfahr [385]. hand [899]. hand-robots [899]. Handbook [203, 230, 521, 391, 498, 116, 219]. handling [919]. Hankel [103]. Hansch [365]. hard [389, 744]. hardback [90]. hardware [880, 879]. Harlow [211]. harmonic [1203, 1030]. harmonics [1201]. harmonique [130]. Harrell [534]. Harrison [23]. Harvard [1]. harvest [899]. harvesting [957]. Harvey [255, 532]. Hausdorff [281]. Hausmann [456]. Hayes [145, 214]. HB [186, 191]. heat [1021, 411, 624, 505, 179, 64]. Heidrich [423]. height [1068]. Heiles [1164]. Heilio [362]. Helein [364]. Helmholtz [1401]. help [1139]. Henon [1164]. Herod [534]. Hessian [463]. Heterogeneity [1277, 739]. heterogeneous [1082, 1331]. heteroscedasticity [336]. Heyman [182]. Hierarchical [637, 898, 74, 1096, 847, 1180]. hierarchy [293]. high [1412, 1341, 1205, 999, 1248, 712, 687, 634, 113, 710, 1346, 1178, 1352]. high-dimensional [687]. high-frequency [999]. high-performance [1341, 1346]. high-speed [113, 1352]. higher [426, 1105, 1270]. higher-order [1105]. highly [1216, 1241]. Hilbert [451]. Hilger [24]. Hill [380, 749]. Hillsdale [536]. hip [1388]. historique [130]. history [118, 183, 774, 123, 353]. Hodge [1115]. hodographic [590]. Hoffman [380]. Holland [382, 116, 208, 230, 353, 530, 531, 524]. Homoclinic [646]. Homogeneity [970]. homogeneous [1168]. homogenization [1397, 138]. homological [979]. homotopy [977]. Honor [86, 234]. honouring [325]. Hopkins [28]. Hordijk [191]. Horn [30, 288]. Horuzhy [195]. Horwood [101, 102, 137, 134]. hot [1280]. household [1069]. Housing [334, 769]. hovering [842]. HP48SX [524]. HRUs [1042]. Hu [493]. Hulanicki [151]. human [1253, 843]. Humi [546]. Hurd [554]. Hurst [1309]. Hurt [4]. Hybrid [867, 1208, 376]. hybridoma [945]. Hydro [713]. hydrochemical [1043]. hydrodynamic [1047]. hydrodynamics [1331]. hydrogenated [1356]. hydrogeological [342]. hydrologic [305]. hydrological [1043, 44, 1042]. hyperbolic [1403, 1370]. hyperelastic [1183]. hyperfunction [489]. HYPERTEXT [392]. hypothesis [761, 1324]. hysteresis [1196].

IBM [497, 361]. ideals [987, 1169].

Identifiability [83, 937]. Identification [167, 860, 1050, 1400, 1219, 1216, 587, 1281, 1286, 441, 675, 169, 1223, 1218, 945, 387, 413]. Identifying [1284]. identity [990, 1024]. Iglehart [86]. II [385, 151, 208, 534, 1277, 1135, 213, 454, 21]. III [151, 208, 133, 424]. iliac [843]. IMACS [235, 237, 704, 81, 154, 37, 106, 153, 171, 242, 239, 240, 238, 231, 236, 290, 302, 343, 399, 437, 458, 466, 372, 477, 503, 510, 515, 537, 563, 572, 581, 591, 614, 625, 643, 661, 669, 701, 725, 741, 783, 810, 824, 848, 862, 882, 902, 924, 934, 958, 1008, 1076, 1093, 1019, 1032, 1119, 1034, 743]. image [4, 554, 898, 872, 597, 618, 892, 452]. imaginary [1184]. imaging [820]. Imai [489]. immersed [1369]. impact [1002, 774]. impacts [1043]. impatient [398]. imperfectly [949, 1148]. implement [1]. Implementation [1367, 880, 413, 419, 1385, 855]. implications [778]. Implicit [1353, 680]. importance [1083, 176]. imprecision [1284]. improve [909]. improved [855]. improvement [471, 1252]. Improving [1375]. impulse [347]. impulses [347, 803]. IMSOR [242]. inactivation [1282]. included [550]. including [252]. inclusion [119]. inclusions [357]. income [1324]. incompressible [139]. Increasing [1249]. incremental [542]. increments [284]. indefinite [204]. independent [284, 1138]. Index [1172, 1225, 1274, 1327, 1362, 1408, 38, 172, 467, 564, 615, 670, 726, 811, 863, 925, 1009, 1077, 768, 1318, 801]. indicators [757]. indices [1143]. indirect [718, 1193]. indistinguishability [83]. individual [242]. induced [1040]. Induction [309, 1198, 1201, 1197, 707, 716, 1187, 1190, 718, 1193]. inductive [150, 1188]. industrial [272, 137, 217, 433, 792, 580, 1124, 1098, 1325, 910, 1283]. industry [663, 1325, 362, 416]. inequalities [496, 560, 460, 35, 488, 352]. Inference [579, 887]. infinite [388, 500, 1151]. infinite-dimensional [388]. inflation [334]. influence [1381]. influenced [1147]. Information [1201, 247, 193, 428, 78, 293, 299, 1025, 337, 1022, 200, 495]. informative [309]. inhomogeneous [1403]. initial [649, 280, 1030, 517, 445]. initial-boundary [649]. initial-value [280, 517]. injection [624]. inner [909]. Input [908, 1154]. Input/output [908]. inputs [1000]. insect [318]. insecticide [1035]. inspection [892]. Inst [529, 555]. Instabilities [424]. instability [650, 831, 778]. instability-driven [650]. installation [1190]. instantaneous [1195]. Institute [363]. instruments [532, 776]. insurance [247]. integrability [970, 827]. Integrable [651, 650]. integral [108, 457, 559, 1110, 496]. integrals [352, 673, 1247, 92]. Integrated [749, 405, 332, 1061, 755, 1062, 55, 301]. Integrating [978, 1117, 600]. integration [548, 692, 751, 1228, 755, 1243, 687, 1055, 750, 1386, 681, 754, 498]. integrations [1235]. integrators [651, 1136]. integro [443]. integro-differential [443]. intelligence [1132]. Intelligent [605, 897, 607, 594, 601, 903, 521]. intensive [1298]. inter [1124]. inter-disciplinary [1124]. interaction [1228, 653, 1383, 1384]. Interactionist [1101]. Interactionist-expectative [1101]. interactions [1121, 1002, 822, 777, 1110]. Interactive [1337, 379, 285, 1097, 177]. Interconnected [668]. Interest [780, 689, 1063, 1065]. interface [1369, 613]. interfaces [144, 829]. interference [1199, 442]. interior [1195]. interlocking [1138]. intermittency [1123]. intermittent [1234]. internal [1050]. International [28, 593, 1071, 778]. interplay [576]. Interpolated [587]. interpolation [69, 1028, 929, 472]. interpretation [857, 818, 1392]. intersection [158]. interval [480, 929, 733, 363]. intra [335, 1324]. intra-daily [335]. intra-year [1324]. Introduction [31, 42, 672, 97, 103, 104, 366, 367, 535, 185, 189, 87, 195, 215, 536]. Introductory [1159]. intrusion [49, 1253]. Intuition [823]. Intuition-based [823]. invariance [147]. invariant [857]. invariants [1342]. Inventory [264]. Inverse [1417, 911, 1402, 1407, 1405, 1406, 657, 1401, 179, 1352]. inversion [541]. inverter [999]. investigating [961]. Investigation [750, 746, 1104]. investigations [588]. investment [1351]. Involution [961]. Involutive [1169, 1170]. involved [1298]. involving [352, 1402]. Iokhvidov [204]. iron [1074]. irregular [127, 1054]. irregularly [1313]. ISBN [254, 267, 272, 257, 260, 273, 264, 269, 252, 265, 262, 268, 259, 261, 253, 258, 255, 256, 270, 266, 263, 271, 22, 27, 25, 31, 34, 5, 6, 9, 23, 30, 32, 24, 28, 29, 4, 8, 26, 35, 33, 7, 36, 381, 384, 386, 383, 396, 391, 393, 388, 392, 380, 394, 389, 390, 385, 382, 397, 398, 387, 487, 489, 499, 500, 497, 501, 490, 485, 494, 498, 495, 496, 486, 492, 491, 520, 488, 502, 493, 88, 92, 101, 115, 126, 133, 136, 152, 122, 91, 94, 95, 105, 121, 144, 147, 148, 150, 96, 90, 102, 129, 137, 139]. ISBN [127, 93, 116, 118, 138, 87, 97, 103, 104, 119, 145, 98, 130, 89, 123, 124, 131, 132, 134, 140, 142, 143, 99, 149, 125, 86, 120, 128, 117, 151, 135, 100, 209, 213, 217, 206, 187, 196, 198, 208, 210, 225, 289, 203, 229, 286, 200, 188, 193, 211, 218, 224, 230, 282, 288, 183, 219, 281, 195, 214, 215, 216, 204, 184, 185, 194, 201, 205, 221, 189, 202, 207, 212, 222, 228, 283, 287, 226, 186, 190, 223, 191, 197, 284]. ISBN [182, 192, 199, 285, 220, 227, 181, 429, 430, 446, 456, 356, 351, 355, 359, 423, 450, 354, 357, 360, 365, 432, 418, 421, 427, 415, 353, 451, 352, 413, 424, 452, 358, 422, 361, 366, 371, 420, 447, 453, 364, 349, 350, 362, 416, 433, 434, 454, 367, 368, 369, 370, 419, 426, 455, 457, 414, 417, 425, 428, 431, 448, 449, 529, 545, 558, 546, 530, 531, 550, 560, 522, 533, 532, 523, 554, 561, 524, 521, 548, 552, 557, 527, 547]. ISBN [553, 525, 562, 526, 549, 551, 535, 555, 556, 559, 10, 15, 16, 17, 19, 11, 20, 12, 14, 13, 18, 21, 534, 536, 395, 141, 146]. ISBN-0-12-594750-X [146]. ISBN-0-12-742687-6 [148]. ISBN-0-12-751955-6 [141]. ISBN-0-7923-1466-2 [395]. ISBN0 [379]. ISBN0-7923-1235-X [379]. island [505]. Isolating [1054]. isolation [629]. Iterated [996, 673]. iterations [805]. Iterative [795, 506, 1380, 1366, 1242, 951, 1261, 1218, 119, 145, 214]. ITI [1102]. IV [124]. IX [100, 17, 20].

J [273, 252, 265, 253, 25, 24, 8, 26, 35, 393, 392, 380, 385, 397, 387, 487, 494, 88, 101, 144, 96, 90, 116, 103, 145, 130, 123, 117, 100, 210, 229, 286, 224, 230, 214, 216, 184, 194, 221, 202, 207, 212, 190, 191, 197, 227, 429, 359, 423, 432, 353, 352, 364, 367, 414, 560, 533, 532, 524, 557, 525, 562, 535, 556, 14, 13, 18, 534, 1277]. J.-M [364]. J.-S [393]. J.-Y [8]. Ja [395]. Jacobi [1264]. Jacoboni [142]. Jacobsen [133]. James [268]. Jameson [992]. Janko [190]. January [1125, 1141, 1020]. Japanese [149, 1273, 1325]. Java [1340]. Jeffries [351]. Job [1111]. Job-shop [1111]. John [6, 88]. joint [916, 1282, 956, 641]. jointly [1278]. Jones [498, 548]. Joseph [216]. Journal [152, 363]. July [1232, 243, 1338]. junction [1269, 1224]. junctional [1224]. June [1226, 107, 582, 883, 1328].

K.-I [386]. Kac [486, 568]. Kahn [392, 132]. Kalman [734, 928]. Kaltofen [121]. Kan [353]. Kaper [529]. Karian [252]. Karlin [86]. Kashiwara [22]. KASTLE [595]. Kawai [22]. KdV [1114, 543]. KdV-like [1114]. KdV-MKdV [543]. Keener [13]. Kehe [226]. Kemper [553]. Kenelly [532]. Kent [129, 545, 546]. Kerr [1406]. Kerr-like [1406]. Kesavan [135]. Keviczky [24]. Kiguradze [558]. Kincaid [145, 214]. kind [823]. Kinematic [914, 901]. Kinetic [945, 944, 1283]. kinetics [1256]. Kirk [550]. Kisielewicz [357]. Klamka [359]. Klatte [350]. Kleijnen [117]. Klimyk [395]. Kluge [553]. Kluwer [254, 272, 260, 264, 261, 253, 258, 256, 270, 271, 5, 9, 32, 24, 4, 35, 7, 36, 381, 384, 383, 396, 388, 390, 395, 379, 397, 489, 490, 496, 492, 493, 136, 150, 127, 138, 149, 151, 217, 187, 198, 188, 193, 282, 281, 195, 185, 201, 190, 191, 284, 199, 285, 429, 430, 456, 356, 359, 423, 450, 357, 360, 418, 421, 427, 451, 352, 413, 424, 452, 422, 371, 447, 364, 362, 416, 433, 368, 369, 370, 455, 417, 428, 431, 448, 529, 558, 560, 523, 521, 557, 555]. Kluwer [559, 20]. Knoerrer [485]. Knowledge [604, 595, 599, 9, 125, 598, 596, 405, 600]. Knowledge-based [604, 598, 596, 600]. known [1148]. Ko [386]. Koch [189]. Kohonen [694]. Kolostori [24]. Korteweg [667, 649, 655, 657, 659]. Kosinski [501]. Kratzel [32]. Kriging [69]. Krizel [211]. KTK [149]. Kuang [522]. Kuiken [433]. Kulisch [502, 350]. Kurzhanski [5]. Kutta [517].

L [269, 253, 24, 379, 499, 490, 88, 133, 116, 145, 89, 131, 143, 86, 120, 128, 230, 214, 189, 228, 191, 285, 181, 446, 456, 355, 436, 431, 435, 530, 533, 554, 17]. laboratory [612, 820]. Lafayette [349]. Lafont [8]. lag [752]. Lagrangian [48, 1259, 340, 474]. lakes [50]. Lakshmikantham [271]. laminar [346, 586]. L’analyse [130]. Land [342, 1045, 321, 749]. lands [749]. landscape [77]. landscapes [747]. Langevin [836]. Langmuir [1121]. language [296, 249, 1180]. Laplace [991]. LaQuey [181]. large [145, 214, 1374, 1288, 294, 698, 750, 1192, 1271]. large-scale [294]. Larson [398]. laser [1377]. LATOR [246]. Lattice [32]. Lau [530]. laws [199, 1370, 3, 1114, 1115, 265]. Layer [1403, 666]. Layer-stripping [1403]. layered [1258, 1397]. LCGs [1085]. leaching [940]. leakages [722]. Learning [939, 1330, 822, 869, 867, 1139, 1101]. least [98]. Lebesgue [548]. Lecture [133, 119, 209, 225, 207, 10, 19]. Lectures [146]. Ledermann [219]. Lee [208, 415, 530, 531]. legacy [95]. Legendre [931]. length [752, 1112]. Lenstra [353]. Lessard [201]. Letter [699, 700]. level [341, 1331, 111, 658, 827, 1178]. Levendorskiĭ [254]. LeVeque [265]. Levine [536]. Levitan [368]. levitation [1191]. Lévy [683]. Lewis [137, 436, 435]. Liapunov [737]. libraries [597]. Library [420, 264, 596, 595, 997, 28, 524]. Lick [96]. lie [395, 674]. life [266, 95, 78, 1324]. life-cycle [1324]. Lifecycle [765, 331]. light [648]. like [1377, 1406, 1114]. likelihood [65, 1309, 759, 337, 306, 1311]. limb [1189]. limit [827, 736]. limitations [880, 941]. limited [53]. limiting [945]. line [1089, 1203, 698, 1130, 928, 922]. Linear [835, 1041, 393, 387, 145, 214, 221, 360, 550, 541, 968, 1026, 731, 1168, 957, 791, 970, 83, 336, 1318, 1159, 1113, 971, 1155, 929, 60, 1079, 1153, 951, 1255, 1317, 84, 1143, 160, 460, 931, 3, 733, 868, 568, 98, 204, 552]. linearization [908]. linearized [1370]. lines [1337, 709, 1344]. link [916, 630, 912, 1156]. linkage [305]. linkages [1090]. Linking [322]. links [664, 1416]. LINPACK [731]. Lions [116, 230]. Liou [435]. Liouville [216, 368]. LISP [384, 269, 30].

LISP-STAT [269]. lithography [1248]. Liu [91]. Livesley [134]. LM [1314, 1191]. load [1137]. load-flow [1137]. local [732, 693, 1353, 246]. locally [1353]. locations [313]. logic [557, 633, 628, 1193, 486, 358, 422]. Logical [301]. Logics [1133]. logistic [1088, 1097]. lognormal [65]. London [207]. long [332, 753, 1059, 756, 800, 1056]. long-distance [800]. long-run [332, 753, 1056]. longer [774]. Longham [211]. loosened [1388]. losses [1201]. low [1282, 712, 1257, 1098, 1016]. low-cost [1098]. low-discrepancy [1257]. lowland [808]. Loxton [207]. Lp [99]. Lp-norm [99]. LSQR [1390]. Lucas [102]. Lugli [142]. lump [1351]. lump-sum [1351]. Lützen [216]. Lyapunov [271]. Lyche [143]. Lyngby [242].

M [257, 273, 262, 261, 263, 271, 22, 34, 24, 26, 392, 397, 487, 497, 494, 486, 488, 121, 137, 139, 119, 146, 149, 100, 198, 203, 211, 224, 185, 202, 212, 283, 190, 197, 199, 227, 456, 357, 421, 352, 452, 364, 350, 362, 367, 368, 370, 363, 529, 546, 560, 533, 554, 561, 547, 549, 535, 555, 559, 16, 17, 14, 534]. MA [393, 392, 394, 398, 500, 494, 498, 495, 502, 125, 286, 183, 184, 223, 446, 432, 358, 361, 453, 434, 454, 419, 414, 545, 546, 550, 522, 533, 532, 554, 561, 524, 548, 552, 527, 547, 525, 562, 526, 549, 551, 535, 556, 391]. MAC [497, 415]. machine [724, 914, 442, 715, 1325, 397]. machinery [1064]. Machines [528, 25, 1200, 721]. Macintosh [497, 203, 415]. Macmillan [257]. macro [361, 1053]. macromolecules [815]. macros [361]. made [1043]. Maeder [223]. magnet [723, 1195]. magnetic [722, 705]. magnetofluid [832]. magnetostatic [1400]. maintenance [710]. major [51]. majorants [1389]. Makhankov [256]. making [575, 580, 749, 576]. Malaysia [1064]. Malaysian [1326].

Malliavin [676]. man [1043, 88]. man-made [1043]. management [80, 44, 314, 1035, 246, 1045, 61, 295, 60, 320, 76, 405, 316, 409]. managerial [576]. Managing [294]. Mancuso [184]. maneuvering [876]. manifolds [646, 501]. manipulation [1288, 888, 641]. manipulator [904, 1156]. manipulators [912, 889]. manual [446]. manufactured [1064]. manufacturing [518, 301, 809]. map [866]. MAPLE [695, 1136, 1135, 1134]. Mapping [956, 390, 587, 897, 177]. mappings [261, 396, 1166, 6]. maps [579]. Marcel [98, 99, 120, 117, 209, 226]. March [1173]. marching [461]. Marcus [488]. Marinov [447]. market [761, 1056]. marketed [90]. Markets [781, 1320, 1319]. Markov [675, 366]. Markovian [1083]. Markowich [192]. Mass [27, 25, 31, 34, 23, 30, 28, 26, 33, 1278, 949, 621]. massively [1340]. Masson [130]. master [216]. Mastering [184]. MATCOM [480]. material [623]. materials [919, 920]. Math [1293, 1300, 1117]. mathematica [223, 533]. Mathematical [4, 207, 1281, 1099, 1379, 804, 1350, 800, 744, 262, 821, 34, 88, 89, 131, 188, 205, 354, 421, 353, 364, 594, 534, 613, 637, 943, 539, 132, 134, 140, 143, 185, 201, 447, 527, 547]. Mathematica(R) [526]. Mathematics [32, 24, 4, 35, 133, 136, 138, 119, 149, 151, 209, 187, 225, 188, 195, 284, 10, 19, 260, 261, 258, 256, 6, 382, 322, 286, 219, 362, 416, 524, 235, 237, 272, 121, 118, 125, 86, 217, 216, 186, 227, 433, 13, 24, 486, 123, 221]. MathLie [1291]. MATLAB(R) [550]. matrices [202, 1184, 1168, 1130]. matrix [1236, 488, 541, 1241, 1292, 1187, 1, 1345, 194]. Matroid [453]. Matrosov [271]. Maxfield [487]. maximising [315]. maximum [65, 675, 759, 337, 1068]. Maxwell [1399]. May [1206, 373, 703, 1086]. McAllester [125]. McCulloch [818]. McCulloch-Pitts [818]. McGraw [380]. McGraw-Hill [380].

MCNP4A [1246]. ME02 [1191]. mean [471, 68, 778]. mean-variance [778]. means [328, 674]. measurements [418, 1400, 1381]. Measures [388]. mechanical [206, 187, 1217, 2, 479]. Mechanics [206, 262, 1366, 1345, 378]. mechanisms [245]. Mechanistic [943, 792]. mechatronic [796, 907, 1212]. media [138, 932, 1406, 1397, 64]. medical [1307, 299]. medicine [821]. medium [1263, 1402]. Meerschaert [547]. meeting [234]. Mellin [492]. melting [837]. Membership [164, 242, 241, 167, 161]. Membership-set [164]. MEMEX [392]. memorization [1252]. memorization-based [1252]. memory [1122, 1059, 854]. memoryless [1152]. Mercier [97]. Merusi [419]. Merzbach [118]. MESFET [1269]. mesh [603]. mesoscale [1040]. Mesterton [34]. Mesterton-Gibbons [34]. metabolism [311]. metallic [1265]. metapopulation [306]. metapopulations [1308]. Metcalf [224]. metering [760]. method [36, 142, 318, 1307, 1108, 1309, 1238, 677, 674, 590, 1241, 1368, 1164, 735, 1023, 988, 1243, 1333, 1337, 698, 1252, 716, 610, 1331, 484, 1, 759, 1067, 570, 1401, 589, 2, 1192, 1289, 63, 112, 482, 346, 666, 1174, 309, 512, 1017, 737, 1311, 928, 474]. method-of-lines [1337]. méthode [682]. Methodology [413, 1081]. Methods [895, 265, 1236, 28, 4, 380, 389, 139, 119, 145, 132, 134, 143, 225, 230, 214, 283, 287, 529, 525, 1231, 1366, 1046, 73, 1027, 1373, 1375, 404, 325, 463, 1281, 506, 1410, 1240, 571, 1353, 66, 1245, 1369, 1246, 1247, 652, 1013, 1066, 1270, 739, 544, 1380, 930, 1158, 687, 1396, 1209, 443, 310, 460, 680, 1229, 517, 1109, 600, 1110, 696, 1018, 568, 65, 268, 97, 228, 367, 131]. metric [204]. Micchelli [198]. mice [804]. microbial [943]. microbiology [955]. microcrystalline [1356]. microscopy [820]. Microsoft [1272]. microtubules [846, 844]. military [1098]. Millar [94]. Miller [546]. millisecond [818]. Minc [488]. Mineola [457, 449]. miniature [1224]. Minimal [1170]. Minimising [315]. Minimum [689]. mining [764]. MINITAB [361]. Miranda [206]. mis [752]. mis-specifying [752]. Misra [26]. missing [1079, 1314]. misspecification [70]. misspecified [1310]. Mitrinovic [35, 352, 560]. mixed [1366, 949, 1157, 371]. mixing [946]. Miyamoto [193]. MKdV [543]. Mlak [451]. MLE [338]. mobile [890, 896, 483, 877, 900]. modal [1397]. mode [916]. Model [904, 1218, 1417, 331, 765, 191, 1147, 724, 1288, 408, 374, 314, 1035, 640, 341, 949, 881, 1091, 635, 664, 1297, 773, 944, 1282, 313, 1360, 336, 846, 1039, 873, 442, 915, 1286, 920, 802, 758, 1329, 720, 307, 1361, 777, 841, 333, 748, 1079, 1387, 348, 1350, 746, 306, 1219, 1138, 1133, 323, 948, 721, 641, 769, 1068, 817, 819, 1319, 941, 1352, 316, 64, 639, 945, 1283, 1, 205]. modèle [690]. Modeling [808, 723, 852, 807, 1188, 853, 1187, 1096, 494, 428, 795, 1097, 109, 947, 1202, 575, 822, 1379, 1177, 788, 1255, 1256, 1355, 800, 1211, 828, 91, 289, 547, 536]. Modelling [954, 57, 1034, 1216, 1280, 51, 1208, 1189, 52, 797, 61, 45, 806, 1381, 1051, 777, 330, 743, 778, 320, 932, 1308, 803, 1124, 1195, 273, 642, 33, 137, 140, 418, 1043, 787, 953, 1288, 1281, 779, 339, 1222, 1196, 66, 1042, 43, 1099, 342, 1215, 78, 77, 56, 1248, 67, 48, 630, 305, 796, 792, 920, 789, 1331, 321, 749, 1209, 75, 889, 637, 804, 1279, 308, 631, 772, 839, 2, 1056, 47, 340, 923, 1110, 744, 917, 794, 1180, 208, 262, 34, 365, 435]. Models [208, 404, 264, 821, 385, 88, 447, 1088, 729, 1046, 167, 332, 1309, 795, 771, 44, 1048, 161, 689, 1061, 1089, 1045, 83, 1052, 58, 730, 763, 829, 1082, 1037, 1197, 1071, 542, 790, 793, 1176, 786, 759, 1067, 1217, 74, 76, 1315, 770, 762, 943, 168, 55, 1258, 112, 176, 1234, 1259, 1096, 710, 1321, 312, 766, 338, 780, 1063, 518, 1223, 937, 623, 878, 1060, 182, 549].

Modern [252, 420]. modes [844]. Modified [176, 1, 659]. Modular [1180, 379, 1116, 1214]. Modulational [831]. modulators [1205]. Moduli [249, 1092]. Moduli-based [249]. molecular [270, 1136, 816]. moments [65, 1307, 835]. Money [99, 758, 764]. monitor [606]. Monte [142, 1031, 1236, 692, 73, 1238, 1239, 1410, 1242, 1241, 1244, 1245, 755, 1066, 1251, 1243, 1252, 687, 1254, 1253, 310, 1255, 1258, 539, 1260, 1261, 1262, 1263, 1235, 1264, 1266, 1265, 338]. morphological [856]. Morrise [553]. Moscow [259]. MOSFET [365]. Moshier [131]. Motion [1000, 656, 683, 378, 639]. Motivation [1209]. motor [714, 711, 1198, 797, 1197, 1187, 1195]. motors [707, 716, 870]. Moving [751]. MPI [1385]. MRI [1372]. MSSA [1034, 743]. MSSA/IMACS [1034, 743]. MTD [933]. Multi [691, 553, 1082, 930, 421, 881, 1189, 957, 970, 838, 1405, 1133, 472]. Multi-agent [1082]. Multi-asset [691]. multi-component [838]. multi-dimensional [1405]. multi-frequency [421]. multi-limb [1189]. multi-model [881]. Multi-parameterized [930]. multi-parametric [970]. multi-period [957]. Multi-Processing [553]. multi-rate [472]. multi-valued [1133]. multiblock [1174]. multibody [376, 997]. Multidimensional [975]. multigrid [1174]. Multilevel [1378]. multilink [664]. Multimedia [1176, 608]. multiobjective [513]. multiple [200, 640, 1247, 59]. multiply [348]. multiprocessor [113]. multiresolution [1397]. multisine [715]. multispecies [1046]. Multistability [1414]. Multistep [1158]. multisummation [971]. Multivariable [640]. multivariate [763, 1130, 124]. MuPAD [1346, 553]. Murdock [268]. Murray [1051]. mushrooms [314]. Muskhelishvili [457]. mutually [774]. mystery [514].

N [25, 4, 383, 395, 95, 138, 360, 413, 454, 457, 449, 940]. N-leaching [940]. NAG [28, 595]. Nakata [152]. NASA [436, 435]. Nash [183]. Nather [523]. NATO [253, 456, 418, 364, 428, 529, 555]. natural [44, 45, 943, 803, 1036]. nature [88, 781, 308, 90]. Naundorf [553]. Navier [584]. Neaga [350]. near [650, 1381, 1165]. near-integrable [650]. Necessary [873]. need [774, 308]. Neittaanmaki [287, 447, 211]. Nematics [364]. Nering [552]. nerve [803]. Nesterov [363]. net [711, 1022]. nets [788, 789, 298, 1124, 1178, 809]. Network [775, 385, 638, 893, 898, 876, 822, 888, 177, 1137, 2, 694, 855, 757]. networks [1092, 860, 891, 908, 1227, 1300, 246, 874, 1416, 1282, 880, 852, 875, 867, 620, 177, 698, 853, 815, 1051, 854, 859, 871, 1355, 1179, 870, 717, 909, 861, 851, 868, 816, 856, 181, 351]. Neumann [202, 418]. Neural [868, 816, 351, 638, 893, 898, 711, 860, 891, 908, 1227, 1300, 876, 874, 1282, 822, 880, 852, 536, 875, 869, 867, 888, 177, 1197, 853, 815, 1051, 854, 879, 859, 871, 1355, 1179, 870, 909, 855, 851, 757, 856]. neural-network [177]. neurofuzzy [887, 1295]. neuromuscular [1224]. neuronal [858]. neurons [818]. neurosciences [821]. neutral [1150, 1143, 1149]. News [37, 106, 153, 171, 231, 236, 290, 302, 343, 399, 437, 458, 466, 477, 503, 510, 515, 537, 563, 572, 581, 591, 614, 625, 643, 661, 669, 701, 725, 741, 783, 810, 824, 848, 862, 882, 902, 924, 934, 958, 1008, 1019, 1032, 1119]. Newton [1293, 642, 36, 1163]. Next [1237]. niche [1361]. niche-model [1361]. Nielsen [144]. nitrogen [807]. nitrous [340]. NJ [385, 536]. no [200, 774]. Node [423]. nodes [1137]. noise [1318, 730, 736]. Nola [9]. nominal [753]. Non [1318, 326, 1136, 73, 724, 1021, 693, 919, 590, 1353, 1376, 1026, 319, 970, 723, 712, 1406, 166, 3, 1263]. non-analog [1263]. non-compliance [319]. non-connected [166]. Non-equilibrium [1136]. non-Gaussian [73]. non-Kerr-like [1406]. Non-linear [1318, 1026, 3]. non-local [1353]. Non-parametric [326]. non-Park-transformable [724]. non-rigid [919]. non-sinewave [723]. non-standard [1376]. non-stationary [693]. non-trivial [590]. non-uniform [1021]. nonautonomous [558]. nonequilibrium [424]. Nonlinear [1215, 839, 271, 646, 1012, 396, 500, 215, 185, 221, 356, 1046, 714, 73, 1080, 647, 677, 1216, 1281, 1407, 465, 1378, 1370, 1145, 652, 1013, 996, 1415, 873, 654, 442, 480, 620, 1406, 697, 513, 179, 1219, 861, 1391, 1016, 85, 636, 1149, 1166, 1146, 937, 969, 394, 99, 149, 562]. Nonlocal [1251]. Nonnegative [202]. nonparametric [71, 1305]. nonrigid [1092]. nonsymmetric [506]. norm [99]. Normal [1162, 1164, 1336]. normality [763]. North [382, 116, 208, 230, 353, 530, 531, 342, 45]. North-Holland [382, 116, 208, 230, 353, 530, 531]. Note [207, 108, 779, 768]. notebooks [21]. Notes [133, 119, 209, 225, 10, 19]. Novak [19]. novel [642, 1204, 641, 1295]. November [626, 644, 742, 784, 812, 1287, 1034, 291, 459, 538, 1394]. Nozickean [781]. nozzles [904]. NSW [1036]. NT [1272]. nuclear [1383]. number [122, 1210, 1085, 347, 29, 207]. numbers [102, 1410, 984, 1139, 222]. numeric [613, 114]. numeric-symbolic [114]. Numerical [1269, 230, 1108, 1377, 1228, 1045, 1248, 1384, 1388, 663, 1171, 950, 85, 659, 1264, 660, 568, 646, 1127, 116, 97, 104, 128, 367, 529, 277, 19, 587, 1183, 476, 999, 571, 1104, 989, 992, 1370, 651, 624, 735, 280, 755, 679, 1268, 1243, 1158, 620, 698, 687, 461, 278, 750, 681, 1190, 932, 113, 570, 657, 1229, 1335, 585, 1290, 1230, 600, 695, 599, 1135, 265, 28, 380, 520, 189, 212, 228, 283, 287, 525, 12]. numerically [1277]. Numerics [519]. Numerik [491]. NY [135, 457, 449]. Nyce [392].

O [258, 36, 120, 128, 553, 17, 242, 1277]. Object [798, 888, 613, 794, 269, 1341, 1288, 605, 793, 114]. Object-oriented [798, 613, 794, 269, 1341, 1288, 605, 793]. objectives [59]. observation [714, 45]. observations [1079, 1314]. observing [1043]. obstacle [893, 483, 570, 922]. Ocean [208, 1, 1244, 461, 348, 262]. October [155, 516, 616, 1363]. ODE [475, 971]. ODEs [1027, 506, 610, 1116, 969]. Oevel [553]. Office [242]. Offshore [53]. OH [436, 435]. oil [1229]. oligopolists [1414]. Olmsted [227]. OLS [755]. on-line [922]. one [677, 1353, 1052, 904, 1396, 1256, 837, 585, 276, 819, 1402]. one- [1052]. one-band [1256]. one-dimensional [677, 1353, 585]. one-step [276]. one-way [1396, 1402]. Ontic [125]. Open [1]. operating [1215]. operation [899, 1100]. operational [1204, 748, 1283]. operations [1097, 1111]. operator [356, 450, 451, 562, 1046, 226]. operators [204, 429, 968, 484, 254, 136, 103, 368]. opportunities [315, 76]. optical [648, 296]. optics [654]. Optimal [760, 411, 957, 1083, 1069, 347, 169, 1223, 331, 357, 318, 1358, 732, 876, 1084, 342, 1152, 1296, 746, 162, 3]. optimality [873]. optimisation [1411, 945]. Optimising [313, 854]. Optimization [1406, 112, 555, 974, 795, 1112, 1213, 1360, 1405, 1175, 59, 1106, 1335, 632, 941, 410, 619, 10]. optimum [539]. option [780]. options [690]. optoelectronic [1356]. Orchestrated [846]. order [1105, 1250, 1270, 915, 1220, 754, 909, 928]. ordered [985]. ordering [838, 844]. orderings [1366]. ordinary [558, 968]. ore [1074]. organisation [774]. Organization [292, 840]. organizational [245]. organizing [866]. orientation [1193]. oriented [269, 1341, 1288, 605, 947, 798, 613, 793, 403, 718, 794]. original [420, 913]. Ortega [212, 367, 535]. Orthogonal [874, 14]. orthopaedy [1372]. Ös.1148 [365]. ÖS686 [192]. Oscar [266]. oscillations [455, 817, 421]. oscillator [278, 1345]. other [73]. outer [1358]. Outliers [779]. output [161, 908, 996, 1155]. output-error [161]. overtraining [851]. OWE [871]. Oxford [273, 387, 224, 221, 528]. oxide [340]. OZCOT [316].

P [23, 30, 8, 381, 392, 398, 496, 520, 122, 147, 150, 116, 130, 132, 142, 117, 151, 188, 211, 230, 282, 205, 287, 182, 192, 432, 413, 361, 447, 434, 533, 554, 562, 13, 1027, 1270, 1277]. p.d.e [677]. PA [212, 90]. Pacific [1048, 1059]. packages [1299]. packs [954]. Pages [82, 174, 40, 107, 155, 275, 303, 243, 291, 344, 400, 469, 373, 459, 438, 504, 566, 478, 538, 516, 511, 573, 592, 662, 671, 582, 626, 644, 616, 727, 825, 703, 742, 784, 812, 849, 883, 935, 1120, 1033, 1020, 1086, 1094, 1181, 1185, 1267, 1294, 1160, 1125, 1141, 1226, 1232, 1173, 1206, 1287, 1275, 1347, 1409, 1418, 1301, 1328, 1338, 1394, 1363, 1357, 101, 94, 95, 105, 96, 102, 93, 97, 103, 104, 10, 15, 16, 17, 19, 11, 20, 12, 14, 13, 18, 21, 98, 99]. pages [100]. Paine [397, 557]. pair [1184]. Panasenko [138]. Pang [393]. Pankov [356]. Paper [28, 148, 487, 486, 488, 90, 205, 446, 457, 449, 550, 533, 527, 536, 1254, 93]. Paperback [30, 265, 389]. Papers [86, 151, 274, 232, 1034, 743, 349]. parabolic [52, 110, 931, 1391]. paradigm [778]. Parallel [1413, 473, 933, 879, 1116, 79, 642, 1236, 87, 423, 535, 1027, 1250, 1122, 901, 708, 709, 1385, 877, 1304, 410, 26]. Parameter [375, 163, 622, 938, 70, 65, 1358, 1307, 1309, 1403, 1282, 480, 159, 1315, 166, 160, 540, 740, 162, 512, 636, 632, 157, 1295]. parameterized [930]. parameters [1404, 1066, 1306, 894, 715, 110, 513, 746, 529]. Parametric [1323, 970, 326, 921, 441, 675, 937]. Parametrization [964]. Parametrized [962, 1000]. Parikh [266]. Paris [130]. parity [1058]. Park [724, 723]. Parker [128]. parking [877]. Parsimonious [630]. part [230, 21]. partial [1127, 492, 199, 456, 529, 545, 546, 1013, 1337, 993, 186, 526]. particle [677, 48, 63, 1259, 696, 1047]. particles [1234]. partiellen [491]. Partington [103]. partners [575]. Parzen [420]. PASCAL [350]. Passekov [188]. path [886, 1072, 909]. paths [921]. pathway [819]. Pattern [841, 856, 52, 875, 654]. patterns [1088, 1054]. PB [186, 191]. PC [252, 540]. PDE [484]. PDE’s [1168, 604, 989, 609]. Pearson [432]. Pecaric [35, 352, 560]. Pedrycz [9]. Peitgen [36]. Percentile [65]. percentiles [70, 324]. perceptive [896]. Performance [1416, 718, 1341, 176, 766, 1346, 946]. Pergamon [387]. period [765, 957]. periodic [138, 356, 1164, 751, 58, 326, 517]. Periodically [1062, 1061]. permanent [723, 1324, 1195]. permanent-magnet [723]. perspective [342, 248]. perspectives [349]. Perturbation [282, 194, 635, 444, 609, 389]. perturbations [1012, 1147, 69, 268]. perturbed [1220, 445]. pest [314]. pests [312]. Peters [126, 554]. Petersburg [363]. Petkovic [119]. petri [1124, 788, 789, 298, 1022, 1178, 809]. Petri-nets [298]. Phase [1227, 1300, 1205, 836, 857, 837, 482, 4]. phase-fitted [482]. phase/single [1205]. phenomena [832, 943]. phenomenological [56]. phenomenology [256]. phenomenon [1196, 851]. Philadelphia [212]. Phillips [28]. Photon [1244]. photosystem [1277]. photovoltaic [66]. physical [366, 364, 305]. physicochemical [805]. physics [845, 1251, 354, 534]. physiology [464]. picture [857]. PID [633, 632]. PIDMaster [631]. Piecewise [84, 1382]. Piecewise-linear [84]. Pier [130]. pig [917]. Pintelon [387]. pitchfork [695]. pitfalls [799]. Pitts [818]. placement [411, 632]. Plamenevskii [136]. planar [1112, 412, 1090]. Planck [569]. plane [988, 1382, 1182]. Planktonic [1121]. Planning [921, 80, 1000, 886]. plant [1383, 74, 306, 910]. plants [806, 950, 794, 1332]. plasma [831, 1190]. plasticity [1373]. plates [1374]. platform [1341]. plural [899]. plus [203, 415]. point [1228, 1310, 982, 737]. points [32]. polar [1198]. pole [632]. pole-placement [632]. policies [329]. pollution [50, 1040]. polycrystalline [1265]. polyhedra [1293, 1163]. polylogarithms [1001]. polynomial [119, 638, 1284, 970, 1169, 1130, 1133, 963, 961]. polynomial-non-linear [970]. polynomials [14, 960, 210, 18]. polytope [159]. polytopes [158]. population [455, 522, 317]. populations [314]. porous [1263, 64]. portfolio [691, 1069]. portion [321]. portrait [1227, 1300]. pose [738]. position [915]. posteriori [1389, 1391]. potential [371, 1285]. pour [690]. Povinelli [436, 435]. Power [941, 1293, 1163, 1205, 693, 1058, 1202, 712, 971, 1383, 1354, 708, 720, 1194, 718, 710, 766, 719]. pp [254, 267, 272, 257, 260, 273, 264, 269, 252, 265, 262, 268, 259, 253, 258, 255, 256, 270, 266, 263, 271, 22, 27, 25, 31, 34, 5, 6, 9, 23, 30, 32, 24, 28, 29, 4, 8, 26, 35, 33, 7, 36, 381, 384, 386, 383, 396, 391, 393, 388, 392, 380, 394, 389, 390, 395, 379, 385, 382, 397, 398, 387, 487, 489, 499, 500, 501, 490, 485, 494, 498, 495, 496, 486, 492, 491, 520, 488, 502, 493, 88, 92, 115, 126, 133, 136, 152, 122, 91, 121, 144, 147, 148, 150, 90, 129, 137, 139, 127, 141, 116, 118, 138, 87]. pp [119, 145, 130, 146, 89, 123, 124, 131, 132, 134, 140, 143, 149, 125, 86, 120, 128, 117, 151, 135, 209, 213, 217, 206, 187, 196, 198, 208, 210, 225, 289, 203, 229, 286, 200, 193, 211, 218, 224, 230, 282, 288, 219, 281, 195, 214, 215, 216, 204, 184, 185, 194, 201, 221, 189, 202, 207, 212, 222, 228, 283, 287, 226, 186, 190, 223, 191, 197, 284, 182, 192, 199, 285, 220, 227, 181, 429, 430, 446, 456, 356, 351, 355, 359]. pp [423, 450, 354, 357, 360, 365, 432, 418, 421, 427, 415, 353, 451, 352, 413, 424, 452, 358, 422, 361, 366, 371, 420, 447, 453, 364, 349, 350, 362, 416, 433, 434, 368, 369, 370, 419, 426, 455, 457, 414, 417, 425, 428, 431, 448, 449, 529, 545, 558, 546, 530, 531, 550, 560, 522, 533, 532, 523, 554, 561, 524, 521, 548, 552, 557, 527, 547, 553, 525, 562, 526, 549, 551, 535, 555, 556, 559, 536, 528]. practical [196, 951, 1223, 128]. practice [28, 980]. practices [313]. practitioners [117]. pre [909]. pre-trained [909]. preceding [700]. precipitation [1037]. precision [1278, 1081]. Preconditioning [1292, 1385]. predatory [781]. Predictable [753]. predicting [324]. Prediction [1105]. Predictive [955, 640, 639]. predictor [680, 482]. Preface [627, 884, 1095, 959, 401, 304, 574, 785, 244, 814]. premia [776]. premixed [346]. Prentice [385]. Prentice-Hall [385]. preprocessor [379]. presence [722, 736, 1314]. Press [267, 273, 255, 266, 22, 29, 8, 391, 393, 392, 394, 389, 387, 499, 500, 497, 501, 485, 494, 495, 520, 502, 92, 152, 91, 94, 95, 105, 144, 147, 148, 137, 139, 141, 93, 87, 103, 145, 146, 89, 140, 143, 86, 196, 203, 200, 224, 214, 194, 221, 207, 283, 197, 181, 446, 354, 432, 415, 366, 453, 434, 454, 367, 419, 414, 425, 522, 533, 532, 561, 524, 552, 527, 547, 525, 562, 549, 551, 535, 534, 125, 189, 1217, 556, 528]. Price [1318]. prices [753]. pricing [692, 1321, 780]. primality [1343]. prime [122]. primer [148]. Principal [1100, 164]. principles [706, 800, 13]. Prioritising [1075]. Pritsker [385, 349]. probabilistic [444, 955, 1266]. probability [420, 729, 1307, 956, 622, 740, 86, 549]. probability-weighted [1307]. probed [1277]. probit [1070]. problem [691, 1299, 608, 1377, 374, 836, 590, 1336, 649, 1000, 62, 1025, 1344, 347, 1137, 738, 589, 733, 601, 249, 1346, 1157, 602, 1144, 393]. problem-solving [601, 249, 602]. problems [211, 371, 546, 1231, 464, 975, 1373, 598, 1376, 1368, 845, 1369, 1407, 1104, 1378, 1405, 280, 1380, 930, 1385, 681, 570, 1406, 697, 513, 931, 179, 1390, 1335, 517, 1290, 1109, 1230, 1402, 568, 665, 383, 396, 287, 418, 427, 552, 433]. procedure [1084, 620, 1261]. procedures [288, 1289]. Proceedings [362, 416]. Process [78, 1417, 322, 57, 1222, 575, 807, 792, 1296, 1355, 663, 472, 945]. processes [577, 138, 187, 284, 366, 1334, 1216, 245, 688, 808, 947, 1099, 838, 1248, 45, 942, 802, 580, 684, 932, 697, 279, 1332, 25, 137]. Processing [553, 578, 1377, 597, 293, 1029, 298, 601, 892, 7]. processor [1167]. processors [933, 708, 709]. produced [1203]. production [944, 1099, 1360, 748, 746, 1283]. professional [246]. profile [1049]. program [26, 379, 285, 1291, 1189, 403]. programming [422, 557, 957, 978, 798, 613, 1354, 542, 60, 951, 1317, 513, 1117, 23, 87, 205, 353, 223]. programs [131, 552, 1130, 479, 539, 599]. progress [251, 190]. project [409]. projecting [1398]. projection [101]. projects [1075]. ProLab2 [897]. prolog [397]. proof [288, 991]. proofs [990, 355, 8]. propagation [476, 648, 1081, 803, 844, 530, 531]. Properties [790, 558, 756, 1361, 1060, 120]. propulsion [1191]. prospective [1209]. protection [781]. proving [1343]. pseudo [429, 1278, 1261]. pseudo-differential [429]. pseudo-experimental [1278]. pseudo-vibration [1261]. pseudodifferential [136]. pseudospectral [589]. public [59, 60]. publication [436, 435]. Publications [262, 206]. publish [93]. published [149]. Publishers [5, 9, 32, 24, 4, 35, 7, 36, 149, 20]. Publishing [27, 25, 23, 28, 26, 349]. pulse [1256]. punishment [330]. purchase [334]. purchasing [1058]. Pure [6, 209, 216, 852]. purpose [528]. purposes [641]. PWM [1205]. PWS [129, 545, 546, 550]. PWS-Kent [129, 545, 546].

quadratic [692, 675]. Qualitative [577, 578, 611, 789, 580, 1209, 1361, 852, 889, 1022, 576]. quality [324, 1052]. quantifier [994]. Quantitative [952, 889, 1022, 576]. Quantum [658, 258, 187, 195, 1238, 1239, 846, 750, 1256]. quantum-mechanical [187]. quasi [1309, 1240, 1382, 707, 687, 1308, 1235, 1264, 719, 1311, 690]. quasi-explicites [690]. quasi-likelihood [1309, 1311]. quasi-Monte [687, 1235]. quasi-Monte-Carlo [1264]. quasi-random [1240]. quasi-resonant [719]. quasi-rotating [707]. quasi-stationary [1308]. quasi-uniform [1382]. quasiconformal [6, 261]. quasigeostrophic [348]. quaternion [738]. Québec [713]. Queensland [321]. questions [54]. queueing [1096, 259]. queues [1083]. quing [91]. Quintic [667].

R [272, 265, 255, 263, 27, 393, 385, 382, 387, 92, 115, 93, 103, 145, 98, 134, 140, 99, 120, 209, 213, 217, 208, 200, 214, 204, 202, 223, 191, 220, 227, 430, 446, 450, 422, 350, 434, 454, 530, 531, 550, 521, 525, 562, 559, 17, 12, 248]. R-technology [248]. Raatz [288]. radar [1404]. radial [1229, 1298]. radiation [1249]. Radical [987]. radicals [487]. Radio [259]. radionuclides [1263]. Rahman [262]. railway [1097, 1138]. rainfall [57, 1312, 1048, 55, 323, 1305]. rains [191]. Ramanujan [21]. random [1092, 417, 431, 1410, 1240, 1244, 1416, 1111, 68, 280, 1210, 164, 1085, 443, 932, 179, 666, 445, 284]. Randomly [1351]. rangeland [307]. rank [754]. Rao [413]. rarefied [1289]. Raster [89]. rate [335, 689, 1087, 1156, 780, 1063, 472, 945]. rates [1065]. ratio [722]. rational [1006, 1315, 1165]. Raton [389, 283]. Ratz [350]. re [323]. re-calibrating [323]. reaction [440]. reactions [805]. reactive [1089]. reactor [1021]. reactors [635]. Reading [27, 25, 31, 34, 23, 30, 28, 26, 33, 398, 286, 183, 184, 223, 358, 361, 526]. Real [708, 984, 386, 987, 604, 1184, 986, 896, 473, 988, 709, 985, 1057, 1139, 1124, 710, 910, 454]. Real-time [708, 896, 473, 709, 710, 910]. real-world [604]. realistic [1245]. realities [88]. reality [823]. realization [484]. really [55]. reasoning [578, 890, 603, 1024, 855, 430]. recipes [520]. reciprocity [1297, 390]. recognition [351, 1344, 114, 856]. recombination [544]. reconfiguration [629]. reconstruction [4, 1403, 1402]. records [122]. Recovery [1404, 1371]. rectangular [411, 544]. rectifier [719]. rectifiers [1203]. recurrent [908, 1354]. Recursive [161, 1006, 72, 642, 1367, 160]. reduce [1074, 384]. reduced [915, 1220, 766]. reduced-order [915]. reduction [1410, 846, 1318, 1211, 998]. reductions [1038]. reductive [454]. redundancy [878]. redundant [912, 885, 639]. Redwood [13]. reference [446, 640, 707]. reflected [686, 684]. reflecting [1410, 685]. Reflections [95]. reflexions [465]. regeneration [1221]. Regenerative [551]. regime [1215]. Regional [1043, 1038, 1042]. regions [992, 1282, 1247, 166, 1182]. regression [1284, 16, 755, 407, 1067, 1041, 197]. Regressive [953]. regressor [755]. regressors [766]. Regular [1131, 490, 1382]. Regularization [734]. Regularizing [735]. Regularizing-numerical-continuation [735]. regulation [870, 1068]. Reid [224]. related [552, 822, 866]. relation [9, 313]. relations [75]. relationship [1222]. relationships [1056]. release [318, 624]. reliability [1252, 1254]. reliable [507, 434]. reluctance [714]. remarks [685]. Remmert [220]. removing [1284]. rendezvous [738]. Renelt [6]. Renormalized [258]. replacement [329]. replacements [1388]. reply [700]. report [1118]. repository [594]. representation [150, 125, 631, 405, 1098, 1090, 395]. representations [408, 568]. reprint [457, 449]. requirements [1222]. resampling [692]. Research [436, 435, 293, 767, 748, 979]. reservoirs [59]. Reshetnyak [261, 127]. residues [975]. resistance [1035]. resistivities [735]. resistivity [1392]. Résolution [682]. Resolved [639]. resolvers [1199]. resonances [713, 1166]. resonant [719]. resource [1111, 321, 747]. resources [1045]. response [1043, 341, 1042, 305, 170, 1383, 715, 835]. responses [332]. restoration [618]. result [502, 319, 519]. resultant [1344]. resultants [1168]. results [624, 60, 1057, 169, 338]. retina [1286]. retrieval [4, 193]. Return [1320]. returns [1404, 1059]. reversal [390]. reversible [736]. Review [254, 267, 272, 257, 260, 273, 264, 269, 252, 265, 262, 268, 259, 261, 253, 258, 255, 256, 270, 266, 263, 271, 22, 27, 25, 31, 34, 5, 6, 9, 23, 30, 32, 24, 28, 29, 4, 8, 26, 35, 33, 7, 381, 384, 386, 383, 396, 391, 393, 388, 392, 380, 394, 389, 390, 395, 379, 385, 382, 397, 398, 487, 489, 499, 500, 497, 501, 490, 485, 494, 498, 495, 496, 486, 492, 491, 520, 488, 502, 493, 88, 92, 101, 115, 126, 133, 136, 152, 122, 91, 94, 95, 105, 121, 144, 147, 148, 150, 96, 90, 102, 129, 137, 139]. Review [127, 141, 93, 116, 118, 138, 87, 97, 103, 104, 119, 145, 98, 130, 146, 89, 123, 124, 131, 132, 134, 140, 142, 143, 99, 149, 125, 86, 120, 128, 117, 151, 135, 100, 209, 213, 217, 206, 187, 196, 198, 208, 210, 225, 289, 203, 229, 286, 200, 188, 193, 211, 218, 224, 230, 282, 288, 183, 219, 281, 195, 214, 215, 216, 204, 184, 185, 194, 201, 205, 221, 189, 202, 207, 212, 222, 228, 283, 287, 226, 186, 190, 223, 191]. Review [197, 284, 182, 192, 199, 285, 220, 227, 181, 429, 430, 446, 456, 356, 351, 355, 359, 423, 436, 450, 354, 357, 360, 365, 432, 418, 421, 427, 415, 353, 451, 352, 413, 424, 452, 358, 422, 361, 366, 371, 420, 447, 453, 364, 349, 350, 362, 416, 433, 434, 454, 367, 368, 369, 370, 419, 426, 455, 457, 414, 417, 425, 428, 431, 448, 449, 435, 529, 545, 558, 546, 530, 531, 550, 560, 522, 533, 532, 523, 554, 561, 524, 521, 548]. Review [552, 557, 527, 547, 553, 525, 562, 526, 549, 551, 535, 555, 556, 559, 10, 15, 16, 17, 19, 11, 20, 12, 14, 13, 18, 21, 534, 536, 377, 1279, 169]. revised [457]. revisited [764]. revolution [263]. Ribenboim [122]. Riccati [1398]. Rice [525]. Richardson [648]. richness [310]. Ridge [90]. Rienstra [433]. Rietman [90]. rigid [1092, 919, 997]. Rigorous [722]. Ringhofer [192]. Rings [1144]. Rinnooy [353]. ripples [1200]. rise [341, 1277]. risk [313, 956]. river [1222, 808, 1043, 1051]. Rivlin [210]. RLS [1081]. road [1411, 745, 1304, 1075, 892]. road-traffic [1304]. Robinson [530, 531]. robot [642, 916, 915, 483, 877, 1156, 922, 894]. robotic [905, 630, 913, 114, 917]. robotics [920, 903]. robots [911, 664, 890, 899, 901, 641, 900, 885, 639]. Robust [58, 1155, 940, 640, 159, 1081, 309, 1295]. rock [1204, 1381]. rods [1407]. Roesser [873]. Rogers [394]. Rogge [456]. rolling [1116]. ROM [524]. Ronan [146]. Ronkin [490]. Roos [491]. root [987, 1053, 1310]. roots [1314]. Rosenbrock [528]. Ross [257, 549]. Rössler [833]. rotating [707]. rotations [1165]. Roubens [190]. rough [521]. roundoff [1091]. routines [731]. Routing [1179, 444, 698]. Ruelle [105]. rule [878]. rules [789, 309]. run [332, 753, 756, 1056]. Runge [517]. runoff [732, 315, 55, 1036]. rural [1070]. Ruscheweyh [225]. Russian [259]. Rycroft [273].

S [254, 257, 259, 253, 270, 271, 31, 9, 35, 381, 393, 388, 495, 486, 520, 121, 102, 141, 87, 131, 128, 135, 225, 193, 288, 183, 195, 215, 204, 185, 201, 430, 352, 433, 434, 368, 426, 560, 549, 551, 536]. S&P [1318]. S.Fr.68 [355]. salinisation [1045]. salinity [342, 1331, 1049]. Samoilenko [187, 421]. sample [1128, 759, 1314]. sample-selection [759]. sampled [1202, 68]. sampling [1246, 1177, 1083, 310, 176]. Samuel [86]. San [22, 29, 499, 497, 501, 485, 92, 91, 139, 141, 87, 89, 86, 203, 415, 367, 425]. Sanchez [9]. sand [341, 52]. Sargsjan [368]. satellite [1167, 1165]. saturation [722, 761]. satureted [721]. sauts [690]. saving [334, 1285]. scalar [1039, 63]. Scale [1044, 1288, 294, 1271]. scanning [164]. Scattering [208, 1299, 530, 461, 657, 1401]. scenarios [249, 1063]. scheduling [1111, 1075, 518, 410]. Scheibe [839]. scheme [676, 881, 461, 686, 1087, 922, 1195]. schemes [992, 679, 867, 1081, 276]. Schempp [124]. Schmeiser [192]. Schoukens [387]. Schrijver [353]. Schrödinger [370, 646, 1012, 1023, 652, 585, 1016, 85, 1017]. Schrödinger-type [1023]. Schulze [553]. Schumaker [499, 143]. Schwarz [426, 12, 1104, 930]. Sci [529, 555]. Science [28, 149, 191, 152, 185, 774, 612, 358, 414]. sciences [744, 273, 394]. Scientific [608, 593, 520, 93, 183, 605, 1082, 507, 502, 367, 535]. scientists [129, 132, 380, 389, 94, 286, 366]. score [336, 112]. Scott [841]. screwworm [318, 317]. SDE’s [685]. sea [341, 49]. search [1031, 610, 619, 417]. Seasonal [1055, 1326]. seasonally [58]. second [270, 263, 31, 212, 457, 737]. section [102]. sector [60]. sectoring [1350]. securities [1273]. Sedgewick [27]. sediment [1047]. sedimentation [808]. segmentation [898]. Seidenberg [974]. Seismo [208]. Seismo-ocean [208]. Selected [1034, 743, 60, 151]. selection [691, 351, 167, 1052, 759, 1298]. Self [575, 774, 866, 632, 840]. Self-explanatory [575]. Self-organization [840]. self-organizing [866]. self-tuning [632]. Selftuning [628]. Semi [1023, 747, 993]. semi-arid [747]. Semi-discrete [1023]. semi-discretization [993]. semialgebraic [964]. semiclassical [836, 827]. semiconducting [1377]. semiconductor [142, 1256, 192]. semivariograms [325]. Sendov [281]. sensitivities [1122]. Sensitivity [69, 1316, 407, 1049, 853, 176, 1157, 780]. Sensor [906]. Sensor-based [906]. separate [1366]. September [1275, 438, 511, 935, 1094, 1357]. sequence [816]. sequences [54, 1257]. sequential [1358]. Ser [529, 555]. serial [912]. serially [1079]. Series [260, 253, 28, 136, 138, 149, 151, 187, 188, 195, 207, 284, 456, 418, 364, 428, 261, 258, 256, 6, 32, 24, 35, 464, 1312, 674, 779, 1054, 326, 971, 756, 1354, 687, 1055, 1053, 72, 931, 1264]. servo [1199]. SESAME [598]. Sessa [9]. Set [151, 351, 167, 161, 164]. set-membership [167]. sets [193, 190, 521, 1358, 964, 45, 375, 160, 460, 309, 963]. settling [1047]. SFr [124, 553]. Sfr.128 [287]. SFR.38 [265]. Sfr.52 [288]. Sfr.78 [386]. Shabbir [435]. Shallow [377, 1015, 53, 474]. Shape [1286, 1411, 856, 555]. shapes [1349]. Shapiro [426]. Shaw [191]. Shedler [551]. sheet [586]. shell [1183, 1376, 602]. Shenton [120, 17]. shielding [1249]. shift [682]. Shih [435]. ship [876]. shop [1111]. Shoreface [341]. Short [1063]. Short-term [1063]. shortest [1072]. Shubin [370]. SIAM [212]. Siewiorek [434]. Sigal [385]. Sigmund [5]. signal [708, 601, 998]. signals [693]. signs [1297]. silicon [1356]. SIMCOS [1180]. Simeonov [496]. simple [323]. simplicity [1073]. simplification [1288, 1131]. simplified [664, 718]. Simul [1293, 1300, 1117]. simulated [1092, 1278, 618, 1049, 1277]. Simulating [829, 1111, 1047]. Simulation [402, 311, 1204, 1034, 711, 1240, 439, 1074, 1356, 944, 49, 712, 1250, 799, 1175, 920, 793, 1224, 412, 55, 1214, 378, 1221, 1040, 410, 619, 623, 252, 765, 117, 1372, 1412, 1108, 404, 795, 325, 1280, 1239, 408, 1097, 54, 999, 314, 1035, 341, 919, 808, 927, 1089, 43, 569, 1378, 1360, 801, 473, 1177, 1030, 730, 683, 406, 788, 1082, 1286, 798, 1313, 942, 407, 806, 177, 1100, 708, 709, 706, 913, 321, 705, 307, 747, 505, 1254, 1253, 1187, 748, 637, 1176, 932, 113, 329, 745, 1304, 1217, 308, 697, 403, 1106, 1303]. simulation [540, 306, 663, 479, 1194, 405, 112, 176, 717, 539, 800, 1260, 1305, 1022, 1263, 376, 1230, 950, 79, 85, 512, 1265, 1178, 997, 316, 660, 794, 64, 743, 257, 142, 551, 379, 285]. simulation-based [405]. Simulations [440, 739, 1348, 836, 1244, 1084, 606, 738, 111, 1174, 659, 1298, 695]. simulator [896, 588, 410]. simulators [575, 1098]. simultaneous [119, 763]. sine [327, 1386]. sinewave [723]. Singapore [1056]. Singh [381]. single [1154]. single-input [1154]. single-phase [1205]. singular [635, 1415, 901, 609, 457]. singularities [583, 369]. singularly [1220, 445]. Sinha [413]. sinks [340]. Sites [446]. Sivasundaram [271]. six [1344]. Size [762, 957]. size-structured [957]. Sketches [965]. sketchpad [1007]. skier [378]. Skill [891]. Skill-based [891]. Skorohod [284]. SLAM [385]. sliding [916]. Slowinski [521]. sludge [807]. small [1400, 45, 1342, 718, 1314]. Smith [283]. Smoluchowski [1333]. Smooth [1182, 921]. smoothed [444]. smoothing [58, 326, 509, 462, 72]. snapshots [432]. Snyder [494]. Sobolev [261]. Societies [242]. Society [207, 242]. socio [777]. socio-economic [777]. soft [744]. SoftLab [612]. Software [203, 1299, 152, 525, 594, 787, 597, 624, 631, 665, 419]. soil [953, 315, 1383, 1279, 1049]. soils [943]. solar [285, 1280, 66]. solenoidal [1398]. Solid [33]. soliton [382, 256]. solitons [660, 149, 270]. Solution [590, 108, 487, 464, 1027, 1021, 732, 1377, 648, 571, 280, 1337, 542, 1137, 738, 1116, 621, 84, 179, 1262, 585, 885, 928, 445]. solutions [456, 356, 558, 647, 1164, 1370, 1013, 971, 655, 912, 278, 1087, 1015, 1317, 1390, 685, 584, 962]. solvers [1366, 475]. Solving [821, 1231, 1376, 1272, 1344, 733, 963, 1127, 608, 667, 989, 960, 735, 1333, 681, 570, 460, 1229, 1261, 601, 249, 1346, 1017, 602, 969, 474]. Some [1027, 54, 845, 465, 620, 1142, 111, 697, 1303, 685, 583, 70, 677, 1375, 1359, 66, 1318, 295, 1060, 178, 279, 1229, 169, 338, 1057]. Sommem [450]. son [130]. Sons [6, 90, 88]. sorghum [313]. Sorkam [313]. Soucek [450]. sounding [1392]. source [720]. sources [1284, 411, 340]. Southampton [262, 206]. Southampton/Springer [262]. Soviet [261, 258, 256, 136, 138, 187, 188, 195, 284]. space [383, 1403, 1201, 590, 791, 656, 1350, 694, 922, 665, 388, 548]. spaced [1313]. spaces [396, 204, 451, 750, 261, 226, 448]. spatial [390, 319, 78, 1052, 321, 75, 317, 55, 1090]. spatial-temporal [55]. special [395]. species [310]. specification [603, 1071, 621]. specified [325]. specifying [752]. Spectral [584, 97, 693, 544]. spectrum [1154, 1184]. Spedicato [101]. speed [1412, 1384, 113, 870, 1352]. Spending [767]. Sperschneider [358]. spherical [325]. Spigler [272, 217]. spindle [817]. spines [917]. spinor [450]. spinor-valued [450]. Spinors [255]. spiral [1332]. spline [381, 667, 462, 69]. splines [58]. split [1046]. splitting [1405, 1399]. spontaneous [1224]. sprachbeschreibung [350]. Springer [262, 115, 133, 122, 121, 96, 97, 119, 123, 142, 128, 100, 213, 206, 225, 289, 215, 216, 222, 228, 192, 220, 227, 365, 350, 10, 19, 18, 21]. Springer-Verlag [96, 97, 100, 10, 19, 18, 21]. squared [471]. squares [98]. Srivastava [559]. St [363]. Stability [1151, 1147, 724, 1148, 679, 513, 994, 1149, 1146, 271, 1150, 640, 992, 1370, 1145, 1025, 1028, 708, 1165, 455]. Stabilization [1153, 1152]. stable [869, 730, 683]. stand [715]. stand-still [715]. standard [1376, 473]. Stanislaw [95]. STAT [269]. State [1113, 1415, 781, 109, 791, 1030, 716, 694]. states [1238, 313]. Static [1183, 1202, 706, 112, 713]. stationary [693, 1308]. statistic [762]. Statistical [1256, 1392, 269, 252, 471, 201, 17, 1251, 407, 117]. Statistics [445, 86, 46, 336, 658, 827, 508]. stator [1200]. status [242]. steady [1030, 716, 474]. steady-state [1030, 716]. steel [663]. Stefan [1377]. Stefan-like [1377]. STENSOR [1003]. step [374, 1060, 276]. stepped [1404]. stepped-frequency [1404]. stepsize [475]. Sterile [318]. sterilisation [954]. Stern [202]. Stewart [194]. Stewartson [660]. stiff [506, 1270]. still [715]. Stillwell [123]. Stimulated [62]. stimulations [803]. Stirling [342]. Stochastic [1048, 1247, 56, 1203, 442, 942, 1197, 1234, 696, 695, 551, 1334, 676, 1147, 677, 19, 674, 54, 66, 673, 440, 679, 1066, 1306, 58, 730, 441, 1333, 686, 1087, 443, 681, 279, 839, 1259, 539, 507, 182]. stock [1320, 1059, 1319]. stocking [957]. stockpiles [1074]. Stokes [584]. Stone [393]. storage [66]. straight [1130]. straight-line [1130]. strange [105]. Strategies [191, 318, 1367, 1351, 634, 310, 746]. strategy [1112, 1193]. stray [1199]. stream [1108, 1049]. stream-tube [1108]. Street [1351]. stress [1336, 1381]. strikes [818]. string [147]. stripping [1403]. strongly [1297, 1382]. structural [1366, 860, 822, 852, 1383, 1062, 762]. structure [89, 167, 404, 1359, 1381, 815, 600, 1110, 887, 1065]. Structured [297, 957, 944]. structures [1012, 185, 424, 1244, 939]. structuring [1026]. Strum [550]. Student [252]. studies [252, 382, 414, 1025, 75, 113, 1195]. Study [1194, 694, 981, 140, 1278, 1203, 336, 894, 1268, 1142, 1037, 1190, 1258, 1290, 736, 340, 1265, 410, 197]. Sturm [368]. Stuttgart [491, 362, 416]. Stuttgart/Kluwer [362, 416]. Sub [1026, 818]. sub-millisecond [818]. Sub-structuring [1026]. subcortical [819]. subject [621]. suboptimal [1220]. substrates [1258]. subsynchronous [713]. subsystems [842]. succession [74]. Suchy [390]. suitable [67]. sum [1351]. summarisation [872]. Summations [1001]. Summit [90]. Sun [194]. supercomputers [126]. supercomputing [530, 208]. superconductor [1248]. Superconvergent [1371]. supervision [577, 580]. supplement [219]. supplied [1202]. supply [840]. Support [245, 484, 521, 1097, 749, 307, 820, 320, 599, 385]. Support-operators [484]. supporting [576]. Surface [462, 91, 1404, 1399, 45, 481, 906, 1110, 1349]. surface-following [906]. surfaces [196, 656, 907, 198]. surfactants [838]. surrogates [1106]. survey [488, 279, 479, 710, 973, 518, 157]. survival [1307]. suspension [1379]. sustainability [321]. sustainable [952]. Svirezhev [188].

Svyaz [259]. Swarz [434]. Sweden [1324]. swept [327]. switched [714, 717]. switching [765]. Sylvester [990]. symbol [578]. Symbolic [1080, 1288, 989, 992, 1013, 993, 997, 1167, 1348, 541, 1340, 1341, 991, 1183, 109, 1164, 978, 613, 484, 1117, 1212, 1346, 998, 114, 1090, 695, 1134]. symbolic-based [541]. Symbolic-numerical [992]. symbolic/numeric [613]. Symmetric [481, 1352, 1368, 1165]. Symmetries [1015, 199]. symmetry [1291, 650]. Symplectic [652, 1386, 1014, 1166]. Synchronization [1123, 1304]. synchronous [1205, 723, 1200, 715, 721, 1195]. Synthesis [1205, 1127, 418, 298, 1090, 452]. synthesizer [858]. synthetic [1098]. System [181, 765, 125, 289, 415, 893, 180, 594, 245, 587, 1280, 598, 836, 247, 246, 1164, 896, 797, 996, 899, 630, 299, 596, 708, 307, 820, 748, 786, 113, 745, 110, 1304, 1387, 585, 169, 819, 114, 910, 737, 250, 660, 409, 497]. systematic [305]. systematically [308]. systematics [796]. Systems [349, 593, 903, 252, 668, 6, 36, 379, 387, 126, 215, 285, 427, 413, 434, 550, 562, 292, 1151, 1080, 404, 1154, 463, 1288, 408, 1281, 44, 860, 1359, 506, 919, 1222, 109, 1026, 1091, 876, 66, 1245, 1208, 731, 1168, 1202, 650, 791, 569, 838, 1370, 1148, 1145, 651, 78, 597, 327, 440, 473, 1159, 799, 1113, 1415, 654, 1270, 995, 796, 294, 441, 1122, 296, 1142, 297, 1152, 298, 1083, 1025, 913, 1155, 750, 1386, 1209, 790, 1220, 1153, 163, 1315, 631, 772, 993, 1143, 540, 833, 835, 2, 479, 1138, 827]. systems [1124, 1214, 1096, 800, 1212, 710, 1171, 312, 938, 600, 1157, 1024, 1149, 1271, 1178, 518, 1146, 997, 250, 963, 445, 809, 607, 961, 1295, 1144, 270, 271, 5, 500, 145, 128, 187, 214, 221, 202, 190, 359, 360, 428]. Szmydt [492].

T [22, 23, 28, 7, 520, 143, 86, 128, 210, 204, 181, 423, 366, 435, 558, 527, 553]. table [953]. tallies [1237]. tandem [1083]. Target [1260]. targets [1265]. Task [600, 410]. tasks [923]. Taylor [8, 464]. TC1 [704]. TC1-Conference [704]. Technical [242, 184, 238]. technique [991, 327, 1029, 297, 179]. techniques [1288, 1244, 1104, 1177, 798, 1279, 419]. Technologies [363, 576]. technology [981, 495, 297, 248, 24, 152]. telerobotics [895]. temperature [1248]. template [403]. template-oriented [403]. temporal [55]. Tensor [1212]. tenure [769]. term [776, 1063, 1065]. termination [1084]. terrain [1044]. test [336, 1188, 1310, 715]. testability [771]. Testing [752, 1061, 763, 1324, 323, 1084, 689, 761]. tests [1059, 1314]. Teubner [491, 362, 416]. Teukolsky [520]. Texas [181, 532]. their [9, 493, 429, 352, 371, 14]. Theorems [227]. theoretical [845, 414]. theories [1129, 1004]. Theory [980, 267, 264, 268, 1167, 29, 381, 386, 499, 488, 133, 127, 124, 207, 226, 450, 421, 371, 420, 447, 521, 562, 1007, 787, 774, 1159, 1171, 973, 1024, 258, 382, 489, 147, 225, 195, 194, 201, 451, 220, 417, 431, 449, 559]. thermal [379, 285, 954, 663, 666]. Thermodynamics [1135, 1134]. thin [1183, 1248, 1349]. Third [593, 30]. thirty [432]. thirty-five [432]. THR [1388]. three [1414, 70, 65, 1307, 1108, 1205, 1403, 1244, 313, 1331, 74, 111, 740, 1192, 1262, 340, 428]. three-dimensional [1108, 1403, 1244, 1331, 1192, 1262, 340, 428]. three-level [111]. three-parameter [70, 65, 1307]. three-phase [1205]. three-phase/single-phase [1205]. TI [532]. TI-85 [532]. tidal [1381]. Tierney [269]. tillage [1279]. timber [907]. Time [1322, 1220, 1393, 765, 390, 413, 1147, 1312, 779, 1026, 896, 1148, 852, 1405, 473, 1113, 1415, 1054, 326, 1142, 756, 1152, 708, 709, 1055, 1155, 793, 1053, 72, 163, 1315, 1289, 717, 710, 1157, 338, 910]. time-delay [1142, 1152, 1155, 1157]. Time-domain [1393, 1405, 1315, 717]. time-series [1054, 72]. Time-suboptimal [1220]. Time-varying [1322, 1148, 1415, 163]. Tirapegui [424]. Toda [149]. Tokyo [149]. tolerance [733]. too [802]. Tool [553, 1412, 631, 1325]. tools [117, 991, 61]. Topics [382, 135]. Topological [448]. topology [988, 1138, 976, 1349]. topology-independent [1138]. torque [714, 1200, 1195]. torsion [1004]. total [1388]. touch [1133]. Tourism [770, 1071]. tourist [1326]. trace [29, 47]. tracer [340]. tracking [1260, 1047]. traffic [1175, 637, 745, 1304, 1072, 1040, 892]. traffic-induced [1040]. trained [909]. training [51, 867, 410]. trajectories [921]. Trajectory [900]. transaction [691]. transfer [1278, 949, 1315, 839]. transfer-function [1315]. transform [693, 1001, 657]. transformable [724]. transformation [492, 722, 723, 1257, 1325]. transformations [1293, 1163, 658]. transformers [1189]. Transient [809, 1204, 1189, 1030]. transistor [1269]. transition [837]. transitions [1250]. translated [204]. Translation [91, 1315, 457, 449, 341]. transmission [1320, 1089, 1188, 709, 928]. transnational [409]. transonic [180]. transparent [397]. transport [1046, 1097, 1244, 650, 840, 1253, 1255, 1263, 1290, 340, 64, 623]. transportation [620]. Transputer [927]. transversal [1081]. traveling [1012]. treated [1244]. Treatise [556]. Treatment [717, 806, 804, 1296, 1109]. trees [688]. Trend [782, 1310]. trends [1058, 1142, 706, 705]. Triangular [1168, 611, 963]. triangulations [1382]. TRIFIT [665]. Trilogy [74]. Trim [545]. trivial [590]. troposphere [47]. Trubowitz [485]. Truemper [453]. trunk [698]. Truss [286]. trust [881]. tube [1108]. Tucker [552]. Tufte [200]. tumor [804]. Tuning [633, 631, 632]. tunneling [834]. tuple [854]. turbine [1335]. turbulence [435, 829]. turbulent [1105, 1234, 1259, 1107]. Turner [428]. tutor [557]. Two [1411, 1259, 917, 1092, 954, 70, 1108, 325, 463, 1353, 916, 1228, 1052, 1050, 831, 857, 1060, 621, 833, 830, 1335, 737, 584, 665]. two- [70, 1108]. Two-dimensional [1411, 917, 1092, 1353, 1052, 621, 830, 1335, 584]. two-joint [916]. Two-particle [1259]. two-phase [857]. two-point [737]. TX [181]. type [492, 1150, 1386, 1230, 1023]. types [583, 8].

U [395, 502, 118, 350]. U.S. [764]. UK [5, 6, 9, 32, 24, 4, 35, 7, 36, 15, 11, 12, 1057]. UK£ [254, 272, 260, 273, 264, 269, 252, 262, 268, 261, 253, 258, 256, 270, 263, 271, 381, 384, 383, 396, 388, 390, 395, 379, 387, 489, 496, 492, 493, 187, 198, 188, 193, 282, 281, 195, 185, 201, 190, 191, 284, 285, 429, 430, 456, 356, 359, 423, 450, 357, 360, 418, 421, 427, 451, 352, 413, 424, 452, 371, 447, 364, 362, 416, 433, 368, 369, 370, 455, 417, 428, 431, 448, 529, 558, 560, 523, 521, 557, 555, 559]. Ulam [486, 95]. Ullman [414]. Ullrich [350]. ultimate [1149]. unanticipated [334]. unbounded [1336, 1247]. uncertain [167, 1028, 1155, 746]. uncertainties [940]. uncertainty [308, 166]. understanding [1002, 403]. underwater [1191, 7]. unemployment [333]. unified [1389]. uniform [1021, 1382]. uniformization [383]. unilateral [1231]. unit [1053, 1310, 1314]. United [313]. units [1043, 1042]. unity [1194, 719]. Univ [520, 140, 224, 207, 528]. University [8, 94, 95, 105, 93, 103, 242, 181]. UNIX [398]. unknown [754]. Unobservable [776]. unpredictable [753]. unreal [266]. unrestricted [1410]. unsteady [346, 1110]. unstructed [611]. unstructured [1104, 1223]. Unsupervised [875]. updating [895]. upper [324]. urban [773, 1039, 1175, 505, 637, 1072, 1068]. US$100 [382]. US$110 [387, 496, 556]. US$112 [254]. US$114 [359]. US$115 [182]. US$116 [433]. US$117.50 [191]. US$118 [284]. US$12.95 [457]. US$122 [424]. US$124 [357]. US$126 [388]. US$129 [149, 188]. US$134 [281, 360, 371]. US$136 [272, 217]. US$139 [381, 195, 362]. US$140 [364]. US$142 [421]. US$144 [418, 417]. US$148 [416]. US$149 [258, 256, 270, 35, 198, 451, 352, 413, 368, 428]. US$150 [555]. US$154 [558]. US$156 [529]. US$16.30 [90]. US$162 [521]. US$174 [253]. US$176 [450]. US$179 [455]. US$189 [264]. US$19.95 [432, 425]. US$193 [383]. US$199 [379, 489]. US$24.95 [361]. US$245 [431]. US$249 [395, 370]. US$27.95 [184, 207]. US$285 [560]. US$29.95 [527]. US$32.95 [266, 420]. US$34.50 [30, 358, 550]. US$34.95 [495, 446]. US$359 [151]. US$39.35 [255]. US$39.75 [27, 26]. US$39.95 [392, 494, 91, 87, 86, 196, 203, 194, 189, 453]. US$40.95 [31]. US$44.50 [366]. US$48 [200]. US$48.50 [34, 13]. US$49 [152]. US$49.50 [394, 286]. US$49.95 [267, 520, 145, 197, 354, 415, 367, 414, 554, 561, 547]. US$5.95 [488]. US$51.30 [186]. US$54.95 [391]. US$55 [282, 557]. US$59 [397]. US$59.95 [22, 393, 426]. US$6.95 [486]. US$64.95 [434]. US$66.50 [447]. US$67 [24]. US$69 [493]. US$69.50 [289]. US$7.95 [487]. US$71 [353]. US$74 [5]. US$75 [454]. US$78 [427]. US$79 [260, 271, 4, 456, 448]. US$8.95 [449]. US$85 [422]. US$87 [190]. US$88 [384, 423]. US$89 [492, 430]. US$96 [285, 559]. US$98 [429]. US$99 [396, 390, 187, 193, 185, 201, 356, 452, 369, 523]. US$99.50 [32]. US133 [261]. usage [854]. Use [859, 583, 324, 760, 1002, 891, 1316, 675, 749, 951, 55, 323, 909, 312]. user [126, 144, 181]. users [596, 599]. uses [1303].

Using [298, 321, 1348, 28, 550, 1372, 1108, 991, 648, 693, 1240, 1003, 949, 908, 1227, 1300, 881, 1042, 1246, 1105, 1282, 313, 966, 822, 1249, 1400, 956, 58, 1175, 1286, 920, 164, 888, 707, 708, 709, 1324, 1344, 633, 1155, 1411, 1051, 505, 60, 889, 1176, 1257, 76, 1355, 1179, 1192, 870, 1096, 861, 1136, 1260, 1346, 892, 1178, 695, 809, 1295]. Utilization [710].

V [261, 256, 271, 35, 388, 500, 115, 147, 97, 213, 188, 283, 284, 450, 360, 358, 371, 369, 431, 448, 363, 20, 534]. vacancies [838]. Vajda [102]. validated [1106]. Validation [624, 167, 462, 786, 1218]. valuation [59]. value [371, 546, 324, 649, 280, 930, 681, 539, 517, 568]. valued [450, 1133, 493]. vapor [64]. var [713]. variability [1074, 1036]. variable [1366, 178, 623]. variables [1003, 777]. Variance [1410, 68, 778]. variant [999]. variates [1249]. variation [675]. variational [211]. variations [852]. variété [678]. Various [990, 336]. varying [1359, 1148, 1322, 1415, 58, 163]. Vector [1335, 1198, 777, 754, 2, 271, 493]. vector-network [2]. Vector-valued [493]. vegetation [75]. vehicle [891, 439, 897, 896, 797, 798]. velocity [885]. verification [731, 519, 502]. Verlag [96, 97, 100, 10, 19, 18, 21]. version [252, 497, 203]. versus [65, 1351, 744]. Vetterling [520]. via [282, 714, 1309, 795, 1144, 635, 1111, 1360, 1215, 327, 473, 1113, 1415, 480, 1254, 1001, 658, 737]. vibration [1204, 1261]. Vichnevetsky [208, 531]. view [1101]. viewpoint [1098]. VII [213, 499]. VIII [99, 16, 556]. Vilenkin [395]. Vinogradov [199]. Vinuesa [14]. virtual [594, 612, 823]. viscoelastic [1108, 1374, 920]. viscous [139]. Visiometrics [828]. Visual [1097]. Visualization [263]. vivo [950]. vocabulary [600]. Vol [22, 151, 10, 19]. Volenec [35]. voltage [999, 712, 707, 720]. voltage-fed [999]. Volterra [108, 443].

Volume [152, 230, 556, 24, 4, 395, 382, 38, 172, 136, 116, 39, 173, 467, 468, 564, 565, 615, 670, 726, 811, 1009, 1077, 1369, 208, 863, 925]. volumes [22, 151]. vortices [830]. vorticity [1262]. Vries [655, 667, 649, 657, 659]. Vvedensky [526].

W [263, 9, 393, 394, 389, 379, 382, 398, 487, 499, 520, 96, 137, 98, 124, 86, 151, 198, 219, 194, 190, 285, 456, 365, 451, 424, 416, 433, 435, 545, 546, 532, 523, 552, 553, 534]. Wacker [416]. wages [1057]. Wald [762]. walk [1240, 1244]. walking [1351, 914]. Wall [1351]. Wallach [454]. Walsh [687, 84]. Walsh-series [687]. Warren [137]. wastewater [806, 1296]. Water [377, 285, 953, 1280, 49, 1045, 1052, 61, 1015, 53, 378, 64, 474]. water-skier [378]. waters [45]. Watt [121]. Wave [1399, 369, 647, 1405, 327, 653, 1014, 1015, 412, 1402]. Wave-splitting [1399, 1405]. waveguides [648, 1396]. wavelet [693]. Wavelets [1334, 1213, 1113, 1415, 100]. waves [262, 1012, 149, 831]. way [1396, 1402]. Weak [1143, 679, 680]. weakening [1195]. weakly [439]. weather [54, 56, 58]. Weibull [1307, 622, 740]. weighted [1307]. Weinman [129]. Weiskamp [148]. welding [904]. welfare [775]. well [1229, 1271]. Wellesley [554]. Wesley [27, 25, 31, 34, 23, 30, 28, 26, 33, 398, 286, 183, 184, 223, 358, 361, 526, 13]. West [349, 251]. Western [1048, 342]. wheat [321, 948]. Who [372]. whose [1184]. wide [1304]. Wien [142, 192, 365]. Wiggins [215]. Wildlife [76]. Wiley [269, 268, 6, 88, 90, 118, 104, 131, 132, 135, 210, 218, 219, 204, 205, 202, 186, 420, 426, 15, 11, 12]. Wilf [141]. Williams [205]. wind [52, 1335]. Windcrest [90]. windings [1192]. Windows [1272]. Winston [30]. within [1043, 307, 747]. without [840]. Wittmeyer [189]. Wittmeyer-Koch [189]. Wiwianka [553]. Wojtaszcyk [151]. Wong [92, 209]. wool [746]. work [432]. workpiece [904]. Workshop [435]. World [235, 237, 604, 1038, 1124]. write [93, 242]. writing [184].

X [389, 96, 191, 527, 14, 267, 381, 379, 387, 500, 520, 488, 88, 144, 147, 146, 143, 149, 151, 356, 450, 451, 370, 426, 558, 524, 553, 21]. xanthan [944]. XI [105, 93, 18, 21]. XII [209, 10]. XIII [98, 210, 212, 12]. XIV [215]. XIX [219, 216]. XSC [350]. XV [11, 13]. XVI [104]. XXII [218]. xxx [214].

Ya [204]. Yang [493, 427]. Yashkov [259]. year [1324]. years [1413, 432]. yield [1323]. yields [314]. York [257, 252, 263, 380, 487, 486, 488, 88, 137, 118, 104, 98, 131, 132, 99, 120, 117, 135, 209, 210, 214, 202, 226, 186, 420, 350, 426, 15, 16, 17, 11, 14, 534]. Yu.G [261, 127]. Yu.I [187]. Yu.L [388]. Yu.M [188]. Yuan [91].

Z [252, 492]. Zariski [266]. Zauderer [186]. Zavialov [258]. Zealand [749]. Zehna [361]. Zelazko [151]. Zeller [124, 424]. zero [4, 987]. zero-dimensional [987]. zeros [119]. Zhe [91]. Zhigljavsky [417]. Zhu [226]. Ziemian [492]. Zimmerman [190]. Zulehner [416]. Zwillinger [391, 498]. Zygmund [151].

References

Milliff:1990:MCM

[1] Ralph F. Milliff. A modified capacitance matrix method to implement coastal boundaries in the Harvard Open Ocean Model. Mathematics and Computers in Simulation, 31 (6):541–564, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090060V.

Richard:1990:VNM

[2] M. J. Richard, R. J. Anderson, and G. C. Andrews. The vector-network method for the modelling of mechanical systems. Mathematics and Computers in Simulation, 31(6):565– 581, February 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090061M.

Rotella:1990:EAA

[3] Frédéric Rotella, Geneviève Dauphin-Tanguy, and Pierre Borne. Evaluation of the accuracy of analytic nonlinear optimal control laws. Mathematics and Computers in Simulation, 31 (6):583–592, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090062N.

Ames:1990:BRPa

[4] W. F. Ames. Book review: Phase retrieval and zero crossings (Mathematical methods in image reconstruction): N. E. Hurt, a volume in Mathematics and its Applications. Kluwer Academic Publishers, Dordrecht, 1989. 304 pp., Dfl 160, US$79, UK £49, ISBN 0-7923-0210-9. Mathematics and Computers in Simulation, 31 (6):593, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090063O.

Ames:1990:BREa

[5] W. F. Ames. Book review: Evolution and control in biological systems: A. B. Kurzhanski and K. Sigmund (Eds), Kluwer Academic Publishers, 1989. 189 pp., Dfl. 140, US$74, UK £44. ISBN 0-7923-0219-2. Mathematics and Computers in Simulation, 31(6):594, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090064P.

Ames:1990:BREb

[6] W. F. Ames. Book review: Elliptic systems and quasiconformal mappings, (Pure and applied mathematics series): H. Renelt, John Wiley & Sons, Chichester, 1988. 146 pp., UK £14.95. ISBN 0-471-9173-1. Mathematics and Computers in Simulation, 31(6):594, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090065Q.

Ames:1990:BRU

[7] W. F. Ames. Book review: Underwater acoustic data processing: Y. T. Chan (Ed.), Kluwer Academic Publishers, Dordrecht, 1989. 674 pp., Dfl 300, UK £97, US $173. ISBN 0-7923-0127-7. Mathematics and Computers in Simulation, 31(6):594, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090066R.

Ames:1990:BRPb

[8] W. F. Ames. Book review: Proofs and types: J.-Y. Girard, P. Taylor and Y. Lafont, Cambridge University Press, 1989. 176 pp., US $34.50. ISBN 0-521-37181-3. Mathematics and Computers in Simulation, 31(6): 594–595, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090067S.

Ames:1990:BRFa

[9] W. F. Ames. Book review: Fuzzy relation equations and their applications to knowledge engineering: A. Di Nola, S. Sessa, W. Pedrycz and E. Sanchez, Kluwer Academic Publishers, Dordrecht, 1989. 278 pp., Dfl. 150, US $79, UK £49. ISBN 0-7923-0307-5. Mathematics and Computers in Simulation, 31(6):595, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090068T.

Brezinski:1990:BRAa

[10] C. Brezinski. Book review: Approximation and optimization: A. Gomez et al. (Eds), Lecture Notes in Mathematics, vol. 1354, Springer-Verlag, Berlin, 1988. XII + 277 pages, DM 42.50. ISBN 3-540-50443-5. Mathematics and Computers in Simulation, 31(6):595, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090069U.

Brezinski:1990:BRE

[11] C. Brezinski. Book review: Elements of Computer algebra with applications: A. G. Akritas, Wiley, New York, 1989. XV + 425 pages, UK £33.75. ISBN 0-471-61163-8. Mathematics and Computers in Simulation, 31(6):596, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090071P.

Brezinski:1990:BRN

[12] C. Brezinski. Book review: Numerical analysis: H. R. Schwarz, Wiley, Chichester, 1989. XIII + 517 pages, UK £15.95. ISBN 0-471-92064-9. Mathematics and Computers in Simulation, 31(6):596, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090072Q.

Brezinski:1990:BRPa

[13] C. Brezinski. Book review: Principles of applied mathematics: J. P. Keener, Addison-Wesley, Redwood City, 1988. XV + 560 pages, US$48.50. ISBN 0-201-15674-1. Mathematics and Computers in Simulation, 31 (6):596, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090070Y.

Brezinski:1990:BRO

[14] C. Brezinski. Book review: Orthogonal polynomials and their applications: J. Vinuesa (Ed.), M. Dekker, New York, 1989. X + 207 pages. ISBN 0-8247-8161-9. Mathematics and Computers in Simulation, 31(6): 596–597, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090073R.

Brezinski:1990:BRAb

[15] C. Brezinski. Book review: Applied functional analysis: D. H. Griffel, Wiley, New York, 1981. 390 pages, UK £18.40. ISBN 0-85312-226-1. Mathematics and Computers in Simulation, 31(6):597, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090074S.

Brezinski:1990:BRAc

[16] C. Brezinski. Book review: Applied regression analysis in econometrics: H. E. Doran, M. Dekker, New York, 1989. VIII + 372 pages, US $89.75. ISBN 0-8247-8049-3. Mathematics and Computers in Simulation, 31(6):597, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090076U.

Brezinski:1990:BRC

[17] C. Brezinski. Book review: Continued fractions in statistical applications: K. O. Bowman and L. R. Shenton, M. Dekker, New York, 1989. IX + 330 pages. ISBN 0-8247-8120-1. Mathematics and Computers in Simulation, 31(6):597, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090075T.

Brezinski:1990:BRPb

[18] C. Brezinski. Book review: Polynomials: E. J. Barbeau, Springer-Verlag, Berlin, 1989. XI + 441 pages. ISBN 3-540-96919-5. Mathematics and Computers in Simulation, 31(6): 597–598, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090077V.

Brezinski:1990:BRD

[19] C. Brezinski. Book review: Deterministic and stochastic error bounds in numerical analysis: E. Novak, Lecture Notes in Mathematics vol. 1349, Springer-Verlag, Berlin, 1988. 113 pages, DM 23. ISBN 3-540-50368-4. Mathematics and Computers in Simulation, 31(6):598, February 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549090078W.

Brezinski:1990:BRG

[20] C. Brezinski. Book review: Geometry and codes: V. D. Goppa, Kluwer Academic Publishers, Dordrecht, 1988. IX + 157 pages, US $64.00. ISBN 90-277-2776-7. Mathematics and Computers in Simulation, 31(6):598, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090079X.

Brezinski:1990:BRR

[21] C. Brezinski. Book review: Ramanujan’s notebooks, part II: B. C. Berndt, Springer-Verlag, Berlin, 1989. XI + 359 pages, DM 168.00. ISBN 3-540-96794-X. Mathematics and Computers in Simulation, 31 (6):598, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475490900803.

Ames:1990:BRAa

[22] W. F. Ames. Book review: Algebraic analysis (2 volumes): M. Kashiwara and T. Kawai (Eds), Academic Press, San Diego, Calif., 1989. Vol. 1:472 pp., Vol. 2:478 pp., US$59.95 each., ISBN 0-12-400465-2 (Vol. 1), ISBN 0-12-400466-0 (Vol. 2). Mathematics and Computers in Simulation, 31(6):599, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090081S.

Ames:1990:BRFb

[23] W. F. Ames. Book review: Functional programming: T. Field and P. Harrison, Addison Wesley Publishing Co., Reading, Mass., 1988. 602 pp., US $37.75. ISBN 0-201-19249-7. Mathematics and Computers in Simulation, 31(6):599, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090083U.

Ames:1990:BRM

[24] W. F. Ames. Book review: Mathematics and control engineering of grinding technology (a volume in Mathematics and its applications — East European series): L. Keviczky, M. Hilger and J. Kolostori, Kluwer Academic Publishers, Dordrecht, 1989. 179 pp., US$67, UK £45, Dfl. 140. ISBN 0-7923-0051-3. Mathematics and Computers in Simulation, 31 (6):599, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090082T.

Ames:1990:BRCa

[25] W. F. Ames. Book review: Computer control of machines and processes: J. G. Bollinger and N. A. Duffie, Addison Wesley Publishing Co., Reading, Mass., 1988. 613 pp., US $53.75. ISBN 0-201-10645-0. Mathematics and Computers in Simulation, 31(6): 599–600, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090084V.

Ames:1990:BRPc

[26] W. F. Ames. Book review: Parallel program design — a foundation: K. M. Chandy and J. Misra, Addison Wesley Publishing Co., Reading, Mass., 1988. 516 pp., US$39.75. ISBN 0-201-05866-9. Mathematics and Computers in Simulation, 31 (6):600, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090085W.

Ames:1990:BRAb

[27] W. F. Ames. Book review: Algorithms (2nd edition): R. Sedgewick, Addison-Wesley Publishing Co., Reading, Mass., 1988. 657 pp, US$39.75. ISBN 0-201-06673-4. Mathematics and Computers in Simulation, 31 (6):601, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090088Z.

Ames:1990:BRNa

[28] W. F. Ames. Book review: Numerical methods in practice — using the NAG Library: T. Hopkins and C. Phillips, International Computer Science Series, Addison-Wesley Publishing Co., Reading Mass., 1988. 308 pp., US $26.95. ISBN 0-201-19248-9 (Paper). Mathematics and Computers in Simulation, 31(6):601, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090086X.

Ames:1990:BRNb

[29] W. F. Ames. Book review: Number theory, trace formulas and discrete groups: K. E. Aubert, E. Bombieri and D. Goldfeld, Academic Press, San Diego, California, 1989. 510 pp., US $69.95. ISBN 0-12-067570-6. Mathematics and Computers in Simulation, 31(6):601, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090087Y.

Ames:1990:BRLa

[30] W. F. Ames. Book review: LISP (third edition): P. H. Winston and B. K. P. Horn, Addison Wesley, Reading, Mass., 1989. 608 pp, Paperback, US$34.50. ISBN 0-201-08319-1. Mathematics and Computers in Simulation, 31(6): 601–602, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475490900892.

Ames:1990:BRCb

[31] W. F. Ames. Book review: Computer Algorithms — Introduction to design and analysis (second edition): S. Baase, Addison Wesley, Reading, Mass., 1988. 415 pp., US$40.95. ISBN 0-201-06035-3. Mathematics and Computers in Simulation, 31 (6):602, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475490900906.

Ames:1990:BRLb

[32] W. F. Ames. Book review: Lattice points (Mathematics and its applications, East European series): E. Kratzel, Kluwer Academic Publishers, Dordrecht, 1988. 320 pp, Dfl. 198, US$99.50, UK £58. ISBN 90-277-2733-3. Mathematics and Computers in Simulation, 31(6): 602, February 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090092W.

Ames:1990:BRS

[33] W. F. Ames. Book review: Solid modelling with DESIGNBASE: H. Chiyokura, Addison Wesley, Reading, Mass., 1988. 300 pp., US $39.75. ISBN 0-201-19245-4. Mathematics and Computers in Simulation, 31 (6):602, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090091V.

Ames:1990:BRCc

[34] W. F. Ames. Book review: A concrete approach to mathematical modelling: M. Mesterton-Gibbons, Addison-Wesley, Reading, Mass., 1989. 597 pp., US$48.50. ISBN 0-201-12910-8. Mathematics and Computers in Simulation, 31(6):603, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090093X.

Ames:1990:BRR

[35] W. F. Ames. Book review: Recent advances in geometric inequalities (Mathematics and its applications, East European series): D. S. Mitrinovic, J. E. Pecaric and V. Volenec, Kluwer Academic Publishers, Dordrecht, 1989. 710 pp., Dfl. 295, US$149, UK £90. ISBN 90-277-2565-9. Mathematics and Computers in Simulation, 31(6):603, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090095Z.

Ames:1990:NMD

[36] W. F. Ames. Newton’s method and dynamical systems: H.-O. Peitgen (Ed.), Kluwer Academic Publishers, Dordrecht, 1989. 226 pp., US $84, Dfl. 150, UK £48, ISBN 0-7923-0113-7. Mathematics and Computers in Simulation, 31(6):603, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090094Y.

Anonymous:1990:NIa

[37] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 31(6):605–622, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/0378475490900962.

Anonymous:1990:AIVa

[38] Anonymous. Author index to volume 31 (1989/90). Mathematics and Computers in Simulation, 31(6): 623–625, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475490900973.

Anonymous:1990:CVa

[39] Anonymous. Contents of volume 31 (1989/90). Mathematics and Computers in Simulation, 31(6):627– 629, February 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475490900984.

Anonymous:1990:PF

[40] Anonymous. Pages 541–630 (February 1990). Mathematics and Computers in Simulation, 31(6):??, February 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1990:EB

[41] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 32(1–2):ii, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475490902092.

Anonymous:1990:I

[42] Anonymous. Introduction. Mathematics and Computers in Simulation, 32(1–2):1–2, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090210A.

Gani:1990:EMS

[43] J. Gani. Epidemic modelling and simulation. Mathematics and Computers in Simulation, 32(1–2):3– 12, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090211Z.

Chapman:1990:CHM

[44] T. G. Chapman. Construction of hydrological models for natural systems management. Mathematics and Computers in Simulation, 32(1–2):

13–37, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475490902122.

Hornberger:1990:MCN

[45] G. M. Hornberger. Modelling complex natural processes with small observation sets: the case of acidification of surface waters in North America and Europe. Mathematics and Computers in Simulation, 32 (1–2):39–47, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475490902133.

Chatwin:1990:CAC

[46] P. C. Chatwin and P. J. Sullivan. Cloud-average concentration statistics. Mathematics and Computers in Simulation, 32(1–2):49– 57, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475490902144.

Taylor:1990:NAM

[47] J. A. Taylor. New approaches to modelling the gobal distribution of trace gases in the troposphere. Mathematics and Computers in Simulation, 32 (1–2):59–64, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475490902155.

Hurley:1990:FML

[48] P. Hurley and W. Physick. Fumigation modelling — a Lagrangian particle approach. Mathematics and Computers in Simulation, 32(1–2): 65–70, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475490902166.

Ghassemi:1990:SSW

[49] F. Ghassemi, A. J. Jakeman, and G. Jacobson. Simulation of sea water intrusion. Mathematics and Computers in Simulation, 32(1–2): 71–76, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475490902177.

Anderssen:1990:DAL

[50] R. S. Anderssen, C. R. Dietrich, and P. A. Green. Designing artificial lakes as pollution control devices. Mathematics and Computers in Simulation, 32(1–2):77–82, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475490902188.

Cuddy:1990:MEE

[51] S. M. Cuddy, P. Laut, J. R. Davis, P. A. Whigham, J. Goodspeed, and T. Duell. Modelling the environmental effects of training on a major Australian army base. Mathematics and Computers in Simulation, 32(1–2):83–88, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475490902199.

Finnigan:1990:MWF

[52] J. J. Finnigan, D. Neil, B. G. Lees, R. J. Croome, and M. Woodgate. Modelling the wind flow pattern around a parabolic sand dune. Mathematics and Computers in Simulation, 32 (1–2):89–94, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090220D.

Noble:1990:OBF

[53] S. Noble and B. Lees. Offshore bar formation in a shallow water, limited fetch environment. Mathematics and Computers in Simulation, 32 (1–2):95–100, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475490902214.

Chia:1990:SFQ

[54] E. Chia. Some fundamental questions in stochastic simulation of weather sequences, with applications to ANU-CLOUD. Mathematics and Computers in Simulation, 32(1–2):101– 106, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475490902225.

Retnam:1990:SRI

[55] M. T. Retnam, B. J. Williams, and G. Kuczera. Simulation of a really integrated spatial-temporal rainfall field for use in rainfall–runoff models. Mathematics and Computers in Simulation, 32(1–2):107–112, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475490902236.

Guenni:1990:SWM

[56] L. Guenni, D. Charles-Edwards, C. W. Rose, R. Braddock, and W. Hogarth. Stochastic weather modelling: a phenomenological approach. Mathematics and Computers in Simulation, 32 (1–2):113–118, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475490902247.

Beer:1990:MRF

[57] T. Beer. Modelling rainfall as a fractal process. Mathematics and Computers in Simulation, 32(1–2): 119–124, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475490902258.

Hutchinson:1990:RCS

[58] M. F. Hutchinson. Robust calibration of seasonally varying stochastic weather models using periodic smoothing splines. Mathematics and Computers in Simulation, 32(1–2): 125–130, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475490902269.

Mahendrarajah:1990:DOM

[59] S. Mahendrarajah. Dynamic optimization with multiple objectives and the valuation of public goods of community reservoirs. Mathematics and Computers in Simulation, 32 (1–2):131–136, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090227A.

McKenney:1990:EAP

[60] D. McKenney and M. Common. The economic analysis of public sector forest management using linear programming: selected results from an Australian application. Mathematics and Computers in Simulation, 32 (1–2):137–142, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090228B.

Henderson-Sellers:1990:MTW

[61] B. Henderson-Sellers and D. R. Gallagher. Modelling tools for water management. Mathematics and Computers in Simulation, 32(1–2): 143–148, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549090229C.

Jellett:1990:SAC

[62] P. M. Jellett. Stimulated annealing for a constrained allocation problem. Mathematics and Computers in Simulation, 32(1–2):149– 154, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090230G.

Roberts:1990:PMS

[63] S. Roberts. A particle method for a scalar advection diffusion equation. Mathematics and Computers in Simulation, 32(1–2):155– 160, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475490902317.

Yang:1990:MCW

[64] J. Z. Yang and I. White. A model of coupled water, water vapor and heat transport in porous media and a simulation analysis of evaporation. Mathematics and Computers in Simulation, 32(1–2):161–166, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475490902328.

Bai:1990:PET

[65] J. Bai, A. J. Jakeman, and J. A. Taylor. Percentile estimation of the threeparameter gamma and lognormal distributions: Methods of moments versus maximum likelihood. Mathematics and Computers in Simulation, 32 (1–2):167–172, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475490902339.

Doran:1990:CSM

[66] D. G. Doran. A comparison of some methods of stochastic modelling of solar energy photovoltaic storage systems. Mathematics and Computers in Simulation, 32(1–2): 173–178, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549090234A.

Haritos:1990:FEF

[67] N. Haritos and D. T. He. A finite element formulation for cables suitable for dynamic modelling. Mathematics and Computers in Simulation, 32 (1–2):179–184, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090235B.

Hansen:1990:EVA

[68] D. V. Hansen, P. J. Dillon, and P. I. Brooker. The estimation variance of the areal mean of a sampled random field. Mathematics and Computers in Simulation, 32(1–2): 185–190, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549090236C.

Dietrich:1990:SKS

[69] C. R. Dietrich. Sensitivity of Kriging and Spline interpolation to data perturbations. Mathematics and Computers in Simulation, 32(1–2):191– 196, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090237D.

Bai:1990:EME

[70] J. Bai, A. J. Jakeman, and M. McAleer. The effects of misspecification in estimating the percentiles of some two- and three-parameter distributions. Mathematics and Computers in Simulation, 32(1–2):197–202, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090238E.

Kohn:1990:NEG

[71] R. Kohn and C. F. Ansley. The nonparametric estimation of growth curves. Mathematics and Computers in Simulation, 32(1–2):203– 208, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090239F.

Ng:1990:RSE

[72] C. N. Ng and P. C. Young. Recursive smoothing of environmental time-series. Mathematics and Computers in Simulation, 32(1–2):209– 214, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090240J.

Benyon:1990:MCO

[73] P. R. Benyon. Monte Carlo and other methods for nonlinear non-Gaussian estimation. Mathematics and Computers in Simulation, 32(1– 2):215–220, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090241A.

Noble:1990:TTH

[74] I. R. Noble, A. D. Moore, and T. J. Montgomery. Trilogy — three hierarchical models of plant succession. Mathematics and Computers in Simulation, 32(1–2):221– 224, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090242B.

Mackey:1990:MFS

[75] B. G. Mackey and T. Bayes. A modelling framework for the spatial extension of ecological relations in vegetation studies. Mathematics and Computers in Simulation, 32 (1–2):225–229, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090243C.

Norton:1990:WMA

[76] T. W. Norton and J. E. Williams. Wildlife management in Australia: New developments and opportunities using computer-based generic models. Mathematics and Computers in Simulation, 32(1–2):231– 235, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090244D.

Green:1990:GAL

[77] D. G. Green, R. E. Reichelt, J. van der Laan, and B. W. Macdonald. A generic approach to landscape modelling. Mathematics and Computers in Simulation, 32(1–2):237– 242, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090245E.

Granger:1990:PMG

[78] K. J. Granger. Process modelling and geographic information systems: breathing life into spatial analysis. Mathematics and Computers in Simulation, 32(1–2):243– 247, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090246F.

Stockwell:1990:PCE

[79] D. R. B. Stockwell and D. G. Green. Parallel computing in ecological simulation. Mathematics and Computers in Simulation, 32(1–2):249– 254, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090247G.

Bowyer:1990:DDC

[80] J. K. Bowyer. Designing a database for coastal planning and management. Mathematics and Computers in Simulation, 32(1–2):255– 260, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549090248H.

Anonymous:1990:ICEa

[81] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 32(1–2):261– 271, April 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549090249I.

Anonymous:1990:PA

[82] Anonymous. Pages 1–272 (April 1990). Mathematics and Computers in Simulation, 32(1–2):??, April 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Godfrey:1990:IIL

[83] Keith R. Godfrey and Michael J. Chapman. Identifiability and indistinguishability of linear compartmental models. Mathematics and Computers in Simulation, 32(3):273– 295, June 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549090185L.

Ohkita:1990:PLA

[84] Masaaki Ohkita and Yasuhiro Kobayashi. Piecewise-linear approximation of solution of linear differential equations by Walsh functions. Mathematics and Computers in Simulation, 32 (3):297–308, June 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090186M.

Taha:1990:NSN

[85] Thiab R. Taha. Numerical simulation of the nonlinear Schrödinger equation. Mathematics and Computers in Simulation, 32(3):309– 312, June 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090187N.

Anonymous:1990:BRPa

[86] Anonymous. Book review: Probability, statistics and mathematics (Papers in Honor of Samuel Karlin): T. W. Anderson, K. B. Athreya and D. L. Iglehart (Eds.), Academic Press, San Diego, 1989. 371 pp., US$39.95, ISBN 0-12-058470-0. Mathematics and Computers in Simulation, 32(3):313, June 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090188O.

Anonymous:1990:BRIa

[87] Anonymous. Book review: Introduction to parallel programming: S. Brawer, Academic Press, San Diego, 1989. 422 pp., US$39.95, ISBN 0-12-128470-0. Mathematics and Computers in Simulation, 32 (3):313–314, June 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090189P.

Anonymous:1990:BRAa

[88] Anonymous. Book review: Alternate realities — mathematical models of nature and man: J. L. Casti, John Wiley & Sons, New York, 1989. 493 pp., GB £25.40, ISBN 0-471-61842-X. Mathematics and Computers in Simulation, 32(3):314, June 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090190T.

Anonymous:1990:BRMa

[89] Anonymous. Book review: The mathematical structure of Raster graphics: E. L. Fiume, Academic Press, San Diego, 1989. 221 pp., US $29.95, ISBN 0-12-257960-7. Mathematics and Computers in Simulation, 32 (3):314–315, June 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090191K.

Anonymous:1990:BRE

[90] Anonymous. Book review: Exploring the geometry of nature: E. Rietman, Windcrest Books, Blue Ridge Summit, Pa. (marketed by J. Wiley & Sons), 1989. 199 pp., US$16.30, ISBN 0-8306-3137-2 (paper), 0-8306-9137-5 (hardback). Mathematics and Computers in Simulation, 32 (3):315, June 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549090192L.

Anonymous:1990:BRCa

[91] Anonymous. Book review: Computational geometry — curve and surface modeling: Su Bu-quing and Liu Ding-Yuan, (Translation of 1979 Chinese Edition by Chang Geng-Zhe). Academic Press, San Diego, 1989. 295 pp., US$39.95, ISBN 0-12-675610-4. Mathematics and Computers in Simulation, 32(3):315–316, June 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090193M.

Anonymous:1990:BRAb

[92] Anonymous. Book review: Asymptotic approximations of integrals: R. Wong, Academic Press, San Diego, 1989. 543 pp. US $69.95, ISBN 0-12-762535-6. Mathematics and Computers in Simulation, 32(3):316, June 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090194N.

Anonymous:1990:BRHa

[93] Anonymous. Book review: How to write and publish a scientific paper: R. A. Day, Cambridge University Press, Cambridge, 1989. ISBN 0-521-36760-3, XI + 211 pages, £7.95. Mathematics and Computers in Simulation, 32(3):316, June 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090195O.

Anonymous:1990:BRCb

[94] Anonymous. Book review: Chambers concise dictionary of scientists:

D. Millar et al., Cambridge University Press, Cambridge, 1989. ISBN 1-85296-354-9, 461 pages, £14.95. Mathematics and Computers in Simulation, 32(3):317, June 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090196P.

Anonymous:1990:BRCc

[95] Anonymous. Book review: From cardinals to chaos. Reflections on the life and legacy of Stanislaw Ulam: N. G. Cooper (ed.), Cambridge University Press, Cambridge, 1989. ISBN 0-521-36734-4, 320 pages, £15. Mathematics and Computers in Simulation, 32(3):317, June 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090197Q.

Anonymous:1990:BRD

[96] Anonymous. Book review: Difference equations from differential equations: W. J. Lick, Springer-Verlag, Berlin, 1989. ISBN 3-540-50739-6, X + 282 pages, DM 68. Mathematics and Computers in Simulation, 32(3):317, June 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090198R.

Anonymous:1990:BRIb

[97] Anonymous. Book review: An introduction to the numerical analysis of spectral methods: B. Mercier, Springer-Verlag, Berlin, 1989. ISBN 3-540-51106-7, V + 154 pages, DM 39. Mathematics and Computers in Simulation, 32(3):317–318, June 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090199S.

Anonymous:1990:BRLa

[98] Anonymous. Book review: Linear least squares computations: R. W. Farebrother, Marcel Dekker, New York, 1988. ISBN 0-8247-7661-5, XIII + 293 pages. Mathematics and Computers in Simulation, 32 (3):318, June 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090201S.

Anonymous:1990:BRNa

[99] Anonymous. Book review: Nonlinear Lp-norm estimation: R. Gonin and A. H. Money, Marcel Dekker, New York, 1989. ISBN 0-8247-8125-2, VIII + 300 pages. Mathematics and Computers in Simulation, 32 (3):318, June 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090202T.

Anonymous:1990:BRW

[100] Anonymous. Book review: Wavelets: J. M. Combes et al. (eds.), Springer-Verlag, Berlin 1989. ISBN 3-540-51159-8, IX + 315 pages. Mathematics and Computers in Simulation, 32(3):318, June 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475490902003.

Anonymous:1990:BRAc

[101] Anonymous. Book review: ABS projection algorithms: J. Abaffy and E. Spedicato, Ellis Horwood, Chichester, 1989. ISBN 0-7458-0423-3, 220 pages, £39.95. Mathematics and Computers in Simulation, 32 (3):319, June 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549090203U.

Anonymous:1990:BRFa

[102] Anonymous. Book review: Fibonacci and Lucas numbers, and the golden section: S. Vajda, Ellis Horwood, Chichester, 1989. ISBN 0-7458-0715-1, 190 pages, £25. Mathematics and Computers in Simulation, 32 (3):319, June 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549090204V.

Anonymous:1990:BRIc

[103] Anonymous. Book review: An introduction to Hankel operators: J. R. Partington, Cambridge University Press, Cambridge, 1989. ISBN 0-521-36791-3, 103 pages, £7.50. Mathematics and Computers in Simulation, 32(3):319, June 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090205W.

Anonymous:1990:BRId

[104] Anonymous. Book review: An introduction to numerical analysis: K. E. Atkinson, Wiley, New York, 1989. ISBN 0-471-50023-2 XVI + 693 pages, 16.50. Mathematics and Computers in Simulation, 32 (3):319, June 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090206X.

Anonymous:1990:BRCd

[105] Anonymous. Book review: Chaotic evolution and strange attractors: D. Ruelle, Cambridge University Press, Cambridge, 1989. ISBN 0-521-36830-8, XI + 96 pages, £8.95. Mathematics and Computers in Simulation, 32(3):320, June 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090207Y.

Anonymous:1990:NIb

[106] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 32(3):321–333, June 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090208Z.

Anonymous:1990:PJ

[107] Anonymous. Pages 273–334 (June 1990). Mathematics and Computers in Simulation, 32(3):??, June 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Abdalkhani:1990:NEV

[108] J. Abdalkhani. A note on examples of Volterra integral equation with exact solution. Mathematics and Computers in Simulation, 32(4): 335–337, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475490901389.

Depiante:1990:SGS

[109] E. V. Depiante. A symbolic generator of state equations for modeling of electrical systems. Mathematics and Computers in Simulation, 32 (4):339–349, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090139A.

Nakano:1990:FDA

[110] Kazushi Nakano, Ichiro Tsurumi, and Setsuo Sagara. Finite-dimensional approach to estimation of functional parameters in a parabolic system. Mathematics and Computers in Simulation, 32(4):351–357, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090140E.

Noye:1990:SET

[111] John Noye. Some explicit threelevel finite-difference simulations of advection. Mathematics and Computers in Simulation, 32(4):359– 372, October 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475490901415.

Rubinstein:1990:OSS

[112] Reuven Y. Rubinstein and Alexander Shapiro. Optimization of static simulation models by the score function method. Mathematics and Computers in Simulation, 32(4):373– 392, October 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475490901426.

Murthy:1990:ASS

[113] C. Siva Ram Murthy and V. Rajaraman. Analytical and simulation studies of a multiprocessor system for highspeed numerical computations. Mathematics and Computers in Simulation, 32(4):393–401, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475490901437.

Tzafestas:1990:NSE

[114] Spyros Tzafestas and George Kechriotis. A numeric-symbolic expert system for 2-D and 3-D object recognition in robotic applications. Mathematics and Computers in Simulation, 32(4):403–418, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475490901448.

Anonymous:1990:BRAd

[115] Anonymous. Book review: Analysis I: R. V. Gamkrelidze (ed.), Springer, Berlin, 1989. 238 pp., DM 128, ISBN 3-540-17008-1. Mathematics and Computers in Simulation, 32(4):419, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475490901459.

Anonymous:1990:BRHb

[116] Anonymous. Book review: Handbook of numerical analysis, volume I: P. G. Ciarlet and J. L. Lions (eds.), North-Holland, Amsterdam, 1990. 652 pp., US $107.25, ISBN 0-444-70366-7. Mathematics and Computers in Simulation, 32(4): 419, October 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090146A.

Anonymous:1990:BRSa

[117] Anonymous. Book review: Statistical tools for simulation practitioners: J. P. C. Kleijnen, Marcel Dekker, New York, 1987. 429 pp., ISBN 0-8247-7333-0. Mathematics and Computers in Simulation, 32(4): 419–420, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090147B.

Anonymous:1990:BRHc

[118] Anonymous. Book review: A history of mathematics: C. B. Boyer and U. C. Merzbach, Wiley, New York, 1989. 762 pp., £14.95, ISBN 0-471-50357-6. Mathematics and Computers in Simulation, 32(4): 420, October 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090149D.

Anonymous:1990:BRIe

[119] Anonymous. Book review: Iterative methods for simultaneous inclusion of polynomial zeros: M. Petkovic, Lecture Notes in Mathematics 1387. Springer, Berlin, 1990. 263 pp., DM 45, ISBN 3-540-51485-6. Mathematics and Computers in Simulation, 32(4):420, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090148C.

Anonymous:1990:BRPb

[120] Anonymous. Book review: Properties of estimators for the gamma function: K. O. Bowman and L. R. Shenton, Marcel Dekker, New York, 1988. 268 pp., ISBN 0-8247-7556-2. Mathematics and Computers in Simulation, 32(4):420, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090150H.

Anonymous:1990:BRCe

[121] Anonymous. Book review: Computers and mathematics: E. Kaltofen and S. M. Watt (eds.), Springer, Berlin, 1989. 400 pp., DM 78, ISBN 3-540-97019-3. Mathematics and Computers in Simulation, 32(4):420– 421, October 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475490901518.

Anonymous:1990:BRB

[122] Anonymous. Book review: The book of prime number records: P. Ribenboim, Springer, Berlin, 1989, 2nd ed. 479 pp., DM 98, ISBN 3-540-97042-8. Mathematics and Computers in Simulation, 32(4):421, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090153A.

Anonymous:1990:BRMb

[123] Anonymous. Book review: Mathematics and its history: J. Stillwell, Springer, Berlin, 1989. 371 pp., DM 98, ISBN 3-540-96981-0. Mathematics and Computers in Simulation, 32(4):421, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475490901529.

Anonymous:1990:BRMc

[124] Anonymous. Book review: Multivariate approximation theory IV: C. K. Chui, W. Schempp and K. Zeller (eds.), Birkhäuser, Basel, 1989. 342 pp., SFr 98, ISBN 3-7643-2384-1. Mathematics and Computers in Simulation, 32(4):421, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090154B.

Anonymous:1990:BRO

[125] Anonymous. Book review: Ontic, a knowledge representation system for mathematics: D. A. McAllester, MIT press, Cambridge, MA, 1989. 158 pp., US $36.95, ISBN 0-262-13235-4. Mathematics and Computers in Simulation, 32(4):421–422, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090155C.

Anonymous:1990:BRAe

[126] Anonymous. Book review: Applications of supercomputers in engineering: algorithms, computer systems and user experience: C. A. Brebbia and A. Peters (eds.), Elsevier, Amsterdam, 1989. 226 pp., US $95, ISBN 0-444-88110-7. Mathematics and Computers in Simulation, 32(4):422, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090158F.

Anonymous:1990:BRGa

[127] Anonymous. Book review: General theory of irregular curves: A. D. Alexandrov and Yu.G. Reshetnyak, Kluwer, Dordrecht, 1989. 288 pp., US $99, ISBN 90277-28119. Mathematics and Computers in Simulation, 32(4):422, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090157E.

Anonymous:1990:BRPc

[128] Anonymous. Book review: Practical numerical algorithms for chaotic systems: T. S. Parker and L. O. Chua, Springer, Berlin, 1989. 348 pp., DM 98, ISBN 3-540-96689-7. Mathematics and Computers in Simulation, 32(4):422, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090156D.

Anonymous:1990:BRFb

[129] Anonymous. Book review: Fortran for scientists and engineers: D. G. Weinman, PWS-Kent, Boston, 1989. 531 pp., US $14.95, ISBN 0-87835-337-2. Mathematics and Computers in Simulation, 32(4): 423, October 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549090160K.

Anonymous:1990:BRLb

[130] Anonymous. Book review: L’analyse harmonique, son développement historique: J. P. Pier, Masson, Paris, 1990. 344 pp., FF 370, ISBN 2-225-81899-1. Mathematics and Computers in Simulation, 32(4): 423, October 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549090159G.

Anonymous:1990:BRMd

[131] Anonymous. Book review: Methods and programs for mathematical functions: S. L. Moshier, Wiley, New York, 1989. 415 pp., £16.95, ISBN 0-74358-0805-0. Mathematics and Computers in Simulation, 32(4): 423, October 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090161B.

Anonymous:1990:BRMe

[132] Anonymous. Book review: Mathematical methods for scientists and engineers: P. B. Kahn, Wiley, New York, 1990. 469 pp., £43.15, ISBN 0-471-62305-9. Mathematics and Computers in Simulation, 32(4): 423, October 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090162C.

Anonymous:1990:BRAf

[133] Anonymous. Book review: Analytic theory of continued fractions III: L. Jacobsen (ed.), Lecture Notes in Mathematics 1406. Springer, Berlin, 1989. 142 pp., DM 25, ISBN 3-540-51830-4. Mathematics and Computers in Simulation, 32(4):424, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090164E.

Anonymous:1990:BRMf

[134] Anonymous. Book review: Mathematical methods for engineers: R. K. Livesley, Ellis Horwood, Chichester, 1989. 462 pp., £15.95, ISBN 0-7458-0714-3. Mathematics and Computers in Simulation, 32(4): 424, October 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090163D.

Anonymous:1990:BRT

[135] Anonymous. Book review: Topics in functional analysis and applications: S. Kesavan, Wiley, New York, 1989, 267 pp., £20.15, ISBN 0-470-21050-8 (Wiley, NY); 81-224-0062-0 (Wiley Eastern). Mathematics and Computers in Simulation, 32(4):424, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090165F.

Anonymous:1990:BRAg

[136] Anonymous. Book review: Algebras of pseudodifferential operators, volume 43: B. A. Plamenevskii, Mathematics and its Applications (Soviet Series). Kluwer, Dordrecht, 1989. 291 pp., Dfl. 210, US $99, £69, ISBN 0-7923-0231-1. Mathematics and Computers in Simulation, 32(4): 424–425, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090166G.

Anonymous:1990:BRFc

[137] Anonymous. Book review: Flow modelling in industrial processes: A. W. Bush, B. A. Lewis and M. D. Warren (eds.), Ellis Horwood, Chichester (Halsted Press, New York), 1989. 256 pp., £35, ISBN 0-7458-0477-2 (Horwood); 0-470-21498-8 (Halsted). Mathematics and Computers in Simulation, 32(4):425, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090167H.

Anonymous:1990:BRHd

[138] Anonymous. Book review: Homogenization: averaging processes in periodic media: N. Bakhvalov and G. Panasenko, Mathematics and its Applications (Soviet Series). Kluwer, Dordrecht, 1989. 366 pp., Dfl. 285, US $139, £94, ISBN 0-7923-0049-1. Mathematics and Computers in Simulation, 32(4):425–426, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090168I.

Anonymous:1990:BRFd

[139] Anonymous. Book review: Finite element methods for viscous incompressible flows: M. D. Gunzburger, Academic Press, San Diego, 1989. 269 pp., US $44.50, ISBN 0-12-307350-2. Mathematics and Computers in Simulation, 32(4):426, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090169J.

Anonymous:1990:BRMg

[140] Anonymous. Book review: Mathematical modelling: a case study approach: R. R. Clements, Cambridge Univ. Press, Cambridge, 1989. 166 pp., US $44.50, ISBN 0-521-34340-2. Mathematics and Computers in Simulation, 32(4):426–427, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090170N.

Anonymous:1990:BRGb

[141] Anonymous. Book review: Generatingfunctionology: H. S. Wilf, Academic Press, San Diego, 1990. 184 pp., US $29.95, ISBN-0-12-751955-6. Mathematics and Computers in Simulation, 32(4):427, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090171E.

Anonymous:1990:BRMh

[142] Anonymous. Book review: The Monte Carlo method for semiconductor device simulation: C. Jacoboni and P. Lugli, Springer, Wien, 1989. Ös 1300, DM 186, ISBN 3-211-82110-4. Mathematics and Computers in Simulation, 32(4):427, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090172F.

Anonymous:1990:BRMi

[143] Anonymous. Book review: Mathematical methods in computer aided geometric design: T. Lyche and L. L. Schumaker (eds.), Academic Press, Boston, 1989. 611 pp., US $49.95, ISBN 0-12-460515-X. Mathematics and Computers in Simulation, 32(4): 427–428, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090173G.

Anonymous:1990:BRCf

[144] Anonymous. Book review: Coordinating user interfaces for consistency: J. Nielsen (ed.), Academic Press, Boston, 1989. 142 pp., US $24.95, ISBN 0-12-518400-X. Mathematics and Computers in Simulation, 32(4):428, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090174H.

Anonymous:1990:BRIf

[145] Anonymous. Book review: Iterative methods for large linear systems: D. R. Kincaid and L. J. Hayes (eds.), Academic Press, Boston, 1990. 319 pp., US$49.95, ISBN 0-12-407475-8. Mathematics and Computers in Simulation, 32(4):428, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090175I.

Anonymous:1990:BRLc

[146] Anonymous. Book review: Lectures on buildings: M. Ronan, Academic Press, Boston, 1989. 201 pp., US $27.95, ISBN-0-12-594750-X. Mathematics and Computers in Simulation, 32(4):428–429, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090176J.

Anonymous:1990:BRCg

[147] Anonymous. Book review: Conformal invariance and string theory: P. Dita and V. Georgescu (eds.), Academic Press, Boston, 1989. 557 pp., US $59.95, ISBN 0-12-218100-X. Mathematics and Computers in Simulation, 32(4):429, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090177K.

Anonymous:1990:BRCh

[148] Anonymous. Book review: The complete C++ primer: K. Weiskamp and B. Flamig, Academic Press, Boston, 1990. 524 pp., US $45, ISBN-0-12-742687-6 (Paper). Mathematics and Computers in Simulation, 32(4): 429, October 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549090179M.

Anonymous:1990:BRNb

[149] Anonymous. Book review: Nonlinear waves and solitons: M. Toda, Mathematics and its Applications (Japanese Series). Kluwer, Dordrecht (co-published with KTK Science Publishers, Tokyo), 1989. 371 pp., Dfl. 260, US$129, £86, ISBN 0-7923-0442-X. Mathematics and Computers in Simulation, 32(4): 429, October 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549090178L.

Anonymous:1990:BRCi

[150] Anonymous. Book review: Change of representation and inductive bias: D. P. Benjamin (ed.), Kluwer, Dordrecht, 1990. 353 pp., Dfl. 180, US $70, £50.50, ISBN 0-7923-9055-5. Mathematics and Computers in Simulation, 32(4):429–430, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090180Q.

Anonymous:1990:BRSb

[151] Anonymous. Book review: Selected papers of Antoni Zygmund: A. Hulanicki, P. Wojtaszcyk and W. Zelazko (eds.), (3 volumes), Mathematics and its Applications (East European Series). Kluwer, Dordrecht, 1989. Vol. I (443 pp.), Vol. II (444 pp.), Vol. III (457 pp.), Dfl. 695, US$359, £229, ISBN (Set) 0-7923-0474-8; Vol. I 0-7923-0471-3; Vol. II 0-7923-0472-1; Vol. III 0-7923-0473-X. Mathematics and Computers in Simulation, 32(4):430, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090181H.

Anonymous:1990:BRAh

[152] Anonymous. Book review: Advances in software science and technology: I. Nakata (Journal Editor), Volume 1 (Journal). Academic Press, Boston, 1990. 215 pp., US$49, ISBN 0-12-037101-4. Mathematics and Computers in Simulation, 32(4): 431, October 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090182I.

Anonymous:1990:NIc

[153] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 32(4):433, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090183J.

Anonymous:1990:ICEb

[154] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 32(4):434– 445, October 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549090184K.

Anonymous:1990:PO

[155] Anonymous. Pages 335–446 (October 1990). Mathematics and Computers in Simulation, 32(4):??, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Walter:1990:F

[156] Eric Walter. Foreword. Mathematics and Computers in Simulation, 32 (5–6):447, December 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090001Y.

Walter:1990:EPB

[157] Eric Walter and Hélène Piet-Lahanier. Estimation of parameter bounds from bounded-error data: a survey. Mathematics and Computers in Simulation, 32(5–6):449–468, December 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549090002Z.

Broman:1990:CAI

[158] V. Broman and M. J. Shensa. A compact algorithm for the intersection and approximation of N-dimensional polytopes. Mathematics and Computers in Simulation, 32(5–6):469– 480, December 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475490900032.

Mo:1990:FRA

[159] S. H. Mo and J. P. Norton. Fast and robust algorithm to compute exact polytope parameter bounds. Mathematics and Computers in Simulation, 32(5– 6):481–493, December 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475490900043.

Piet-Lahanier:1990:ERC

[160] H. Piet-Lahanier and E. Walter. Exact recursive characterization of feasible parameter sets in the linear case. Mathematics and Computers in Simulation, 32(5–6):495– 504, December 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475490900054.

Clement:1990:RMS

[161] Thierry Clement and Sylviane Gentil. Recursive membership set estimation for output-error models. Mathematics and Computers in Simulation, 32(5– 6):505–513, December 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475490900065.

Rao:1990:RDO

[162] Ashok K. Rao and Yih-Fang Huang. Recent developments in optimal bounding ellipsoidal parameter estimation. Mathematics and Computers in Simulation, 32(5–6):515– 526, December 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475490900076.

Norton:1990:PBT

[163] J. P. Norton and S. H. Mo. Parameter bounding for time-varying systems. Mathematics and Computers in Simulation, 32(5–6):527– 534, December 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475490900087.

Keesman:1990:MSE

[164] Karel Keesman. Membership-set estimation using random scanning and principal component analysis. Mathematics and Computers in Simulation, 32(5–6):535–543, December 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475490900098.

Smit:1990:BEA

[165] M. K. Smit and J. W. Verhoof. A bounded-error approach to accuracy analysis in ellipsometry. Mathematics and Computers in Simulation, 32(5– 6):545–551, December 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549090010G.

Piet-Lahanier:1990:CNC

[166] H. Piet-Lahanier and E. Walter. Characterization of non-connected parameter uncertainty regions. Mathematics and Computers in Simulation, 32(5– 6):553–560, December 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475490900117.

Belforte:1990:ISS

[167] G. Belforte, B. Bona, and V. Cerone. Identification, structure selection and validation of uncertain models with set-membership error description. Mathematics and Computers in Simulation, 32(5–6):561– 569, December 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475490900128.

Pronzato:1990:EDB

[168] Luc Pronzato and Eric Walter. Experiment design for bounded-error models. Mathematics and Computers in Simulation, 32(5–6):571– 584, December 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475490900139.

Tempo:1990:OAS

[169] R. Tempo and A. Vicino. Optimal algorithms for system identification: a review of some recent results. Mathematics and Computers in Simulation, 32(5–6):585– 595, December 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549090014A.

Kouvaritakis:1990:BED

[170] B. Kouvaritakis and M. S. Trimboli. Bounded-error data, and frequency response design. Mathematics and Computers in Simulation, 32(5– 6):597–607, December 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090015B.

Anonymous:1990:NId

[171] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 32(5–6):609–627, December 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/037847549090016C.

Anonymous:1990:AIVb

[172] Anonymous. Author index to volume 32 (1990). Mathematics and Computers in Simulation, 32(5–6): 629–633, December 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549090017D.

Anonymous:1990:CVb

[173] Anonymous. Contents of volume 32 (1990). Mathematics and Computers in Simulation, 32(5–6):635– 638, December 1990. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549090018E.

Anonymous:1990:PD

[174] Anonymous. Pages 447–638 (December 1990). Mathematics and Computers in Simulation, 32(5–6):??, December 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1991:EB

[175] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 33(1):ii, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190019Y.

Rubinstein:1991:MIS

[176] Reuven Y. Rubinstein. Modified importance sampling for performance evaluation and sensitivity analysis of computer simulation models. Mathematics and Computers in Simulation, 33(1):1–22, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475491900204.

Korn:1991:DFG

[177] Granino A. Korn. Design of functiongenerating mapping networks by interactive neural-network simulation.

Mathematics and Computers in Simulation, 33(1):23–31, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190021T.

Midy:1991:CSE

[178] P. Midy and Y. Yakovlev. Computing some elementary functions of a complex variable. Mathematics and Computers in Simulation, 33(1): 33–49, July 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549190022U.

Riganti:1991:STR

[179] R. Riganti. A solution technique for random and nonlinear inverse heat conduction problems. Mathematics and Computers in Simulation, 33(1):51–64, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190023V.

Barron:1991:DTC

[180] R. M. Barron and R. K. Naeem. 2-D transonic calculations on a flowbased grid system. Mathematics and Computers in Simulation, 33(1): 65–67, July 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549190024W.

Anonymous:1991:BRU

[181] Anonymous. Book review: The user’s directory of computer networks: T.

L. LaQuey, Digital Press of the University of Texas System, Austin, TX, 1990. 630 pp., ISBN 1-55558-047-5. Mathematics and Computers in Simulation, 33(1):69, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190025X.

Anonymous:1991:BRSa

[182] Anonymous. Book review: Stochastic models: D. P. Heyman, Elsevier, Amsterdam, 1990. 734 pp., US$115, Dfl.295, ISBN 0-444-87473-9. Mathematics and Computers in Simulation, 33(1):69–70, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190026Y.

Anonymous:1991:BRHa

[183] Anonymous. Book review: A history of scientific computing: S. G. Nash (ed.), Addison-Wesley, Reading, MA, 1990. ISBN 0-201-50814-1. Mathematics and Computers in Simulation, 33(1):70, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190027Z.

Anonymous:1991:BRMa

[184] Anonymous. Book review: Mastering technical writing: J. C. Mancuso, Addison-Wesley, Reading, MA, 1990. 249 pp., US$27.95, ISBN 0-201-52350-7. Mathematics and Computers in Simulation, 33(1):70, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475491900282.

Anonymous:1991:BRMb

[185] Anonymous. Book review: Mathematical structures of nonlinear science — an introduction: M. S. Berger, Kluwer, Dordrecht, 1990. 421 pp., Dfl.220, US$99, UK£79, ISBN 0-7923-0728-3. Mathematics and Computers in Simulation, 33(1):70–71, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475491900293.

Anonymous:1991:BRPa

[186] Anonymous. Book review: Partial differential equations of applied mathematics: E. Zauderer, Wiley, New York, 1989, 2nd ed. 891 pp., US$51.30 (HB), $29.75 (PB), ISBN 0-471-61298-7. Mathematics and Computers in Simulation, 33(1):71, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475491900307.

Anonymous:1991:BRCa

[187] Anonymous. Book review: Control of quantum-mechanical processes and systems: A. G. Butkovskiy and Yu.I. Samoilenko, Mathematics and its Applications (Soviet Series), Kluwer, Dordrecht, 1990. 232 pp., Dfl.175, US$99, UK£65, ISBN 0-7923-0689-9. Mathematics and Computers in Simulation, 33(1):71–72, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549190031W.

Anonymous:1991:BRFa

[188] Anonymous. Book review: Fundamentals of mathematical evolutionary genetics: Yu.M. Svirezhev and V. P. Passekov, Mathematics and its Applications (Soviet Series), Kluwer, Dordrecht, 1990. 395pp., Dfl.240, US$129, UK£74, ISBN 90-277-2772-4. Mathematics and Computers in Simulation, 33(1):72, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190032X.

Anonymous:1991:BRNa

[189] Anonymous. Book review: Numerical analysis — an introduction: L. Elden and L. Wittmeyer-Koch, Academic press, Boston, 1990. 347 pp., US$39.95, ISBN 0-12-236430-9. Mathematics and Computers in Simulation, 33(1):72, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190033Y.

Anonymous:1991:BRPb

[190] Anonymous. Book review: Progress in fuzzy sets and systems: W. H. Janko, M. Roubens and H. J. Zimmerman (eds.), Kluwer, Dordrecht, 1990. 188 pp., Dfl.150, US$87, UK£54, ISBN 0-7923-0730-5. Mathematics and Computers in Simulation, 33 (1):73, July 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475491900363.

Anonymous:1991:BRRa

[191] Anonymous. Book review: The rains model of acidification: J. Alcamo, R. Shaw and L. Hordijk (eds.), (Science and Strategies in Europe). Kluwer, Dordrecht, 1990. 402 pp., Dfl.200, US$117.50, UK£72, ISBN 07923-0781-X (HB); 07923-0782-8 (PB). Mathematics and Computers in Simulation, 33(1):73, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549190034Z.

Anonymous:1991:BRSb

[192] Anonymous. Book review: Semiconductor equations: P. A. Markowich, C. A. Ringhofer and C. Schmeiser, Springer, Wien, 1990. 248 pp., ÖS686, DM98, ISBN 3-211-82157-0. Mathematics and Computers in Simulation, 33(1):73, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475491900352.

Anonymous:1991:BRFb

[193] Anonymous. Book review: Fuzzy sets in information retrieval and cluster analysis: S. Miyamoto, Kluwer, Dordrecht, 1990. 259 pp., Dfl.159, US$99, UK£64, ISBN 0-7923-0721-6. Mathematics and Computers in Simulation, 33(1):73–74, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475491900374.

Anonymous:1991:BRMc

[194] Anonymous. Book review: Matrix perturbation theory: G. W. Stewart and J. Sun, Academic Press, Boston, 1990. 365 pp., US$39.95, ISBN 0-12-670230-6. Mathematics and Computers in Simulation, 33(1):74, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475491900385.

Anonymous:1991:BRIa

[195] Anonymous. Book review: Introduction to algebraic quantum field theory: S. S. Horuzhy, Mathematics and its Applications (Soviet Series). Kluwer, Dordrecht, 1990. 301 pp., Dfl.260, US$139, UK£84, ISBN 90-277-2722-8. Mathematics and Computers in Simulation, 33(1):74–75, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475491900396.

Anonymous:1991:BRCb

[196] Anonymous. Book review: Curves and surfaces for computer-aided geometric design — a practical guide: G. Farin, Academic Press, Boston, 1990, 2nd ed. 444 pp., US$39.95, ISBN 0-12-249051-7. Mathematics and Computers in Simulation, 33(1):75, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190040A.

Anonymous:1991:BRRb

[197] Anonymous. Book review: Regression estimators — a comparative study: M. H. J. Gruber, Academic Press, Boston, 1990. 347 pp., US$49.95, ISBN 0-12-304752-8. Mathematics and Computers in Simulation, 33(1):75–76, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190041Z.

Anonymous:1991:BRCc

[198] Anonymous. Book review: Computation of curves and surfaces: W. Dahmen, M. Gasca and C. A. Micchelli (eds.), Kluwer, Dordrecht, 1989. 536 pp., Dfl.265, US$149, UK£95, ISBN 0-7923-0724-0. Mathematics and Computers in Simulation, 33 (1):76, July 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475491900422.

Anonymous:1991:BRSc

[199] Anonymous. Book review: Symmetries of partial differential equations — conservation laws, applications, algorithms: A. M. Vinogradov, (ed.), Kluwer, Dordrecht, 1989. 451 pp., ISBN 0-7923-0594-9. Mathematics and Computers in Simulation, 33(1):76, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475491900433.

Anonymous:1991:BRE

[200] Anonymous. Book review: Envisioning information: E. R. Tufte, Graphics Press, Cheshire, CT, 1990. 126 pp., US$48 (multiple copies $45), no ISBN given. Mathematics and Computers in Simulation, 33 (1):77, July 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475491900455.

Anonymous:1991:BRMd

[201] Anonymous. Book review: Mathematical and statistical developments of evolutionary theory: S. Lessard (ed.), Kluwer, Dordrecht, 1990. 363 pp., Dfl.195, US$99, UK£67, ISBN 0-7923-0595-7. Mathematics and Computers in Simulation, 33 (1):77, July 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475491900444.

Anonymous:1991:BRNb

[202] Anonymous. Book review: Nonnegative matrices in dynamic systems: A. Berman, M. Neumann and R. J. Stern, Wiley, New York, 1989. 167 pp., £29.85, ISBN 0-471-62074-2. Mathematics and Computers in Simulation, 33(1):77, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475491900466.

Anonymous:1991:BRDa

[203] Anonymous. Book review: The desktop fractal design (Macintosh version): M. Barnsley, Academic Press, San Diego, 1990. Software plus 40 pp. Handbook, US$39.95, ISBN 0-12-079064-5. Mathematics and Computers in Simulation, 33(1):78, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475491900488.

Anonymous:1991:BRL

[204] Anonymous. Book review: Linear operators in spaces with an indefinite metric: T. Ya. Azizov and I. S. Iokhvidov (translated by E. R. Dawson), Wiley, Chichester, 1989. 304 pp., £39.95, ISBN 0-471-92129-7. Mathematics and Computers in Simulation, 33(1):78, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475491900477.

Anonymous:1991:BRMe

[205] Anonymous. Book review: Model building in mathematical programming: H. P. Williams, Wiley, Chichester, 1990, 3rd ed. paper £14.50, ISBN 0-471-92581-0; cloth £29.95, ISBN 0-471-92580-2. Mathematics and Computers in Simulation, 33(1): 78–79, July 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475491900499.

Anonymous:1991:BRB

[206] Anonymous. Book review: Boundary elements in mechanical and electrical engineering: C. A. Brebbia and A. Chaudouet-Miranda (eds.), Computational Mechanics Publications, Southampton, and Springer, Berlin, 1990. 569 pp., ISBN 1-85312-062-6 (Computational Mechanics), ISBN 3-540-52645-5 (Springer). Mathematics and Computers in Simulation, 33 (1):79, July 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475491900514.

Anonymous:1991:BRNc

[207] Anonymous. Book review: Number theory and cryptography: J. H. Loxton, (ed.), London Mathematical Society Lecture Note Series 154. Cambridge Univ. Press, Cambridge, 1990. 235 pp., US$27.95, ISBN 0-521- 39877-0. Mathematics and Computers in Simulation, 33(1):79, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190050D.

Anonymous:1991:BRCd

[208] Anonymous. Book review: Computational acoustics: D. Lee, A. Cakmac and R. Vichnevetsky (eds.), North-Holland, Amsterdam, 1990. Volume I. Ocean Acoustic Models and Supercomputing. 276 pp., ISBN 0-444-88720-2. Volume II. Scattering, Gaussian Beams and Aeroacoustics. 322 pp., ISBN 0-444-88721-0. Volume III. Seismo-ocean acoustics and Modelling. 343 pp., ISBN 0-444-88722-9. Mathematics and Computers in Simulation, 33(1):79–80, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475491900525.

Anonymous:1991:BRAa

[209] Anonymous. Book review: Asymptotic and computational analysis: R.

Wong (ed.), Lecture Notes in Pure and Applied Mathematics 124. Marcel Dekker, New York, 1990. XII + 755 pp., ISBN 0-8247-8347-6. Mathematics and Computers in Simulation, 33(1):80, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475491900569.

Anonymous:1991:BRCe

[210] Anonymous. Book review: Chebyshev polynomials: T. J. Rivlin, Wiley, New York, 1990, 2nd ed. XIII + 249 pp., ISBN 0-471-62896-4. Mathematics and Computers in Simulation, 33(1):80, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475491900536.

Anonymous:1991:BRFc

[211] Anonymous. Book review: Finite element approximation of variational problems and applications: M. Krizel and P. Neittaanmäki, Longham, Harlow, 1990. 239 pp., £36.00, ISBN 0-582-05666-7. Mathematics and Computers in Simulation, 33 (1):80, July 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475491900558.

Anonymous:1991:BRNd

[212] Anonymous. Book review: Numerical analysis, a second course: J. M. Ortega, SIAM, Philadelphia, PA, 1990. XIII + 201 pp., ISBN 0-89871-250-5. Mathematics and Computers in Simulation, 33(1):80, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475491900547.

Anonymous:1991:BRAb

[213] Anonymous. Book review: Analysis II: R. V. Gamkrelidze (ed.), Springer, Berlin, 1990. VII + 255 pp., DM128.00, ISBN 3-540-18179-2. Mathematics and Computers in Simulation, 33(1):81, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190057A.

Anonymous:1991:BRIb

[214] Anonymous. Book review: Iterative methods for large linear systems: D. R. Kincaid and L. J. Hayes (eds.), Academic Press, New York, 1990. xxx + 319 pp., $49.95, ISBN 0-12-407475-8. Mathematics and Computers in Simulation, 33(1):81, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190058B.

Anonymous:1991:BRIc

[215] Anonymous. Book review: Introduction to applied nonlinear dynamical systems and chaos: S. Wiggins, Springer, Berlin, 1990. XIV + 672 pp., DM98, ISBN 3-540-97003-7. Mathematics and Computers in Simulation, 33(1):81, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190060G.

Anonymous:1991:BRJ

[216] Anonymous. Book review: Joseph Liouville, 1809–1882, master of pure and applied mathematics: J. Lützen, Springer, Berlin, 1990. XIX + 884 pp., DM228, ISBN 0-387-97180-7. Mathematics and Computers in Simulation, 33(1):81, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549190059C.

Anonymous:1991:BRAc

[217] Anonymous. Book review: Applied and industrial mathematics: R. Spigler (ed.), Kluwer, Dordrecht, 1990. 374 pp., US$136, ISBN 0-7923-0521-3. Mathematics and Computers in Simulation, 33(1):82, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475491900639.

Anonymous:1991:BRFd

[218] Anonymous. Book review: Fractal geometry: K. Falconer, Wiley, Chichester, 1990. XXII + 288 pp., £19.95, ISBN 0-471- 92287-0. Mathematics and Computers in Simulation, 33(1):82, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475491900617.

Anonymous:1991:BRHb

[219] Anonymous. Book review: Handbook of applicable mathematics, supplement: W. Ledermann (ed.), Wiley, Chichester, 1990. XIX + 479 pp., £52.50, ISBN 0-471-91825-3. Mathematics and Computers in Simulation, 33(1):82, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475491900628.

Anonymous:1991:BRTa

[220] Anonymous. Book review: Theory of complex functions: R. Remmert, Springer, Berlin, 1990. 440 pp., DM118, ISBN 3-540-97195-5. Mathematics and Computers in Simulation, 33(1):82, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190064A.

Anonymous:1991:BRMf

[221] Anonymous. Book review: Mathematics of linear and nonlinear systems: D. J. Bell, Clarendon Press, Oxford, 1990. 304 pp., £40.00, ISBN 0-19-856332-9. Mathematics and Computers in Simulation, 33(1):83, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190067D.

Anonymous:1991:BRNe

[222] Anonymous. Book review: Numbers: H.-D. Ebbinghaus et al., Springer, Berlin, 1990. 395 pp., DM98, ISBN 3-540-97202-1. Mathematics and Computers in Simulation, 33 (1):83, July 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549190065B.

Anonymous:1991:BRPc

[223] Anonymous. Book review: Programming in mathematica: R. Maeder, Addison-Wesley, Reading, MA, 1990. 267 pp., ISBN 0-201-51002-2. Mathematics and Computers in Simulation, 33(1):83, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549190066C.

Anonymous:1991:BRFe

[224] Anonymous. Book review: Fortran 90 explained: M. Metcalf and J. Reid, Oxford Univ. Press, Oxford, 1990. 294 pp., £14.95, ISBN 0-19-853772-7. Mathematics and Computers in Simulation, 33(1):83–84, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549190068E.

Anonymous:1991:BRCf

[225] Anonymous. Book review: Computational methods and function theory: S. Ruscheweyh et al. (eds.), Lecture Notes in Mathematics 1435. Springer, Berlin, 1990. 211 pp., DM37, ISBN 3-540-52768-0. Mathematics and Computers in Simulation, 33 (1):84, July 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549190071A.

Anonymous:1991:BRO

[226] Anonymous. Book review: Operator theory in function spaces: Kehe Zhu, Marcel Dekker, New York, 1990.

258 pp., ISBN 0-8247-8411-1. Mathematics and Computers in Simulation, 33(1):84, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190069F.

Anonymous:1991:BRTb

[227] Anonymous. Book review: Theorems and counterexamples in mathematics: B. R. Gelbaum and J. M. H. Olmsted, Springer, Berlin, 1990. 345 pp., DM84, ISBN 3-540-97342-7. Mathematics and Computers in Simulation, 33(1):84, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190070J.

Anonymous:1991:BRNf

[228] Anonymous. Book review: Numerical continuation methods: E. L. Allogower and K. Georg, Springer, Berlin, 1990. 388 pp., DM128, ISBN 3-540-12760-7. Mathematics and Computers in Simulation, 33(1):84–85, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190072B.

Anonymous:1991:BRDb

[229] Anonymous. Book review: The discipline of curiosity: J. Groen et al. (eds.), Elsevier, Amsterdam, 1990. 156 pp., ISBN 0-444-88861-6. Mathematics and Computers in Simulation, 33(1):85, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549190073C.

Anonymous:1991:BRFf

[230] Anonymous. Book review: Finite element methods (part 1): P. G. Ciarlet and J. L. Lions (eds.), Handbook of Numerical Analysis, Volume 2. North-Holland, Amsterdam, 1991. 928 pp., ISBN 0-444-70365-9. Mathematics and Computers in Simulation, 33(1):85, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549190074D.

Anonymous:1991:NIa

[231] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 33(1):87, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549190075E.

Anonymous:1991:FAC

[232] Anonymous. First announcement and call for papers. Mathematics and Computers in Simulation, 33(1): 88–89, July 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549190076F.

Anonymous:1991:FA

[233] Anonymous. First announcement. Mathematics and Computers in Simulation, 33(1):90, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190077G.

Anonymous:1991:MHG

[234] Anonymous. A meeting to honor Garrett Birkhoff on his 80th birthday. Mathematics and Computers in Simulation, 33(1):91, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190078H.

Anonymous:1991:IWCa

[235] Anonymous. 13th IMACS World Congress on Computation and Applied Mathematics. Mathematics and Computers in Simulation, 33(1): 92–93, July 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549190079I.

Anonymous:1991:NIb

[236] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 33(1):94–95, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190080M.

Anonymous:1991:IWCb

[237] Anonymous. 13th IMACS World Congress on Computation and Applied Mathematics. Mathematics and Computers in Simulation, 33 (1):96, July 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549190081D.

Anonymous:1991:ITC

[238] Anonymous. IMACS technical commitees. Mathematics and Computers in Simulation, 33(1):97–99, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/037847549190082E.

Anonymous:1991:ICEa

[239] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 33(1):100– 107, July 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549190083F.

Anonymous:1991:ICEb

[240] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 33(1):108, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/037847549190084G.

Anonymous:1991:AM

[241] Anonymous. Application for membership. Mathematics and Computers in Simulation, 33(1):109, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/037847549190085H.

Anonymous:1991:AIA

[242] Anonymous. Application for individual or associate membership: (to apply for Affiliated Society status, write to the IMACS Affiliated Societies Coordinating Office, c/o IM-SOR, Bldg 321, Technical University of Denmark, 2800 Lyngby, Denmark). Mathematics and Computers in Simulation, 33(1):110, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190086I.

Anonymous:1991:PJ

[243] Anonymous. Pages 1–110 (July 1991). Mathematics and Computers in Simulation, 33(1):??, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Venets:1991:P

[244] V. I. Venets. Preface. Mathematics and Computers in Simulation, 33(2):111, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190155V.

Burkov:1991:SSP

[245] V. N. Burkov, S. P. Kolesnikov, and A. V. Schepkin. Support system for processes of designing organizational mechanisms. Mathematics and Computers in Simulation, 33 (2):113–120, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190156W.

Demetrovics:1991:LPD

[246] J. Demetrovics, Gy. Gyepesi, L. Hannák, T. Remzsö, and F. Urbánski. LATOR — professional database management system for local networks. Mathematics and Computers in Simulation, 33 (2):121–128, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549190157X.

Demetrovics:1991:ISI

[247] J. Demetrovics, L. Hannák, A. Heppes, T. Remzsö, and F. Urbánszki. Information system for insurance companies (ABLAK). Mathematics and Computers in Simulation, 33(2):129– 137, August 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549190158Y.

McHenry:1991:RTC

[248] William K. McHenry. R-technology and case: analysis and perspective. Mathematics and Computers in Simulation, 33(2):139–150, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/037847549190159Z.

Simov:1991:MBL

[249] G. Simov and I. Popchev. Modulibased language for building problemsolving scenarios. Mathematics and Computers in Simulation, 33(2):151– 164, August 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475491901605.

Wajs:1991:CAC

[250] W. Wajs and M. Capik. Computeraided control system design for drilling systems. Mathematics and Computers in Simulation, 33(2):165– 172, August 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549190161U.

Wolcott:1991:APW

[251] Peter Wolcott. Ada: progress in the West and the East. Mathematics and Computers in Simulation, 33 (2):173–174, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190162V.

A:1991:BRM

[252] W. F. A. Book review: Modern statistical systems, and GPSS simulation (the first course): Z. A. Karian and E. J. Dudewicz, including a disk: Student version of GPSS/PC and case studies, Freeman, New York, 1991. 468 pp., UK£39.95, ISBN 0-7167-8232-4. Mathematics and Computers in Simulation, 33(2):175, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190164X.

A:1991:BRRa

[253] W. F. A. Book review: Recent advances in Fourier analysis and its applications: J. S. Byrnes and J. L. Byrnes (eds.), (NATO ASI Series). Kluwer, Dordrecht, 1990. 686 pp., Dfl.300, US$174, UK£108, ISBN 0-7923-0875-1. Mathematics and Computers in Simulation, 33(2):175, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190163W.

A:1991:BRAa

[254] W. F. A. Book review: Asymptotic distribution of eigenvalues of differential operators: S. Levendorskiĭ, Kluwer, Dordrecht, 1990. 279 pp., Dfl.220, US$112, UK£74, ISBN 0-7923-0539-6. Mathematics and Computers in Simulation, 33 (2):175–176, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190165Y.

A:1991:BRSa

[255] W. F. A. Book review: Spinors and calibrations: F. R. Harvey, Academic Press, Boston, 1990. 323 pp., US$39.35, ISBN 0-12-329650-1. Mathematics and Computers in Simulation, 33(2):176, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190166Z.

A:1991:BRSb

[256] W. F. A. Book review: Soliton phenomenology: V. G. Makhankov, mathematics and its applications (Soviet series), Kluwer, Dordrecht, 1990. 452 pp., Dfl.260, US$149, UK£91, ISBN 90-277-2830-5. Mathematics and Computers in Simulation, 33(2): 176, August 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475491901672.

A:1991:BRCa

[257] W. F. A. Book review: A course in simulation: S. M. Ross, Macmillan, New York, 1990. 202 pp., ISBN 0-02-403891-1. Mathematics and Computers in Simulation, 33(2): 177, August 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475491901683.

A:1991:BRRb

[258] W. F. A. Book review: Renormalized quantum field theory: O. I. Zavialov, mathematics and its applications (Soviet series), Kluwer, Dordrecht, 1990. 524 pp., Dfl.295, US$149, UK£92, ISBN 90-277-2758-9. Mathematics and Computers in Simulation, 33(2):177, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475491901694.

A:1991:BRQa

[259] W. F. A. Book review: Queueing analysis for computers: S. Yashkov, (in Russian), Radio i Svyaz, Moscow, 1989, 216 pp., ISBN 5-256-00304-6. Mathematics and Computers in Simulation, 33(2):177–178, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475491901708.

A:1991:BRCb

[260] W. F. A. Book review: Convolutional calculus: I. H. Dimovski, mathematics and its applications (East European Series). Kluwer, Dordrecht, 1990. 184 pp., Dfl.150, US$79, UK£53, ISBN 0-7923-0623-6. Mathematics and Computers in Simulation, 33(2): 178, August 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549190171X.

A:1991:BRQb

[261] W. F. A. Book review: Quasiconformal mappings and Sobolev spaces: V. M. Gol’dshte and Yu.G. Reshetnyak, mathematics and its applications (Soviet series), Kluwer, Dordrecht, Dfl.260, US133, UK£85, ISBN 0-7923-0543-4. Mathematics and Computers in Simulation, 33(2):178– 179, August 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549190172Y.

A:1991:BRO

[262] W. F. A. Book review: Ocean waves mechanics, computational fluid dynamics, and mathematical modelling: M. Rahman (ed.), Computational Mechanics Publications, Southampton/Springer, Berlin, 1990. 954 pp., UK£145, ISBN 1-85312-065-0. Mathematics and Computers in Simulation, 33(2):179, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190173Z.

A:1991:BRVa

[263] W. F. A. Book review: Visualization — The second computer revolution: R. M. Friedhoff and W. Benzon, Freeman, New York, 1991. 215 pp., UK£17.95, ISBN 0-7167-2231-3. Mathematics and Computers in Simulation, 33(2):179, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475491901742.

A:1991:BRI

[264] W. F. A. Book review: Inventory models (theory and decision library): A. Chikan (ed.), Kluwer, Dordrecht, 1990, 418 pp., Dfl.325, US$189, UK£115, ISBN 0-7923-0494-2. Mathematics and Computers in Simulation, 33(2):180, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475491901775.

A:1991:BRN

[265] W. F. A. Book review: Numerical methods for conservation laws: R. J. LeVeque, Birkhäuser, Basel, 1990. 214 pp. (paperback), SFR.38, ISBN 3-7643-2464-3. Mathematics and Computers in Simulation, 33(2): 180, August 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475491901753.

A:1991:BRU

[266] W. F. A. Book review: The unreal life of Oscar Zariski: C. Parikh, Academic Press, Boston, 1990, 264 pp., US$32.95, ISBN 0-12-545030-3. Mathematics and Computers in Simulation, 33(2):180, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475491901764.

A:1991:BRAb

[267] W. F. A. Book review: Approximation theory and functional analysis: C. K. Chui (ed.), Academic Press, Boston, 1990. 247 pp., US$49.95, ISBN 0-12-174583-X. Mathematics and Computers in Simulation, 33(2): 181, August 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475491901797.

A:1991:BRP

[268] W. F. A. Book review: Perturbations — theory and methods: James A. Murdock, Wiley, Chichester, 1991, 509 pp., UK£43.65, ISBN 0-471-61294-4. Mathematics and Computers in Simulation, 33(2):181, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475491901786.

A:1991:BRL

[269] W. F. A. Book review: LISP-STAT (an object-oriented environment for statistical computing and dynamic graphics): L. Tierney, Wiley, Chichester, 1990. 397 pp., UK£32.15, ISBN 0-471-50916-7. Mathematics and Computers in Simulation, 33 (2):181–182, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190180B.

A:1991:BRSc

[270] W. F. A. Book review: Solitons in molecular systems: A. S. Davydov, (second edition). Kluwer, Dordrecht, 1991. 413 pp., Dfl.250, US$149, UK£87, ISBN 0-7923-1029-2. Mathematics and Computers in Simulation, 33(2):182, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475491901812.

A:1991:BRVb

[271] W. F. A. Book review: Vector Lyapunov functions and stability analysis of nonlinear systems: V. Lakshmikantham, V. M. Matrosov and S. Sivasundaram, Kluwer, Dordrecht, 1991. 172 pp., Dfl.120, US$79, UK£41, ISBN 0-7923-1152-3. Mathematics and Computers in Simulation, 33(2): 182, August 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475491901823.

A:1991:BRAc

[272] W. F. A. Book review: Applied and industrial mathematics: R. Spigler (ed.), Kluwer, Dordrecht, 1990. 374 pp., Dfl.220, US$136, UK£77, ISBN 0-7923-0521-3. Mathematics and Computers in Simulation, 33 (2):182–183, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475491901834.

A:1991:BRCc

[273] W. F. A. Book review: Computer modelling in the environmental sciences: D. G. Farmer and M. J. Rycroft, Clarendon Press, Oxford, 1991. 379 pp., UK£45, ISBN 0-19-853394-2. Mathematics and Computers in Simulation, 33(2):183, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475491901845.

Anonymous:1991:ACP

[274] Anonymous. Announcement and call for papers. Mathematics and Computers in Simulation, 33(2):185– 195, August 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475491901856.

Anonymous:1991:PA

[275] Anonymous. Pages 111–195 (August 1991). Mathematics and Computers in Simulation, 33(2):??, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Singh:1991:CDE

[276] Arindama Singh. Comments on decouplings and Euler’s one-step schemes. Mathematics and Computers in Simulation, 33(3):197–203, November 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549190118M.

Ballesteros:1991:NAE

[277] F. Ballesteros. A numerical approach to evaluate the describing function (DF). Mathematics and Computers in Simulation, 33(3):205– 208, November 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549190119N.

Liu:1991:ECN

[278] I-Shih Liu and I. Suliciu. Energy control of the numerical solutions of an elastic oscillator. Mathematics and Computers in Simulation, 33(3): 209–221, November 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190120R.

Popescu:1991:SCG

[279] Ileana Popescu and Ion Vaduva. A survey on computer generation of some classes of stochastic processes. Mathematics and Computers in Simulation, 33(3):223–241, November 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549190121I.

Harlow:1991:NSR

[280] D. G. Harlow and T. J. Delph. The numerical solution of random initialvalue problems. Mathematics and Computers in Simulation, 33(3):243– 258, November 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549190122J.

Anonymous:1991:BRHc

[281] Anonymous. Book review: Hausdorff approximations: Bl. Sendov, Kluwer, Dordrecht, 1990. 364 pp., Dfl.230, US$134, UK£81, ISBN 0-7923-0901-4. Mathematics and Computers in Simulation, 33(3):259, November 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549190123K.

Anonymous:1991:BRGa

[282] Anonymous. Book review: Gradient Estimation via Perturbation Analysis: P. Glasserman, Kluwer, Dordrecht, 1991. 221 pp., Dfl.116, US$55, UK£32.75, ISBN 0-7923-9095-4. Mathematics and Computers in Simulation, 33(3):259–260, November 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549190124L.

Anonymous:1991:BRNg

[283] Anonymous. Book review: Numerical methods for engineers: D. V. Griffiths and I. M. Smith, CRC Press, Boca Raton, 1991. 330 pp., ISBN 0-8493-8610-1. Mathematics and Computers in Simulation, 33 (3):260, November 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190127O.

Anonymous:1991:BRRc

[284] Anonymous. Book review: Random processes with independent increments: A. V. Skorohod, Mathematics and its Applications (Soviet Series). Kluwer, Dordrecht, 1991. 279 pp., Dfl.190, US$118, UK£65, ISBN 0-7923-0340-7. Mathematics and Computers in Simulation, 33 (3):260, November 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190125M.

Anonymous:1991:BRSd

[285] Anonymous. Book review: Simulation of water based thermal solar systems (EURSOL — An interactive program): W. L. Dutre, Kluwer, Dordrecht, 1991. 256 pp., Dfl.150, US$96, UK£52, ISBN 0-7923-1236-8. Mathematics and Computers in Simulation, 33(3):260, November 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190126N.

Anonymous:1991:BRDc

[286] Anonymous. Book review: Discrete mathematics for computer scientists: J. K. Truss, Addison-Wesley, Reading, MA, 1991. 565 pp., US$49.50, ISBN 0-201-17564-9. Mathematics and Computers in Simulation, 33 (3):261, November 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190129Q.

Anonymous:1991:BRNh

[287] Anonymous. Book review: Numerical methods for free boundary problems: P. Neittaanmaki (ed.), Birkhäuser, Basel, 1991. 439 pp., Sfr.128, ISBN 3-7643-2641-7 (Basel); 0-8176-2641-7 (Boston). Mathematics and Computers in Simulation, 33 (3):261, November 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549190128P.

Anonymous:1991:BRGb

[288] Anonymous. Book review: Graphbased proof procedures for Horn clauses: S. Raatz, Birkhäuser, Boston, 1990. 147 pp., Sfr.52, ISBN 0-8176-3530-0 (Boston), 3-7643-3530-0 (Basel). Mathematics and Computers in Simulation, 33(3):261–262, November 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549190130U.

Anonymous:1991:BRCg

[289] Anonymous. Book review: Continuous system modeling: François E. Cellier, Springer, Berlin, 1991. 755 pp., US$69.50, ISBN 0-387-97502-0. Mathematics and Computers in Simulation, 33(3):262, November 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549190131L.

Anonymous:1991:NIc

[290] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 33(3):263–274, November 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/037847549190132M.

Anonymous:1991:PN

[291] Anonymous. Pages 197–274 (November 1991). Mathematics and Computers in Simulation, 33(3):??, November 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Artamonov:1991:OCA

[292] Eugeniy Artamonov. Organization of computer-aided design systems. Mathematics and Computers in Simulation, 33(4):275–279, December 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549190105C.

Gudas:1991:FRI

[293] S. Gudas. A framework for research of information processing hierarchy in enterprise. Mathematics and Computers in Simulation, 33(4): 281–285, December 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190106D.

Karibskii:1991:MDL

[294] A. V. Karibskii. Managing the development of large-scale systems. Mathematics and Computers in Simulation, 33 (4):287–293, December 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190107E.

Kerekfy:1991:SAF

[295] Pál Kerékfy and Tibor Remzśmó. On some applications of form management. Mathematics and Computers in Simulation, 33(4):295–302, December 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549190108F.

Kistlerov:1991:CAD

[296] V. L. Kistlerov, P. I. Kitsul, and B. M. Miller. Computer-aided design of the optical devices control systems based on the language of algebraic computations FLAC. Mathematics and Computers in Simulation, 33 (4):303–307, December 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190109G.

Kovacs:1991:SAT

[297] Vilmos Kovács and George L. Kovács. Structured analysis technique and technology in control systems design. Mathematics and Computers in Simulation, 33(4):309–316, December 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549190110O.

Kulba:1991:UPN

[298] V. V. Kulba and A. R. Shvetsov. Using Petri-nets for data processing systems analysis and synthesis. Mathematics and Computers in Simulation, 33 (4):317–321, December 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190111F.

Kulmagambetov:1991:DMI

[299] A. R. Kulmagambetov and I. R. Kulmagambetov. Design of medical information system for dynamic control. Mathematics and Computers in Simulation, 33(4):323–329, December 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549190112G.

Uzhastov:1991:CAD

[300] I. A. Uzhastov. Computer-aided design of distributed databases. Mathematics and Computers in Simulation, 33 (4):331–338, December 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549190113H.

Yuditskiy:1991:LCF

[301] S. A. Yuditskiy and S. S. Pokalev. Logical control of flexible integrated manufacturing. Mathematics and Computers in Simulation, 33(4):339– 347, December 1991. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549190114I.

Anonymous:1991:NId

[302] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 33(4):349–356, December 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/037847549190115J.

Anonymous:1991:PD

[303] Anonymous. Pages 275–356 (December 1991). Mathematics and Computers in Simulation, 33(4):??, December 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Mayer:1992:P

[304] D. G. Mayer and A. J. Jakeman. Preface. Mathematics and Computers in Simulation, 33(5–6): 357, April 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549290121V.

Jakeman:1992:SAM

[305] A. J. Jakeman, G. M. Hornberger, I. G. Littlewood, P. G. Whitehead, J. W. Harvey, and K. E. Bencala. A systematic approach to modelling the dynamic linkage of climate, physical catchment descriptors and hydrologic response components. Mathematics and Computers in Simulation, 33 (5–6):359–366, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290122W.

Possingham:1992:MSM

[306] H. P. Possingham, I. Davies, I. R. Noble, and T. W. Norton. A metapopulation simulation model for assessing the likelihood of plant and animal extinctions. Mathematics and Computers in Simulation, 33(5–6): 367–372, April 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290123X.

Ludwig:1992:ERS

[307] J. A. Ludwig, R. E. Sinclair, and I. R. Noble. Embedding a rangeland simulation model within a decision support system. Mathematics and Computers in Simulation, 33 (5–6):373–378, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290124Y.

Norton:1992:HMS

[308] T. W. Norton and J. E. Williams. Habitat modelling and simulation for nature conservation: a need to deal systematically with uncertainty. Mathematics and Computers in Simulation, 33 (5–6):379–384, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290125Z.

Stockwell:1992:ISR

[309] David R. B. Stockwell and Ian R. Noble. Induction of sets of rules from animal distribution data: a robust and informative method of data analysis. Mathematics and Computers in Simulation, 33(5–6):385– 390, April 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475492901262.

Neave:1992:EFS

[310] H. M. Neave, R. B. Cunningham, T. W. Norton, and H. A. Nix. Evaluation of field sampling strategies for estimating species richness by Monte Carlo methods. Mathematics and Computers in Simulation, 33 (5–6):391–396, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475492901273.

Barford:1992:SAC

[311] J. P. Barford, P. J. Phillips, and C. Harbour. Simulation of animal cell metabolism. Mathematics and Computers in Simulation, 33(5– 6):397–402, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475492901284.

Shen:1992:DUM

[312] Zuo R. Shen. The development and use of models and expert systems to aid the control of crop pests in China. Mathematics and Computers in Simulation, 33(5–6):403–413, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475492901295.

Gerik:1992:OCP

[313] T. J. Gerik, L. J. Wade, W. D. Rosenthal, and R. L. Vanderlip. Optimising cultural practices for grain sorghum in relation to climatic risk at three locations in the United States, using the Sorkam model. Mathematics and Computers in Simulation, 33 (5–6):415–419, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475492901309.

Clift:1992:SME

[314] A. D. Clift and M. A. Terras. A simulation model to estimate effects of farm management and pest populations on yields of cultivated mushrooms. Mathematics and Computers in Simulation, 33(5–6):421–426, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290131Y.

Carroll:1992:MSE

[315] C. Carroll, M. Littleboy, and M. Halpin. Minimising soil erosion and runoff by maximising cropping opportunities. Mathematics and Computers in Simulation, 33(5–6):427– 432, April 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290132Z.

Wells:1992:OCC

[316] A. T. Wells and A. B. Hearn. OZ-COT: a cotton crop simulation model for management. Mathematics and Computers in Simulation, 33(5–6): 433–438, April 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475492901332.

Mayer:1992:ACS

[317] D. G. Mayer, M. G. Atzeni, and D. G. Butler. Adaptation of climex for spatial screwworm fly population dynamics. Mathematics and Computers in Simulation, 33(5–6): 439–444, April 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475492901343.

Atzeni:1992:SIR

[318] M. G. Atzeni, D. G. Mayer, and D. G. Butler. Sterile insect release method — optimal strategies for eradication of screwworm fly. Mathematics and Computers in Simulation, 33 (5–6):445–450, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475492901354.

Die:1992:DSC

[319] David J. Die and Reg A. Watson. Dissipation of spatial closure benefits as a result of non-compliance. Mathematics and Computers in Simulation, 33 (5–6):451–456, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475492901365.

Monypenny:1992:MDM

[320] Richard Monypenny. Modelling of dynamic management for decision support. Mathematics and Computers in Simulation, 33(5–6):457– 462, April 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475492901376.

Littleboy:1992:USL

[321] M. Littleboy, M. J. Grundy, M. J. Bryant, D. O. Gooding, and B. W. Carey. Using spatial land resource data and computer simulation modelling to evaluate sustainability of wheat cropping for a portion of the eastern Darling Downs, Queensland. Mathematics and Computers in Simulation, 33 (5–6):463–468, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475492901387.

Anderssen:1992:LMA

[322] Bob Anderssen. Linking mathematics with applications: the comparative assessment process. Mathematics and Computers in Simulation, 33 (5–6):469–475, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475492901398.

Robinson:1992:TRC

[323] J. Brett Robinson. Testing and recalibrating a simple model of daily rainfall for use in eastern Australia. Mathematics and Computers in Simulation, 33(5–6):477–482, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290140C.

Bai:1992:UEV

[324] J. Bai, A. J. Jakeman, and M. McAleer. On the use of extreme value distributions for predicting the upper percentiles of environmental quality data. Mathematics and Computers in Simulation, 33(5–6):483– 488, April 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475492901413.

Brooker:1992:CMS

[325] P. I. Brooker, G. C. Cock, and M. A. Stewart. Comparison of methods for simulation of two dimensional data honouring specified spherical semivariograms. Mathematics and Computers in Simulation, 33(5– 6):489–494, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475492901424.

Hutchinson:1992:NPS

[326] M. F. Hutchinson. Non-parametric smoothing of almost annually periodic time series. Mathematics and Computers in Simulation, 33(5–6): 495–500, April 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475492901435.

Haritos:1992:CDS

[327] N. Haritos. The characteristics of dynamic systems via the swept sine wave technique. Mathematics and Computers in Simulation, 33(5–6):

501–506, April 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475492901446.

Benyon:1992:DDE

[328] P. R. Benyon. Derivation of dynamic estimation equations by means of the Dirac delta function. Mathematics and Computers in Simulation, 33 (5–6):507–512, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475492901457.

Murthy:1992:SAA

[329] D. N. P. Murthy, B. P. Iskandar, and R. J. Wilson. A simulation approach to analysis of free replacement policies. Mathematics and Computers in Simulation, 33(5–6):513– 518, April 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475492901468.

McAleer:1992:MED

[330] Michael McAleer. Modelling in econometrics: the deterrent effect of capital punishment. Mathematics and Computers in Simulation, 33(5–6): 519–532, April 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475492901479.

Aiyoshi:1992:AOC

[331] Eitaro Aiyoshi and Atsushi Maki. An application of optimal control analysis to the lifecycle model. Mathematics and Computers in Simulation, 33 (5–6):533–538, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290148A.

Bewley:1992:ELR

[332] Ronald Bewley and Denzil G. Fiebig. Estimation of long-run responses in dynamic models with integrated data. Mathematics and Computers in Simulation, 33(5–6):539– 544, April 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290149B.

McAleer:1992:BEN

[333] Michael McAleer and Jeremy Smith. Bootstrap estimates of a new classical model of unemployment. Mathematics and Computers in Simulation, 33 (5–6):545–550, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290150F.

Takase:1992:HPS

[334] Mitsuo Takase. Housing purchase and saving behavior under unanticipated inflation. Mathematics and Computers in Simulation, 33(5–6): 551–556, April 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475492901516.

Bewley:1992:DID

[335] Ronald Bewley, Philip Lowe, and Robert Trevor. On the distribution of intra-daily exchange rate changes. Mathematics and Computers in Simulation, 33(5–6):557– 562, April 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475492901527.

Hall:1992:SVS

[336] A. D. Hall. A study of various score test statistics for heteroscedasticity in the general linear model. Mathematics and Computers in Simulation, 33 (5–6):563–568, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475492901538.

Oya:1992:DFI

[337] Kosuke Oya and Kimio Morimune. The distribution of the full information maximum likelihood estimator. Mathematics and Computers in Simulation, 33(5–6):569–574, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475492901549.

Tse:1992:MSC

[338] Y. K. Tse. MLE of some continuous time financial models: Some Monte Carlo results. Mathematics and Computers in Simulation, 33(5– 6):575–580, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290155A.

Common:1992:EMA

[339] M. S. Common. Economic modelling and Australian carbon dioxide emissions. Mathematics and Computers in Simulation, 33(5–6):581– 596, April 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549290156B.

Taylor:1992:GTD

[340] John A. Taylor. A global threedimensional Lagrangian tracer transport modelling study of the sources and sinks of nitrous oxide. Mathematics and Computers in Simulation, 33 (5–6):597–602, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290157C.

Cowell:1992:STM

[341] P. J. Cowell, P. S. Roy, and R. A. Jones. Shoreface translation model: Computer simulation of coastal-sand-body response to sea level rise. Mathematics and Computers in Simulation, 33(5–6):603–608, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290158D.

Gomboso:1992:GMO

[342] J. Gomboso and F. Ghassemi. Groundwater modelling and optimal salinity control in the North Stirling Land Conservation District, Western Australia: a hydrogeological and economic perspective. Mathematics and Computers in Simulation, 33(5–6): 609–614, April 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290159E.

Anonymous:1992:NIa

[343] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 33(5–6):615–624, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290160I.

Anonymous:1992:PAa

[344] Anonymous. Pages 357–624 (April 1992). Mathematics and Computers in Simulation, 33(5–6):??, April 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1992:EB

[345] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 34(1):iv, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290093V.

Sod:1992:FDM

[346] Gary A. Sod. A flame dictionary method for an unsteady, premixed laminar flame. Mathematics and Computers in Simulation, 34 (1):1–21, May 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549290094W.

Miller:1992:OIC

[347] B. M. Miller and E. Ja. Rubinovich. Optimal impulse control problem with constrained number of impulses. Mathematics and Computers in Simulation, 34(1):23–49, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290095X.

Ozsoy:1992:CBQ

[348] Emin Özsoy, Carlos J. Lozano, and Allan R. Robinson. A consistent baroclinic quasigeostrophic ocean model in multiply connected ocean domains. Mathematics and Computers in Simulation, 34(1):51– 79, May 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549290096Y.

Anonymous:1992:BRPa

[349] Anonymous. Book review: Papers — experiences — perspectives: A. A. B. Pritsker, Systems Publishing Corporation, West Lafayette, IN, 1990. 577 pp., ISBN 0-938974-03-3. Mathematics and Computers in Simulation, 34(1):81, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290097Z.

Anonymous:1992:BRPb

[350] Anonymous. Book review: PASCAL XSC — sprachbeschreibung mit beispielen (fundamentals with examples): R. Klatte, U. Kulisch, M. Neaga, D. Ratz and Ch. Ullrich, Springer, Berlin, 1991. 345 pp., ISBN 3-540-53714-7 (Berlin), 0-387-53714-7 (New York). Mathematics and Computers in Simulation, 34(1):81, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475492900982.

Anonymous:1992:BRCa

[351] Anonymous. Book review: Code recognition and set selection with neural networks: C. Jeffries, Birkhäuser, Boston, Basel, Berlin, 1991. 166 pp., ISBN 0-8176-3585-8 or 3-7643-3585-8. Mathematics and Computers in Simulation, 34(1):82, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290100U.

Anonymous:1992:BRIa

[352] Anonymous. Book review: Inequalities involving functions and their integrals and derivatives: D. S. Mitrinovic, J. E. Pecaric and A. M. Fink, Kluwer, Dordrecht, 1991. 587 pp., Dfl.280, US$149, UK£95, ISBN 0-7923-1330-5. Mathematics and Computers in Simulation, 34(1):82, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475492900993.

Anonymous:1992:BRHa

[353] Anonymous. Book review: History of mathematical programming: J. K. Lenstra, A. H. G. Rinnooy Kan and A. Schrijver, North-Holland, Amsterdam, 1991. 142 pp., US$71, Dfl.124, ISBN 0-444-88818-7. Mathematics and Computers in Simulation, 34(1):82–83, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290101L.

Anonymous:1992:BRDa

[354] Anonymous. Book review: Differential equations and mathematical physics: C. Bennewitz (ed.), Academic Press, Boston, 1991. 365 pp., US$49.95, ISBN 0-12-089040-2. Mathematics and Computers in Simulation, 34(1):83, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290102M.

Anonymous:1992:BRCb

[355] Anonymous. Book review: Canonical equational proofs: L. Bachmair, Birkhäuser, Boston, Basel, Berlin, 1991. 135 pp., S.Fr.68, DM.78, ISBN 0-8176-3555-6 or 3-7643-3555-6. Mathematics and Computers in Simulation, 34(1):83–84, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290103N.

Anonymous:1992:BRB

[356] Anonymous. Book review: Bounded and almost periodic solutions of nonlinear operator differential equations: A. A. Pankov, Kluwer, Dordrecht, 1990. 221 pp., Dfl.170, US$99, UK£59, ISBN 0-7923-0585-X. Mathematics and Computers in Simulation, 34 (1):84, May 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290104O.

Anonymous:1992:BRDb

[357] Anonymous. Book review: Differential inclusions and optimal control: M. Kisielewicz, Kluwer, Dordrecht, 1991. 240 pp., Dfl.210, US$124, UK£77, ISBN 0-7923-0675-9. Mathematics and Computers in Simulation, 34(1):84, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290105P.

Anonymous:1992:BRLa

[358] Anonymous. Book review: LOGIC — a foundation for computer science: V. Sperschneider and G. Antoniou, Addison-Wesley, Reading, MA, 1991. 495 pp., US$34.50, ISBN 0-201-56514-5. Mathematics and Computers in Simulation, 34(1): 84–85, May 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290106Q.

Anonymous:1992:BRCc

[359] Anonymous. Book review: Controllability of dynamical systems: J. Klamka, Kluwer, Dordrecht, 1991. 248 pp., Dfl.195, US$114, UK£72, ISBN 0-7923-0822-0. Mathematics and Computers in Simulation, 34 (1):85, May 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549290107R.

Anonymous:1992:BRDc

[360] Anonymous. Book review: Discrete linear control systems: V. N. Fomin, Kluwer, Dordrecht, 1991. 302 pp., Dfl.220, US$134, UK£75, ISBN 0-7923-1248-1. Mathematics and Computers in Simulation, 34 (1):85, May 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549290108S.

Anonymous:1992:BRMa

[361] Anonymous. Book review: A MINITAB companion — with macros: P. Zehna, Addison-Wesley, Reading, MA, 1992. 400 pp., US$24.95, ISBN 0-201-55592-1 with IBM 5 1/4in macro disk; 0-201-55580-8 with IBM 312in macro disk. Mathematics and Computers in Simulation, 34 (1):86, May 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475492901103.

Anonymous:1992:BRPc

[362] Anonymous. Book review: Proceedings of the fifth European conference on mathematics in industry: M. Heilio (ed.), Teubner, Stuttgart/Kluwer, Dordrecht, 1991. 400 pp., Dfl.260, US$139, UK£89, ISBN 3-519-02176-5 (Teubner); 0-7923-1317-8 (Kluwer). Mathematics and Computers in Simulation, 34(1):86, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290109T.

Anonymous:1992:ICN

[363] Anonymous. Interval computations: New journal: (Editor-in-Chief: V. M. Nesterov, Institute For New Technologies, St. Petersburg) 1991. Mathematics and Computers in Simulation, 34(1):86, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290111S.

Anonymous:1992:BRN

[364] Anonymous. Book review: Nematics — mathematical and physical aspects: J.-M. Coron, J.-M., Chidaglia and F. Helein (eds.), NATO ASI Series. Kluwer, Dordrecht, 1991. 428 pp., Dfl.215, US$140, UK£74, ISBN 0-7923-1113-2. Mathematics and Computers in Simulation, 34(1): 86–87, May 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290112T.

Anonymous:1992:BRDd

[365] Anonymous. Book review: The drift diffusion equation and its application in MOSFET modelling: W. Hansch, Springer, Wien, 1991. 271 pp., Ös.1148, DM.164, ISBN 3-211-82222-4. Mathematics and Computers in Simulation, 34(1):87, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290113U.

Anonymous:1992:BRMb

[366] Anonymous. Book review: Markov processes — an introduction for physical scientists: D. T. Gillespie, Academic Press, Boston, 1992. 565 pp., US$44.50, ISBN 0-12-283955-2. Mathematics and Computers in Simulation, 34(1):87, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290114V.

Anonymous:1992:BRSa

[367] Anonymous. Book review: Scientific computing and differential equations — an introduction of numerical methods: G. H. Golub and J. M. Ortega, Academic Press, San Diego, 1992. US$49.95, ISBN 0-12-289255-0. Mathematics and Computers in Simulation, 34(1):87, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290115W.

Anonymous:1992:BRSb

[368] Anonymous. Book review: Sturm– Liouville and Dirac operators: B. M. Levitan and I. S. Sargsjan, Kluwer, Dordrecht, 1991. 350 pp., Dfl.240, US$149, UK£83, ISBN 0-7923-0992-8. Mathematics and Computers in Simulation, 34(1):88, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290116X.

Anonymous:1992:BRSc

[369] Anonymous. Book review: Singularities of caustics and wave fronts: V.

I. Arnold, Kluwer, Dordrecht, 1990. 259 pp., Dfl.160, US$99, UK£56, ISBN 0-7923-1038-1. Mathematics and Computers in Simulation, 34 (1):88, May 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290117Y.

Anonymous:1992:BRSd

[370] Anonymous. Book review: The Schrödinger equation: F. A. Berezin and M. A. Shubin, Kluwer, Dordrecht, 1991. 555 pp., Dfl. 395, US$249, UK£136, ISBN 0-7923-1218-X. Mathematics and Computers in Simulation, 34(1):88–89, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290118Z.

Anonymous:1992:BRMc

[371] Anonymous. Book review: Mixed boundary value problems of potential theory and their applications in engineering: V. I. Fabrikant, Kluwer, Dordrecht, 1991. 451 pp., Dfl.270, US$134, UK£89, ISBN 0-7923-1157-4. Mathematics and Computers in Simulation, 34(1):89, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475492901192.

Anonymous:1992:WWI

[372] Anonymous. Who is who in IMACS. Mathematics and Computers in Simulation, 34(1):91–99, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/0378475492901206.

Anonymous:1992:PM

[373] Anonymous. Pages 1–100 (May 1992). Mathematics and Computers in Simulation, 34(1):??, May 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Choi:1992:FSA

[374] Y. S. Choi and Kwong-Yu Chan. Fractional step algorithm for a model problem in electrochemistry. Mathematics and Computers in Simulation, 34 (2):101–112, August 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290047K.

Moore:1992:PSB

[375] Ramon Moore. Parameter sets for bounded-error data. Mathematics and Computers in Simulation, 34 (2):113–119, August 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290048L.

Sreenath:1992:HCE

[376] N. Sreenath. A hybrid computation environment for multibody simulation. Mathematics and Computers in Simulation, 34(2):121– 140, August 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549290049M.

Neta:1992:AFE

[377] B. Neta. Analysis of finite elements and finite differences for shallow water equations: a review. Mathematics and Computers in Simulation, 34 (2):141–161, August 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290050Q.

Silverberg:1992:SMW

[378] Larry Silverberg and John Richard Gardiner. Simulation of the mechanics of water-skier motion. Mathematics and Computers in Simulation, 34 (2):163–181, August 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290051H.

Ames:1992:BRSa

[379] W. F. Ames and C. Brezinski. Book review: Simulation of thermal systems — a modular program with an interactive preprocessor (EMGP3): W. L. Dutre, Kluwer, Dordrecht, 1991. 600 pp., Dfl.320, US$199, UK£110, ISBN0-7923-1235-X. Mathematics and Computers in Simulation, 34(2): 183, August 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290052I.

Ames:1992:BRNa

[380] W. F. Ames and C. Brezinski. Book review: Numerical methods for Engineers and Scientists: J. D. Hoffman, McGraw-Hill, New York, 1992. 825 pp., ISBN 0-0702-9213-2. Mathematics and Computers in Simulation, 34 (2):183–184, August 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290053J.

Ames:1992:BRA

[381] W. F. Ames and C. Brezinski. Book review: Approximation theory, spline functions and applications: S. P. Singh (ed.), Kluwer, Dordrecht, 1992. 479 pp., Dfl.240, US$139, UK£82, ISBN 0-7923-1574-X. Mathematics and Computers in Simulation, 34(2): 184, August 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290054K.

Ames:1992:BRTa

[382] W. F. Ames and C. Brezinski. Book review: Topics in soliton theory: R. W. Carroll, North-Holland mathematics studies volume 167. North-Holland, Amsterdam, 1991. 428 pp., Dfl.195, US$100, ISBN 0-4448-8869-1. Mathematics and Computers in Simulation, 34(2):184, August 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290055L.

Ames:1992:BRD

[383] W. F. Ames and C. Brezinski. Book review: Discrete groups in space and uniformization problems: B. N. Apanasov, Kluwer, Dordrecht, 1991. 482 pp., Dfl.295, UK£99, US$193, ISBN 0-7923-0216-8. Mathematics and Computers in Simulation, 34 (2):184–185, August 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290056M.

Ames:1992:BRCa

[384] W. F. Ames and C. Brezinski. Book review: Computer Algebra with LISP and REDUCE: F. Brackx and D. Constales, Kluwer, Dordrecht, 1991. 264 pp., Dfl.160, US$88, UK£54, ISBN 0-7923-1441-7. Mathematics and Computers in Simulation, 34(2): 185, August 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290058O.

Ames:1992:BRSb

[385] W. F. Ames and C. Brezinski. Book review: SLAM II: network models for decision support: A. A. B. Pritsker, C. E. Sigal and R. D. J. Hammesfahr, Prentice-Hall, Englewood Cliffs, NJ, 1989. 558 pp., ISBN 0-1381-2819-7. Mathematics and Computers in Simulation, 34(2): 185, August 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290057N.

Ames:1992:BRCb

[386] W. F. Ames and C. Brezinski. Book review: Complexity theory of real functions: K.-I Ko, Birkhäuser, Basel, 1991. 308 pp., Sfr.78, DM94, ISBN 0-8176-3586-6; 3-7643-3586-6. Mathematics and Computers in Simulation, 34(2):186, August 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290059P.

Ames:1992:ILS

[387] W. F. Ames and C. Brezinski. Identification of linear systems: J. Schoukens and R. Pintelon, Pergamon Press, Oxford, 1991. 317 pp., UK£55, US$110, ISBN 0-0804-0734-X. Mathematics and Computers in Simulation, 34(2): 186, August 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549290060T.

Ames:1992:BRMa

[388] W. F. Ames and C. Brezinski. Book review: Measures and differential equations in infinite-dimensional space: Yu.L. Dalecky and S. V. Fomin, Kluwer, Dordrecht, 1991. 337 pp., Dfl.220, US$126, UK£75, ISBN 0-7923-1517-0. Mathematics and Computers in Simulation, 34 (2):186–187, August 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290061K.

Ames:1992:BRP

[389] W. F. Ames and C. Brezinski. Book review: Perturbation methods for Engineers and Scientists: A. W. Bush, CRC Press, Boca Raton, FL, 1992. 303 pp., ISBN 0-8493-8608-X (hard cover); 0-8493-8614-4 (paperback). Mathematics and Computers in Simulation, 34(2):187, August 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290062L.

Ames:1992:BRRa

[390] W. F. Ames and C. Brezinski. Book review: Reciprocity, spatial mapping and time reversal in electromagnetics: C. Altman and K. Suchy, Kluwer, Dordrecht, 1991. 286 pp., Dfl.185, US$99, UK£62, ISBN 0-7923-1339-9. Mathematics and Computers in Simulation, 34(2):187–188, August 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290063M.

Ames:1992:BRH

[391] W. F. Ames and C. Brezinski. Book review: Handbook of differential equations: D. Zwillinger, Academic Press, Boston, Ma, 2nd ed., 1992. 787 pp., US$54.95, ISBN 0-1278-4391-4. Mathematics and Computers in Simulation, 34(2):188, August 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290065O.

Ames:1992:BRMb

[392] W. F. Ames and C. Brezinski. Book review: From MEMEX to HYPER-TEXT: J. M. Nyce and P. Kahn (eds.), Academic Press, Boston, MA, 1991. 367 pp., US$39.95, ISBN 0-1252-3270-5. Mathematics and Computers in Simulation, 34(2): 188, August 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549290064N.

Ames:1992:BRL

[393] W. F. Ames and C. Brezinski. Book review: The linear complementarity problem: R. W. Cottle, J.-S. Pang and R. E. Stone, Academic Press, Boston, MA, 1992. 762 pp., US$59.95, ISBN 0-1219-2350-9. Mathematics and Computers in Simulation, 34 (2):188–189, August 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290066P.

Ames:1992:BRNb

[394] W. F. Ames and C. Brezinski. Book review: Nonlinear equations in the applied sciences: W. F. Ames and C. Rogers (eds.), Academic Press, Boston, MA, 1992. 474 pp., US$49.50, ISBN 0-1205-6752-0. Mathematics and Computers in Simulation, 34(2): 189, August 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549290067Q.

Ames:1992:BRRb

[395] W. F. Ames and C. Brezinski. Book review: Representation of lie groups and special functions, volume 1: N. Ja Vilenkin and A. U. Klimyk, Kluwer, Dordrecht, 1991. 608 pp., Dfl.460, US$249, UK£155, ISBN-0-7923-1466-2. Mathematics and Computers in Simulation, 34(2):189– 190, August 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290068R.

Ames:1992:BRG

[396] W. F. Ames and C. Brezinski. Book review: Geometry of Banach spaces, duality mappings and nonlinear problems: I. Cioranescu, Kluwer, Dordrecht, 1990. 260 pp, Dfl.175, US$99, UK£62, ISBN 0-7923-0910-3. Mathematics and Computers in Simulation, 34(2):190, August 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290069S.

Ames:1992:BRTb

[397] W. F. Ames and C. Brezinski. Book review: The transparent prolog machine: M. Eisenstadt, M. Brayshaw and J. Paine, Kluwer, Dordrecht, 1991. 176 pp., Dfl.110, US$59, ISBN 0-7923-1447-6. Mathematics and Computers in Simulation, 34(2): 190, August 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290070W.

Ames:1992:BRU

[398] W. F. Ames and C. Brezinski. Book review: UNIX for the impatient: P. W. Abrahams and B. A. Larson, Addison-Wesley, Reading, MA, 1992. 559 pp., ISBN 0-2015-5703-7. Mathematics and Computers in Simulation, 34(2):191, August 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290071N.

Anonymous:1992:NIb

[399] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 34(2):193–203, August 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290072O.

Anonymous:1992:PAb

[400] Anonymous. Pages 101–204 (August 1992). Mathematics and Computers in Simulation, 34(2):??, August 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Javor:1992:P

[401] András Jávor. Preface. Mathematics and Computers in Simulation, 34(3– 4):205–206, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290001W.

Ameling:1992:SDF

[402] Walter Ameling. Simulation of distributed functions. Mathematics and Computers in Simulation, 34(3–4): 207–218, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290002X.

Oren:1992:TOA

[403] T. I. Ören, R. N. Wendt, D. G. King, O. Abou-Rabia, and L. G. Birta. A template-oriented approach for simulation program understanding. Mathematics and Computers in Simulation, 34(3–4):219– 229, September 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290003Y.

Breitenecker:1992:MME

[404] Felix Breitenecker. Models, methods and experiments — a new structure for simulation systems. Mathematics and Computers in Simulation, 34(3– 4):231–260, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290004Z.

Rozenblit:1992:IKR

[405] Jerzy W. Rozenblit and Jhyfang Hu. Integrated knowledge representation and management in simulation-based design generation. Mathematics and Computers in Simulation, 34(3–4): 261–282, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475492900052.

Javor:1992:DCS

[406] András Jávor. Demon controlled simulation. Mathematics and Computers in Simulation, 34(3–4):283– 296, September 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475492900063.

Kleijnen:1992:SAS

[407] Jack P. C. Kleijnen. Sensitivity analysis of simulation experiments: regression analysis and statistical design. Mathematics and Computers in Simulation, 34(3–4):297–315, September 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475492900074.

Ceric:1992:DRC

[408] Vlatko Ceric and Ray J. Paul. Diagrammatic representations of the conceptual simulation model for discrete event systems. Mathematics and Computers in Simulation, 34(3– 4):317–324, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475492900085.

Yolles:1992:DTP

[409] Maurice Yolles and Mohamed Pirani. A design for a transnational project management system. Mathematics and Computers in Simulation, 34(3– 4):325–339, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475492900096.

Wen:1992:SSA

[410] Chuan yuan Wen and Zheng she Liu. Simulation study on the architecture optimization and task scheduling of a parallel flight training simulator. Mathematics and Computers in Simulation, 34(3–4):341–350, September 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290010E.

Cahlon:1992:OPH

[411] B. Cahlon, I. E. Schochetman, and M. Shillor. Optimal placement of heat sources on a rectangular grid. Mathematics and Computers in Simulation, 34(3–4):351–364, September 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475492900115.

Pistella:1992:SPC

[412] Francesca Pistella and Vincenzo Casulli. Simulation of a planar combustion wave. Mathematics and Computers in Simulation, 34(3–4):365– 373, September 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475492900126.

Anonymous:1992:BRIb

[413] Anonymous. Book review: Identification of continuous-time systems (Methodology and computer implementation): N. K. Sinha and G. P. Rao (eds.), Kluwer, Dordrecht, 1991. 637 pp., Dfl.275, US$149, UK£92, ISBN 0-7923-1336-4. Mathematics and Computers in Simulation, 34(3– 4):375, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475492900137.

Anonymous:1992:BRTa

[414] Anonymous. Book review: Theoretical studies in computer science: J. D. Ullman (ed.), Academic Press, Boston, MA, 1992. 339 pp., US$49.95, ISBN 0-1270-8240-9. Mathematics and Computers in Simulation, 34(3– 4):375, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475492900148.

Anonymous:1992:BRF

[415] Anonymous. Book review: Fractal attraction — a fractal design system for the Macintosh: K. D. Lee and Y. Cohen, Academic Press, San Diego, CA, 1992. 67 pp. plus a Mac Disk, US$49.95, ISBN 0-1244-0740-4. Mathematics and Computers in Simulation, 34(3–4):376, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475492900159.

Anonymous:1992:BRPd

[416] Anonymous. Book review: Proceedings of the fourth European conference on mathematics in industry: H. Wacker and W. Zulehner (eds.), Teubner, Stuttgart/Kluwer, Dordrecht, 1990. 433 pp., Dfl.240, US$148, UK£83, ISBN 3-5190-2175-7 (Teubner), 0-7923-1036-5 (Kluwer). Mathematics and Computers in Simulation, 34(3– 4):376, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290016A.

Anonymous:1992:BRTb

[417] Anonymous. Book review: Theory of global random search: A. A. Zhigljavsky, Kluwer, Dordrecht, 1991.

341 pp., Dfl.220, US$144, UK£76, ISBN 0-7923-1122-1. Mathematics and Computers in Simulation, 34(3– 4):376–377, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290017B.

Anonymous:1992:BREa

[418] Anonymous. Book review: Electromagnetic modelling and measurements for analysis and synthesis problems: B. de Neumann (ed.), NATO ASI Series. Kluwer, Dordrecht, 1991. 447 pp., Dfl.260, US$144, UK£87, ISBN 0-7923-1265-1. Mathematics and Computers in Simulation, 34(3– 4):377, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290018C.

Anonymous:1992:BRSe

[419] Anonymous. Book review: Software implementation techniques: D. E. Merusi, Digital Press, Burlington, MA, 1992. 743 pp., ISBN 1-5555-8090-4. Mathematics and Computers in Simulation, 34(3–4):377, September 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290019D.

Anonymous:1992:BRMd

[420] Anonymous. Book review: Modern probability theory and its applications: E. Parzen, Wiley Classics Library. Wiley, New York, 1992 (original 1960). 464 pp., US$32.95, ISBN 0-4715-7278-0. Mathematics and Computers in Simulation, 34(3– 4):377–378, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290020H.

Anonymous:1992:BREb

[421] Anonymous. Book review: Elements of the mathematical theory of multi-frequency oscillations: A. M. Samoilenko, Kluwer, Dordrecht, 1991. 313 pp., Dfl.260, US$142, UK£88, ISBN 0-7923-1438-7. Mathematics and Computers in Simulation, 34(3– 4):378, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475492900218.

Anonymous:1992:BRLb

[422] Anonymous. Book review: Logic programming — new frontiers: D. R. Brough (ed.), Kluwer, Dordrecht, 1992. 310 pp., Dfl.150, US$85, ISBN 0-7923-1546-4. Mathematics and Computers in Simulation, 34(3–4): 378, September 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475492900229.

Anonymous:1992:BRCd

[423] Anonymous. Book review: Computing with T. Node parallel architecture: D. Heidrich and J. C. Grossetie (eds.), Kluwer, Dordrecht, 1991. 264 pp., Dfl.160, US$88, UK£54, ISBN 0-7923-1483-2. Mathematics and Computers in Simulation, 34(3– 4):378–379, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290023A.

Anonymous:1992:BRIc

[424] Anonymous. Book review: Instabilities and nonequilibrium structures III: E. Tirapegui and W. Zeller (eds.), Kluwer, Dordrecht, 1991. 370 pp., Dfl.220, US$122, UK£74, ISBN 0-7923-1153-1. Mathematics and Computers in Simulation, 34(3–4):379, September 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290024B.

Anonymous:1992:BRTc

[425] Anonymous. Book review: TEX by example — a beginner’s guide: A. Borde, Academic Press, San Diego, CA, 1992. 169 pp., US$19.95, ISBN 0-1211-7650-9. Mathematics and Computers in Simulation, 34(3–4): 379, September 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290025C.

Anonymous:1992:BRSf

[426] Anonymous. Book review: The Schwarz function and its generalization to higher dimensions: H. S. Shapiro, Wiley, New York, 1992. 108 pp., US$59.95, ISBN 0-4715-7127-X. Mathematics and Computers in Simulation, 34(3–4):379– 380, September 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549290026D.

Anonymous:1992:BREc

[427] Anonymous. Book review: Exterior differential systems and equivalence problems: K. Yang, Kluwer, Dordrecht, 1992. 196 pp., Dfl.130, US$78, UK£45, ISBN 0-7923-1593-6. Mathematics and Computers in Simulation, 34(3–4):380, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290027E.

Anonymous:1992:BRTd

[428] Anonymous. Book review: Threedimensional modeling with geoscientific information systems: A. K. Turner (ed.), NATO ASI Series. Kluwer, Dordrecht, 1992. 443 pp., Dfl.260, US$149, UK£88, ISBN 0-7923-1550-2. Mathematics and Computers in Simulation, 34(3–4):380, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549290028F.

Anonymous:1992:BRAa

[429] Anonymous. Book review: Analytic pseudo-differential operators and their applications: J. A. Dubinskii, Kluwer, Dordrecht, 1991. 252 pp., Dfl.175, US$98, UK£59, ISBN 0-7923-1296-1. Mathematics and Computers in Simulation, 34(3–4):380– 381, September 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290029G.

Anonymous:1992:BRAb

[430] Anonymous. Book review: Automated reasoning: R. S. Boyer (ed.), Kluwer, Dordrecht, 1991. 365 pp., Dfl.160, US$89, UK£54, ISBN 0-7923-1409-3. Mathematics and Computers in Simulation, 34(3–4):381, September 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290030K.

Anonymous:1992:BRTe

[431] Anonymous. Book review: Theory of random determinants: V. L. Girko, Kluwer, Dordrecht, 1990. 677 pp., Dfl. 420, US$245, UK£47, ISBN 0-7923-0233-8. Mathematics and Computers in Simulation, 34(3– 4):381, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290031B.

Anonymous:1992:BRDe

[432] Anonymous. Book review: Digital at work — snapshots from the first thirtyfive years: J. P. Pearson (ed.), Digital Press, Burlington, MA, 1992. 212 pp., US$19.95, ISBN 0-1321-3489-6. Mathematics and Computers in Simulation, 34(3–4):382, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290034E.

Anonymous:1992:BRPe

[433] Anonymous. Book review: Problems in applied, industrial and engineering mathematics: H. K. Kuiken and S. W. Rienstra (eds.), Kluwer, Dordrecht, 1992. 265 pp., Dfl.195, US$116, UK£68, ISBN 0-7923-1680-0. Mathematics and Computers in Simulation, 34(3–4):382, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290032C.

Anonymous:1992:BRRa

[434] Anonymous. Book review: Reliable computer systems — design and evaluation: D. P. Siewiorek and R. S. Swarz, Digital Press, Burlington, MA, 2nd ed., 1992. 916 pp., US$64.95, ISBN 1-5555-8075-0. Mathematics and Computers in Simulation, 34(3– 4):382, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290033D.

Anonymous:1992:BRW

[435] Anonymous. Book review: Workshop on engineering turbulence modelling: L. A. Povinelli, W. W. Liou, A. Shabbir and T. H. Shih (eds.), NASA conference publication 10088. NASA Lewis Research Center, Cleveland, OH, 1992. Mathematics and Computers in Simulation, 34(3–4): 383, September 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290036G.

Anonymous:1992:BRCe

[436] Anonymous. Book review: Computational fluid dynamics: L. A. Povinelli (ed.), NASA conference publication 10078. NASA Lewis Research Center, Cleveland, OH, 1992. Mathematics and Computers in Simulation, 34 (3–4):383, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290035F.

Anonymous:1992:NIc

[437] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 34(3–4):385–396, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/037847549290037H.

Anonymous:1992:PS

[438] Anonymous. Pages 205–396 (September 1992). Mathematics and Computers in Simulation, 34(3–4):??, September 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Drescher:1992:SWC

[439] G. Drescher and B. Fellenberg. Simulation of weakly correlated fields with an application to vehicle dynamics. Mathematics and Computers in Simulation, 34(5):397–409, November 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290073P.

Hebert:1992:SSR

[440] D. J. Hebert. Simulations of stochastic reaction–diffusion systems. Mathematics and Computers in Simulation, 34 (5):411–432, November 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290074Q.

Kazimierczyk:1992:ECF

[441] Piotr Kazimierczyk. Explicit correction formulae for parametric identification of stochastic differential systems. Mathematics and Computers in Simulation, 34(5):433–450, November 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290075R.

Jaiswal:1992:SAN

[442] M. P. Jaiswal, C. L. Sharma, and Karmeshu. Stochastic analysis of a nonlinear machine interference model. Mathematics and Computers in Simulation, 34(5):451– 458, November 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290076S.

Makroglou:1992:CMS

[443] Athena Makroglou. Collocation methods for stochastic Volterra integrodifferential equations with random forcing functions. Mathematics and Computers in Simulation, 34(5):459– 466, November 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290077T.

Gong:1992:ASP

[444] Wei-Bo Gong and Henning Schulzrinne. Application of smoothed perturbation analysis to probabilistic routing. Mathematics and Computers in Simulation, 34(5):467–485, November 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290078U.

Zavattaro:1992:SSS

[445] M. G. Zavattaro and R. Riganti. Statistics of the solution of singularly perturbed systems with random initial conditions. Mathematics and Computers in Simulation, 34(5): 487–499, November 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290079V.

Anonymous:1992:BRAc

[446] Anonymous. Book review: Alpha architecture reference manual: R. L. Sites (ed.), Digital Press, Burlington, MA, 1992. 600 pp., paper, US$34.95, ISBN 0-1303-3663-7. Mathematics and Computers in Simulation, 34 (5):501, November 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290080Z.

Anonymous:1992:BRMe

[447] Anonymous. Book review: Mathematical models in electrical circuits: theory applications: C. A. Marinov and P. Neittaanmaki, Kluwer, Dordrecht, 1991. 160 pp., Dfl.120, US$66.50, UK£40, ISBN 0-7923-1155-8. Mathematics and Computers in Simulation, 34(5):501–502, November 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290081Q.

Anonymous:1992:BRTf

[448] Anonymous. Book review: Topological function spaces: A. V. Arkhangel’skii, Kluwer, Dordrecht, 1991. 205 pp., Dfl.140, US$79, UK£48, ISBN 0-7923-1531-6. Mathematics and Computers in Simulation, 34 (5):502, November 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290082R.

Anonymous:1992:BRTg

[449] Anonymous. Book review: Theory of approximation: N. I. Achieser, Dover, Mineola, NY, 1992 (reprint of the 1956 translation). 317 pp., paper, US$8.95, ISBN 0-4866-7129-1. Mathematics and Computers in Simulation, 34(5):502, November 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290083S.

Anonymous:1992:BRCf

[450] Anonymous. Book review: Clifford algebra and spinor-valued functions (a function theory for the Dirac operator): R. Delanghe, F. Sommem and V. Soucek, Kluwer, Dordrecht, 1992. 485 pp., Dfl.295, US$176, UK£102, ISBN 0-7923-0229-X. Mathematics and Computers in Simulation, 34 (5):503, November 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290085U.

Anonymous:1992:BRHb

[451] Anonymous. Book review: Hilbert spaces and operator theory: W. Mlak, Kluwer, Dordrecht, 1991. 290 pp., Dfl.235, US$149, UK£81, ISBN 0-7923-1042-X. Mathematics and Computers in Simulation, 34(5):503, November 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290084T.

Anonymous:1992:BRId

[452] Anonymous. Book review: Image synthesis: M. Bret, Kluwer, Dordrecht, 1991. 289 pp., Dfl.180, US$99, UK£61, ISBN 0-7923-1488-3. Mathematics and Computers in Simulation, 34(5): 503–504, November 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290086V.

Anonymous:1992:BRMf

[453] Anonymous. Book review: Matroid decomposition: K. Truemper, Academic Press, Boston, MA, 1992, 398 pp., US$39.95, ISBN 0-1270-1225-7. Mathematics and Computers in Simulation, 34(5):504, November 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290087W.

Anonymous:1992:BRRb

[454] Anonymous. Book review: Real reductive groups II: N. R. Wallach, Academic Press, Boston, MA, 1992. US$75, ISBN 0-1273-2961-7. Mathematics and Computers in Simulation, 34(5): 504–505, November 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290088X.

Anonymous:1992:BRSg

[455] Anonymous. Book review: Stability and oscillations in delay differential equations of population dynamics: K. Gopalsamy, Kluwer, Dordrecht, 1992. 501 pp., Dfl.295, US$179, UK£103, ISBN 0-7923-1594-4. Mathematics and Computers in Simulation, 34(5):505, November 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290089Y.

Anonymous:1992:BRAd

[456] Anonymous. Book review: Approximation by solutions of partial differential equations: B. Fuglede, M. Goldstein, W. Hausmann and L. Rogge, NATO ASI Series. Kluwer, Dordrecht, 1992. 201 pp., Dfl.135, US$79, UK£47, ISBN 0-7923-1700-9. Mathematics and Computers in Simulation, 34(5): 505–506, November 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475492900904.

Anonymous:1992:BRSh

[457] Anonymous. Book review: Singular integral equations: N. I. Muskhelishvili, Dover, Mineola, NY, 1991 (reprint of the 1953 translation of the second revised edition of 1946). 447 pp., paper, US$12.95, ISBN 0-4866-6893-2. Mathematics and Computers in Simulation, 34(5):506, November 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290091T.

Anonymous:1992:NId

[458] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 34(5):507–514, November 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/037847549290092U.

Anonymous:1992:PN

[459] Anonymous. Pages 397–514 (November 1992). Mathematics and Computers in Simulation, 34(5):??, November 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Piet-Lahanier:1992:CMS

[460] Hélène Piet-Lahanier, Sándor M. Veres, and Eric Walter. Comparison of methods for solving sets of linear inequalities in the boundederror context. Mathematics and Computers in Simulation, 34(6):515– 524, December 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290038I.

Lee:1992:NMS

[461] Ding Lee, Martin H. Schultz, William L. Siegmann, and Donald F. St. Mary. A numerical marching scheme to compute scattering in the ocean. Mathematics and Computers in Simulation, 34 (6):525–540, December 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290039J.

Lu:1992:SAS

[462] Hongmin Lu and Frank H. Mathis. Surface approximation by spline smoothing and generalized crossvalidation. Mathematics and Computers in Simulation, 34(6):541–549, December 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290040N.

Burger:1992:CTM

[463] Jacques Burger, Jean-Sébastien Le Brizaut, and Marc Pogu. Comparison of two methods for the calculation of the gradient and of the Hessian of cost functions associated with differential systems. Mathematics and Computers in Simulation, 34 (6):551–562, December 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549290041E.

Asaithambi:1992:TSS

[464] N. S. Asaithambi and J. B. Garner. Taylor series solution of a class of diffusion problems in physiology. Mathematics and Computers in Simulation, 34(6):563–570, December 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290042F.

Fossard:1992:SRA

[465] A. J. Fossard. Some reflexions about nonlinear control. Mathematics and Computers in Simulation, 34(6):571– 583, December 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290043G.

Anonymous:1992:NIe

[466] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 34(6):585–591, December 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/037847549290044H.

Anonymous:1992:AIV

[467] Anonymous. Author index to volume 34 (1992). Mathematics and Computers in Simulation, 34(6):593– 595, December 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549290045I.

Anonymous:1992:CV

[468] Anonymous. Contents of volume 34 (1992). Mathematics and Computers in Simulation, 34(6):597–599, December 1992. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549290046J.

Anonymous:1992:PD

[469] Anonymous. Pages 515–599 (December 1992). Mathematics and Computers in Simulation, 34(6):??, December 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1993:EB

[470] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 35(1):ii, February 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549390032P.

Angus:1993:CAI

[471] John E. Angus. Computer-assisted improvement of mean squared error in statistical estimation. Mathematics and Computers in Simulation, 35 (1):1–13, February 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549390033Q.

Verriest:1993:CDP

[472] Erik I. Verriest. The continuization of a discrete process and applications in interpolation and multirate control. Mathematics and Computers in Simulation, 35(1):15– 31, February 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549390034R.

Ho:1993:PSR

[473] Yu-Chi Ho, C. G. Cassandras, and M. Makhlouf. Parallel simulation of real-time systems via the Standard Clock approach. Mathematics and Computers in Simulation, 35 (1):33–41, February 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390035S.

Yang:1993:GLM

[474] J.-Y. Yang, C. A. Hsu, and W. H. Hui. A generalized Lagrangian method for solving the steady shallow water equations. Mathematics and Computers in Simulation, 35(1):43– 61, February 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549390036T.

Birta:1993:AAS

[475] L. G. Birta, M. Yang, and O. Abou-Rabia. An adaptive approach to stepsize control in ODE solvers. Mathematics and Computers in Simulation, 35(1):63–78, February 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390037U.

Calzada:1993:NAF

[476] Maria Calzada and Yu Song. A numerical algorithm for flame advection and propagation. Mathematics and Computers in Simulation, 35 (1):79–92, February 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549390038V.

Anonymous:1993:NIa

[477] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 35(1):93–101, February 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549390039W.

Anonymous:1993:PF

[478] Anonymous. Pages 1–102 (February 1993). Mathematics and Computers in Simulation, 35(1):??, February 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Richard:1993:SSP

[479] Marc J. Richard and Clément M. Gosselin. A survey of simulation programs for the analysis of mechanical systems. Mathematics and Computers in Simulation, 35(2):103– 121, April 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549390007H.

Jaulin:1993:GNP

[480] Luc Jaulin and Eric Walter. Guaranteed nonlinear parameter estimation from bounded-error data via interval analysis [MATCOM 932]. Mathematics and Computers in Simulation, 35(2):123–137, April 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390008I.

Mittelmann:1993:SCS

[481] Hans D. Mittelmann. Symmetric capillary surface in a cube. Mathematics and Computers in Simulation, 35 (2):139–152, April 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390009J.

Simos:1993:PCP

[482] T. E. Simos. A predictor–corrector phase-fitted method for y′′ = f(x, y). Mathematics and Computers in Simulation, 35(2):153–159, April 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/037847549390010R.

Miyata:1993:AOA

[483] Hitoshi Miyata and Masaaki Ohkita. An algorithm of obstacle avoidance for an autonomous mobile robot. Mathematics and Computers in Simulation, 35(2):161–171, April 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390011I.

Liska:1993:SOM

[484] Richard Liska, Mikhail Yu. Shashkov, and Andrey V. Solovjov. Supportoperators method for PDE discretization: symbolic algorithms and realization. Mathematics and Computers in Simulation, 35(2):173– 183, April 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549390012J.

Ames:1993:BRGa

[485] W. F. Ames and C. Brezinski. Book review: The geometry of algebraic Fermi curves: D. Gieseker, H. Knoerrer and E. Trubowitz, Academic Press, San Diego, CA, 1993. 236 pp., ISBN 0-122-82620-5. Mathematics and Computers in Simulation, 35(2):185, April 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390013K.

Ames:1993:BRMa

[486] W. F. Ames and C. Brezinski. Book review: Mathematics and logic: M. Kac and S. M. Ulam, Dover, New York, 1992. 170 pp., paper, US$6.95, ISBN 0-486-67085-6. Mathematics and Computers in Simulation, 35 (2):185, April 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549390014L.

Ames:1993:BRAa

[487] W. F. Ames and C. Brezinski. Book review: Abstract algebra and solution by radicals: J. E. Maxfield and M. W. Maxfield, Dover, New York, 1992. 209 pp., paper, US$7.95, ISBN 0-486-67121-6. Mathematics and Computers in Simulation, 35 (2):186, April 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549390015M.

Ames:1993:BRSa

[488] W. F. Ames and C. Brezinski. Book review: A survey of matrix theory and matrix inequalities: M. Marcus and H. Minc, Dover, New York, 1992. 180 pp., paper, US$5.95, ISBN 0-486-67102-X. Mathematics and Computers in Simulation, 35(2):186, April 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390016N.

Ames:1993:BRAb

[489] W. F. Ames and C. Brezinski. Book review: Applied hyperfunction theory: I. Imai, Kluwer, Dordrecht, 1992. 438 pp., Dfl.360, US$199, UK£123, ISBN 0-792-31507-3. Mathematics and Computers in Simulation, 35 (2):187, April 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549390018P.

Ames:1993:BRF

[490] W. F. Ames and C. Brezinski. Book review: Functions of completely regular growth: L. I. Ronkin, Kluwer, Dordrecht, 1992. 392 pp., ISBN 0-792-31677-0. Mathematics and Computers in Simulation, 35 (2):187, April 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549390017O.

Ames:1993:BRNa

[491] W. F. Ames and C. Brezinski. Book review: Numerik partiellen differentialgleichungen: C. Grossman and H.- G. Roos, Teubner, Stuttgart, 1992. 477 pp., ISBN 3-519-02089-0. Mathematics and Computers in Simulation, 35(2):187–188, April 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549390019Q.

Ames:1993:BRMb

[492] W. F. Ames and C. Brezinski. Book review: The Mellin transformation and Fuchsian type partial differential equations: Z. Szmydt and B. Ziemian, Kluwer, Dordrecht, 1992. 223 pp., Dfl. 150, US$89, UK£52, ISBN 0-792-31683-5. Mathematics and Computers in Simulation, 35 (2):188, April 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549390021L.

Ames:1993:BRV

[493] W. F. Ames and C. Brezinski. Book review: Vector-valued functions and their applications: C.-G. Hu and C.- c. Yang, Kluwer, Dordrecht, 1992. 160 pp., Dfl.120, US$69, UK£42, ISBN 0-792-31605-3. Mathematics and Computers in Simulation, 35 (2):188, April 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549390020U.

Ames:1993:BRGb

[494] W. F. Ames and C. Brezinski. Book review: Generative modeling for computer graphics and CAD: J. M. Snyder, Academic Press, Boston, MA, 1992. 311 pp., US$39.95, ISBN 0-126-54040-3. Mathematics and Computers in Simulation, 35(2):189, April 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390023N.

Ames:1993:BRIa

[495] W. F. Ames and C. Brezinski. Book review: Information in the Enterprise: (It’s more than technology): G. Darnton and S. Giacoletto, Digital Press, Burlington, MA, 1992. 318 pp., US$34.95, ISBN 0-131-76173-0. Mathematics and Computers in Simulation, 35(2):189, April 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390022M.

Ames:1993:BRIb

[496] W. F. Ames and C. Brezinski. Book review: Integral inequalities and applications: D. Bainov and P. Simeonov, Kluwer, Dordrecht, 1992. 245 pp., Dfl.185, US$110, UK£64, ISBN 0-792-31714-9. Mathematics and Computers in Simulation, 35(2):189– 190, April 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549390024O.

Ames:1993:BRDa

[497] W. F. Ames and C. Brezinski. Book review: The desktop fractal design system: M. F. Barnsley, (Macintosh version 2.0, IBM Version 2.0). Academic Press, San Diego, CA, 1992. ISBN 0-120-79065-3 (MAC), 0-120-79066-1 (IBM). Mathematics and Computers in Simulation, 35 (2):190, April 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549390025P.

Ames:1993:BRH

[498] W. F. Ames and C. Brezinski. Book review: Handbook of integration: D. Zwillinger, Jones and Bartlett, Boston, MA, 1992. 367 pp., ISBN 0-867-20293-9. Mathematics and Computers in Simulation, 35(2):190, April 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549390026Q.

Ames:1993:BRAc

[499] W. F. Ames and C. Brezinski. Book review: Approximation theory VII: E. W. Cheney, C. K. Chui and L. L. Schumaker (eds.), Academic Press, San Diego, CA, 1993. 249 pp., ISBN 0-121-74589-9. Mathematics and Computers in Simulation, 35(2):191, April 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549390027R.

Ames:1993:BRAd

[500] W. F. Ames and C. Brezinski. Book review: Analysis and control of nonlinear infinite dimensional systems: V. Barbu, Academic Press, Boston, MA, 1993. 476 pp., ISBN 0-120-78145-X. Mathematics and Computers in Simulation, 35(2):191, April 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390028S.

Ames:1993:BRDb

[501] W. F. Ames and C. Brezinski. Book review: Differential manifolds: A. A. Kosinski, Academic Press, San Diego, CA, 1993. 248 pp., ISBN 0-124-21850-4. Mathematics and Computers in Simulation, 35(2):191– 192, April 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549390029T.

Ames:1993:BRSb

[502] W. F. Ames and C. Brezinski. Book review: Scientific computing with automatic result verification: E. Adams and U. Kulisch, Academic Press, Boston, MA, 1993. 612 pp., ISBN 0-120-44210-8. Mathematics and Computers in Simulation, 35(2):192, April 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390030X.

Anonymous:1993:NIb

[503] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 35(2):193–202, April 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549390031O.

Anonymous:1993:PA

[504] Anonymous. Pages 103–202 (April 1993). Mathematics and Computers in Simulation, 35(2):??, April 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Mann:1993:CSU

[505] Kenneth J. Mann. Computer simulation of an urban heat island using finite elements. Mathematics and Computers in Simulation, 35(3): 203–209, September 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549390001B.

Chronopoulos:1993:IMN

[506] A. T. Chronopoulos and C. T. Pedro. Iterative methods for nonsymmetric systems in DAEs and stiff ODEs codes. Mathematics and Computers in Simulation, 35(3):211–232, September 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549390002C.

Vignes:1993:SAR

[507] J. Vignes. A stochastic arithmetic for reliable scientific computation. Mathematics and Computers in Simulation, 35(3):233–261, September 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549390003D.

Sullivan:1993:FCC

[508] P. J. Sullivan and H. Ye. Further comments on “cloud-averaged” concentration statistics. Mathematics and Computers in Simulation, 35(3): 263–269, September 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390004E.

Loe:1993:EBS

[509] Kia-Fock Loe. Energy-based smoothing of data. Mathematics and Computers in Simulation, 35(3):271–277, September 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549390005F.

Anonymous:1993:NIc

[510] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 35(3):279–301, September 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/037847549390006G.

Anonymous:1993:PS

[511] Anonymous. Pages 203–302 (September 1993). Mathematics and Computers in Simulation, 35(3):??, September 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Tali-Maamar:1993:ACM

[512] N. Tali-Maamar, T. Damak, J.-P. Babary, and M. T. Nihtilä. Application of a collocation method for simulation of distributed parameter bioreactors. Mathematics and Computers in Simulation, 35(4):303– 319, October 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549390061X.

Osman:1993:SMN

[513] Mohamed S. Osman and Abou-Zaid H. El-Banna. Stability of multiobjective nonlinear programming problems with fuzzy parameters. Mathematics and Computers in Simulation, 35 (4):321–326, October 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390062Y.

Franksen:1993:BCM

[514] Ole Immanuel Franksen. Babbage and cryptography. Or, the mystery of Admiral Beaufort’s cipher. Mathematics and Computers in Simulation, 35 (4):327–367, October 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390063Z.

Anonymous:1993:NId

[515] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 35(4):369–382, October 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493900642.

Anonymous:1993:PO

[516] Anonymous. Pages 303–386 (October 1993). Mathematics and Computers in Simulation, 35(4):??, October 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Simos:1993:ERK

[517] T. E. Simos. Embedded Runge– Kutta methods for periodic initialvalue problems. Mathematics and Computers in Simulation, 35(5):387– 395, November 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475493900402.

Tzafestas:1993:DSC

[518] Spyros Tzafestas and Alekos Triantafyllakis. Deterministic scheduling in computing and manufacturing systems: a survey of models and algorithms. Mathematics and Computers in Simulation, 35(5):397– 434, November 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549390041R.

Kulisch:1993:NAR

[519] Ulrich Kulisch and L. B. Rall. Numerics with automatic result verification. Mathematics and Computers in Simulation, 35(5):435–450, November 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549390042S.

Ames:1993:BRNb

[520] W. F. Ames and C. Brezinski. Book review: Numerical recipes in Fortran (The art of scientific computing): W. H. Press, S. A. Teukolsky, W. T. Vetterling and B. P. Flannery, Cambridge Univ. Press, Cambridge, 2nd ed., 1992. 963 pp., US$49.95, ISBN 0-521-43064-X. Mathematics and Computers in Simulation, 35(5): 451–452, November 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390043T.

Anonymous:1993:BRI

[521] Anonymous. Book review: Intelligent decision support (Handbook of applications and advances of the rough sets theory): R. Slowinski (ed.), Kluwer, Dordrecht, 1992. 471 pp., Dfl.260, US$162, UK£88, ISBN 0-792- 3192-0. Mathematics and Computers in Simulation, 35(5):452, November 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549390044U.

Anonymous:1993:BRDa

[522] Anonymous. Book review: Delay differential equations with applications in population dynamics: Y. Kuang, Academic Press, Boston, MA, 1993. 398 pp., ISBN 0-124-27610-5. Mathematics and Computers in Simulation, 35 (5):452–453, November 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390045V.

Anonymous:1993:BRFa

[523] Anonymous. Book review: Fuzzy data analysis: H. Bandermer and W. Nather, Kluwer, Dordrecht, 1992. 341 pp., Dfl.180, US$99, UK£62, ISBN 0-792-31772-6. Mathematics and Computers in Simulation, 35 (5):453, November 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390047X.

Anonymous:1993:BRH

[524] Anonymous. Book review: HP48SX engineering mathematics library: J. F. Holland, (with ROM card). Academic Press, Boston, MA, 1992. 632 pp., ISBN 0-123-52380-X. Mathematics and Computers in Simulation, 35 (5):453, November 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390046W.

Anonymous:1993:BRNa

[525] Anonymous. Book review: Numerical methods, software and analysis: J. R. Rice, Academic Press, Boston, MA, 2nd ed., 1993. 720 pp., ISBN 0-125-87755-2. Mathematics and Computers in Simulation, 35(5): 453–454, November 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390048Y.

Anonymous:1993:BRPa

[526] Anonymous. Book review: Partial differential equations with Mathematica(R): D. Vvedensky, Addison-Wesley, Reading, MA, 1992. 465 pp., ISBN 0-201-54409-1. Mathematics and Computers in Simulation, 35 (5):454, November 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549390049Z.

Anonymous:1993:BRMa

[527] Anonymous. Book review: Mathematical T_{E}X by example: A. Borde, Academic Press, Boston, MA, 1993. 352 pp., paper, US$29.95, ISBN 0-121-17645-2. Mathematics and Computers in Simulation, 35(5):455, November 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475493900505.

Tzafestas:1993:MPH

[528] Spyros Tzafestas. Machines with a purpose: H. Rosenbrock, Oxford Univ. Press, Oxford, 1990. 223 + x pp. Mathematics and Computers in Simulation, 35(5):455–456, November 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549390051U.

Anonymous:1993:BRAa

[529] Anonymous. Book review: Asymptotic and numerical methods for partial differential equations with critical parameters: H. G. Kaper and M. Garbey (eds.), NATO Adv. Sci. Inst. Ser. Kluwer, Dordrecht, 1993. 372 pp., Dfl.245, US$156, UK£95, ISBN 0-792-32061-1. Mathematics and Computers in Simulation, 35 (5):456, November 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390054X.

Anonymous:1993:BRCa

[530] Anonymous. Book review: Computational acoustics — scattering, supercomputing and propagation: R. L. Lau, D. Lee and A. R. Robinson, North-Holland, Amsterdam, 1993. 456 pp., ISBN 0-444-89761-5. Mathematics and Computers in Simulation, 35 (5):456, November 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390052V.

Anonymous:1993:BRCb

[531] Anonymous. Book review: Computational acoustics — acoustic propagation: D. Lee, R. Vichnevetsky and A. R. Robinson (eds.), North-Holland, Amsterdam, 1993. 457 pp., ISBN 0-444-89762-3. Mathematics and Computers in Simulation, 35 (5):456, November 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390053W.

Anonymous:1993:BRE

[532] Anonymous. Book review: Explorations with the Texas instruments TI-85: J. C. Harvey and J.

W. Kenelly (eds.), Academic Press, Boston, MA, 1993. 349 pp., ISBN 0-123-29070-8. Mathematics and Computers in Simulation, 35(5):456– 457, November 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549390055Y.

Anonymous:1993:BRDb

[533] Anonymous. Book review: Differential equations with mathematica: M. L. Abell and J. P. Braselton, Academic Press, Boston, MA, 1992. 631 pp., paper, ISBN 0-120-41538-0. Mathematics and Computers in Simulation, 35(5):457, November 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475493900572.

Chafee:1993:BRD

[534] Nathaniel Chafee. Book review: Differential equations with applications to mathematical physics: W. F. Ames, E. M. Harrell II and J. V. Herod (eds.), Academic Press, New York, 1993. ISBN 0-120-56740-7. Mathematics and Computers in Simulation, 35(5):457, November 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549390056Z.

Anonymous:1993:BRSa

[535] Anonymous. Book review: Scientific computing, an introduction with parallel computing: G. Golub and J. M. Ortega, Academic Press, Boston, MA, 1993. 442 pp., ISBN 0-122-89253-4. Mathematics and Computers in Simulation, 35(5):458, November 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475493900583.

Gustafson:1993:BRI

[536] Karl Gustafson. Book review: Introduction to neural and cognitive modeling: D. S. Levine, Erlbaum, Hillsdale, NJ, 1991. 452 pp., ISBN 0-805-80267-3 or 0-805-80268-1 (paper). Mathematics and Computers in Simulation, 35 (5):458–460, November 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493900594.

Anonymous:1993:NIe

[537] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 35(5):461–467, November 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475493900608.

Anonymous:1993:PN

[538] Anonymous. Pages 387–468 (November 1993). Mathematics and Computers in Simulation, 35(5):??, November 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Sarper:1993:MCS

[539] Hüseyin Sarper. Monte Carlo simulation for analysis of the optimum value distribution in stochastic mathematical programs. Mathematics and Computers in Simulation, 35(6):469– 480, December 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475493900653.

Point:1993:PES

[540] N. Point, A. Vande Wouwer, M. Remy, and M. Zeitz. PC environment for simulation and parameter estimation of distributed parameter systems. Mathematics and Computers in Simulation, 35(6):481–491, December 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475493900664.

Arsham:1993:LSB

[541] Hossein Arsham, Darush Davani, and Jae B. Yu. A linear symbolic-based approach to matrix inversion. Mathematics and Computers in Simulation, 35 (6):493–500, December 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493900675.

Mahmoud:1993:AIA

[542] F. F. Mahmoud, A. K. Al-Saffar, and K. A. Hassan. An adaptive incremental approach for the solution of convex programming models. Mathematics and Computers in Simulation, 35 (6):501–508, December 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493900686.

Taha:1993:DDE

[543] Thiab R. Taha. A differentialdifference equation for a KdV-MKdV equation. Mathematics and Computers in Simulation, 35(6):509–512, December 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475493900697.

Karageorghis:1993:EBB

[544] Andreas Karageorghis. On the equivalence between basis recombination and boundary bordering formulations for spectral collocation methods in rectangular domains. Mathematics and Computers in Simulation, 35(6): 513–524, December 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390070B.

Anonymous:1993:BRAb

[545] Anonymous. Book review: Applied partial differential equations: D. W. Trim, PWS-Kent, Boston, MA, 1990. 485 pp., ISBN 0-534-92134-5. Mathematics and Computers in Simulation, 35(6):525, December 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493900723.

Anonymous:1993:BRB

[546] Anonymous. Book review: Boundary value problems and partial differential equations: M. Humi and W. B. Miller, PWS-Kent, Boston, MA, 1992. 342 pp., ISBN 0-534-92880-3. Mathematics and Computers in Simulation, 35(6):525, December 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493900712.

Anonymous:1993:BRMb

[547] Anonymous. Book review: Mathematical modeling: M. M. Meerschaert, Academic Press, Boston, MA, 1993. 287 pp., US$49.95, ISBN 0-124-87650-1. Mathematics and Computers in Simulation, 35(6):525– 526, December 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475493900734.

Anonymous:1993:BRLa

[548] Anonymous. Book review: Lebesgue integration on Euclidean space: F. Jones,Jones and Bartlett, Boston, MA, 1993. 588 pp., ISBN 0-867-20203-3. Mathematics and Computers in Simulation, 35(6):526, December 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475493900745.

Anonymous:1993:BRPb

[549] Anonymous. Book review: Probability models: S. M. Ross, Academic Press, Boston, MA 5th ed., 1993. 556 pp., ISBN 0-125-98455-3. Mathematics and Computers in Simulation, 35 (6):526, December 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493900756.

Anonymous:1993:BRCc

[550] Anonymous. Book review: Contemporary linear systems using MAT-LAB(R) (computer disks included): R. D. Strum and D. E. Kirk, PWS, Boston, MA, 1994. 673 pp. (paper), US$34.50, ISBN 0-534-93273-8. Mathematics and Computers in Simulation, 35(6):527, December 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493900767.

Anonymous:1993:BRR

[551] Anonymous. Book review: Regenerative stochastic simulation: G. S. Shedler, Academic Press, Boston, MA, 1993. 400 pp., ISBN 0-126-39360-5. Mathematics and Computers in Simulation, 35(6):527, December 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475493900778.

Anonymous:1993:BRLb

[552] Anonymous. Book review: Linear programs and related problems: E. D. Nering and A. W. Tucker, Academic Press, Boston, MA, 1993. 584 pp. ISBN 0-125-15440-2. Mathematics and Computers in Simulation, 35 (6):528, December 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390079A.

Anonymous:1993:BRMc

[553] Anonymous. Book review: MuPAD (Multi-Processing Algebra Data Tool (Benutzerhandbuch)): B. Fuchssteiner, W. Wiwianka, K. Gottheil, A. Kemper, O. Kluge, K. Morrise, H. Naundorf, G. Oevel and T. Schulze, Birkhäuser, Basel, 1993. 414 pp., Sfr. 62, ISBN 3-764-32872-X. Mathematics and Computers in Simulation, 35 (6):528, December 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493900789.

Anonymous:1993:BRFb

[554] Anonymous. Book review: Fractal image compression: M. F. Barnsley and L. P. Hurd, Peters, Wellesley, MA, 1993. 244 pp., US$49.95, ISBN 1-568-81000-8. Mathematics and Computers in Simulation, 35(6): 528–529, December 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390080E.

Anonymous:1993:BRSb

[555] Anonymous. Book review: Shape optimization and free boundaries: M. C. Delfour (ed.), NATO Adv. Sci. Inst. Ser. Kluwer, Dordrecht, 1992. 462 pp., Dfl.255, US$150, UK£86, ISBN 0-792-31944-3. Mathematics and Computers in Simulation, 35 (6):529, December 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493900826.

Anonymous:1993:BRTa

[556] Anonymous. Book review: Treatise on analysis, Volume 10-VIII:

J. Dieudonné (ed.), Academic Press, Boston, MA, 1993. 356 pp., US$110, ISBN 0-122-15508-4. Mathematics and Computers in Simulation, 35 (6):529, December 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493900815.

Anonymous:1993:BRLc

[557] Anonymous. Book review: The logic programming tutor: J. Paine, Kluwer, Dordrecht, 1992. 365 pp., Dfl.100, US$55, UK£34, ISBN 0-792-31448-4. Mathematics and Computers in Simulation, 35(6):529–530, December 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475493900837.

Anonymous:1993:BRAc

[558] Anonymous. Book review: Asymptotic properties of solutions of nonautonomous ordinary differential equations: I. T. Kiguradze and T. A. Chanturia, Kluwer, Dordrecht, 1993. 331 pp., Dfl.230, US$154, UK£92.50, ISBN 0-792-32059-X. Mathematics and Computers in Simulation, 35 (6):530, December 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493900859.

Anonymous:1993:BRTb

[559] Anonymous. Book review: Theory and applications of convolution integral equations: H. M. Srivastava and R. G. Buschman, Kluwer, Dordrecht, 1992. 239 pp., Dfl.165, US$96, UK£56, ISBN 0-792-31891-9. Mathematics and Computers in Simulation, 35(6):530, December 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475493900848.

Anonymous:1993:BRCd

[560] Anonymous. Book review: Classical and new inequalities in analysis: D. S. Mitrinovic, J. E. Pecaric and A. M. Fink, Kluwer, Dordrecht, 1992. 725 pp., Dfl.425, US$285, UK£170.50, ISBN 0-792-32064-6. Mathematics and Computers in Simulation, 35 (6):531, December 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549390086A.

Anonymous:1993:BRFc

[561] Anonymous. Book review: Fractals everywhere: M. F. Barnsley, Academic Press, Boston, MA, end ed., 1993. 531 pp., US$49.95, ISBN 0-120-79061-0. Mathematics and Computers in Simulation, 35(6):531, December 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549390087B.

Anonymous:1993:BRNb

[562] Anonymous. Book review: Nonlinear feedback control systems (An operator theory approach): R. J. P. Figueredo and G. Chen, Academic Press, Boston, MA, 1993. 220 pp., ISBN 0-122-08630-9. Mathematics and Computers in Simulation, 35 (6):531, December 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549390088C.

Anonymous:1993:NIf

[563] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 35(6):533–539, December 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/037847549390089D.

Anonymous:1993:AIV

[564] Anonymous. Author index to volume 35 (1993). Mathematics and Computers in Simulation, 35(6):541– 543, December 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549390090H.

Anonymous:1993:CV

[565] Anonymous. Contents of volume 35 (1993). Mathematics and Computers in Simulation, 35(6):545–546, December 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475493900918.

Anonymous:1993:PD

[566] Anonymous. Pages 469–546 (December 1993). Mathematics and Computers in Simulation, 35(6):??, December 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1994:EBa

[567] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 36(1):ii, April 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494900442.

deCursi:1994:NML

[568] J. E. Souza de Cursi. Numerical methods for linear boundary value problems based on Feyman–Kac representations. Mathematics and Computers in Simulation, 36(1):1–16, April 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494900450.

Garcia-Olivares:1994:FPE

[569] Antonio García-Olivares and Antonio Muñoz. Fokker–Planck equations in the simulation of complex systems. Mathematics and Computers in Simulation, 36(1):17–48, April 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494900469.

Noor:1994:NMS

[570] Muhammad Aslam Noor and Rafi Ahmad Ashrafi. On a numerical method for solving obstacle problems. Mathematics and Computers in Simulation, 36(1):49–55, April 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494900477.

Das:1994:CNS

[571] Bigyani Das, Stanly Steinberg, Danjie Zhang, and Thomas Robey. Comparisons of numerical solution methods for differential equations with discontinuous coefficients. Mathematics and Computers in Simulation, 36(1):57–75, April 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900485.

Anonymous:1994:NIa

[572] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 36(1):77–87, April 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900493.

Anonymous:1994:PAa

[573] Anonymous. Pages 1–88 (April 1994). Mathematics and Computers in Simulation, 36(1):??, April 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Singh:1994:P

[574] Madan G. Singh and Nuria Piera-Carrete. Preface. Mathematics and Computers in Simulation, 36(2): 89–90, June 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549490023X.

Forbus:1994:SES

[575] Kenneth D. Forbus. Self-explanatory simulators: making computers partners in the modeling process. Mathematics and Computers in Simulation, 36(2):91–101, June 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900248.

Singh:1994:DTS

[576] Madan G. Singh. Decision technologies for supporting the interplay between qualitative and quantitative aspects of managerial decision making. Mathematics and Computers in Simulation, 36(2):103–114, June 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900256.

Aguilar-Martin:1994:QCD

[577] Joseph Aguilar-Martin. Qualitative control, diagnostic and supervision of complex processes. Mathematics and Computers in Simulation, 36 (2):115–127, June 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900264.

Bartoli:1994:QRC

[578] J. A. Bartoli and J. L. Le Moigne. Qualitative reasoning and complex symbol processing. Mathematics and Computers in Simulation, 36(2):129– 136, June 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475494900272.

Wellman:1994:ICM

[579] Michael P. Wellman. Inference in cognitive maps. Mathematics and Computers in Simulation, 36(2):137– 148, June 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475494900280.

Leyval:1994:QAD

[580] Lydie Leyval, Jacky Montmain, and Sylviane Gentil. Qualitative analysis for decision making in supervision of industrial continuous processes. Mathematics and Computers in Simulation, 36(2):149–163, June 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900299.

Anonymous:1994:NIb

[581] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 36(2):165–172, June 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900302.

Anonymous:1994:PJ

[582] Anonymous. Pages 89–172 (June 1994). Mathematics and Computers in Simulation, 36(2):??, June 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Attili:1994:UBE

[583] Basem S. Attili and Yaqoub Shehadeh. The use of block elimination for the calculation of some types of singularities efficiently. Mathematics and Computers in Simulation, 36 (3):173–184, August 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900027.

Warsi:1994:SSN

[584] Z. U. A. Warsi and G. P. Koomullil. Spectral solutions of the Navier– Stokes equations in arbitrary twodimensional domains. Mathematics and Computers in Simulation, 36 (3):185–193, August 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900035.

Simos:1994:ESN

[585] T. E. Simos and G. V. Mitsou. An expert system for the numerical solution of the one-dimensional Schrödinger equation. Mathematics and Computers in Simulation, 36 (3):195–202, August 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900043.

Villasenor:1994:FSC

[586] R. Villaseñor. A flame sheet calculation of a confined buoyancy laminar diffusion flame. Mathematics and Computers in Simulation, 36 (3):203–208, August 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900051.

Bursal:1994:IMS

[587] Faruk H. Bursal and Benson H. Tongue. Interpolated mapping system identification as a numerical algorithm. Mathematics and Computers in Simulation, 36(3):209– 220, August 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549490006X.

Kaiser:1994:ESI

[588] Mark J. Kaiser, Kenneth L. Kaiser, and Walter L. Weeks. Electrohydrodynamic simulator investigations. Mathematics and Computers in Simulation, 36(3):221–240, August 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900078.

Razzaghi:1994:PCM

[589] Mohsen Razzaghi and Gamal N. Elnagar. A pseudospectral collocation method for the brachistochrone problem. Mathematics and Computers in Simulation, 36(3):241– 246, August 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475494900086.

Dantas:1994:SNT

[590] J. F. Dantas, J. A. Cuminato, G. F. Leal Ferreira, and M. T. Figueiredo. Solution of a non-trivial space charge problem by the hodographic method. Mathematics and Computers in Simulation, 36(3):247– 259, August 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475494900094.

Anonymous:1994:NIc

[591] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 36(3):261–267, August 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494900108.

Anonymous:1994:PAb

[592] Anonymous. Pages 173–268 (August 1994). Mathematics and Computers in Simulation, 36(3):??, August 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Houstis:1994:TIC

[593] E. N. Houstis, J. R. Rice, and R. Vichnevetsky. Third international conference on expert systems for scientific computing. Mathematics and Computers in Simulation, 36(4–6):267–503, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). Special triple issue.

Boisvert:1994:AIV

[594] Ronald F. Boisvert. The architecture of an intelligent virtual mathematical software repository system. Mathematics and Computers in Simulation, 36 (4–6):269–279, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494900620.

Sastry:1994:KBF

[595] Venkat V. S. S. Sastry and John C. Mason. Knowledge based front-end to NAG library — KASTLE. Mathematics and Computers in Simulation, 36 (4–6):281–292, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900639.

Laug:1994:DKB

[596] Patrick Laug. DOMINO: a knowledgebased system for the users of a finite element library. Mathematics and Computers in Simulation, 36(4– 6):293–301, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900647.

Grimm:1994:AES

[597] Felix Grimm, Horst Bunke, and Jürg Hählen. An approach to expert systems for image processing software libraries. Mathematics and Computers in Simulation, 36(4–6):303– 313, October 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475494900655.

Carpraux:1994:SKB

[598] Jean-François Carpraux and Jocelyne Erhel. SESAME: a knowledgebased system for eigenvalue problems. Mathematics and Computers in Simulation, 36(4–6):315– 325, October 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475494900663.

vanZuylen:1994:KBS

[599] Henk J. van Zuylen. Knowledge based support of users of numerical programs. Mathematics and Computers in Simulation, 36(4–6):327– 336, October 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475494900671.

Sundaram:1994:TSV

[600] Ananthapadmanaban Sundaram, James K. McDowell, and Martin C. Hawley. Task structure: a vocabulary for integrating numerical methods and knowledge-based systems. Mathematics and Computers in Simulation, 36 (4–6):337–346, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549490068X.

Shekhar:1994:TIP

[601] Chandra Shekhar, Sabine Moisan, and Monique Thonnat. Towards an intelligent problem-solving environment for signal processing. Mathematics and Computers in Simulation, 36(4– 6):347–359, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494900698.

Willamowski:1994:DSC

[602] Jutta Willamowski, François Chevenet, and François Jean-Marie. A development shell for cooperative problemsolving environments. Mathematics and Computers in Simulation, 36(4– 6):361–379, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900701.

Hurley:1994:CBR

[603] Neil Hurley. A case based reasoning approach to mesh specification for adaptive finite element analysis. Mathematics and Computers in Simulation, 36 (4–6):381–388, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549490071X.

Cahill:1994:KBA

[604] Eamonn Cahill. Knowledge-based algorithm construction for real-world engineering PDEs. Mathematics and Computers in Simulation, 36(4–6): 389–400, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900728.

Cuyt:1994:IOO

[605] Annie Cuyt, Brigitte Verdonk, and Jan Verelst. Intelligent object-oriented scientific computation. Mathematics and Computers in Simulation, 36(4– 6):401–411, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900736.

Labrie:1994:EAM

[606] R. Labrie, C. Thilloy, P. A. Tanguy, and G. H. Moll. An expert assistant to monitor finite element simulations. Mathematics and Computers in Simulation, 36(4–6):413– 422, October 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475494900744.

Zhao:1994:ICA

[607] Feng Zhao. Intelligent computing about complex dynamical systems. Mathematics and Computers in Simulation, 36(4–6):423– 432, October 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475494900752.

Anupam:1994:SPS

[608] Vinod Anupam, Chandrajit Bajaj, Fausto Bernardini, Steve Cutchin, Jindon Chen, Daniel Schikore, Guoliang Xu, Peinan Zhang, and Weiping Zhang. Scientific problem solving in a distributed and collaborative multimedia environment. Mathematics and Computers in Simulation, 36 (4–6):433–442, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494900760.

Li:1994:DDS

[609] Peiyuan Li and Richard L. Peskin. Domain decomposition for singular perturbation PDEs. Mathematics and Computers in Simulation, 36(4–6): 443–455, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900779.

Li:1994:NSM

[610] Peiyuan Li and Richard L. Peskin. A new search method for domain decomposition for ODEs. Mathematics and Computers in Simulation, 36(4– 6):457–466, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900787.

Froncioni:1994:QFD

[611] A. M. Froncioni and R. L. Peskin. Qualitative flow descriptors for unstructed triangular grids. Mathematics and Computers in Simulation, 36 (4–6):467–477, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900795.

Hoffman:1994:SVL

[612] Christoph M. Hoffman, Elias N. Houstis, John R. Rice, Ann Christine Catlin, Margaret Gaitatzes, Sanjiva Weerawarana, N-H. Linda Wang, Christos G. Takoudis, and David G. Taylor. SoftLab — a virtual laboratory for computational science. Mathematics and Computers in Simulation, 36 (4–6):479–491, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900809.

Lambe:1994:OOM

[613] Larry Lambe and Richard Luczak. Object-oriented mathematical programming and symbolic/numeric interface. Mathematics and Computers in Simulation, 36(4–6):493– 503, October 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475494900817.

Anonymous:1994:NId

[614] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 36(4–6):505–516, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/0378475494900825.

Anonymous:1994:AIVa

[615] Anonymous. Author index to volume (1994). Mathematics and Computers in Simulation, 36(4–6):517– 519, October 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475494900833.

Anonymous:1994:PO

[616] Anonymous. Pages 269–520 (October 1994). Mathematics and Computers in Simulation, 36(4–6):??, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1994:EBb

[617] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 37(1):ii, November 21, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900531.

Lamotte:1994:CSA

[618] J. L. Lamotte and R. Alt. Comparison of simulated annealing algorithms for image restoration. Mathematics and Computers in Simulation, 37(1): 1–15, November 21, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549490054X.

Yunker:1994:SOG

[619] James M. Yunker and Jeffrey D. Tew. Simulation optimization by genetic search. Mathematics and Computers in Simulation, 37(1):17– 28, November 21, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900558.

Klempous:1994:SNA

[620] Ryszard Klempous, Jerzy Kotowski, and Ewa Szlachcic. Some numerical aspects of the gradient procedure for design of nonlinear transportation networks. Mathematics and Computers in Simulation, 37(1):29– 35, November 21, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900566.

Noye:1994:EST

[621] B. J. Noye and M. Dehghan. Explicit solution of two-dimensional diffusion subject to specification of mass. Mathematics and Computers in Simulation, 37(1):37–45, November 21, 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475494900574.

Qiao:1994:PEW

[622] Hongzhu Qiao and Chris P. Tsokos. Parameter estimation of the Weibull probability distribution. Mathematics and Computers in Simulation, 37(1): 47–55, November 21, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900582.

Zenger:1994:SVD

[623] Kai Zenger and Raimo Ylinen. Simulation of variable delays in material transport models. Mathematics and Computers in Simulation, 37(1): 57–72, November 21, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900590.

Guilain:1994:VNS

[624] S. Guilain and J. Vignes. Validation of numerical software results — Application to the computation of apparent heat release in directinjection diesel engines. Mathematics and Computers in Simulation, 37(1): 73–92, November 21, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900604.

Anonymous:1994:NIe

[625] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 37(1):93–98, November 21, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475494900612.

Anonymous:1994:PNa

[626] Anonymous. Pages 1–98 (21 November 1994). Mathematics and Computers in Simulation, 37(1):??, November 21, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1994:P

[627] Anonymous. Preface. Mathematics and Computers in Simulation, 37(2– 3):99, November 30, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494001308.

Pfeiffer:1994:SCC

[628] Bernd-Markus Pfeiffer and Rolf Isermann. Selftuning of classical controllers with fuzzy-logic. Mathematics and Computers in Simulation, 37(2–3): 101–110, November 30, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900116.

Kobi:1994:FDI

[629] A. Kobi, S. Nowakowski, and J. Ragot. Fault detection-isolation and control reconfiguration. Mathematics and Computers in Simulation, 37(2–3):111– 117, November 30, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900124.

Jadot:1994:PML

[630] F. Jadot, D. Galardini, and R. Gorez. Parsimonious modelling of link flexibility in a robotic system. Mathematics and Computers in Simulation, 37(2–3): 119–123, November 30, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900132.

Ostroversnik:1994:CRS

[631] M. Ostroveršnik, Z. Šehić, B. Zupančič, and M. Šega. Concept representation of the software tool PID-Master for systems modelling and controllers tuning. Mathematics and Computers in Simulation, 37(2–3):125– 131, November 30, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900140.

Tzafestas:1994:PST

[632] Spyros Tzafestas and George Kapsiotis. PID self-tuning control combining pole-placement and parameter optimization features. Mathematics and Computers in Simulation, 37(2–3):133– 142, November 30, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000026.

Li:1994:TCP

[633] M. X. Li, P. M. Bruijn, and H. B. Verbruggen. Tuning cascade PID controllers using fuzzy logic. Mathematics and Computers in Simulation, 37(2–3): 143–151, November 30, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000034.

Luttmann:1994:DCS

[634] Reiner Luttmann, Gerhard Bitzer, and Jens Hartkopf. Development of control strategies for high cell density cultivations. Mathematics and Computers in Simulation, 37(2–3):153– 164, November 30, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000042.

Dochain:1994:ADM

[635] D. Dochain and B. Bouaziz. Approximation of the dynamical model of fixed bed reactors via a singular perturbation approach. Mathematics and Computers in Simulation, 37(2–3):165– 172, November 30, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000050.

Tali-Maamar:1994:CND

[636] N. Tali-Maamar, M. T. Nihtilä, and J. P. Babary. On control of nonlinear distributed parameter bioreactors. Mathematics and Computers in Simulation, 37(2–3):173–181, November 30, 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475494900159.

Mikhailov:1994:HCC

[637] Lyudmil Mikhailov and Raymond Hanus. Hierarchical control of congested urban traffic — mathematical modelling and simulation. Mathematics and Computers in Simulation, 37(2–3): 183–188, November 30, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494900167.

Balestrino:1994:NAP

[638] A. Balestrino and F. Bini Verona. A new adaptive polynomial neural network. Mathematics and Computers in Simulation, 37(2–3):189– 194, November 30, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494900175.

Zagorianos:1994:RMM

[639] A. Zagorianos, E. Kontoyannis, and S. Tzafestas. Resolved motion model based predictive control of redundant robots. Mathematics and Computers in Simulation, 37(2–3):195– 205, November 30, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494900183.

Codron:1994:MGP

[640] P. Codron and P. Boucher. Multivariable generalized predictive control with new multiple reference model: a robust stability analysis. Mathematics and Computers in Simulation, 37(2–3): 207–219, November 30, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900191.

Tahboub:1994:NMM

[641] K. A. Tahboub and P. C. Müller. A novel model manipulation of elastic-joint robots for control purposes. Mathematics and Computers in Simulation, 37(2–3):221–225, November 30, 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475494900205.

Abdalla:1994:NPR

[642] E. Abdalla, H. J. Pu, M. Müller, A. A. Tantawy, L. Abdelatif, and H. Nour Eldin. A novel parallel recursive Newton–Euler algorithm for modelling and computation of robot dynamics. Mathematics and Computers in Simulation, 37(2–3):227– 240, November 30, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900213.

Anonymous:1994:NIf

[643] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 37(2–3):241–246, November 30, 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475494900221.

Anonymous:1994:PNb

[644] Anonymous. Pages 99–246 (30 November 1994). Mathematics and Computers in Simulation, 37(2–3):??, November 30, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Taha:1994:F

[645] Thiab R. Taha, Mark J. Ablowitz, Jerry L. Bona, and Alan C. Newell. Foreword. Mathematics and Computers in Simulation, 37(4–5):247, December 5, 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475494900426.

Ablowitz:1994:HMN

[646] Mark J. Ablowitz and Constance M. Schober. Homoclinic manifolds and numerical chaos in the nonlinear Schrödinger equation. Mathematics and Computers in Simulation, 37(4–5): 249–264, December 5, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494000182.

Bona:1994:ABS

[647] Jerry Bona, Keith Promislow, and Gene Wayne. On the asymptotic behavior of solutions to nonlinear, dispersive, dissipative wave equations. Mathematics and Computers in Simulation, 37(4–5):265–277, December 5, 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475494000190.

Chinni:1994:ASL

[648] V. R. Chinni, C. R. Menyuk, and P. K. A. Wai. Accurate solution of light propagation in optical waveguides using Richardson extrapolation. Mathematics and Computers in Simulation, 37(4–5):279– 291, December 5, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000204.

Fokas:1994:IBV

[649] A. S. Fokas and A. R. Its. An initial-boundary value problem for the Korteweg–de Vries equation. Mathematics and Computers in Simulation, 37(4–5):293–321, December 5, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475494000212.

Forest:1994:ESI

[650] M. Gregory Forest, Christopher G. Goedde, and Amarendra Sinha. Effects of symmetry on instabilitydriven chaotic energy transport in near-integrable, Hamiltonian systems. Mathematics and Computers in Simulation, 37(4–5):323–339, December 5, 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475494000220.

Grammaticos:1994:IDS

[651] B. Grammaticos and B. Dorizzi. Integrable discrete systems and numerical integrators. Mathematics and Computers in Simulation, 37(4–5):341– 352, December 5, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494000239.

Herbst:1994:SMN

[652] B. M. Herbst, F. Varadi, and M. J. Ablowitz. Symplectic methods for the nonlinear Schrödinger equation. Mathematics and Computers in Simulation, 37(4–5):353– 369, December 5, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494000247.

Hirota:1994:DWI

[653] Ryogo Hirota, Tatsuya Imai, Satoshi Tsujimoto, and Yasuhiro Ohta. Discrete 2N-wave interaction. Mathematics and Computers in Simulation, 37(4–5):371–383, December 5, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/0378475494000255.

Indik:1994:CPF

[654] R. A. Indik and A. C. Newell. Computing for pattern forming systems in nonlinear optics. Mathematics and Computers in Simulation, 37(4–5):385– 403, December 5, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000263.

Karakashian:1994:ASG

[655] Ohannes Karakashian and William McKinney. On the approximation of solutions of the generalized Korteweg–de Vries–Burgers equation. Mathematics and Computers in Simulation, 37(4–5): 405–416, December 5, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000271.

Nakayama:1994:MSD

[656] Kazuaki Nakayama and Miki Wadati. Motion of surfaces in 3-dimensional space. Mathematics and Computers in Simulation, 37(4–5):417–430, December 5, 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549400028X.

Osborne:1994:AAN

[657] A. R. Osborne. Automatic algorithm for the numerical inverse scattering transform of the Korteweg– de Vries equation. Mathematics and Computers in Simulation, 37(4–5):431– 450, December 5, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000298.

Ramani:1994:QHL

[658] A. Ramani, B. Grammaticos, and E. Caurier. Quantum Hamiltonians from level statistics via dressing transformations. Mathematics and Computers in Simulation, 37(4–5):451– 459, December 5, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494000301.

Taha:1994:NSC

[659] Thiab R. Taha. Numerical simulations of the complex modified Korteweg– de Vries equation. Mathematics and Computers in Simulation, 37(4–5):461– 467, December 5, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549400031X.

White:1994:NSS

[660] P. W. White and J. A. C. Weideman. Numerical simulation of solitons and dromions in the Davey– Stewartson system. Mathematics and Computers in Simulation, 37(4–5):469– 479, December 5, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494000328.

Anonymous:1994:NIg

[661] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 37(4–5):481–486, December 5, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/0378475494900434.

Anonymous:1994:PDa

[662] Anonymous. Pages 247–486 (5 December 1994). Mathematics and Computers in Simulation, 37(4–5):??, December 5, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Rafaralahy:1994:NST

[663] H. Rafaralahy, M. Darouach, and G. Krzakala. Numerical simulation of thermal process in steel industry. Mathematics and Computers in Simulation, 37(6):487–490, December 31, 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475494900507.

Ficola:1994:SFE

[664] Antonio Ficola, Michele La Cava, and Pietro Muraca. A simplified finite element model of multilink robots with flexible links. Mathematics and Computers in Simulation, 37(6):491– 503, December 31, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000565.

vanderWijngaart:1994:CTF

[665] Rob F. van der Wijngaart. Concepts of TRIFIT, a flexible software environment for problems in two space dimensions. Mathematics and Computers in Simulation, 37(6):505– 525, December 31, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000573.

Sod:1994:REM

[666] Gary A. Sod. A random element method for a thermal boundary layer. Mathematics and Computers in Simulation, 37(6):527–538, December 31, 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475494000999.

El-Zoheiry:1994:QSS

[667] H. El-Zoheiry, L. Iskandar, and B. El-Naggar. The Quintic spline for solving the Korteweg–de Vries equation. Mathematics and Computers in Simulation, 37(6):539–549, December 31, 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475494001014.

Alastruey:1994:ISD

[668] Carlos F. Alastruey and JoséR. González de Mendívil. Interconnected systems with delays. Mathematics and Computers in Simulation, 37(6):551– 569, December 31, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494001022.

Anonymous:1994:NIh

[669] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 37(6):571–579, December 31, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/0378475494900515.

Anonymous:1994:AIVb

[670] Anonymous. Author index to volume 37 (1994). Mathematics and Computers in Simulation, 37(6):581– 584, December 31, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494900523.

Anonymous:1994:PDb

[671] Anonymous. Pages 487–584 (31 December 1994). Mathematics and Computers in Simulation, 37(6):??, December 31, 1994. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Bouleau:1995:I

[672] N. Bouleau. Introduction. Mathematics and Computers in Simulation, 38(1–3):1–5, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495900217.

Gaines:1995:BIS

[673] J. G. Gaines. A basis for iterated stochastic integrals. Mathematics and Computers in Simulation, 38(1–3):7–11, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493E00619.

Castell:1995:EAM

[674] Fabienne Castell and Jessica Gaines. An efficient approximation method for stochastic differential equations by means of the exponential Lie series. Mathematics and Computers in Simulation, 38(1–3):13– 19, May 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475493E0062A.

Kazimierczyk:1995:PIM

[675] Piort Kazimierczyk. To the parametric identification of Markov diffusions; the use of the maximum quadratic variation functional. Mathematics and Computers in Simulation, 38(1–3):21–33, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475493E0063B.

Bally:1995:ESS

[676] Vlad Bally and Denis Talay. The Euler scheme for stochastic differential equations: error analysis with Malliavin calculus. Mathematics and Computers in Simulation, 38(1–3): 35–41, May 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475493E0064C.

Bossy:1995:SPM

[677] Mireille Bossy and Denis Talay. A stochastic particle method for some one-dimensional nonlinear p.d.e. Mathematics and Computers in Simulation, 38(1–3):43–50, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475493E0065D.

Grorud:1995:EDA

[678] Axel Grorud and Monique Pontier. Equation différentielle anticipative sur une variété et approximations. Mathematics and Computers in Simulation, 38(1–3):51–61, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493E0066E.

Hofmann:1995:SWN

[679] Norbert Hofmann. Stability of weak numerical schemes for stochastic differential equations. Mathematics and Computers in Simulation, 38(1–3): 63–68, May 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475493E0067F.

Platen:1995:WIP

[680] Eckhard Platen. On weak implicit and predictor–corrector methods. Mathematics and Computers in Simulation, 38(1–3):69–76, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493E0068G.

Milshtein:1995:SBV

[681] G. N. Milshtein. The solving of boundary value problems by numerical integration of stochastic equations. Mathematics and Computers in Simulation, 38(1–3):77–85, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493E0069H.

BenAlaya:1995:REE

[682] Mohamed Ben Alaya. Résolution des équations elliptiques par la méthode du shift. Mathematics and Computers in Simulation, 38(1–3):87– 96, May 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475493E0070L.

Janicki:1995:CSD

[683] Aleksander Janicki. Computer simulation of diffusions driven by αstable Lévy motion. Mathematics and Computers in Simulation, 38 (1–3):97–101, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493E0071C.

Liu:1995:DCR

[684] Yingjie Liu. Discretization of a class of reflected diffusion processes. Mathematics and Computers in Simulation, 38(1–3):103–108, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493E0072D.

Slominski:1995:SRA

[685] Leszek Słomiński. Some remarks on approximation of solutions of SDE’s with reflecting boundary conditions. Mathematics and Computers in Simulation, 38(1–3):109–117, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493E0073E.

Lepingle:1995:ESR

[686] D. Lépingle. Euler scheme for reflected stochastic differential equations. Mathematics and Computers in Simulation, 38(1–3):119–126, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493E0074F.

Larcher:1995:NIH

[687] G. Larcher and W. Ch. Schmid. On the numerical integration of highdimensional Walsh-series by quasi-Monte Carlo methods. Mathematics and Computers in Simulation, 38 (1–3):127–134, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493E0075G.

Chauvin:1995:BPT

[688] Brigitte Chauvin. Branching processes, trees and the Boltzmann equation. Mathematics and Computers in Simulation, 38(1–3):135– 141, May 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475493E0076H.

Fournie:1995:MDE

[689] Eric Fournié. Minimum distance estimation and testing for interest rate models. Mathematics and Computers in Simulation, 38(1–3):143– 150, May 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475493E0077I.

Zhang:1995:FQE

[690] Xiaolan Zhang. Formules quasiexplicites pour les options américaines dans un modèle de diffusion avec sauts. Mathematics and Computers in Simulation, 38(1–3):151– 161, May 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475493E0078J.

Akian:1995:MAP

[691] Marianne Akian, Jose Luis Menaldi, and Agnès Sulem. Multi-asset portfolio selection problem with transaction costs. Mathematics and Computers in Simulation, 38(1–3):163– 172, May 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475493E0079K.

Barraquand:1995:MCI

[692] Jérôme Barraquand. Monte Carlo integration, quadratic resampling, and asset pricing. Mathematics and Computers in Simulation, 38(1–3):173– 182, May 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475493E0080O.

Clouet:1995:ELP

[693] J. F. Clouet, J. P. Fouque, and M. Postel. Estimation of local power spectral densities for non-stationary signals using wavelet transform. Mathematics and Computers in Simulation, 38(1–3):183–188, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493E0081F.

Thiran:1995:SKN

[694] Patrick Thiran and Martin Hasler. Study of the Kohonen network with a discrete state space. Mathematics and Computers in Simulation, 38 (1–3):189–197, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493E0082G.

Xu:1995:SPB

[695] Kedai Xu. Stochastic pitchfork bifurcation: numerical simulations and symbolic calculations using MAPLE. Mathematics and Computers in Simulation, 38(1–3):199– 209, May 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475493E0083H. Prob-

abilités numériques (Paris, 1992).

Wagner:1995:SPM

[696] Wolfgang Wagner. Stochastic particle methods and approximation of the Boltzmann equation. Mathematics and Computers in Simulation, 38 (1–3):211–216, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475493E0084I.

Ogawa:1995:SPS

[697] Shigeyoshi Ogawa. Some problems in the simulation of nonlinear diffusion processes. Mathematics and Computers in Simulation, 38(1–3): 217–223, May 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475493E0085J.

Kushner:1995:ENM

[698] Harold J. Kushner and Jichuan Yang. An effective numerical method for controlled routing in large trunk line networks. Mathematics and Computers in Simulation, 38(1–3):225– 239, May 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475493E0086K.

Hornung:1995:LE

[699] Ulrich Hornung. Letter to the editor. Mathematics and Computers in Simulation, 38(1–3):241, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/0378475495900233.

Mittelmann:1995:ARP

[700] Hans D. Mittelmann. Author’s reply to the preceding letter. Mathematics and Computers in Simulation, 38(1–3):241, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495900551.

Anonymous:1995:NIa

[701] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 38(1–3):243–253, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549590025X.

Anonymous:1995:EBa

[702] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 38(1–3):ifc, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495900195.

Anonymous:1995:PM

[703] Anonymous. Pages 1–253 (May 1995). Mathematics and Computers in Simulation, 38(1–3):??, May 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Olivier:1995:FIT

[704] Guy Olivier and Réal-Paul Bouchard. The Fourth IMACS TC1-Conference. Mathematics and Computers in Simulation, 38(4–6):255, August 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475495900136.

Lowther:1995:NTS

[705] D. A. Lowther. New trends in the simulation of magnetic devices. Mathematics and Computers in Simulation, 38 (4–6):257–262, August 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549500035V.

LeDoeuff:1995:GPN

[706] R. Le Doeuff and M. F. Benkhoris. General principles and new trends in the simulation of static converters and drives. Mathematics and Computers in Simulation, 38(4– 6):263–270, August 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549500036W.

Kutkut:1995:AVC

[707] Nasser H. Kutkut, Hassan M. Cherradi, and Thomas A. Lipo. Analysis of voltage controlled induction motors using quasi-rotating reference frame. Mathematics and Computers in Simulation, 38(4–6):271– 281, August 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549500037X.

Lavoie:1995:RTS

[708] M. Lavoie, V. Qué-Do, J. L. Houle, and J. Davidson. Real-time simulation of power system stability using parallel digital signal processors. Mathematics and Computers in Simulation, 38 (4–6):283–292, August 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549400038L.

Le-Huy:1995:DRT

[709] Hoang Le-Huy and Jean-Claude Soumagne. Digital real-time simulation of transmission lines using parallel processors. Mathematics and Computers in Simulation, 38(4–6): 293–301, August 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549500039Z.

Schneider:1995:URT

[710] H. Schneider, M. Metz, H. Piquet, H. Foch, J. Courault, and O. Lapierre. Utilization of real-time models for the survey and maintenance of high power electrotechnical systems. Mathematics and Computers in Simulation, 38 (4–6):303–310, August 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000405.

BorgesdaSilva:1995:SNN

[711] Luiz Eduardo Borges da Silva, Germano Lambert Torres, Ernesto Castillo Saturno, Alexandre P. Alves da Silva, and Xuan Dai Do. Simulation of a neural net controller for motor drives. Mathematics and Computers in Simulation, 38(4–6):311– 322, August 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549500041U.

Guimaraes:1995:SNC

[712] Carlos Guimarães, Guy Olivier, and Georges-Émile April. Simulation of a non conventional high current low voltage power converter with EMTP. Mathematics and Computers in Simulation, 38(4–6):323– 333, August 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549500042V.

Sybille:1995:ESR

[713] G. Sybille, P. Giroux, and L. Gérin-Lajoie. Effect of subsynchronous resonances on Hydro–Québec static var compensators. Mathematics and Computers in Simulation, 38(4–6):335– 344, August 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549500043W.

BenAmor:1995:ANT

[714] L. Ben Amor, L. A. Dessaint, and O. Akhrif. Adaptive nonlinear torque control of a switched reluctance motor via flux observation. Mathematics and Computers in Simulation, 38 (4–6):345–358, August 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549400044K.

Mpanda-Mabwe:1995:ESM

[715] B. Mpanda-Mabwe, K. Beya, M. Crappe, and M. Delhaye. Estimation of synchronous machine parameters from multisine stand-still frequency response test data. Mathematics and Computers in Simulation, 38(4–6):359– 367, August 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549400045L.

Laporte:1995:CMI

[716] B. Laporte, G. Vinsard, and J. C. Mercier. A computation method for induction motors in steadystate. Mathematics and Computers in Simulation, 38(4–6):369– 376, August 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549500046Z.

Sana:1995:TDT

[717] Abdou-Rahmani Sana, Jean Mahseredjian, Xuan Dai-Do, and Hermann Dommel. Treatment of discontinuities in time-domain simulation of switched networks. Mathematics and Computers in Simulation, 38(4– 6):377–387, August 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000472.

Robyns:1995:PAS

[718] B. Robyns, F. Labrique, and H. Buyse. Performance analysis of a simplified indirect field oriented control algorithm for small power induction actuators. Mathematics and Computers in Simulation, 38(4–6):389– 398, August 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000483.

Tou:1995:ADC

[719] M. Tou, R. Chaffaï, K. Al-Haddad, and V. Rajagopalan. Analysis and design consideration of a unity power factor quasi-resonant rectifier. Mathematics and Computers in Simulation, 38 (4–6):399–413, August 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549400049P.

Lott:1995:NAC

[720] C. Lott, J. H. Xu, S. Saadate, and B. Davat. A new approach to control by model of a voltage source GTO active power filter. Mathematics and Computers in Simulation, 38 (4–6):415–422, August 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000508.

Tahan:1995:GMS

[721] S.-A. Tahan, I. Kamwa, P. Viarouge, and M. Ferfra. A generalized model of satureted synchronous machines. Mathematics and Computers in Simulation, 38(4–6):423– 435, August 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549500051X.

Garrido:1995:RDT

[722] M. Garrido, N. Janssens, J. Ph. Conard, E. Dejaeger, and E. Matagne. Rigorous definition of the transformation ratio in the presence of magnetic saturation and leakages. Mathematics and Computers in Simulation, 38 (4–6):437–443, August 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549500052Y.

Grenier:1995:MCN

[723] Damien Grenier and Jean-Paul Louis. Modeling for control of non-sinewave permanent-magnet synchronous drives by extending Park’s transformation. Mathematics and Computers in Simulation, 38(4–6):445– 452, August 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549500053Z.

Bhathena:1995:SAN

[724] Firdaus Bhathena, George C. Verghese, and Michel Poloujadoff. Stability analysis of a non-Park-transformable electrical machine model. Mathematics and Computers in Simulation, 38 (4–6):453–463, August 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000542.

Anonymous:1995:NIb

[725] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 38(4–6):465–474, August 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475495900152.

Anonymous:1995:AIVa

[726] Anonymous. Author index to volume 38 (1995). Mathematics and Computers in Simulation, 38(4–6): 475–478, August 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495900179.

Anonymous:1995:PA

[727] Anonymous. Pages 255–478 (August 1995). Mathematics and Computers in Simulation, 38(4–6):??, August 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1995:EBb

[728] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 39(1–2):ii, November 8, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495900314.

Bai:1995:EME

[729] Jun Bai, Anthony J. Jakeman, and Michael McAleer. Empirical models for evaluating errors in fitting extremes of a probability distribution. Mathematics and Computers in Simulation, 39 (1–2):1–7, November 8, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495001046.

Janicki:1995:CSA

[730] Aleksander Janicki and Aleksander Weron. Computer simulation of attractors in stochastic models with α-stable noise. Mathematics and Computers in Simulation, 39(1–2): 9–19, November 8, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549500132H.

FalcoKorn:1995:ELV

[731] C. FalcóKorn and C. P. Ullrich. Extending LINPACK by verification routines for linear systems. Mathematics and Computers in Simulation, 39(1–2): 21–37, November 8, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494001335.

Brummer:1995:CSO

[732] Johannes Brummer. A constructive solution for optimal local control of runoff events. Mathematics and Computers in Simulation, 39(1– 2):39–52, November 8, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494001346.

Shary:1995:SLI

[733] Sergey P. Shary. Solving the linear interval tolerance problem. Mathematics and Computers in Simulation, 39(1–2): 53–85, November 8, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549500135K.

Miller:1995:RGK

[734] B. M. Miller and E. Ya. Rubinovich. Regularization of a generalized Kalman filter. Mathematics and Computers in Simulation, 39(1–2):87– 108, November 8, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549500024R.

Han:1995:RNC

[735] B. Han, H. Y. Yin, and J. Q. Liu. Regularizing-numerical-continuation method for solving the resistivities of the earth. Mathematics and Computers in Simulation, 39(1–2):109– 114, November 8, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549500025S.

Tambe:1995:DBL

[736] S. S. Tambe, S. R. Inamdar, and B. D. Kulkarni. Diffusive broadening of limit cycle in presence of noise: a case study of reversible Brusselator. Mathematics and Computers in Simulation, 39(1–2): 115–124, November 8, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549500026T.

Vanualailai:1995:CAT

[737] Jito Vanualailai, Shin ichi Nakagiri, and Jun-Hong Ha. Collision avoidance in a two-point system via Liapunov’s second method. Mathematics and Computers in Simulation, 39(1–2): 125–141, November 8, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549500027U.

Mukundan:1995:QSP

[738] R. Mukundan and K. R. Ramakrishnan. A quaternion solution to the pose determination problem for rendezvous and docking simulations. Mathematics and Computers in Simulation, 39(1–2):

143–153, November 8, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549500028V.

Kaipio:1995:SHE

[739] J. P. Kaipio, J. Tervo, and M. Vauhkonen. Simulations of the heterogeneity of environments by finite element methods. Mathematics and Computers in Simulation, 39(1–2):155– 172, November 8, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495952124.

Qiao:1995:ETP

[740] Hongzhu Qiao and Chris P. Tsokos. Estimation of the three parameter Weibull probability distribution. Mathematics and Computers in Simulation, 39(1–2):173–185, November 8, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495952135.

Anonymous:1995:NIc

[741] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 39(1–2):187–193, November 8, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475495900330.

Anonymous:1995:PNa

[742] Anonymous. Pages 1–193 (8 November 1995). Mathematics and Computers in Simulation, 39(1–2):??, November 8, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

McAleer:1995:SPM

[743] M. McAleer, A. J. Jakeman, and B. Henderson-Sellers. Selected papers of the MSSA/IMACS 10th Biennial Conference on Modelling and Simulation. Mathematics and Computers in Simulation, 39(3–4):195, November 24, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475495900276.

Vichnevetsky:1995:MMS

[744] Robert Vichnevetsky. Mathematical modelling: the soft versus the hard sciences. Mathematics and Computers in Simulation, 39(3–4):197– 206, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495000590.

Nakanishi:1995:DRT

[745] Toshio Nakanishi, Kunio Yikai, Jun ichi Satoh, Isao Miyoshi, Akira Satoh, and Michiya Takahashi. The development of a road traffic simulation system in broad areas. Mathematics and Computers in Simulation, 39(3–4):207– 212, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495000607.

Pleasants:1995:IOG

[746] A. B. Pleasants, S. C. Hendy, G. C. Wake, and D. G. McCall. Investigation of optimal grazing strategies for wool production when model parameters are uncertain. Mathematics and Computers in Simulation, 39(3–4):213– 218, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000619.

Ludwig:1995:SRD

[747] J. A. Ludwig and S. G. Marsden. A simulation of resource dynamics within degraded semi-arid landscapes. Mathematics and Computers in Simulation, 39(3–4):219–224, November 24, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000620.

McCown:1995:AAP

[748] R. L. McCown, G. L. Hammer, J. N. G. Hargreaves, D. Holzworth, and N. I. Huth. APSIM: an agricultural production system simulation model for operational research. Mathematics and Computers in Simulation, 39(3–4):225– 231, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000632.

Luckman:1995:IEB

[749] P. G. Luckman, N. A. Trustrum, L. J. Brown, and J. R. Dymond. Integrated economic-biophysical modelling to support land use decisionmaking in eroding New Zealand hill lands. Mathematics and Computers in Simulation, 39(3–4):233–238, November 24, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475495000644.

Lobanov:1995:ILQ

[750] Yu. Yu. Lobanov. Investigation of the large quantum systems by the numerical integration in functional spaces. Mathematics and Computers in Simulation, 39(3–4):239–244, November 24, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475494000650.

Franses:1995:MAF

[751] Philip Hans Franses and Richard Paap. Moving average filters and periodic integration. Mathematics and Computers in Simulation, 39(3–4):245– 249, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495000668.

Bewley:1995:TCE

[752] Ronald Bewley and Minxian Yang. Testing for cointegration: the effects of mis-specifying the lag length. Mathematics and Computers in Simulation, 39(3–4):251–255, November 24, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549500067X.

Juselius:1995:PUC

[753] Katarina Juselius. Predictable and unpredictable components of the long-run growth in nominal prices. Mathematics and Computers in Simulation, 39(3–4): 257–263, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000681.

Morimune:1995:ERC

[754] Kimio Morimune and Akihisa Mantani. Estimating the rank of co-integration when the order of a vector autoregression is unknown. Mathematics and Computers in Simulation, 39(3–4):265– 271, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000698.

Hisamatsu:1995:EDO

[755] H. Hisamatsu, J. L. Knight, and K. Maekawa. The exact distribution of the OLS and GLS estimators in regression with an integrated regressor and correlated errors — comparison of numerical and Monte Carlo integration. Mathematics and Computers in Simulation, 39(3–4):273– 277, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000704.

Korosi:1995:ALR

[756] Gábor Kőrösi, László Lovrics, and László Mátyás. Aggregation and the long run properties of economic time series. Mathematics and Computers in Simulation, 39(3–4):279– 286, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494000716.

Vishwakarma:1995:NNF

[757] Keshav P. Vishwakarma. A neural network to forecast business cycle indicators. Mathematics and Computers in Simulation, 39(3–4):287– 291, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494000728.

Lim:1995:CEC

[758] Lee Kian Lim. Cointegration and an error correction model of money demand for Australia. Mathematics and Computers in Simulation, 39(3–4):293– 297, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549400073X.

Nawata:1995:ESS

[759] Kazumitsu Nawata. Estimation of sample-selection models by the maximum likelihood method. Mathematics and Computers in Simulation, 39(3–4): 299–303, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475494000741.

Bartels:1995:ODE

[760] Robert Bartels and Denzil G. Fiebig. Optimal design in end-use metering experiments. Mathematics and Computers in Simulation, 39(3–4):305– 309, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000753.

George:1995:ABA

[761] Peter J. George, Ernest H. Oksanen, and Michael R. Veall. Analytic and bootstrap approaches to testing a market saturation hypothesis. Mathematics and Computers in Simulation, 39(3–4): 311–315, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000765.

Oya:1995:SCW

[762] Kosuke Oya. Size corrections for the Wald statistic in structural equation models. Mathematics and Computers in Simulation, 39(3–4):317– 322, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000777.

Jarque:1995:TMN

[763] C. M. Jarque and C. R. McKenzie. Testing for multivariate normality in simultaneous equations models. Mathematics and Computers in Simulation, 39(3–4):323–328, November 24, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475494000789.

McAleer:1995:DME

[764] Michael McAleer and Michael R. Veall. Data mining and the con in econometrics: the U.S. demand for money revisited. Mathematics and Computers in Simulation, 39(3–4):329– 333, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000796.

Aiyoshi:1995:LSB

[765] Eitaro Aiyoshi and Atsushi Maki. Lifecycle simulation based on the feedback system and the choice of switching time for the period model. Mathematics and Computers in Simulation, 39(3–4): 335–341, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000807.

Smith:1995:PAE

[766] Jeremy Smith and Michael McAleer. The performance of alternative estimators in models with generated regressors when the expectations equation has reduced explanatory power. Mathematics and Computers in Simulation, 39(3–4):343–346, November 24, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000814.

Marinova:1995:SRD

[767] Dora Marinova. Spending on research and development and economic growth: a cointegration approach. Mathematics and Computers in Simulation, 39(3–4): 347–352, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000826.

Hu:1995:NCG

[768] Baiding Hu. A note on calculating the Gini index. Mathematics and Computers in Simulation, 39(3–4):353– 358, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000838.

Takase:1995:EEM

[769] Mitsuo Takase. Econometric estimation for a model with dynamic housing tenure choice. Mathematics and Computers in Simulation, 39(3–4):359– 365, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000844.

Ong:1995:TDM

[770] Christine Ong. Tourism demand models: a critique. Mathematics and Computers in Simulation, 39(3–4):367– 372, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000851.

Brindley:1995:EGM

[771] Damian Brindley. Endogenous growth models: a critique of empirical testability. Mathematics and Computers in Simulation, 39(3–4):373–377, November 24, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000863.

Oxley:1995:ESA

[772] L. T. Oxley. An expert systems approach to econometric modelling. Mathematics and Computers in Simulation, 39(3–4):379–383, November 24, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549400087X.

Fukushima:1995:CEE

[773] Takashi Fukushima and Yukiko Shigeno. Congestion externality and economic efficiency in a model of urban economy. Mathematics and Computers in Simulation, 39(3–4):385–391, November 24, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475494000881.

Foster:1995:ISO

[774] John Foster. The impact of the self organisation approach on economic science: why economic theory and history need no longer be mutually exclusive domains. Mathematics and Computers in Simulation, 39(3–4):393– 398, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000893.

Hayashi:1995:NDC

[775] Toshihiko Hayashi. Network dynamics: competition and welfare. Mathematics and Computers in Simulation, 39(3–4): 399–402, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549400090X.

McDonald:1995:UCC

[776] A. David McDonald and A. Stan Hurn. Unobservable cyclical components in term premia of fixed-term financial instruments. Mathematics and Computers in Simulation, 39(3–4):403– 409, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000911.

Masih:1995:MDI

[777] Rumi Masih. Modelling the dynamic interactions among crime, deterrence and socio-economic variables: evidence from a vector error-correction model. Mathematics and Computers in Simulation, 39(3–4):411–416, November 24, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475494000923.

Mitchell:1995:MII

[778] Jason Mitchell. Modelling implications of the instability of the meanvariance paradigm in international finance. Mathematics and Computers in Simulation, 39(3–4):417–423, November 24, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475494000935.

Chan:1995:OFT

[779] Wai sum Chan. Outliers and financial time series modelling: a cautionary note. Mathematics and Computers in Simulation, 39(3–4):425– 430, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000947.

Tse:1995:IRM

[780] Y. K. Tse. Interest rate models and option pricing: a sensitivity analysis. Mathematics and Computers in Simulation, 39(3–4):431– 436, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000959.

Coram:1995:MPP

[781] Bruce Talbot Coram. Markets for protection and predatory behaviour in a Nozickean state of nature. Mathematics and Computers in Simulation, 39(3–4): 437–440, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494000960.

El-Shaarawi:1995:TDE

[782] A. H. El-Shaarawi. Trend detection and estimation with environmental applications. Mathematics and Computers in Simulation, 39(3–4):441– 447, November 24, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000978.

Anonymous:1995:NId

[783] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 39(3–4):449–455, November 24, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495900292.

Anonymous:1995:PNb

[784] Anonymous. Pages 195–455 (24 November 1995). Mathematics and Computers in Simulation, 39(3–4):??, November 24, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Troch:1995:P

[785] I. Troch and F. Breitenecker. Preface. Mathematics and Computers in Simulation, 39(5–6):457–458, November 30, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495900578.

Murray-Smith:1995:EED

[786] D. J. Murray-Smith. Enhanced environments for the development and validation of dynamic system models. Mathematics and Computers in Simulation, 39(5–6):459–464, November 30, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475494001059.

Borutzky:1995:ABG

[787] W. Borutzky, G. Dauphin-Tanguy, and J. U. Thoma. Advances in bond graph modelling: theory, software, applications. Mathematics and Computers in Simulation, 39(5–6):465– 475, November 30, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001066.

Javor:1995:PNA

[788] András Jávor. Petri nets and AI in modeling and simulation. Mathematics and Computers in Simulation, 39(5–6): 477–484, November 30, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001078.

Kluwe:1995:QMB

[789] M. Kluwe, V. Krebs, J. Lunze, and H. Richter. Qualitative modelling based on rules, Petri nets, and differential equations. Mathematics and Computers in Simulation, 39(5–6):485– 489, November 30, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549500108X.

Maschke:1995:PDS

[790] B. M. Maschke and M. Villarroya. Properties of descriptor systems arising from bond graph models. Mathematics and Computers in Simulation, 39(5–6): 491–497, November 30, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001091.

Franke:1995:LSS

[791] Dieter Franke. A linear state space approach to a class of discrete-event systems. Mathematics and Computers in Simulation, 39(5–6):499– 503, November 30, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494001102.

Juslin:1995:EMM

[792] K. Juslin. Experience on mechanistic modelling of industrial process dynamics with APROS. Mathematics and Computers in Simulation, 39(5–6): 505–511, November 30, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549500111X.

Mattsson:1995:SOO

[793] Sven Erik Mattsson. Simulation of object-oriented continuous time models. Mathematics and Computers in Simulation, 39(5–6):513–518, November 30, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475494001126.

Wollhaf:1995:OOM

[794] Konrad Wöllhaf and Sebastian Engell. Object-oriented modelling and simulation of batch plants. Mathematics and Computers in Simulation, 39(5–6): 519–525, November 30, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495001133.

Breitner:1995:IDE

[795] M. H. Breitner, B. Koslik, O. von Stryk, and H. J. Pesch. Iterative design of economic models via simulation, optimization and modeling. Mathematics and Computers in Simulation, 39(5–6): 527–532, November 30, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495001145.

Junker:1995:SMM

[796] Frank Junker. A systematics of modelling mechatronic systems. Mathematics and Computers in Simulation, 39(5–6):533–539, November 30, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475495001157.

Gissinger:1995:MMV

[797] G. L. Gissinger, Y. Chamaillard, and T. Stemmelen. Modelling a motor vehicle and its braking system. Mathematics and Computers in Simulation, 39(5–6):541–548, November 30, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495001169.

Kecskemethy:1995:OOP

[798] Andrés Kecskeméthy and Manfred Hiller. Object-oriented programming techniques in vehicle dynamics simulation. Mathematics and Computers in Simulation, 39(5–6):549–558, November 30, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495001170.

Houbak:1995:SES

[799] Niels Houbak. Simulation of energy systems — guidelines and pitfalls. Mathematics and Computers in Simulation, 39(5–6):559–564, November 30, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495001182.

Schlacher:1995:MMC

[800] Kurt Schlacher and Andreas Kugi. Mathematical modeling and computational principles for the analysis and simulation of long-distance energy systems. Mathematics and Computers in Simulation, 39(5–6):565– 572, November 30, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475494001199.

Gunther:1995:DIE

[801] Michael Günther and Uwe Feldmann. The DAE-index in electric circuit simulation. Mathematics and Computers in Simulation, 39(5–6):573– 582, November 30, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001200.

King:1995:BPT

[802] Rudibert King. Are biological processes too complex to model? Mathematics and Computers in Simulation, 39(5–6): 583–588, November 30, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001212.

Rattay:1995:MEP

[803] Frank Rattay. Modelling of the excitation and the propagation of nerve impulses by natural and artificial stimulations. Mathematics and Computers in Simulation, 39(5–6):589– 595, November 30, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001224.

Miklavcic:1995:MMT

[804] Damijan Miklavčič, Tomaž Jarm, Rihard Karba, and Gregor Serša. Mathematical modelling of tumor growth in mice following electrotherapy and bleomycin treatment. Mathematics and Computers in Simulation, 39(5–6):597– 602, November 30, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001236.

Gontar:1995:CID

[805] V. Gontar. Calculus of iterations and dynamics of physicochemical reactions. Mathematics and Computers in Simulation, 39(5–6):603–608, November 30, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495001248.

Koehne:1995:MSF

[806] M. Koehne, K. Hoen, and M. Schuhen. Modelling and simulation of final clarifiers in wastewater treatment plants. Mathematics and Computers in Simulation, 39(5–6):609–616, November 30, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549500125X.

Isaacs:1995:MND

[807] S. H. Isaacs, H. Zhao, H. Søeberg, and M. Kümmel. Modeling the nitrogen dynamics in an alternating activated sludge process. Mathematics and Computers in Simulation, 39(5–6): 617–626, November 30, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001261.

Engelhardt:1995:MSS

[808] Christof Engelhardt, Dieter Prochnow, and Heinz Bungartz. Modeling and simulation of sedimentation processes in a lowland river. Mathematics and Computers in Simulation, 39(5–6):627– 633, November 30, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495001273.

Zerhouni:1995:TAM

[809] N. Zerhouni, M. Ferney, and A. El Moudni. Transient analysis of manufacturing systems using continuous Petri nets. Mathematics and Computers in Simulation, 39(5–6):635– 639, November 30, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495001285.

Anonymous:1995:NIe

[810] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 39(5–6):641–647, November 30, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475495900594.

Anonymous:1995:AIVb

[811] Anonymous. Author index to volume 39 (1995). Mathematics and Computers in Simulation, 39(5–6):649– 653, November 30, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495900616.

Anonymous:1995:PNc

[812] Anonymous. Pages 457–653 (30 November 1995). Mathematics and Computers in Simulation, 39(5–6):??, November 30, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1995:EBc

[813] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 40(1–2):ii, December 5, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495900470.

Witten:1995:P

[814] Matthew Witten. Preface. Mathematics and Computers in Simulation, 40 (1–2):1–3, December 5, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495900497.

Liebman:1995:ANN

[815] Michael N. Liebman. Application of neural networks to the analysis of structure and function in biologically active macromolecules. Mathematics and Computers in Simulation, 40(1– 2):5–22, December 5, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000152.

Wu:1995:NNM

[816] Cathy Wu, Sailaja Shivakumar, Hsu-Ping Lin, Srinivas Veldurti, and Yugesh Bhatikar. Neural networks for molecular sequence classification. Mathematics and Computers in Simulation, 40(1–2):23–33, December 5, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475495000164.

Thomas:1995:CMS

[817] Elizabeth Thomas and Robert E. Wyatt. A computational model of spindle oscillations. Mathematics and Computers in Simulation, 40(1–2): 35–69, December 5, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495000176.

Softky:1995:MPS

[818] William Softky. McCulloch-Pitts strikes back: a biophysical interpretation of cortical neurons as submillisecond binary devices. Mathematics and Computers in Simulation, 40(1–2):71–79, December 5, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/0378475495000188.

Travis:1995:CMO

[819] Bryan J. Travis. A computational model of one pathway in the cat subcortical auditory system. Mathematics and Computers in Simulation, 40(1–2): 81–99, December 5, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549500019X.

Mason:1995:EIS

[820] Patrick A. Mason. Establishing an imaging system to support a microscopy laboratory. Mathematics and Computers in Simulation, 40(1–2):101– 106, December 5, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000206.

Adomian:1995:SMM

[821] George Adomian. Solving the mathematical models of neurosciences and medicine. Mathematics and Computers in Simulation, 40(1–2):107– 114, December 5, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000218.

Gonzalez-Lima:1995:ANN

[822] F. Gonzalez-Lima and A. R. McIntosh. Analysis of neural network interactions related to associative learning using structural equation modeling. Mathematics and Computers in Simulation, 40(1–2):115–140, December 5, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549500022X.

Mikiten:1995:IBC

[823] Terry M. Mikiten. Intuition-based computing: a new kind of ‘virtual reality’. Mathematics and Computers in Simulation, 40(1–2):141–147, December 5, 1995. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000231.

Anonymous:1995:NIf

[824] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 40(1–2):149–154, December 5, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/0378475495900519.

Anonymous:1995:PD

[825] Anonymous. Pages 1–154 (5 December 1995). Mathematics and Computers in Simulation, 40(1–2):??, December 5, 1995. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Mosekilde:1996:F

[826] Erik Mosekilde and Peter Leth Christiansen. Foreword. Mathematics and Computers in Simulation, 40(3– 4):155–158, April 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475496900075.

Robnik:1996:ELS

[827] Marko Robnik. The energy level statistics of Hamiltonian systems between integrability and chaos: the semiclassical limit. Mathematics and Computers in Simulation, 40(3–4): 159–179, April 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475495000305.

Zabusky:1996:VMC

[828] N. J. Zabusky, D. Silver, and V. Fernandez. Visiometrics and modeling in computational fluid dynamics. Mathematics and Computers in Simulation, 40(3–4):181–191, April 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000313.

Jensen:1996:SMT

[829] Mogens H. Jensen. Simulating models of turbulence and interfaces. Mathematics and Computers in Simulation, 40(3–4):193–206, April 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000321.

Rasmussen:1996:DVT

[830] J. Juul Rasmussen, J. S. Hesthaven, J. P. Lynov, A. H. Nielsen, and M. R. Schmidt. Dipolar vortices in two-dimensional flows. Mathematics and Computers in Simulation, 40 (3–4):207–221, April 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549500033X.

Karpman:1996:MIP

[831] V. I. Karpman, J. P. Lynov, P. K. Michelsen, and J. Juul Rasmussen. Modulational instability of plasma waves in two dimensions. Mathematics and Computers in Simulation, 40 (3–4):223–234, April 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000348.

Feudel:1996:BPM

[832] Fred Feudel, Norbert Seehafer, and Olaf Schmidtmann. Bifurcation phenomena of the magnetofluid equations. Mathematics and Computers in Simulation, 40(3–4):235– 245, April 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475495000356.

Rasmussen:1996:BTC

[833] J. Rasmussen, E. Mosekilde, and C. Reick. Bifurcations in two coupled Rössler systems. Mathematics and Computers in Simulation, 40 (3–4):247–270, April 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495000364.

Lazarides:1996:GAT

[834] N. Lazarides and M. P. Sørensen. Gap anisotropy and tunneling currents. Mathematics and Computers in Simulation, 40(3–4):271– 280, April 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475495000372.

Reick:1996:LRF

[835] Christian H. Reick. Linear response functions of chaotic systems and equilibrium moments. Mathematics and Computers in Simulation, 40(3–4): 281–295, April 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475495000380.

Cruzeiro-Hansson:1996:CPP

[836] Leonor Cruzeiro-Hansson and V. M. Kenkre. Comment on the phase problem in the semiclassical Langevin simulations of the Davydov system. Mathematics and Computers in Simulation, 40(3–4):297–304, April 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000399.

Peyrard:1996:DMP

[837] Michel Peyrard and Thierry Dauxois. DNA melting: a phase transition in one dimension. Mathematics and Computers in Simulation, 40 (3–4):305–318, April 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000402.

Gilhoj:1996:EVS

[838] Henriette Gilhøj, Mohamed Laradji, Bernd Dammann, Claus Jeppesen, Ole G. Mouritsen, Søren Toxværd, and Martin J. Zuckermann. Effect of vacancies and surfactants on the dynamics of ordering processes in multi-component systems. Mathematics and Computers in Simulation, 40 (3–4):319–337, April 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000410.

Rasmussen:1996:NSM

[839] K. Ø. Rasmussen, Yu. B. Gaididei, O. Bang, and P. L. Christiansen. Nonlinear and stochastic modelling of energy transfer in Scheibe aggregates. Mathematics and Computers in Simulation, 40(3–4):339– 358, April 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475495000429.

Kessler:1996:CST

[840] John O. Kessler. Consumption, supply and transport: Self-organization without direct communication. Mathematics and Computers in Simulation, 40 (3–4):359–370, April 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495000437.

Mazin:1996:PFB

[841] W. Mazin, K. E. Rasmussen, E. Mosekilde, P. Borckmans, and G. Dewel. Pattern formation in the bistable Gray–Scott model. Mathematics and Computers in Simulation, 40(3–4): 371–396, April 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475495000445.

Gustafson:1996:BDS

[842] Karl Gustafson. Biological dynamical subsystems of hovering flight. Mathematics and Computers in Simulation, 40(3–4):397–410, April 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495000453.

Thomsen:1996:CDD

[843] J. S. Thomsen, Li. Mosekilde, J. Barlach, C. H. Søgaard, and E. Mosekilde. Computerized determination of 3-D connectivity density in human iliac crest bone biopsies. Mathematics and Computers in Simulation, 40 (3–4):411–423, April 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000461.

Tuszynski:1996:CMG

[844] J. A. Tuszyński, D. Sept, and B. Trpisová. The cell’s microtubules: Growth dynamics, ordering of dipoles and modes of energy propagation. Mathematics and Computers in Simulation, 40(3–4): 425–442, April 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549500047X.

Eilbeck:1996:SAC

[845] J. C. Eilbeck and V. Z. Enolśkii. Some applications of computer algebra to problems in theoretical physics. Mathematics and Computers in Simulation, 40(3–4):443–452, April 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000488.

Hameroff:1996:ORQ

[846] Stuart Hameroff and Roger Penrose. Orchestrated reduction of quantum coherence in brain microtubules: a model for consciousness. Mathematics and Computers in Simulation, 40 (3–4):453–480, April 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475496804769.

Scott:1996:HEC

[847] Alwyn Scott. The hierarchical emergence of consciousness. Mathematics and Computers in Simulation, 40 (3–4):481–489, April 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549500050X.

Anonymous:1996:NIa

[848] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 40(3–4):491–504, April 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475496900099.

Anonymous:1996:PA

[849] Anonymous. Pages 155–504 (April 1996). Mathematics and Computers in Simulation, 40(3–4):??, April 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Tzafestas:1996:E

[850] Spyros G. Tzafestas. Editorial. Mathematics and Computers in Simulation, 40(5–6):505–506, May 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475496900154.

Tzafestas:1996:OPB

[851] S. G. Tzafestas, P. J. Dalianis, and G. Anthopoulos. On the overtraining phenomenon of backpropagation neural networks. Mathematics and Computers in Simulation, 40(5–6): 507–521, May 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000038.

Grujic:1996:MQA

[852] Ljubomir T. Grujić and Anthony N. Michel. Modeling and qualitative analysis of continuous-time neural networks under pure structural variations. Mathematics and Computers in Simulation, 40(5–6):523–533, May 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000046.

Lamy:1996:MSA

[853] D. Lamy. Modeling and sensitivity analysis of neural networks. Mathematics and Computers in Simulation, 40(5–6):535–548, May 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000054.

Mitchell:1996:OMU

[854] R. J. Mitchell, J. M. Bishop, and P. R. Minchinton. Optimising memory usage in n-tuple neural networks. Mathematics and Computers in Simulation, 40(5–6):549–563, May 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000062.

Tzafestas:1996:INN

[855] S. G. Tzafestas and G. B. Stamou. An improved neural network for fuzzy reasoning implementation. Mathematics and Computers in Simulation, 40 (5–6):565–576, May 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000070.

Yu:1996:PRB

[856] P. Yu, V. Anastassopoulos, and A. N. Venetsanopoulos. Pattern recognition based on morphological shape analysis and neural networks. Mathematics and Computers in Simulation, 40 (5–6):577–595, May 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000089.

Kontoravdis:1996:TPC

[857] D. Kontoravdis, S. Kollias, and A. Stafylopatis. A two-phase connectionist approach to invariant picture interpretation. Mathematics and Computers in Simulation, 40(5–6): 597–613, May 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000097.

Scordilis:1996:NFS

[858] Michael S. Scordilis. A neuronal formant synthesizer. Mathematics and Computers in Simulation, 40 (5–6):615–622, May 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000100.

Papadourakis:1996:UAN

[859] G. Papadourakis, M. Vourkas, S. Micheloyannis, and B. Jervis. Use of artificial neural networks for clinical diagnosis. Mathematics and Computers in Simulation, 40(5–6): 623–635, May 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475496000110.

Chassiakos:1996:ISS

[860] Anastassios G. Chassiakos and Sami F. Masri. Identification of structural systems by neural networks. Mathematics and Computers in Simulation, 40 (5–6):637–656, May 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000127.

Sbarbaro:1996:DNC

[861] D. Sbarbaro, K. J. Hunt, and P. J. Gawthrop. Designing nonlinear controllers using connectionist networks. Mathematics and Computers in Simulation, 40(5–6):657– 663, May 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000135.

Anonymous:1996:NIb

[862] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 40(5–6):665–676, May 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475496900178.

Anonymous:1996:AIVa

[863] Anonymous. Author index of Volume 40 (1996). Mathematics and Computers in Simulation, 40(5–6): 677–679, May 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475496900191.

Anonymous:1996:EBa

[864] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 41(1–2):ii, June 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475496900270.

Borne:1996:E

[865] P. Borne, S. G. Tzafestas, and T. Fukuda. Editorial. Mathematics and Computers in Simulation, 41 (1–2):1, June 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475496900294.

Kangas:1996:DAS

[866] Jari Kangas and Teuvo Kohonen. Developments and applications of the self-organizing map and related algorithms. Mathematics and Computers in Simulation, 41(1–2): 3–12, June 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475496882231.

Karayiannis:1996:HLS

[867] Nicolaos B. Karayiannis. Hybrid learning schemes for fast training of feedforward neural networks. Mathematics and Computers in Simulation, 41 (1–2):13–28, June 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000550.

Warwick:1996:NNL

[868] Kevin Warwick. Neural networks for linear control: an analysis. Mathematics and Computers in Simulation, 41(1–2):29–37, June 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000569.

Hepworth:1996:ANC

[869] S. J. Hepworth and A. L. Dexter. Adaptive neural control with stable learning. Mathematics and Computers in Simulation, 41(1–2): 39–51, June 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000577.

Rovithakis:1996:DAR

[870] G. A. Rovithakis and M. A. Christodoulou. Direct adaptive regulation using dynamic neural networks: Application to dc motors speed control. Mathematics and Computers in Simulation, 41(1–2):53–62, June 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000585.

Pican:1996:HOA

[871] Nicolas Pican and Frédéric Alexandre. How OWE architectures encode contextual effects in artificial neural networks. Mathematics and Computers in Simulation, 41(1–2): 63–74, June 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475495000593.

Esteban:1996:CAI

[872] Luis Esteban and Doan Hoang. The connection array for image summarisation. Mathematics and Computers in Simulation, 41(1–2):75– 86, June 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475496000602.

Idczak:1996:NOC

[873] Dariusz Idczak. Necessary optimality conditions for a nonlinear continuous n-dimensional Roesser model. Mathematics and Computers in Simulation, 41(1–2):87–94, June 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495000615.

Francois:1996:OCD

[874] B. François. Orthogonal considerations in the design of neural networks for function approximation. Mathematics and Computers in Simulation, 41(1–2):95–108, June 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000623.

Hamad:1996:UPC

[875] D. Hamad, C. Firmin, and J.-G. Postaire. Unsupervised pattern classification by neural networks. Mathematics and Computers in Simulation, 41(1–2):109–116, June 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000631.

Djouani:1996:FON

[876] K. Djouani and Y. Hamam. Feedback optimal neural network controller for dynamic systems — a ship maneuvering example. Mathematics and Computers in Simulation, 41 (1–2):117–127, June 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549500064X.

Miyata:1996:CAM

[877] Hitoshi Miyata, Makoto Ohki, Yasuyuki Yokouchi, and Masaaki Ohkita. Control of the autonomous mobile robot DREAM-1 for a parallel parking. Mathematics and Computers in Simulation, 41(1–2):129– 138, June 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000658.

Zhao:1996:GAE

[878] J. Zhao, R. Gorez, and V. Wertz. Genetic algorithms for the elimination of redundancy and/or rule contribution assessment in fuzzy models. Mathematics and Computers in Simulation, 41(1–2):139–148, June 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000666.

Neusser:1996:PDN

[879] S. Neusser and B. Höfflinger. Parallel digital neural hardware for controller design. Mathematics and Computers in Simulation, 41(1–2): 149–160, June 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000674.

Goser:1996:IAN

[880] Karl F. Goser. Implementation of artificial neural networks into hardware: Concepts and limitations. Mathematics and Computers in Simulation, 41 (1–2):161–171, June 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000682.

Delmotte:1996:GSM

[881] François Delmotte, Laurent Dubois, and Pierre Borne. A general scheme for multi-model controller using trust. Mathematics and Computers in Simulation, 41(1–2):173– 186, June 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000690.

Anonymous:1996:NIc

[882] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 41(1–2):187–197, June 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475496900312.

Anonymous:1996:PJ

[883] Anonymous. Pages 1–197 (June 1996). Mathematics and Computers in Simulation, 41(1–2):??, June 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Borne:1996:P

[884] P. Borne, S. C. Tzafestas, and T. Fukuda. Preface. Mathematics and Computers in Simulation, 41(3–4):199, July 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496900063.

Tzafestas:1996:SVC

[885] S. Tzafestas, A. Zagorianos, and T. Pimenides. A solution to the velocity control of redundant robots. Mathematics and Computers in Simulation, 41(3–4):201–208, July 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000704.

Oustaloup:1996:CPP

[886] A. Oustaloup and H. Linarès. The CRONE path planning. Mathematics and Computers in Simulation, 41 (3–4):209–217, July 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000712.

Tzafestas:1996:FNC

[887] Spyros Tzafestas, Spyros Raptis, and George Stamou. A flexible neurofuzzy cell structure for general fuzzy inference. Mathematics and Computers in Simulation, 41(3–4):219– 233, July 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000720.

Kiguchi:1996:OMU

[888] Kazuo Kiguchi and Toshio Fukuda. Object manipulation under disturbance using fuzzy-neural network. Mathematics and Computers in Simulation, 41(3–4):235– 244, July 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000739.

Meslin:1996:DCM

[889] Jean-Marc Meslin and Philippe Coiffet. Dynamic control of manipulators using qualitative and quantitative modelling. Mathematics and Computers in Simulation, 41(3–4): 245–255, July 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000747.

Fourie:1996:CMR

[890] Cornelius J. Fourie and Leon A. Deist. Control of mobile robots with fuzzy reasoning. Mathematics and Computers in Simulation, 41 (3–4):257–262, July 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000755.

Daxwanger:1996:SBV

[891] Wolfgang Daxwanger and Günther Schmidt. Skill-based vehicle guidance by use of artificial neural networks. Mathematics and Computers in Simulation, 41(3–4):263– 271, July 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000763.

Taktak:1996:AIR

[892] Rached Taktak, Michel Dufaut, Didier Wolf, and René Husson. Analysis and inspection of road traffic using image processing. Mathematics and Computers in Simulation, 41 (3–4):273–283, July 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000771.

Ball:1996:ANN

[893] N. R. Ball. Application of a neural network based classifier system to AGV obstacle avoidance. Mathematics and Computers in Simulation, 41 (3–4):285–296, July 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/037847549500078X.

Ibrahim:1996:CSE

[894] M. Yousef Ibrahim. Computer study on the effect of an articulated robot’s parameters on its dynamic characteristics under different balancing conditions. Mathematics and Computers in Simulation, 41(3–4):297– 306, July 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475495000798.

Triboulet:1996:MUE

[895] Jean Triboulet, Edie Camille N’zi, and Florent Chavand. Methods for updating the environment’s geometric database in telerobotics. Mathematics and Computers in Simulation, 41 (3–4):307–320, July 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495000801.

Gharapetian:1996:RTS

[896] VahéDer Gharapetian and Michèle Rombaut. A real-time simulator of the perceptive system of a mobile vehicle. Mathematics and Computers in Simulation, 41(3–4):321– 328, July 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/037847549500081X.

Eleter:1996:IMP

[897] B. Eleter and M. Rombaut. Intelligent mapping of the ProLab2 vehicle dynamic environment. Mathematics and Computers in Simulation, 41 (3–4):329–336, July 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000828.

Bhandarkar:1996:HNN

[898] Suchendra M. Bhandarkar, Jean Koh, and Minsoo Suk. A hierarchical neural network and its application to image segmentation. Mathematics and Computers in Simulation, 41 (3–4):337–355, July 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000836.

Inoue:1996:COP

[899] Seiichi Inoue, Takeo Ojika, Michiko Harayama, Takahiro Kobayashi, and Takashi Imai. Cooperated operation of plural hand-robots for automatic harvest system. Mathematics and Computers in Simulation, 41 (3–4):357–365, July 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000844.

Tounsi:1996:TGM

[900] M. Tounsi and J. F. Le Corre. Trajectory generation for mobile robots. Mathematics and Computers in Simulation, 41(3–4):367– 376, July 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475495000852.

Khalil:1996:KAS

[901] W. Khalil and D. Murareci. Kinematic analysis and singular configurations of a class of parallel robots. Mathematics and Computers in Simulation, 41 (3–4):377–390, July 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495000860.

Anonymous:1996:NId

[902] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 41(3–4):391–401, July 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475496900087.

Khodabandehloo:1996:ERI

[903] Koorosh Khodabandehloo and Sanja Vraneš. European Robotics and Intelligent Systems Conference. Mathematics and Computers in Simulation, 41 (5–6):403–406, August 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495000879.

Lauridsen:1996:MBC

[904] J. K. Lauridsen, O. Madsen, H. Holm, I. Hafsteinsson, and J. Boelskifte. Model based control of a one degree of freedom workpiece manipulator for welding of nozzles. Mathematics and Computers in Simulation, 41 (5–6):407–417, August 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000887.

Davies:1996:AER

[905] Brian Davies, S. Harris, A. Razban, and J. Efstathiou. Application experience of a robotic cell for automated adhesive dispensing. Mathematics and Computers in Simulation, 41 (5–6):419–427, August 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000895.

Nunes:1996:SBD

[906] Urbano Nunes, Pedro Faia, and A. T. de Almeida. Sensor-based 3-D autonomous surface-following control. Mathematics and Computers in Simulation, 41(5–6):429– 444, August 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000909.

Parkin:1996:MAA

[907] R. M. Parkin and M. R. Jackson. A mechatronic approach for analysing timber surfaces. Mathematics and Computers in Simulation, 41(5–6): 445–450, August 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000917.

Delgado:1996:IOL

[908] A. Delgado, C. Kambhampati, and K. Warwick. Input/output linearization using dynamic recurrent neural networks. Mathematics and Computers in Simulation, 41(5–6):451– 460, August 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475495000925.

Sanders:1996:ANN

[909] D. A. Sanders, B. P. Haynes, G. E. Tewkesbury, and I. J. Stott. The addition of neural networks to the inner feedback path in order to improve on the use of pre-trained feed forward estimators. Mathematics and Computers in Simulation, 41(5– 6):461–472, August 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495000933.

Tzafestas:1996:RTE

[910] Spyros G. Tzafestas and Paraskevas J. Dalianis. A real-time expert data filtering system for industrial plant environments. Mathematics and Computers in Simulation, 41(5–6):473– 484, August 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475495000941.

Carrera:1996:IDF

[911] Eliodoro Carrera and Miguel A. Serna. Inverse dynamics of flexible robots. Mathematics and Computers in Simulation, 41(5–6):485– 508, August 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549500095X.

Kauschke:1996:CFS

[912] Michael Kauschke. Closed form solutions applied to redundant serial link manipulators. Mathematics and Computers in Simulation, 41(5–6): 509–516, August 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000968.

Lefeber:1996:OFS

[913] Dirk Lefeber. An original O(N) formulation for simulation of robotic systems. Mathematics and Computers in Simulation, 41(5–6):517– 524, August 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000976.

Howard:1996:KAW

[914] D. Howard, S. J. Zhang, and D. J. Sanger. Kinematic analysis of a walking machine. Mathematics and Computers in Simulation, 41(5–6): 525–538, August 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000984.

Jen:1996:ROD

[915] C. W. Jen, D. A. Johnson, and R. Gorez. A reduced-order dynamic model for end-effector position control of a flexible robot arm. Mathematics and Computers in Simulation, 41 (5–6):539–558, August 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495000992.

Ficola:1996:SMC

[916] Antonio Ficola and Michele La Cava. A sliding mode controller for a two-joint robot with an elastic link. Mathematics and Computers in Simulation, 41 (5–6):559–569, August 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475496811359.

Wadie:1996:TDM

[917] I. H. C. Wadie, G. L. Purnell, and K. Khodabandehloo. Two-dimensional modelling of pig carcass spines for robotic evisceration. Mathematics and Computers in Simulation, 41(5– 6):571–578, August 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001018.

Griesser:1996:BBF

[918] M. T. Grießer and P. M. Taylor. The bending behaviour of fabrics: an energy approach. Mathematics and Computers in Simulation, 41(5– 6):579–586, August 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001026.

Cugini:1996:DSN

[919] U. Cugini, P. Denti, and C. Rizzi. Design and simulation of non-rigid materials handling systems. Mathematics and Computers in Simulation, 41 (5–6):587–593, August 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001034.

Karakerezis:1996:SMR

[920] A. Karakerezis and K. Khodabandehloo. Simulation modelling for robotics grasping of flexible viscoelastic materials using a discrete element model. Mathematics and Computers in Simulation, 41(5–6):595– 614, August 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495001042.

Jamhour:1996:PST

[921] E. Jamhour and P. J. André. Planning smooth trajectories along parametric paths. Mathematics and Computers in Simulation, 41(5–6):615– 626, August 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495001050.

Yu:1996:LCS

[922] J. S. Yu and P. C. Müller. An online cartesian space obstacle avoidance scheme for robot arms. Mathematics and Computers in Simulation, 41 (5–6):627–637, August 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495001069.

Thomas:1996:AMA

[923] J. P. Thomas and N. Nissanke. An algebra for modelling assembly tasks. Mathematics and Computers in Simulation, 41(5–6):639– 659, August 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475495001077.

Anonymous:1996:NIe

[924] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 41(5–6):661–674, August 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475496900592.

Anonymous:1996:AIVb

[925] Anonymous. Author index of Volume 41 (1996). Mathematics and Computers in Simulation, 41(5–6): 675–678, August 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475496900610.

Anonymous:1996:EBb

[926] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 42(1):ii, September 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475496900439.

Fung:1996:TSD

[927] K. K. Fung and S. Y. R. Hui. Transputer simulation of decoupled electrical circuits. Mathematics and Computers in Simulation, 42(1):1– 13, September 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495000518.

Woolfson:1996:ATL

[928] M. S. Woolfson and S. Y. R. Hui. Application of the transmission line method to the solution of the continuous Kalman filter equations of general order. Mathematics and Computers in Simulation, 42(1):15– 33, September 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495001093.

Markov:1996:LIU

[929] S. Markov, E. Popova, U. Schneider, and J. Schulze. On linear interpolation under interval data. Mathematics and Computers in Simulation, 42 (1):35–45, September 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001107.

Kim:1996:MPS

[930] S.-B. Kim, A. Hadjidimos, E. N. Houstis, and J. R. Rice. Multiparameterized Schwarz alternating methods for elliptic boundary value problems. Mathematics and Computers in Simulation, 42(1):47–76, September 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475495001115.

Razzaghi:1996:ALS

[931] Mohsen Razzaghi and Mohammad Habibi. Application of Legendre series to the control problems governed by linear parabolic equations. Mathematics and Computers in Simulation, 42(1):77–84, September 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495001336.

Morante:1996:MNS

[932] A. Belleni Morante, G. Lauro, and R. Monaco. Modelling and numerical simulation of contamination processes in random media. Mathematics and Computers in Simulation, 42 (1):85–95, September 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475496000043.

Kabakchiev:1996:PAA

[933] Christo Kabakchiev and Vera Behar. Parallel architecture of adaptive MTD processors. Mathematics and Computers in Simulation, 42(1): 97–105, September 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475496000584.

Anonymous:1996:NIf

[934] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 42(1):107–121, September 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475496900452.

Anonymous:1996:PS

[935] Anonymous. Pages 1–121 (September 1996). Mathematics and Computers in Simulation, 42(1):??, September 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Hanus:1996:E

[936] R. J. L. Hanus. Editorial. Mathematics and Computers in Simulation, 42(2–3):123, October 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475496900336.

Walter:1996:IDN

[937] Eric Walter and Luc Pronzato. On the identifiability and distinguishability of nonlinear parametric models. Mathematics and Computers in Simulation, 42(2–3):125– 134, October 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495001239.

Stortelder:1996:PED

[938] Walter J. H. Stortelder. Parameter estimation in dynamic systems. Mathematics and Computers in Simulation, 42 (2–3):135–142, October 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001174.

Hojsgaard:1996:LSD

[939] Søren Højsgaard. Learning structures from data and experts. Mathematics and Computers in Simulation, 42(2– 3):143–152, October 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001220.

Paul:1996:RAL

[940] W. Paul and E. Witte. Robust approaches to N-leaching under uncertainties. Mathematics and Computers in Simulation, 42(2–3):153– 157, October 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495001247.

VanImpe:1996:PLM

[941] Jan F. M. Van Impe. Power and limitations of model based bioprocess optimization. Mathematics and Computers in Simulation, 42(2–3): 159–169, October 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549500128X.

Kinder:1996:SSB

[942] Michael Kinder and Wolfgang Wiechert. Stochastic simulation of biotechnical processes. Mathematics and Computers in Simulation, 42(2–3):171– 178, October 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495001301.

Panikov:1996:MMM

[943] N. S. Panikov. Mechanistic mathematical models of microbial growth in bioreactors and in natural soils: Explanation of complex phenomena. Mathematics and Computers in Simulation, 42(2–3):179–186, October 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495001271.

Garcia-Ochoa:1996:SXG

[944] F. García-Ochoa, V. E. Santos, and A. Alcón. Simulation of xanthan gum production by a chemically structured kinetic model. Mathematics and Computers in Simulation, 42(2– 3):187–195, October 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001131.

daSilva:1996:KMH

[945] A. Lourenço da Silva, A. Marc, J. M. Engasser, and J. L. Goergen. Kinetic model of hybridoma cultures for the identification of rate limiting factors and process optimisation. Mathematics and Computers in Simulation, 42 (2–3):197–205, October 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475496837037.

Viesturs:1996:CMF

[946] U. E. Viesturs and J. J. Vanags. Correlation of mixing and fermentation performance. Mathematics and Computers in Simulation, 42(2–3): 207–211, October 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475495001204.

Farkas:1996:BOM

[947] I. Farkas and Z. Rendik. Block oriented modeling of drying processes. Mathematics and Computers in Simulation, 42(2–3):213– 219, October 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475496000031.

Szoke:1996:NCM

[948] Sz. Szöke and E. Persoons. A new conceptual model to compute the drying front in wheat. Mathematics and Computers in Simulation, 42(2– 3):221–232, October 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475496837049.

DeMoor:1996:BGB

[949] M. De Moor and D. Berckmans. Building a grey box model to model the energy and mass transfer in an imperfectly mixed fluid by using experimental data. Mathematics and Computers in Simulation, 42(2–3): 233–244, October 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001263.

Stirbet:1996:NSV

[950] Alexandrina D. Stirbet and Reto J. Strasser. Numerical simulation of the in vivo fluorescence in plants. Mathematics and Computers in Simulation, 42 (2–3):245–253, October 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549500114X.

Munford:1996:UIL

[951] Alan G. Munford. The use of iterative linear programming in practical applications of animal diet formulation. Mathematics and Computers in Simulation, 42(2–3):255– 261, October 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495001158.

Richter:1996:QAS

[952] O. Richter and R. Seppelt. Quantitative aspects of sustainable agriculture. Mathematics and Computers in Simulation, 42(2–3):263– 269, October 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495001166.

Boucneau:1996:RMS

[953] G. Boucneau, M. Van Meirvenne, J. Desmet, and G. Hofman. Regressive modelling of soil water table fluctuations. Mathematics and Computers in Simulation, 42(2–3): 271–277, October 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001255.

Avila:1996:MTS

[954] Isabel M. L. B. Ávila, Maria da Conceição Manso, and Cristina L. M. Silva. Modelling the thermal sterilisation of foods inside packs with two divisions. Mathematics and Computers in Simulation, 42(2–3):279– 285, October 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/0378475495001190.

Nicolai:1996:PFM

[955] B. M. Nicolaï and J. F. Van Impe. Predictive food microbiology: a probabilistic approach. Mathematics and Computers in Simulation, 42(2–3): 287–292, October 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001298.

Hollinger:1996:MDR

[956] Steven E. Hollinger and Leszek Kuchar. Mapping disease risk using a joint probability distribution. Mathematics and Computers in Simulation, 42 (2–3):293–298, October 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/037847549500131X.

Forsberg:1996:OSH

[957] Odd Inge Forsberg. Optimal stocking and harvesting of size-structured farmed fish: a multi-period linear programming approach. Mathematics and Computers in Simulation, 42(2– 3):299–305, October 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475495001328.

Anonymous:1996:NIg

[958] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 42(2–3):307–318, October 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/037847549690035X.

Jacob:1996:P

[959] Gérard Jacob. Preface. Mathematics and Computers in Simulation, 42(4– 6):319–320, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496900002.

Gerdt:1996:CAS

[960] V. Gerdt. Computer algebra and solving polynomials. Mathematics and Computers in Simulation, 42(4– 6):321, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497844102.

Zharkov:1996:IAI

[961] A. Yu. Zharkov and Yu. A. Blinkov. Involution approach to investigating polynomial systems. Mathematics and Computers in Simulation, 42(4–6): 323–332, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496000067.

Winkler:1996:PSA

[962] Franz Winkler. Parametrized solutions of algebraic equations. Mathematics and Computers in Simulation, 42(4– 6):333–338, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496000079.

Wang:1996:SPE

[963] Dongming Wang. Solving polynomial equations: Characteristic sets and triangular systems. Mathematics and Computers in Simulation, 42(4– 6):339–351, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496000080.

Gonzalez-Lopez:1996:PSS

[964] M. J. González-López, T. Recio, and F. Santos. Parametrization of semialgebraic sets. Mathematics and Computers in Simulation, 42(4–6): 353–362, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496000092.

Duval:1996:SCF

[965] Dominique Duval and Jean-Claude Reynaud. Sketches and computations over fields. Mathematics and Computers in Simulation, 42(4–6): 363–373, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496000109.

Gomez-Diaz:1996:EUD

[966] Teresa Gómez-Díaz. Examples of using dynamic constructible closure. Mathematics and Computers in Simulation, 42(4–6):375–383, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475496800015.

MacCallum:1996:CAD

[967] M. A. H. MacCallum. Computer algebra and differential equations. Mathematics and Computers in Simulation, 42(4–6):385–386, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475496000122.

Bronstein:1996:FLO

[968] Manuel Bronstein. On the factorisation of linear ordinary differential operators. Mathematics and Computers in Simulation, 42(4–6):387– 389, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475496000134.

Wolf:1996:ASN

[969] Thomas Wolf and Andreas Brand. Approaches to solving nonlinear ODEs. Mathematics and Computers in Simulation, 42(4–6):391– 398, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475496000146.

Gerdt:1996:HIC

[970] V. P. Gerdt. Homogeneity of integrability conditions for multi-parametric families of polynomial-non-linear evolution equations. Mathematics and Computers in Simulation, 42(4–6): 399–408, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475496000158.

Jung:1996:AMF

[971] F. Jung, F. Naegele, and J. Thomann. An algorithm of multisummation of formal power series, solutions of linear ODE equations. Mathematics and Computers in Simulation, 42(4– 6):409–425, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847549600016X.

Giusti:1996:EC

[972] M. Giusti. Elimination and complexity. Mathematics and Computers in Simulation, 42(4–6):427, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497844114.

Solerno:1996:CBE

[973] Pablo Solernó. Complexity bounds in elimination theory — a survey. Mathematics and Computers in Simulation, 42(4–6):429–438, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475496000171.

Boulier:1996:OSE

[974] F. Boulier. An optimization of Seidenberg’s elimination algorithm in differential algebra. Mathematics and Computers in Simulation, 42(4–6): 439–448, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496000183.

Berenstein:1996:MRC

[975] C. A. Berenstein and A. Yger. Multidimensional residues and complexity problems. Mathematics and Computers in Simulation, 42(4–6):449– 457, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475496000195.

Sergeraert:1996:GAT

[976] F. Sergeraert. Geometry and algebraic topology. Mathematics and Computers in Simulation, 42(4–6): 459, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475496000201.

Real:1996:ACH

[977] Pedro Real. An algorithm computing homotopy groups. Mathematics and Computers in Simulation, 42(4– 6):461–465, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475496000213.

Garcia:1996:IFP

[978] Julio Jesús Rubio García. Integrating functional programming and symbolic computation. Mathematics and Computers in Simulation, 42(4–6): 467–473, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475496000225. See

erratum [1096].

Roos:1996:CAR

[979] Jan-Erik Roos. On computerassisted research in homological algebra. Mathematics and Computers in Simulation, 42(4–6):475–490, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475496000237.

Trevisan:1996:TPA

[980] V. Trevisan. Theory and practice of algebraic algorithms. Mathematics and Computers in Simulation, 42 (4–6):491, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496000250.

Abell:1996:ASD

[981] Martha L. Abell and James P. Braselton. Advancing the study of differential equations with technology. Mathematics and Computers in Simulation, 42(4–6):493– 508, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475496000262.

Shirayanagi:1996:FPG

[982] Kiyoshi Shirayanagi. Floating point Gröbner bases. Mathematics and Computers in Simulation, 42(4–6): 509–528, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496000274.

Winkler:1996:ACG

[983] Franz Winkler. Algebraic computation in geometry. Mathematics and Computers in Simulation, 42(4–6): 529–537, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496000286.

Recio:1996:RAN

[984] T. Recio. Real algebraic numbers. Mathematics and Computers in Simulation, 42(4–6):539–540, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475496000304.

Ligatsikas:1996:GCR

[985] Zenon Ligatsikas, Renaud Rioboo, and Marie Françoise Roy. Generic computation of the real closure of an ordered field. Mathematics and Computers in Simulation, 42(4–6): 541–549, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475496000316.

Duval:1996:DER

[986] D. Duval and L. González-Vega. Dynamic evaluation and real closure. Mathematics and Computers in Simulation, 42(4–6):551–560, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475496000328.

Becker:1996:RCZ

[987] E. Becker and T. Wörmann. Radical computations of zero-dimensional ideals and real root counting. Mathematics and Computers in Simulation, 42(4– 6):561–569, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847549600033X.

Hong:1996:EMA

[988] Hoon Hong. An efficient method for analyzing the topology of plane real algebraic curves. Mathematics and Computers in Simulation, 42(4– 6):571–582, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496000341.

Ganzha:1996:SDN

[989] V. Ganzha. Symbolic devices and numerical solving of PDEs. Mathematics and Computers in Simulation, 42 (4–6):583, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497844126.

Akritas:1996:VPS

[990] Alkiviadis G. Akritas, Evgenia K. Akritas, and Genadii I. Malaschonok. Various proofs of Sylvester’s (determinant) identity. Mathematics and Computers in Simulation, 42(4–6): 585–593, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496000353.

Bungartz:1996:PCC

[991] H. Bungartz, M. Griebel, D. Röschke, and C. Zenger. A proof of convergence for the combination technique for the Laplace equation using tools of symbolic computation. Mathematics and Computers in Simulation, 42(4– 6):595–605, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496000365.

Ganzha:1996:SNC

[992] V. G. Ganzha and E. V. Vorozhtsov. Symbolic-numerical computation of the stability regions for Jameson’s schemes. Mathematics and Computers in Simulation, 42(4–6):607– 615, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475496000377.

Pfeiffer:1996:SSD

[993] B.-M Pfeiffer and W. Marquardt. Symbolic semi-discretization of partial differential equation systems. Mathematics and Computers in Simulation, 42(4–6):617–628, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475496000389.

Steinberg:1996:SAQ

[994] Stanly Steinberg and Richard Liska. Stability analysis by quantifier elimination. Mathematics and Computers in Simulation, 42(4–6):629– 638, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475496000390.

Jacob:1996:DS

[995] G. Jacob. Dynamical systems. Mathematics and Computers in Simulation, 42(4–6):639, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497844138.

Hespel:1996:IDO

[996] Christiane Hespel. Iterated derivatives of the output of a nonlinear dynamic system and Faà di Bruno formula. Mathematics and Computers in Simulation, 42(4–6):641– 657, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475496000407.

Villard:1996:SDL

[997] D. Villard and B. Arnaldi. Symbolic differentiation library for simulation of multibody rigid systems. Mathematics and Computers in Simulation, 42(4– 6):659–673, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496000419.

Szymkat:1996:ASR

[998] Maciej Szymkat. Algorithms for symbolic reduction of signal flow graphs. Mathematics and Computers in Simulation, 42(4–6):675– 684, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475496000420.

Chindris:1996:VNS

[999] Mircea Chindriş and Tiberiu Coloşi. A variant for numerical simulation of a high-frequency converter with voltage-fed inverter. Mathematics and Computers in Simulation, 42(4–6): 685–692, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496004326.

Guyon:1996:MPP

[1000] C. Guyon. Motion planning problem: Parametrized inputs and drift problem. Mathematics and Computers in Simulation, 42(4–6):693– 706, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475496000444.

Minh:1996:SPE

[1001] Hoang Ngoc Minh. Summations of polylogarithms via evaluation transform. Mathematics and Computers in Simulation, 42(4–6):707–728, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475496000456.

Caprasse:1996:ICA

[1002] H. Caprasse. The impact of computer algebra on the use and understanding of fundamental interactions. Mathematics and Computers in Simulation, 42(4–6):729, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847549784414X.

Daftardar-Gejji:1996:EAV

[1003] Varsha Daftardar-Gejji. Evaluation of Ashtekar variables using STEN-SOR. Mathematics and Computers in Simulation, 42(4–6):731–738, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475496000468.

Fonseca-Neto:1996:ACT

[1004] Joel B. Fonseca-Neto, M. J. Reboucas, and M. A. H. MacCallum. Algebraic computing in torsion theories of gravitation. Mathematics and Computers in Simulation, 42(4–6): 739–748, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847549600047X.

Jacob:1996:GCD

[1005] G. Jacob. Graphs and curves drawing. Mathematics and Computers in Simulation, 42(4–6):749, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497844151.

Fiorot:1996:RCD

[1006] J. C. Fiorot, P. Jeannin, and B. Sucher.

Recursive calculation of the derivatives of rational curves in ( br ) form. Mathematics and Computers in Simulation, 42(4–6):751– 764, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475496000481.

Baudon:1996:CGS

[1007] O. Baudon and J. M. Laborde.

Cabri-graph, a sketchpad for graph theory. Mathematics and Computers in Simulation, 42(4–6):765– 774, November 1996. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475496000493.

Anonymous:1996:NIh

[1008] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 42(4–6):775–786, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475496900026.

Anonymous:1996:AIVc

[1009] Anonymous. Author index of volume 42 (1996). Mathematics and Computers in Simulation, 42(4–6): 787–790, November 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847549690004X.

Anonymous:1997:EBa

[1010] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 43(1):ii, January 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/0378475497900070.

Taha:1997:F

[1011] Thiab R. Taha. Foreword. Mathematics and Computers in Simulation, 43(1):1, January 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847549600050X.

Ablowitz:1997:NSE

[1012] M. J. Ablowitz, B. M. Herbst, and C. M. Schober. The nonlinear Schrödinger equation: Asymmetric perturbations, traveling waves and chaotic structures. Mathematics and Computers in Simulation, 43(1):3– 12, January 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475496000511.

Hereman:1997:SMC

[1013] Willy Hereman and Ameina Nuseir.

Symbolic methods to construct exact solutions of nonlinear partial differential equations. Mathematics and Computers in Simulation, 43(1):13– 27, January 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475496000535.

Lu:1997:SAW

[1014] Xiaowu Lu and Rudolf Schmid.

A symplectic algorithm for wave equations. Mathematics and Computers in Simulation, 43(1):29–38, January 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475496000523.

Mansfield:1997:SES

[1015] Elizabeth L. Mansfield and Peter A.

Clarkson. Symmetries and exact solutions for a 2 + 1-dimensional shallow water wave equation. Mathematics and Computers in Simulation, 43 (1):39–55, January 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475496000547.

Sulem:1997:FNS

[1016] C. Sulem and P. L. Sulem. The focusing nonlinear Schrödinger equation: Effect of the coupling to a low frequency field. Mathematics and Computers in Simulation, 43 (1):57–66, January 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475496000559.

Twizell:1997:FDM

[1017] E. H. Twizell, A. G. Bratsos, and J. C. Newby. A finitedifference method for solving the cubic Schrödinger equation. Mathematics and Computers in Simulation, 43 (1):67–75, January 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475496000560.

Weideman:1997:FDM

[1018] J. A. C. Weideman and B. M. Herbst. Finite difference methods for an AKNS eigenproblem. Mathematics and Computers in Simulation, 43 (1):77–88, January 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475496000572.

Anonymous:1997:NIa

[1019] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 43(1):89–100, January 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/0378475497900094.

Anonymous:1997:PJ

[1020] Anonymous. Pages 1–100 (January 1997). Mathematics and Computers in Simulation, 43(1):??, January 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Bogaerts:1997:ASN

[1021] Ph. Bogaerts, J. Castillo, and R. Hanus. Analytical solution of the non-uniform heat exchange in a reactor cooling coil with constant fluid flow. Mathematics and Computers in Simulation, 43(2):101– 113, February 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475496000596.

Shibata:1997:PNB

[1022] H. Shibata, T. Ushio, S. Kumagai, and S. Kodama. Petri net based qualitative simulation with quantitative information. Mathematics and Computers in Simulation, 43(2):115– 122, February 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497816357.

He:1997:SDM

[1023] Qiming He. Semi-discrete method for generalized Schrödinger-type equations. Mathematics and Computers in Simulation, 43(2):123–138, February 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475496000614.

Tchamova:1997:ERT

[1024] A. Tchamova. Evidence reasoning theory with application to the identity estimation and data association systems. Mathematics and Computers in Simulation, 43(2):139– 142, February 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475496000626.

Kuznetsov:1997:PIS

[1025] N. Kuznetsov and P. Kloeden. The problem of information stability in computer studies of continuous systems. Mathematics and Computers in Simulation, 43(2):143–158, February 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475496000638.

Devenish:1997:SSN

[1026] D. G. Devenish and B. J. Stone. Substructuring non-linear dynamic systems in the time domain. Mathematics and Computers in Simulation, 43 (2):159–170, February 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847549600064X.

Birta:1997:SPE

[1027] Louis G. Birta and Liming Yang. Some P(EC)^{m}E methods for the parallel solution of ODEs. Mathematics and Computers in Simulation, 43 (2):171–182, February 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496000651.

Lamotte:1997:SIA

[1028] J. L. Lamotte and F. Delay. On the stability of the 2D interpolation algorithms with uncertain data. Mathematics and Computers in Simulation, 43 (2):183–201, February 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496000663.

Kabakchiev:1997:ADP

[1029] Chr. Kabakchiev, B. Vassileva, and I. Miteva. Adaptive Doppler processing technique. Mathematics and Computers in Simulation, 43(2):203– 208, February 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475496000675.

Jacobson:1997:EIT

[1030] Sheldon H. Jacobson. The effect of initial transient on the steadystate simulation harmonic analysis gradient estimators. Mathematics and Computers in Simulation, 43(2): 209–221, February 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475496000687.

Albeanu:1997:MCA

[1031] Grigore Albeanu. A Monte Carlo approach for controlled search. Mathematics and Computers in Simulation, 43(2):223–228, February 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475496000699.

Anonymous:1997:NIb

[1032] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 43(2):229–240, February 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475497816369.

Anonymous:1997:PF

[1033] Anonymous. Pages 101–240 (February 1997). Mathematics and Computers in Simulation, 43(2):??, February 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Binning:1997:SPM

[1034] P. Binning, H. Bridgman, M. McAleer, and B. Williams. Selected papers of the MSSA/IMACS 11th Biennial Conference on Modelling and Simulation, November 1995. Mathematics and Computers in Simulation, 43 (3–6):241, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000463.

Clift:1997:DSM

[1035] A. Clift, G. Herron, and M. A. Terras. DEMANIR, a simulation model of insecticide resistance development and management. Mathematics and Computers in Simulation, 43 (3–6):243–250, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000062.

Ye:1997:ANV

[1036] W. Ye, D. P. Hansen, A. J. Jakeman, P. Sharma, and R. Cooke. Assessing the natural variability of runoff: Clarence Basin catchments, NSW, Australia. Mathematics and Computers in Simulation, 43(3–6): 251–260, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000074.

Kothavala:1997:EPE

[1037] Zavareh Kothavala. Extreme precipitation events and the applicability of global climate models to the study of floods and droughts. Mathematics and Computers in Simulation, 43 (3–6):261–268, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000086.

Filar:1997:RAW

[1038] J. A. Filar and P. S. Gaertner. A regional allocation of world CO _{2} emission reductions. Mathematics and Computers in Simulation, 43(3– 6):269–275, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000098.

Herbert:1997:CSD

[1039] Jan M. Herbert, Glenn T. Johnson, and A. John Arnfield. Coupling of a scalar dispersion and an urban canyon energy budget model. Mathematics and Computers in Simulation, 43(3– 6):277–283, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000104.

Sydow:1997:STI

[1040] Achim Sydow, Thomas Lux, Peter Mieth, and Ralf-Peter Schäfer. Simulation of traffic-induced air pollution for mesoscale applications. Mathematics and Computers in Simulation, 43 (3–6):285–290, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000116.

Singh:1997:LRB

[1041] K. P. Singh, A. McDaniel, A. A. Bartolucci, T. M. Smoot, and R. Holmes. Linear regression and back-calculation analyses with applications to AIDS cases in Alabama. Mathematics and Computers in Simulation, 43(3– 6):291–295, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000128.

Flugel:1997:CGR

[1042] Wolfgan-Albert Flügel. Combining GIS with regional hydrological modelling using hydrological response units (HRUs): an application from Germany. Mathematics and Computers in Simulation, 43(3–6):297– 304, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847549700013X.

Bende:1997:RHM

[1043] Ulrike Bende. Regional hydrochemical modelling by delineation of chemical hydrological response units (CHRUs) within a GIS: an approach of observing man-made impacts in the Broel River catchment (Germany). Mathematics and Computers in Simulation, 43(3–6):305–312, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000141.

Gallant:1997:SDT

[1044] John C. Gallant and Michael F.

Hutchinson. Scale dependence in terrain analysis. Mathematics and Computers in Simulation, 43(3–6):313– 321, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000153.

Ghassemi:1997:NMM

[1045] Fereidoun Ghassemi, Andrew Close, and Jim R. Kellett. Numerical models for the management of land and water resources salinisation. Mathematics and Computers in Simulation, 43 (3–6):323–329, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000165.

Barry:1997:ASO

[1046] D. A. Barry, C. T. Miller, P. J. Culligan, and K. Bajracharya. Analysis of split operator methods for nonlinear and multispecies groundwater chemical transport models. Mathematics and Computers in Simulation, 43 (3–6):331–341, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000177.

Willgoose:1997:HPT

[1047] G. Willgoose. A hydrodynamic particle tracking algorithm for simulating settling of sediment. Mathematics and Computers in Simulation, 43 (3–6):343–349, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000189.

Chapman:1997:SMD

[1048] Tom G. Chapman. Stochastic models for daily rainfall in the Western Pacific. Mathematics and Computers in Simulation, 43(3–6):351– 358, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475497000190.

Mroczkowski:1997:SSS

[1049] Marek Mroczkowski and George Kuczera. Sensitivity of simulated stream salinity to soil chloride profile. Mathematics and Computers in Simulation, 43(3–6):359–366, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000207.

Hansen:1997:IIF

[1050] D. P. Hansen, A. J. Jakeman, C. Kendall, and G. Weizu. Identification of internal flow dynamics in two experimental catchments. Mathematics and Computers in Simulation, 43 (3–6):367–375, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000219.

Maier:1997:MCB

[1051] H. R. Maier and G. C. Dandy. Modelling cyanobacteria (blue-green algae) in the River Murray using artificial neural networks. Mathematics and Computers in Simulation, 43 (3–6):377–386, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000220.

Hamilton:1997:CSS

[1052] D. P. Hamilton, G. C. Hocking, and J. C. Patterson. Criteria for selection of spatial dimension in the application of one- and two-dimensional water quality models. Mathematics and Computers in Simulation, 43(3– 6):387–393, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000232.

Morimune:1997:AAU

[1053] Kimio Morimune and Kenji Miyazaki. ARIMA approach to the unit root analysis of macro economic time series. Mathematics and Computers in Simulation, 43(3–6):395– 403, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475497000244.

Hurn:1997:ICP

[1054] A. S. Hurn and A. D. McDonald. Isolating cyclical patterns in irregular time-series data. Mathematics and Computers in Simulation, 43(3– 6):405–412, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000256.

Leong:1997:SIE

[1055] Kenneth Leong. Seasonal integration in economic time series. Mathematics and Computers in Simulation, 43 (3–6):413–419, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000268.

Sequeira:1997:EML

[1056] John M. Sequeira. Econometric modelling of long-run relationships in the Singapore currency futures market. Mathematics and Computers in Simulation, 43(3–6):421– 427, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847549700027X.

Oxley:1997:CGP

[1057] Les Oxley and David Greasley. Convergence in GDP per capita and real wages: Some results for Australia and the UK. Mathematics and Computers in Simulation, 43(3– 6):429–436, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000281.

Enders:1997:CTG

[1058] Walter Enders and Stan Hurn. Common trends and generalized purchasing power parity. Mathematics and Computers in Simulation, 43(3–6): 437–443, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000293.

Koong:1997:TLM

[1059] C. S. Koong, Albert K. Tsui, and W. S.

Chan. On tests for long memory in Pacific Basin stock returns. Mathematics and Computers in Simulation, 43 (3–6):445–449, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847549700030X.

McKenzie:1997:PST

[1060] C. R. McKenzie. The properties of some two step estimators of ARMA Models. Mathematics and Computers in Simulation, 43(3–6):451– 456, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000311.

Franses:1997:TPI

[1061] Philip Hans Franses and Michael McAleer. Testing periodically integrated autoregressive models. Mathematics and Computers in Simulation, 43(3–6):457–465, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000323.

Maekawa:1997:PIA

[1062] Koichi Maekawa. Periodically integrated autoregression with a structural break. Mathematics and Computers in Simulation, 43(3–6):467– 473, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000335.

Tse:1997:STI

[1063] Y. K. Tse. Short-term interest rate models and generation of interest rate scenarios. Mathematics and Computers in Simulation, 43(3– 6):475–480, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000347.

Siddique:1997:DMM

[1064] M. A. B. Siddique. Demand for machinery and manufactured goods in Malaysia. Mathematics and Computers in Simulation, 43(3–6):481– 486, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000359.

Watkins:1997:TSI

[1065] Clinton Watkins. The term structure of interest rates and economic activity: an empirical critique. Mathematics and Computers in Simulation, 43 (3–6):487–493, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000360.

Hurn:1997:EPS

[1066] A. Stan Hurn and K. A. Lindsay. Estimating the parameters of stochastic differential equations by Monte Carlo methods. Mathematics and Computers in Simulation, 43(3–6): 495–501, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000372.

Nawata:1997:EGR

[1067] K. Nawata. Estimation of generalised regression models by the grouping method. Mathematics and Computers in Simulation, 43(3–6):503– 510, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475497000384.

Takase:1997:ARM

[1068] Mitsuo Takase. Architectural regulation on the maximum height of buildings and urban economic model. Mathematics and Computers in Simulation, 43(3–6):511–517, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000396.

Maki:1997:OHP

[1069] Atsushi Maki and Eitaro Aiyoshi. Optimal household portfolio behavior. Mathematics and Computers in Simulation, 43(3–6):519– 525, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475497000402.

Hu:1997:PAC

[1070] Baiding Hu and Michael McAleer. A probit analysis of consumer behaviour in rural China. Mathematics and Computers in Simulation, 43 (3–6):527–534, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000414.

Lim:1997:FSI

[1071] Christine Lim. The functional specification of international tourism demand models. Mathematics and Computers in Simulation, 43(3–6):535– 543, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000426.

Salim:1997:CSP

[1072] V. Salim. A comparison of shortest path algorithms in urban traffic assignment. Mathematics and Computers in Simulation, 43(3–6):545– 551, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000438.

Keuzenkamp:1997:CS

[1073] Hugo A. Keuzenkamp and Michael McAleer. The complexity of simplicity. Mathematics and Computers in Simulation, 43(3–6):553– 561, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847549700044X.

Everett:1997:SRV

[1074] J. E. Everett. Simulation to reduce variability in iron ore stockpiles. Mathematics and Computers in Simulation, 43(3–6):563–568, March 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000451.

Qiu:1997:PSR

[1075] Min Qiu. Prioritising and scheduling road projects by genetic algorithm. Mathematics and Computers in Simulation, 43(3–6):569– 574, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475497827619.

Anonymous:1997:ICEa

[1076] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 43(3–6):575– 588, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0378475497900057.

Anonymous:1997:AIV

[1077] Anonymous. Author index of volume 43 (1997). Mathematics and Computers in Simulation, 43(3–6): 589–592, March 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475497827620.

Anonymous:1997:EBb

[1078] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 44(1):ii, May 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497900112.

McKenzie:1997:ELM

[1079] C. R. McKenzie and C. A. Kapuscinski. Estimation in a linear model with serially correlated errors when observations are missing. Mathematics and Computers in Simulation, 44(1):1–9, May 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000025.

Besancon:1997:SCN

[1080] Gildas Besançon and Guy Bornard.

Symbolic computation in nonlinear control systems analysis: the example of bilinearization. Mathematics and Computers in Simulation, 44(1):11–27, May 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000037.

Papaodysseus:1997:EAF

[1081] C. Papaodysseus, E. Koukoutsis, G. Stavrakakis, and C. C. Halkias. Exact analysis of the finite precision error generation and propagation in the FAEST and the fast transversal algorithms: a general methodology for developing robust RLS schemes. Mathematics and Computers in Simulation, 44(1):29–41, May 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000049.

Joshi:1997:MAS

[1082] Anupam Joshi, Tzvetan Drashansky, John Rice, Sanjiva Weerawarana, and Elias Houstis. Multi-agent simulation of complex heterogeneous models in scientific computing. Mathematics and Computers in Simulation, 44(1):43–59, May 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000487.

Kuruganti:1997:OIS

[1083] I. Kuruganti and S. Strickland. Optimal importance sampling for Markovian systems with applications to tandem queues. Mathematics and Computers in Simulation, 44(1):61– 79, May 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475497000499.

Feldman:1997:OTT

[1084] Phillip Feldman, Adviti Muni, and Glen Swindle. An optimal termination testing procedure for discrete event simulations. Mathematics and Computers in Simulation, 44(1):81–98, May 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000517.

LEcuyer:1997:RNG

[1085] Pierre L’Ecuyer and Terry H. Andres. A random number generator based on the combination of four LCGs. Mathematics and Computers in Simulation, 44(1):99–107, May 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000529.

Anonymous:1997:PM

[1086] Anonymous. Pages 1–107 (May 1997). Mathematics and Computers in Simulation, 44(1):??, May 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Mackevicius:1997:CRE

[1087] Vigirdas Mackevičius. Convergence rate of Euler scheme for stochastic differential equations: Functionals of solutions. Mathematics and Computers in Simulation, 44(2):109– 121, September 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000475.

Aracil:1997:BPL

[1088] J. Aracil, E. Ponce, and L. Pizarro.

Behavior patterns of logistic models with a delay. Mathematics and Computers in Simulation, 44(2):123– 141, September 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000608.

Fung:1997:DTL

[1089] K. K. Fung and S. Y. R. Hui.

Duality of transmission line models for simulation of reactive circuit components. Mathematics and Computers in Simulation, 44(2):143– 154, September 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000682.

Wu:1997:CFS

[1090] Tzong-Mou Wu and Cha’o-Kuang Chen. Closed form synthesis of planar and spatial linkages function generation with symbolic representation. Mathematics and Computers in Simulation, 44(2):155–162, September 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000694.

Diamond:1997:MRC

[1091] P. Diamond, P. E. Kloeden, V. S. Kozyakin, and A. V. Pokrovskii. A model for roundoff and collapse in computation of chaotic dynamical systems. Mathematics and Computers in Simulation, 44(2):163–185, September 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000700.

Albrecht:1997:CEM

[1092] A. Albrecht, S. K. Cheung, K. S. Leung, and C. K. Wong. Computing elastic moduli of two-dimensional random networks of rigid and nonrigid bonds by simulated annealing. Mathematics and Computers in Simulation, 44(2): 187–215, September 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000724.

Anonymous:1997:ICEb

[1093] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 44(2):217–224, September 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/037847549790001X.

Anonymous:1997:PS

[1094] Anonymous. Pages 109–224 (September 1997). Mathematics and Computers in Simulation, 44(2):??, September 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Ceric:1997:P

[1095] Vlatko Čerić. Preface. Mathematics and Computers in Simulation, 44(3):231, October 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847549781453X.

Sargent:1997:MQS

[1096] R. G. Sargent. Modeling queueing systems using hierarchical control flow graph models. Mathematics and Computers in Simulation, 44 (3):233–249, October 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000542.

Ceric:1997:VIM

[1097] Vlatko Čerić. Visual interactive modeling and simulation as a decision support in railway transport logistic operations. Mathematics and Computers in Simulation, 44(3):251– 261, October 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000554.

Slamic:1997:FVR

[1098] Miroslav Slamić, Ivo Penzar, Dražen Penzar, and Jadranko Stjepanović.

Fixed-viewpoint representation of synthetic environment for the low-cost industrial and military simulators. Mathematics and Computers in Simulation, 44(3):263–269, October 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000566.

Georgiev:1997:MMF

[1099] Tz. Georgiev, Al. Ratkov, and St. Tzonkov. Mathematical modelling of fed-batch fermentation processes for amino acid production. Mathematics and Computers in Simulation, 44 (3):271–285, October 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000591.

Kurtanjek:1997:PCA

[1100] Zelimir Kurtanjek. Principal component analysis of bioreactor fedbatch operation by computer simulation. Mathematics and Computers in Simulation, 44(3):287– 294, October 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847549700058X.

Stojanov:1997:IEV

[1101] Georgi Stojanov, Stevo Bozinovski, and Goran Trajkovski. Interactionistexpectative view on agency and learning. Mathematics and Computers in Simulation, 44(3):295– 310, October 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000578.

Anonymous:1997:I

[1102] Anonymous. (ITI ’96). Mathematics and Computers in Simulation, 44 (3):??, October 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Deville:1997:F

[1103] M. O. Deville and S. Gavrilakis. Foreword. Mathematics and Computers in Simulation, 44(4):311, November 14, 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000967.

Formaggia:1997:NIS

[1104] Luca Formaggia, Alan Scheinine, and Alfio Quarteroni. A numerical investigation of Schwarz domain decomposition techniques for elliptic problems on unstructured grids. Mathematics and Computers in Simulation, 44(4): 313–330, November 14, 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000621.

Gatski:1997:PCF

[1105] Thomas B. Gatski. Prediction of complex flow fields using higher-order turbulent closures. Mathematics and Computers in Simulation, 44(4):331– 345, November 14, 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000633.

Otto:1997:BVC

[1106] John Otto, Marius Paraschivoiu, Serhat Yesilyurt, and Anthony T. Patera. Bayesian-validated computersimulation surrogates for optimization and design: Error estimates and applications. Mathematics and Computers in Simulation, 44(4):347– 367, November 14, 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847549700061X.

Wesseling:1997:CTF

[1107] P. Wesseling, M. Zijlema, A. Segal, and C. G. M. Kassels. Computation of turbulent flow in general domains. Mathematics and Computers in Simulation, 44(4):369–385, November 14, 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475497000645.

Bereaux:1997:NST

[1108] Yves Béreaux and Jean-Robert Clermont. Numerical simulation of twoand three-dimensional complex flows of viscoelastic fluids using the streamtube method. Mathematics and Computers in Simulation, 44(4):387–400, November 14, 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475497000657.

Sprossig:1997:TFP

[1109] Wolfgang Sprössig and Klaus Gürlebeck. On the treatment of fluid problems by methods of Clifford analysis. Mathematics and Computers in Simulation, 44(4):401–413, November 14, 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000669.

Tong:1997:NAM

[1110] R. P. Tong. A new approach to modelling an unsteady free surface in boundary integral methods with application to bubble-structure interactions. Mathematics and Computers in Simulation, 44(4):415–426, November 14, 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000670.

Golenko-Ginzburg:1997:JSR

[1111] Dimitri Golenko-Ginzburg and Aharon Gonik. Job-shop resource scheduling via simulating random operations. Mathematics and Computers in Simulation, 44(5):427–440, December 1, 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000748.

Castillo:1997:GLS

[1112] Jose E. Castillo and James S. Otto.

A generalized length strategy for direct optimization in planar grid generation. Mathematics and Computers in Simulation, 44(5):441–456, December 1, 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475497000797.

Hsiao:1997:SAL

[1113] Chun-Hui Hsiao. State analysis of linear time delayed systems via Haar wavelets. Mathematics and Computers in Simulation, 44(5):457–470, December 1, 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847549700075X.

Sanders:1997:CCL

[1114] Jan A. Sanders and Jing Ping Wang. Classification of conservation laws for KdV-like equations. Mathematics and Computers in Simulation, 44(5):471– 481, December 1, 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000761.

Sanders:1997:HDC

[1115] Jan A. Sanders and Jing Ping Wang. Hodge decomposition and conservation laws. Mathematics and Computers in Simulation, 44(5):483–493, December 1, 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475497000773.

Nordling:1997:PSM

[1116] Patrik Nordling and Anders Sjö. Parallel solution of modular ODEs with application to rolling bearing dynamics. Mathematics and Computers in Simulation, 44(5):495–504, December 1, 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000785.

RubioGarcia:1997:EIF

[1117] Julio Jesús Rubio García. Erratum to “Integrating functional programming and symbolic computation”: Math. Comp. Simul. 42 (1996) 467–473. Mathematics and Computers in Simulation, 44(5):505–511, December 1, 1997. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497001304. See

[957].

Troch:1997:CR

[1118] Inge Troch. Conference report. Mathematics and Computers in Simulation, 44(5):513–516, December 1, 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475497800018.

Anonymous:1997:NIc

[1119] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 44(5):517–526, December 1, 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/0378475497900033.

Anonymous:1997:PD

[1120] Anonymous. Pages 427–526 (1 December 1997). Mathematics and Computers in Simulation, 44(5):??, December 1, 1997. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Bees:1998:PIC

[1121] M. A. Bees, I. Mezic, and J. McGlade. Planktonic interactions and chaotic advection in Langmuir circulation. Mathematics and Computers in Simulation, 44(6):527–544, January 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475497000979.

Keeping:1998:DMP

[1122] B. R. Keeping and C. C. Pantelides. A distributed memory parallel algorithm for the efficient computation of sensitivities of differentialalgebraic systems. Mathematics and Computers in Simulation, 44(6):545– 558, January 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000992.

Bischi:1998:SIC

[1123] Gian-Italo Bischi, Luciano Stefanini, and Laura Gardini. Synchronization, intermittency and critical curves in a duopoly game. Mathematics and Computers in Simulation, 44(6): 559–585, January 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001006.

Rodrigo:1998:MRW

[1124] Renuka D. Rodrigo and Miles G. Nicholls. Modelling real world industrial systems: an inter-disciplinary approach based on petri nets. Mathematics and Computers in Simulation, 44 (6):587–597, January 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001602.

Anonymous:1998:PJa

[1125] Anonymous. Pages 527–618 (1 January 1998). Mathematics and Computers in Simulation, 44(6):??, January 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1998:Ea

[1126] Anonymous. Editorial. Mathematics and Computers in Simulation, 45(1– 2):1–2, January 31, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497100465.

Akers:1998:ASN

[1127] Robert L. Akers, Paul Baffes, Elaine Kant, Curt Randall, Stanly Steinberg, and Robert L. Young. Automatic synthesis of numerical codes for solving partial differential equations. Mathematics and Computers in Simulation, 45(1–2):3–22, January 31, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000827.

Nakagawa:1998:CAA

[1128] Shigekazu Nakagawa, Naoto Niki, and Hiroki Hashiguchi. Computer algebra application to the distribution of sample correlation coefficient. Mathematics and Computers in Simulation, 45(1– 2):23–32, January 31, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000839.

Calmet:1998:CAF

[1129] J. Calmet and W. M. Seiler. Computer algebra and field theories. Mathematics and Computers in Simulation, 45(1– 2):33–37, January 31, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000840.

Llovet:1998:CCP

[1130] Juan Llovet, Bonifacio Castaño, and Raquel Martínez. Computing the characteristic polynomial of multivariate polynomial matrices given by straight-line programs. Mathematics and Computers in Simulation, 45(1– 2):39–57, January 31, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000852.

Ortiz:1998:RES

[1131] Alejandro A. R. Trejo Ortiz and Guillermo Fernández Anaya. Regular expression simplification. Mathematics and Computers in Simulation, 45(1– 2):59–71, January 31, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497000864.

Calmet:1998:CAA

[1132] Jacques Calmet. Computer algebra and artificial intelligence. Mathematics and Computers in Simulation, 45(1– 2):73–82, January 31, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000876.

Roanes-Lozano:1998:PMM

[1133] Eugenio Roanes-Lozano, Luis M. Laita, and Eugenio Roanes-Macías. A polynomial model for multi-valued Logics with a touch of Algebraic Geometry and Computer Algebra. Mathematics and Computers in Simulation, 45(1– 2):83–99, January 31, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000888.

Taylor:1998:TMS

[1134] R. Taylor. Thermodynamics with Maple: I — Symbolic Computation. Mathematics and Computers in Simulation, 45(1–2):101–119, January 31, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847549700089X. Non-

standard applications of computer algebra (Linz/Wailea, HI, 1996/1997).

Taylor:1998:TMI

[1135] Ross Taylor. Thermodynamics with Maple: II — Numerical and graphical applications. Mathematics and Computers in Simulation, 45(1–2):121– 146, January 31, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000906. Non-

standard applications of computer algebra (Linz/Wailea, HI, 1996/1997).

Searles:1998:NEM

[1136] Debra J. Searles, Dennis J. Isbister, and Denis J. Evans. Non-equilibrium molecular dynamics integrators using Maple. Mathematics and Computers in Simulation, 45(1–2):147– 162, January 31, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000918. Non-

standard applications of computer algebra (Linz/Wailea, HI, 1996/1997).

Montes:1998:ASL

[1137] Antonio Montes. Algebraic solution of the load-flow problem for a 4nodes electrical network. Mathematics and Computers in Simulation, 45(1–2): 163–174, January 31, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847549700092X.

Roanes-Lozano:1998:ATI

[1138] Eugenio Roanes-Lozano and Luis M. Laita. An applicable topologyindependent model for railway interlocking systems. Mathematics and Computers in Simulation, 45(1–2):175– 183, January 31, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000931.

Recio:1998:DCA

[1139] T. Recio and M. J. González-López. Does Computer Algebra help at all learning about real numbers? Mathematics and Computers in Simulation, 45(1–2):185–195, January 31, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497000943.

Bohm:1998:DAD

[1140] Josef Böhm. Dimensional Analysis with DERIVE. Mathematics and Computers in Simulation, 45(1–2):197– 205, January 31, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497000955.

Anonymous:1998:PJb

[1141] Anonymous. Pages 1–218 (31 January 1998). Mathematics and Computers in Simulation, 45(1–2):??, January 31, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Kolmanovskii:1998:SNT

[1142] V. B. Kolmanovskii and J.-P Richard.

Some new trends in the study of time-delay systems. Mathematics and Computers in Simulation, 45(3–4): 219–221, February 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001018.

Picard:1998:WCC

[1143] P. Picard, O. Sename, and J. F.

Lafay. Weak controllability and controllability indices for linear neutral systems. Mathematics and Computers in Simulation, 45(3–4):223– 233, February 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847549700102X.

Conte:1998:DFD

[1144] G. Conte, A. M. Perdon, and A. Lombardo. Dynamic Feedback Decoupling Problem for Delay-differential Systems via Systems over Rings. Mathematics and Computers in Simulation, 45(3–4): 235–244, February 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001031.

Goubet-Bartholomeus:1998:DDS

[1145] A. Goubet-Bartholoméüs, M. Dambrine, and J. P. Richard. Delay-dependent stability domains of nonlinear delay systems. Mathematics and Computers in Simulation, 45(3–4):245– 256, February 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001043.

Verriest:1998:SND

[1146] Erik I. Verriest and Woihida Aggoune. Stability of nonlinear differential delay systems. Mathematics and Computers in Simulation, 45(3–4): 257–267, February 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001055.

Beretta:1998:SEM

[1147] Edoardo Beretta, Vladimir Kolmanovskii, and Leonid Shaikhet. Stability of epidemic model with time delays influenced by stochastic perturbations. Mathematics and Computers in Simulation, 45(3–4):269– 277, February 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001067.

Goodall:1998:SCF

[1148] David P. Goodall. Stability criteria for feedback-controlled, imperfectly known, bilinear systems with timevarying delay. Mathematics and Computers in Simulation, 45(3–4):279– 289, February 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001079.

Tchangani:1998:SAD

[1149] A. P. Tchangani, M. Dambrine, and J. P. Richard. Stability, attraction domains, and ultimate boundedness for nonlinear neutral systems. Mathematics and Computers in Simulation, 45(3–4):291–298, February 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497001080.

Arino:1998:SCN

[1150] Ovide Arino and Valery R. Nosov.

On stability of a class of neutral type functional differential equations. Mathematics and Computers in Simulation, 45(3–4):299–307, February 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497001092.

Baker:1998:SGS

[1151] Christopher T. H. Baker and Arsalang Tang. Stability of grazing systems with infinite delays. Mathematics and Computers in Simulation, 45(3–4): 309–317, February 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001109.

Kubo:1998:ESS

[1152] T. Kubo and E. Shimemura. Exponential stabilization of systems with time-delay by optimal memoryless feedback. Mathematics and Computers in Simulation, 45(3–4):319– 328, February 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001110.

Mounier:1998:SCL

[1153] Hugues Mounier. Stabilization of a class of linear delay systems. Mathematics and Computers in Simulation, 45(3–4):329–338, February 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497001122.

Brethe:1998:EAF

[1154] D. Brethé and J. J. Loiseau. An effective algorithm for finite spectrum assignment of single-input systems with delays. Mathematics and Computers in Simulation, 45(3–4): 339–348, February 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001134.

Li:1998:RGC

[1155] Huaizhong Li, Silviu-Iulian Niculescu, Luc Dugard, and Jean-Michel Dion. Robust guaranteed cost control of uncertain linear time-delay systems using dynamic output feedback. Mathematics and Computers in Simulation, 45(3–4): 349–358, February 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001146.

Rodkina:1998:RCL

[1156] Alexandra Rodkina. On rate control of n-link manipulator robot. Mathematics and Computers in Simulation, 45(3–4): 359–364, February 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001158.

Takahashi:1998:DMS

[1157] Kazukiyo Takahashi and Keiji Watanabe. 3 Disk mixed sensitivity problem of time-delay systems. Mathematics and Computers in Simulation, 45(3–4): 365–376, February 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847549700116X.

Kim:1998:MNM

[1158] A. V. Kim and V. G. Pimenov. Multistep numerical methods for functional differential equations. Mathematics and Computers in Simulation, 45(3–4): 377–384, February 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497001171.

Hotzel:1998:LSF

[1159] Richard Hotzel and Michel Fliess. On linear systems with a fractional derivation: Introductory theory and examples. Mathematics and Computers in Simulation, 45(3–4):385– 395, February 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497001183.

Anonymous:1998:PF

[1160] Anonymous. Pages 219–406 (1 February 1998). Mathematics and Computers in Simulation, 45(3–4):??, February 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1998:Eb

[1161] Anonymous. Editorial. Mathematics and Computers in Simulation, 45(5– 6):409–411, March 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497001195.

Bruno:1998:NF

[1162] A. D. Bruno. Normal forms. Mathematics and Computers in Simulation, 45 (5–6):413–427, March 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001201.

Bruno:1998:NPP

[1163] A. D. Bruno. Newton polyhedra and power transformations. Mathematics and Computers in Simulation, 45(5– 6):429–443, March 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001213. See

erratum [1293].

Edneral:1998:SAP

[1164] Victor F. Edneral. A symbolic approximation of periodic solutions of the Henon–Heiles system by the normal form method. Mathematics and Computers in Simulation, 45(5–6): 445–463, March 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001225.

Sadov:1998:FDS

[1165] Sergey Sadov. Functions that determine stability of rational rotations of a near symmetric satellite. Mathematics and Computers in Simulation, 45 (5–6):465–484, March 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001237.

Todesco:1998:ENR

[1166] E. Todesco, M. Gemmi, and M. Giovannozzi. Evaluation of nonlinear resonances in 4D symplectic mappings. Mathematics and Computers in Simulation, 45(5–6):485–496, March 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475497001249.

Abad:1998:ASP

[1167] A. Abad, A. Elipe, J. Palacián, and J. F. San-Juan. ATESAT: a symbolic processor for artificial satellite theory. Mathematics and Computers in Simulation, 45(5–6):497– 510, March 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475497001250.

Ferro:1998:TMD

[1168] Giuseppa Carra’ Ferro. Triangular matrices, differential resultants and systems of linear homogeneous PDE’s. Mathematics and Computers in Simulation, 45(5–6):511– 518, March 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475497001262.

Gerdt:1998:IBP

[1169] Vladimir P. Gerdt and Yuri A. Blinkov. Involutive bases of polynomial ideals. Mathematics and Computers in Simulation, 45(5–6):519– 541, March 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497001274.

Gerdt:1998:MIB

[1170] Vladimir P. Gerdt and Yuri A. Blinkov.

Minimal involutive bases. Mathematics and Computers in Simulation, 45(5– 6):543–560, March 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001286.

Seiler:1998:NAC

[1171] Werner M. Seiler. Numerical analysis of constrained Hamiltonian systems and the formal theory of differential equations. Mathematics and Computers in Simulation, 45(5–6): 561–576, March 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001298.

Anonymous:1998:Ia

[1172] Anonymous. Index. Mathematics and Computers in Simulation, 45(5– 6):577–579, March 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498000949.

Anonymous:1998:PMa

[1173] Anonymous. Pages 409–590 (1 March 1998). Mathematics and Computers in Simulation, 45(5–6):??, March 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Spitaleri:1998:MMG

[1174] Rosa Maria Spitaleri and Valentina Micacchi. A multiblock multigrid grid generation method for complex simulations. Mathematics and Computers in Simulation, 46(1):1– 12, April 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497001535.

Javor:1998:SOU

[1175] A. Jávor and G. Szűcs. Simulation and optimization of urban traffic using AI. Mathematics and Computers in Simulation, 46(1):13– 21, April 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497001547.

Molnar:1998:MUS

[1176] István Molnár and Pham Quang Huy. Multimedia using simulation models. Mathematics and Computers in Simulation, 46(1):23– 33, April 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497001559.

Hur:1998:DES

[1177] Youngmin Hur and Stephen A. Szygenda. Design error simulation based on error modeling and sampling techniques. Mathematics and Computers in Simulation, 46(1):35– 46, April 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497001560.

Tricas:1998:DCS

[1178] Fernando Tricas and Javier Martínez.

Distributed control systems simulation using high level Petri nets. Mathematics and Computers in Simulation, 46(1):47–55, April 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001572.

Redey:1998:RBC

[1179] Á. L. Rédey, M. J. Morse, and A. Platt. Routing in broadband communication networks using neural computations. Mathematics and Computers in Simulation, 46(1):57– 65, April 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497001584.

Zupancic:1998:MHM

[1180] B. Zupančič. Modular hierarchical modelling with SIMCOS language. Mathematics and Computers in Simulation, 46(1):67–76, April 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497001596.

Anonymous:1998:PAa

[1181] Anonymous. Pages 1–86 (1 April 1998).

Mathematics and Computers in Simulation, 46(1):??, April 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Tinoco-Ruiz:1998:SCG

[1182] J. G. Tinoco-Ruiz and P. Barrera-Sánchez. Smooth and convex grid generation over general plane regions. Mathematics and Computers in Simulation, 46(2):87–102, April 30, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498000743.

Cagnol:1998:SEH

[1183] John Cagnol and Jean-Paul Marmorat. Static equilibrium of hyperelastic thin shell: symbolic and numerical computation. Mathematics and Computers in Simulation, 46(2): 103–115, April 30, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847549800072X.

Coleman:1998:CRC

[1184] R. Coleman. On the construction of real canonical forms of Hamiltonian matrices whose spectrum is an imaginary pair. Mathematics and Computers in Simulation, 46(2):117– 155, April 30, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498000731.

Anonymous:1998:PAb

[1185] Anonymous. Pages 87–172 (30 April 1998). Mathematics and Computers in Simulation, 46(2):??, April 30, 1998.

CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1998:Ec

[1186] Anonymous. Editorial. Mathematics and Computers in Simulation, 46 (3–4):173, May 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847549700133X.

Matsuo:1998:MSM

[1187] Takayoshi Matsuo, Steffen Bernet, R. Stephen Colby, and Thomas A. Lipo. Modeling and simulation of matrix converter/induction motor drive. Mathematics and Computers in Simulation, 46(3–4):175– 195, May 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497001341.

Jufer:1998:MTC

[1188] M. Jufer, N. Macabrey, and M. Perrottet. Modeling and test of contactless inductive energy transmission. Mathematics and Computers in Simulation, 46(3–4):197– 211, May 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497001353.

Enright:1998:MML

[1189] W. Enright, J. Arrillaga, N. R. Watson, and J. Zavahir. Modelling multilimb transformers with an electromagnetic transient program. Mathematics and Computers in Simulation, 46 (3–4):213–223, May 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001365.

Mimoune:1998:NSA

[1190] S. M. Mimoune, J. Fouladgar, and G. Develey. 3D numerical study of the annular induction plasma installation. Mathematics and Computers in Simulation, 46(3–4):225– 238, May 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497001377.

Yoshida:1998:DCL

[1191] K. Yoshida, H. Takami, and L. Shi. Decoupled-control of levitation and propulsion in underwater LM car ME02. Mathematics and Computers in Simulation, 46(3–4):239– 256, May 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475497001389.

Richard:1998:CEF

[1192] N. Richard. Calculation of electromagnetic forces on large generator end-windings under fault conditions using a three-dimensional finite element method. Mathematics and Computers in Simulation, 46(3– 4):257–263, May 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001390.

Robyns:1998:FLB

[1193] B. Robyns, H. Buyse, and F. Labrique.

Fuzzy logic based field orientation in an indirect FOC strategy of an induction actuator. Mathematics and Computers in Simulation, 46(3–4): 265–274, May 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001407.

Richardeau:1998:SSC

[1194] F. Richardeau, Y. Chéron, and L. Marroyo. Study and simulation of a cycloconverter-active filter device with a unity power factor. Mathematics and Computers in Simulation, 46(3– 4):275–287, May 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001419.

Zhong:1998:MES

[1195] L. Zhong, M. F. Rahman, and K. W.

Lim. Modelling and experimental studies of an instantaneous torque and field weakening control scheme for an interior permanent magnet synchronous motor drive. Mathematics and Computers in Simulation, 46(3– 4):289–300, May 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001420.

Debruyne:1998:CMH

[1196] H. Debruyne, S. Clenet, and F. Piriou.

Characterisation and modelling of hysteresis phenomenon. Mathematics and Computers in Simulation, 46(3– 4):301–311, May 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001432.

Kruger:1998:SNM

[1197] L. Krüger, D. Naunin, and C. Garbrecht. Stochastic and neural models of an induction motor. Mathematics and Computers in Simulation, 46 (3–4):313–324, May 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001444.

Bouscayrol:1998:CCV

[1198] Alain Bouscayrol, Maria Pietrzak-David, and Bernard de Fornel. Comparison of cartesian vector control and polar vector control for induction motor drives. Mathematics and Computers in Simulation, 46(3–4): 325–337, May 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001456.

Binder:1998:EIB

[1199] A. Binder. Electromagnetic interference of brake stray fields on resolvers in brushless d.c. servo drives. Mathematics and Computers in Simulation, 46 (3–4):339–347, May 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001468.

Khezzar:1998:AFT

[1200] A. Khezzar and B. Davat. Active filtering of torque ripples in double stator synchronous machines. Mathematics and Computers in Simulation, 46 (3–4):349–359, May 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847549700147X.

Cheny:1998:ILD

[1201] Cyrille Cheny and Jean-Marie Kauffmann. Information losses in decoupling space harmonics effects for an induction drive. Mathematics and Computers in Simulation, 46(3–4): 361–372, May 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001481.

Flieller:1998:GSD

[1202] Damien Flieller, Jean-Paul Louis, and Jens Barrenscheen. General sampled data modeling of power systems supplied by static converter with digital and analog controller. Mathematics and Computers in Simulation, 46(3– 4):373–385, May 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475497001493.

Guerin:1998:SSL

[1203] P. Guerin, M. Machmoum, and R. Le Doeuff. Stochastic study of line harmonic currents produced by rectifiers. Mathematics and Computers in Simulation, 46(3–4):387– 396, May 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847549700150X.

Biechl:1998:STO

[1204] H. Biechl, A. Koch, and A. Matt. Simulation of the transient operational behaviour of a novel electromechanical vibration exciter for rock drills. Mathematics and Computers in Simulation, 46(3–4):397– 411, May 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475497001511.

Cambronne:1998:SDS

[1205] J. P. Cambronne and X. Pierre. Synthesis of different synchronous modulators for high power three-phase/single-phase PWM converters. Mathematics and Computers in Simulation, 46 (3–4):413–423, May 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475497001523.

Anonymous:1998:PMb

[1206] Anonymous. Pages 173–440 (1 May 1998). Mathematics and Computers in Simulation, 46(3–4):??, May 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1998:Ed

[1207] Anonymous. Editorial. Mathematics and Computers in Simulation, 46(5– 6):441–443, June 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498000755.

Engell:1998:MAH

[1208] Sebastian Engell. Modelling and analysis of hybrid systems. Mathematics and Computers in Simulation, 46(5– 6):445–464, June 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498000767.

Lunze:1998:QMD

[1209] Jan Lunze. Qualitative modelling of dynamical systems: Motivation, methods, and prospective applications. Mathematics and Computers in Simulation, 46(5–6):465– 483, June 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498000779.

Hellekalek:1998:GRN

[1210] P. Hellekalek. Good random number generators are (not so) easy to find. Mathematics and Computers in Simulation, 46(5–6):485–505, June 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL http://random.mat.sbg.ac.
at/results/peter/A19final.pdf;
https://www.sciencedirect.com/science/
article/pii/S0378475498000780.

Setnes:1998:CRF

[1211] M. Setnes, R. Babuška, and H. B.

Verbruggen. Complexity reduction in fuzzy modeling. Mathematics and Computers in Simulation, 46(5–6):

507–516, June 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498000792.

Schlacher:1998:TAB

[1212] Kurt Schlacher, Andreas Kugi, and Rudolf Scheidl. Tensor analysis based symbolic computation for mechatronic systems. Mathematics and Computers in Simulation, 46(5–6):517– 525, June 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475498000809.

Fedorova:1998:WOA

[1213] A. N. Fedorova and M. G. Zeitlin. Wavelets in optimization and approximations. Mathematics and Computers in Simulation, 46(5–6):527– 534, June 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475498000810.

Rukgauer:1998:SMD

[1214] A. Rükgauer and W. Schiehlen. Simulation of modular dynamic systems. Mathematics and Computers in Simulation, 46(5–6):535–542, June 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475498000822.

Gottsche:1998:NDM

[1215] Th. H. Göttsche, K. J. Hunt, and T. A. Johansen. Nonlinear dynamics modelling via operating regime decomposition. Mathematics and Computers in Simulation, 46(5–6):543– 550, June 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498000834.

Boutayeb:1998:MIH

[1216] M. Boutayeb, M. Darouach, and P. M.

Frank. Modelling and identification for highly nonlinear processes. Mathematics and Computers in Simulation, 46 (5–6):551–557, June 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498000846.

Neumann:1998:CSD

[1217] M. Neumann and H. Hahn. Computer simulation and dynamic analysis of a mechanical press based on different engineer models. Mathematics and Computers in Simulation, 46(5– 6):559–574, June 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498000858.

Zele:1998:MVI

[1218] Mina Žele and – Dani Juričić. Model validation in iterative identification and controller design. Mathematics and Computers in Simulation, 46(5– 6):575–583, June 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847549800086X.

Rivera:1998:INC

[1219] C. J. Rivera and J. V. R. Prasad. Identification of a nonlinear compressor model. Mathematics and Computers in Simulation, 46(5–6):585– 592, June 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475498000871.

Mikhailov:1998:TSC

[1220] S. A. Mikhailov and P. C. Müller. Time-suboptimal control design of singularly perturbed systems by reduced order feedback design. Mathematics and Computers in Simulation, 46(5– 6):593–600, June 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498000883.

Stocklmayer:1998:SRD

[1221] Ch Stöcklmayer and W. Höflinger. Simulation of the regeneration of dust filters. Mathematics and Computers in Simulation, 46(5–6):601– 609, June 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475498000895.

Cuno:1998:RBC

[1222] Bernd Cuno and Stephan Theobald. The relationship between control requirements, process complexity and modelling effort in the design process of river control systems. Mathematics and Computers in Simulation, 46 (5–6):611–619, June 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498000901.

Versyck:1998:OED

[1223] K. J. Versyck, J. E. Claes, and J. F.

Van Impe. Optimal experimental design for practical identification of unstructured growth models. Mathematics and Computers in Simulation, 46 (5–6):621–629, June 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498000913.

Naka:1998:SAE

[1224] Takashi Naka. Simulation analysis of the effects of the junctional folds on spontaneous generation of the miniature endplate current at neuromuscular junction. Mathematics and Computers in Simulation, 46(5– 6):631–639, June 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498000925.

Anonymous:1998:Ib

[1225] Anonymous. Index. Mathematics and Computers in Simulation, 46(5– 6):657–660, June 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001359.

Anonymous:1998:PJc

[1226] Anonymous. Pages 441–660 (1 June 1998). Mathematics and Computers in Simulation, 46(5–6):??, June 1, 1998.

CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Delgado:1998:PPA

[1227] Alberto Delgado. Phase portrait approximation using dynamic neural networks. Mathematics and Computers in Simulation, 47(1):1– 11, July 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475497001614. See

erratum [1300].

Gammal:1998:NII

[1228] A. Gammal. Numerical integration of the interaction energy of two point charges. Mathematics and Computers in Simulation, 47(1):13– 16, July 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475498001177.

Savioli:1998:SNM

[1229] Gabriela B. Savioli, Pablo M. Jacovkis, and M. Susana Bidner. On some numerical methods for solving 2D radial flow towards an oil well. Mathematics and Computers in Simulation, 47(1):17–36, July 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498001165.

Stefanovic:1998:AEA

[1230] Dragoslav L. Stefanovic and Heinz G. Stefan. An accurate and efficient algorithm for numerical simulation of conduction-type problems. Mathematics and Computers in Simulation, 47 (1):37–46, July 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001190.

Arnold:1998:SPU

[1231] Martin Arnold and Kurt Frischmuth.

Solving problems with unilateral constraints by DAE methods. Mathematics and Computers in Simulation, 47(1):47–67, July 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001189.

Anonymous:1998:PJd

[1232] Anonymous. Pages 1–82 (1 July 1998).

Mathematics and Computers in Simulation, 47(1):??, July 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1998:Ee

[1233] Anonymous. Editorial. Mathematics and Computers in Simulation, 47 (2–5):83, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001384.

Sabelfeld:1998:SMC

[1234] K. K. Sabelfeld. Stochastic models for coagulation of aerosol particles in intermittent turbulent flows. Mathematics and Computers in Simulation, 47 (2–5):85–101, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498000950.

Sobol:1998:QMC

[1235] I. M. Sobol. On quasi-Monte Carlo integrations. Mathematics and Computers in Simulation, 47(2–5):103– 112, August 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475498000962.

Alexandrov:1998:EPM

[1236] V. N. Alexandrov. Efficient parallel Monte Carlo methods for matrix computations. Mathematics and Computers in Simulation, 47(2–5): 113–122, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498000974.

Beikert:1998:NEE

[1237] G. J. Beikert. Next event estimation of cell tallies. Mathematics and Computers in Simulation, 47(2–5): 123–132, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498001207.

Blume:1998:ESQ

[1238] D. Blume, M. Lewerenz, and K. B. Whaley. Excited states by quantum Monte Carlo method. Mathematics and Computers in Simulation, 47(2– 5):133–141, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498000986.

Cerf:1998:MCS

[1239] N. J. Cerf and S. E. Koonin. Monte Carlo simulation of quantum computation. Mathematics and Computers in Simulation, 47(2–5):143– 152, August 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498000998.

Coulibaly:1998:SDU

[1240] Ibrahim Coulibaly and Christian Lécot.

Simulation of diffusion using quasirandom walk methods. Mathematics and Computers in Simulation, 47(2– 5):153–163, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001001.

Dimov:1998:NHC

[1241] I. T. Dimov and V. N. Alexandrov. A new highly convergent Monte Carlo method for matrix computations. Mathematics and Computers in Simulation, 47(2–5):165– 181, August 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498001013.

Dimov:1998:ECC

[1242] I. T. Dimov and T. V. Gurov. Estimates of the computational complexity of iterative Monte Carlo algorithm based on Green’s function approach. Mathematics and Computers in Simulation, 47(2–5):183– 199, August 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498001025.

Karaivanova:1998:EAA

[1243] Aneta Karaivanova and Ivan Dimov. Error analysis of an adaptive Monte Carlo method for numerical integration. Mathematics and Computers in Simulation, 47(2–5): 201–213, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001037.

Doyle:1998:PTT

[1244] J. P. Doyle and H. Rief. Photon transport in three-dimensional structures treated by random walk techniques: Monte Carlo benchmark of ocean colour simulations. Mathematics and Computers in Simulation, 47(2– 5):215–241, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001219.

Dubi:1998:AAM

[1245] A. Dubi. Analytic approach & Monte Carlo methods for realistic systems analysis. Mathematics and Computers in Simulation, 47(2–5):243– 269, August 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498001220.

Gallmeier:1998:CCM

[1246] Franz Gallmeier and Werner Gaubatz.

Criticality calculations with MCNP4A using correlated sampling methods. Mathematics and Computers in Simulation, 47(2–5):271–285, August 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498001049.

Genz:1998:SMM

[1247] Alan Genz. Stochastic methods for multiple integrals over unbounded regions. Mathematics and Computers in Simulation, 47(2–5):287–298, August 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498001050.

Gueorguiev:1998:NMP

[1248] Y. Gueorguiev, K. Vutova, and G. Mladenov. Numerical modelling of the processes of exposure and development in electron beam lithography on high temperature superconductor thin films. Mathematics and Computers in Simulation, 47(2–5): 299–307, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001347.

Halton:1998:IER

[1249] John H. Halton and Pradip K. Sarkar.

Increasing the efficiency of radiation shielding calculations by using antithetic variates. Mathematics and Computers in Simulation, 47(2–5): 309–318, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001074.

Handscomb:1998:SOT

[1250] David Handscomb. Simulation of order transitions on a parallel computer. Mathematics and Computers in Simulation, 47(2–5):319–327, August 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498001086.

Janke:1998:NMC

[1251] Wolfhard Janke. Nonlocal Monte Carlo algorithms for statistical physics applications. Mathematics and Computers in Simulation, 47(2–5):329– 346, August 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498001098.

Labeau:1998:CBM

[1252] P. E. Labeau and E. Zio. The cellto-boundary method in the frame of memorization-based Monte Carlo algorithms. A new computational improvement in dynamic reliability. Mathematics and Computers in Simulation, 47(2–5):347–360, August 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498001104.

Marseguerra:1998:MCS

[1253] Marzio Marseguerra and Enrico Zio.

Monte Carlo simulation of the effects of human intrusion on groundwater contaminant transport. Mathematics and Computers in Simulation, 47(2– 5):361–369, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001116.

Marseguerra:1998:CPD

[1254] M. Marseguerra, E. Zio, J. Devooght, and P. E. Labeau. A concept paper on dynamic reliability via Monte Carlo simulation. Mathematics and Computers in Simulation, 47(2–5): 371–382, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001128.

Nedjalkov:1998:CMC

[1255] M. Nedjalkov and I. Dimov. Convergency of the Monte Carlo algorithms for linear transport modeling. Mathematics and Computers in Simulation, 47(2–5):383–390, August 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847549800113X.

Nedjalkov:1998:SMP

[1256] M. Nedjalkov, T. Gurov, and I. Dimov. Statistical modeling of pulse excited electron quantum kinetics in a one-band semiconductor. Mathematics and Computers in Simulation, 47(2– 5):391–402, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498001141.

Ninomiya:1998:CNC

[1257] Syoiti Ninomiya. Constructing a new class of low-discrepancy sequences by using the β-adic transformation. Mathematics and Computers in Simulation, 47(2–5):403– 418, August 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475498001153.

Robert:1998:MCS

[1258] C. Robert, B. Gruzza, L. Bideux, and P. Bondot. A Monte Carlo study for electron elastic backscattering based on layered models for substrates. Mathematics and Computers in Simulation, 47(2–5):419– 427, August 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475498001232.

Sabelfeld:1998:TPS

[1259] K. K. Sabelfeld and O. Kurbanmuradov. Two-particle stochastic Eulerian–Lagrangian models of turbulent dispersion. Mathematics and Computers in Simulation, 47(2–5): 429–440, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498001244.

Semerdjiev:1998:TTU

[1260] Tz. A. Semerdjiev, V. P. Jilkov, and D. S. Angelova. Target tracking using Monte Carlo simulation. Mathematics and Computers in Simulation, 47(2– 5):441–447, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001256.

Shalimova:1998:MCI

[1261] I. A. Shalimova. A Monte Carlo iterative procedure for solving the pseudovibration elastic equation. Mathematics and Computers in Simulation, 47(2–5):449–453, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475498001268.

Simonov:1998:MCS

[1262] Nikolai A. Simonov. Monte Carlo solution to three-dimensional vorticity equation. Mathematics and Computers in Simulation, 47(2–5):455– 459, August 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847549800127X.

Smidts:1998:NAM

[1263] O. F. Smidts and J. Devooght. A non-analog Monte Carlo simulation of transport of radionuclides in a porous medium. Mathematics and Computers in Simulation, 47(2–5): 461–472, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498001281.

Tichy:1998:NAJ

[1264] Robert F. Tichy. Numerical analysis of Jacobi series — a quasi-Monte-Carlo approach. Mathematics and Computers in Simulation, 47(2–5): 473–481, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498001293.

Tombuyses:1998:CSE

[1265] B. Tombuyses and A. Dubus. Contribution to the study of elastic electron backscattering from polycrystalline metallic targets by Monte Carlo simulation. Mathematics and Computers in Simulation, 47(2–5):483– 491, August 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847549800130X.

Tombuyses:1998:BMC

[1266] B. Tombuyses, P. R. DeLuca, and C. Smidts. Backward Monte Carlo for probabilistic dynamics. Mathematics and Computers in Simulation, 47(2– 5):493–505, August 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498001311.

Anonymous:1998:PAc

[1267] Anonymous. Pages 83–518 (1 August 1998). Mathematics and Computers in Simulation, 47(2–5):??, August 1, 1998.

CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Ismail:1998:NSC

[1268] M. S. Ismail and T. R. Taha. A numerical study of compactons. Mathematics and Computers in Simulation, 47(6): 519–530, September 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001323.

Abouchabaka:1998:NAF

[1269] J. Abouchabaka, R. Aboulaïch, and A. Souissi. Numerical approach of a free boundary in the junction field effect transistor — MES-FET. Mathematics and Computers in Simulation, 47(6):531–539, September 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498001335.

Ismail:1998:NHO

[1270] Gamal A. F. Ismail and Iman H.

Ibrahim. A new higher order effective P–C methods for stiff systems. Mathematics and Computers in Simulation, 47(6):541–552, September 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498001360.

Toumi:1998:WAA

[1271] Ahmed Toumi. A well adapted approach to block-diagonalization of large scale systems. Mathematics and Computers in Simulation, 47(6):553– 570, September 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498001372.

Korn:1998:SDE

[1272] Granino A. Korn. Solving 5997 differential equations under Microsoft Windows 95 and NT. Mathematics and Computers in Simulation, 47(6):571– 582, September 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498001396.

Nowman:1998:CJB

[1273] K. Ben Nowman and Ghulam Sorwar. Computation of Japanese bonds and derivative securities. Mathematics and Computers in Simulation, 47(6): 583–588, September 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498001402.

Anonymous:1998:Ic

[1274] Anonymous. Index. Mathematics and Computers in Simulation, 47(6):601– 603, September 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475498001591.

Anonymous:1998:PS

[1275] Anonymous. Pages 519–604 (1 September 1998). Mathematics and Computers in Simulation, 47(6):??, September 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1998:Ef

[1276] Anonymous. Editorial. Mathematics and Computers in Simulation, 48 (1):1, November 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001499.

Strasser:1998:HPI

[1277] Reto J. Strasser and Alexandrina D.

Stirbet. Heterogeneity of photosystem II probed by the numerically simulated chlorophyll a fluorescence rise (O–J–I–P). Mathematics and Computers in Simulation, 48(1): 3–9, November 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001505.

Azevedo:1998:SAP

[1278] I. C. A. Azevedo, F. A. R. Oliveira, and M. C. Drumond. A study on the accuracy and precision of external mass transfer and diffusion coefficients jointly estimated from pseudo-experimental simulated data. Mathematics and Computers in Simulation, 48(1):11–22, November 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498001517.

Mouazen:1998:RFE

[1279] A. M. Mouazen and M. Neményi. A review of the finite element modelling techniques of soil tillage. Mathematics and Computers in Simulation, 48 (1):23–32, November 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001529.

Buzas:1998:MSA

[1280] J. Buzás, I. Farkas, A. Biró, and R. Németh. Modelling and simulation aspects of a solar hot water system. Mathematics and Computers in Simulation, 48(1):33–46, November 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498001530.

Chalabi:1998:MMM

[1281] Z. S. Chalabi. Mathematical methods for modelling and identification of nonlinear agricultural systems. Mathematics and Computers in Simulation, 48 (1):47–52, November 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001542.

Geeraerd:1998:EMP

[1282] Annemie H. Geeraerd, Carl H. Herremans, Linda R. Ludikhuyze, Marc E. Hendrickx, and Jan F. Van Impe. Evaluation of model parameter accuracy by using joint confidence regions: application to low complexity neural networks to describe enzyme inactivation. Mathematics and Computers in Simulation, 48(1):53– 64, November 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001487.

deAndres-Toro:1998:KMB

[1283] B. de Andrés-Toro, J. M. Girón-Sierra, J. A. López-Orozco, C. Fernández-Conde, J. M. Peinado, and F. García-Ochoa. A kinetic model for beer production under industrial operational conditions. Mathematics and Computers in Simulation, 48(1):65– 74, November 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001475.

Brauner:1998:IRS

[1284] Neima Brauner and Mordechai Shacham.

Identifying and removing sources of imprecision in polynomial regression. Mathematics and Computers in Simulation, 48(1):75–91, November 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498001463.

Tantau:1998:ESP

[1285] Hans-Juergen Tantau. Energy saving potential of greenhouse climate control. Mathematics and Computers in Simulation, 48(1):93–101, November 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498001451.

Kanali:1998:SIU

[1286] Christopher Kanali, Haruhiko Murase, and Nobuo Honami. Shape identification using a charge simulation retina model. Mathematics and Computers in Simulation, 48(1):103–118, November 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847549800144X.

Anonymous:1998:PN

[1287] Anonymous. Pages 1–132 (1 November 1998). Mathematics and Computers in Simulation, 48(1):??, November 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Carpanzano:1998:SMT

[1288] Emanuele Carpanzano and Claudio Maffezzoni. Symbolic manipulation techniques for model simplification in object-oriented modelling of large scale continuous systems. Mathematics and Computers in Simulation, 48(2):133– 150, December 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001414.

Rjasanow:1998:TCP

[1289] Sergej Rjasanow and Wolfgang Wagner. On time counting procedures in the DSMC method for rarefied gases. Mathematics and Computers in Simulation, 48(2):151–176, December 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498001426.

Spee:1998:NSG

[1290] E. J. Spee, J. G. Verwer, P. M.

de Zeeuw, J. G. Blom, and W. Hundsdorfer. A numerical study for global atmospheric transport–chemistry problems. Mathematics and Computers in Simulation, 48(2):177–204, December 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498001554.

Baumann:1998:MPD

[1291] G. Baumann. MathLie a program of doing symmetry analysis. Mathematics and Computers in Simulation, 48(2): 205–223, December 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475498001438.

Juncu:1998:PAG

[1292] Gh. Juncu and C. Popa. Preconditioning by approximations of the Gram matrix for convection–diffusion equations. Mathematics and Computers in Simulation, 48(2):225–233, December 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498001566.

Bruno:1998:ENP

[1293] A. D. Bruno. Erratum to “Newton polyhedra and power transformations”[Math. Comp. Simul. 45 (1998) 429–443]. Mathematics and Computers in Simulation, 48(2):235, December 1, 1998. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475498001621. See

[1163].

Anonymous:1998:PD

[1294] Anonymous. Pages 133–250 (1 December 1998). Mathematics and Computers in Simulation, 48(2):??, December 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

daSilva:1999:NAR

[1295] Ivan N. da Silva, Lucia V. R. de Arruda, and Wagner C. do Amaral. A novel approach to robust parameter estimation using neurofuzzy systems. Mathematics and Computers in Simulation, 48(3):251– 268, February 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847549800161X.

Nejjari:1999:EOC

[1296] F. Nejjari, G. Roux, B. Dahhou, and A. Benhammou. Estimation and optimal control design of a biological wastewater treatment process. Mathematics and Computers in Simulation, 48(3):269– 280, February 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847549800158X.

Friedrich:1999:DRB

[1297] J. Friedrich. A dual reciprocity boundary element model for the degradation of strongly eroded archaeological signs. Mathematics and Computers in Simulation, 48(3):281– 293, February 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475498001608.

Valdes:1999:GSC

[1298] J. L. Valdés, R. Biscay, and J. C.

Jimenez. Geometric selection of centers for radial basis function approximations involved in intensive computer simulations. Mathematics and Computers in Simulation, 48(3):295– 306, February 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499000026.

Albertsen:1999:CFS

[1299] Niels Christian Albertsen, Jean-Marie Chesneaux, Søren Christiansen, and Armand Wirgin. Comparison of four software packages applied to a scattering problem. Mathematics and Computers in Simulation, 48(3):307– 317, February 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499000038.

Delgado:1999:EPP

[1300] Alberto Delgado. Erratum to “Phase portrait approximation using dynamic neural networks” [Math. Comp. Simul. 47 (1998) 1–11]. Mathematics and Computers in Simulation, 48 (3):319, February 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000075. See

[1227].

Anonymous:1999:PF

[1301] Anonymous. Pages 251–338 (February 1999). Mathematics and Computers in Simulation, 48(3):??, February 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1999:Ea

[1302] Anonymous. Editorial. Mathematics and Computers in Simulation, 48 (4–6):339–340, June 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000142.

Pagan:1999:SUS

[1303] Adrian Pagan. Some uses of simulation in econometrics. Mathematics and Computers in Simulation, 48 (4–6):341–349, June 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000130.

Namekawa:1999:CSA

[1304] Mitsuhiro Namekawa, Akira Satoh, Hideki Mori, Kunio Yikai, and Toshio Nakanishi. Clock synchronization algorithm for parallel road-traffic simulation system in a wide area. Mathematics and Computers in Simulation, 48(4–6):351–359, June 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000154.

Sharma:1999:NAD

[1305] Ashish Sharma and Upmanu Lall.

A nonparametric approach for daily rainfall simulation. Mathematics and Computers in Simulation, 48(4–6): 361–371, June 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499000166.

Hurn:1999:EPS

[1306] A. S. Hurn and K. A. Lindsay. Estimating the parameters of stochastic differential equations. Mathematics and Computers in Simulation, 48 (4–6):373–384, June 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000178.

Bartolucci:1999:AMS

[1307] Alfred A. Bartolucci, Karan P. Singh, Anne D. Bartolucci, and Sejong Bae. Applying medical survival data to estimate the three-parameter Weibull distribution by the method of probabilityweighted moments. Mathematics and Computers in Simulation, 48(4–6): 385–392, June 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847549900018X.

Pollett:1999:MQS

[1308] P. K. Pollett. Modelling quasistationary behaviour in metapopulations. Mathematics and Computers in Simulation, 48(4–6):393– 405, June 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475499000191.

Biondini:1999:EHP

[1309] Riccardo Biondini and Yan-Xia Lin. Estimating the Hurst parameter in fractional ARIMA(p, d, q) models via the quasi-likelihood method. Mathematics and Computers in Simulation, 48(4–6):407–416, June 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000208.

Morimune:1999:DTU

[1310] Kimio Morimune and Mitsuru Nakagawa. The discontinuous trend unit root test when the break point is misspecified. Mathematics and Computers in Simulation, 48(4–6): 417–427, June 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847549900021X.

Wang:1999:QLM

[1311] You-Gan Wang, Yan-Xia Lin, and M. D. E. Haywood. A quasi-likelihood method for fractal-dimension estimation. Mathematics and Computers in Simulation, 48(4–6):429– 436, June 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475499000324.

Breslin:1999:FDR

[1312] M. C. Breslin and J. A. Belward.

Fractal dimensions for rainfall time series. Mathematics and Computers in Simulation, 48(4–6):437– 446, June 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499000233.

Kentwell:1999:GCF

[1313] D. J. Kentwell, L. M. Bloom, and G. A. Comber. Geostatistical conditional fractal simulation with irregularly spaced data. Mathematics and Computers in Simulation, 48 (4–6):447–456, June 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000245.

Toda:1999:LTU

[1314] Hiro Y. Toda and C. R. McKenzie. LM tests for unit roots in the presence of missing observations: small sample evidence. Mathematics and Computers in Simulation, 48 (4–6):457–468, June 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000257.

Norton:1999:TBT

[1315] J. P. Norton. Translation of bounds on time-domain behaviour of dynamical systems into parameter bounds for discrete-time rational transferfunction models. Mathematics and Computers in Simulation, 48(4–6):

469–478, June 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475499000269.

Fiebig:1999:SBU

[1316] Denzil G. Fiebig and Pierre-François Uldry. Sensitivity bounds for use with flawed data. Mathematics and Computers in Simulation, 48(4–6): 479–486, June 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475499000270.

Nishimura:1999:ECS

[1317] Kazuo Nishimura and Makoto Yano. On the existence of chaotic solutions in dynamic linear programming. Mathematics and Computers in Simulation, 48(4–6):487–496, June 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000282.

Harrison:1999:NLN

[1318] Robert G. Harrison, Dejin Yu, Les Oxley, Weiping Lu, and Donald George. Non-linear noise reduction and detecting chaos: some evidence from the S&P Composite Price Index. Mathematics and Computers in Simulation, 48 (4–6):497–502, June 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000294.

Tsui:1999:CCC

[1319] Albert K. Tsui and Qiao Yu. Constant conditional correlation in a bivariate GARCH model: evidence from the stock markets of China. Mathematics and Computers in Simulation, 48(4–6):503–509, June 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000300.

Chan:1999:RTA

[1320] W. S. Chan, Harry W. C. Lo, and S. H.

Cheung. Return transmission among stock markets of Greater China. Mathematics and Computers in Simulation, 48(4–6):511–518, June 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000312.

Sequeira:1999:EAP

[1321] John M. Sequeira, Michael McAleer, and Ying-Foon Chow. Estimation of alternative pricing models for currency futures contracts. Mathematics and Computers in Simulation, 48 (4–6):519–530, June 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000348.

Groenewold:1999:TVE

[1322] Nicolaas Groenewold and Patricia Fraser. Time-varying estimates of CAPM betas. Mathematics and Computers in Simulation, 48(4–6): 531–539, June 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475499000336.

Hall:1999:PFA

[1323] A. D. Hall. Parametric forecasts of Australian yield curves. Mathematics and Computers in Simulation, 48 (4–6):541–549, June 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847549900035X.

Leong:1999:TLC

[1324] Kenneth Leong and Michael McAleer. Testing the life-cycle permanent income hypothesis using intra-year data for Sweden. Mathematics and Computers in Simulation, 48(4–6):551– 560, June 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475499000361.

Tsuji:1999:EAI

[1325] Masatsugu Tsuji, Shoichi Miyahara, and Mineo Ishikawa. An empirical analysis of industrial transformation in the Japanese machine tool industry. Mathematics and Computers in Simulation, 48(4–6):561–572, June 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000373.

Lim:1999:SAM

[1326] Christine Lim and Michael McAleer. A seasonal analysis of Malaysian tourist arrivals to Australia. Mathematics and Computers in Simulation, 48 (4–6):573–583, June 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000385.

Anonymous:1999:Ia

[1327] Anonymous.  Index.  Mathemat-

ics and Computers in Simulation, 48 (4–6):597–600, June 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000889.

Anonymous:1999:PJa

[1328] Anonymous. Pages 339–600 (June 1999). Mathematics and Computers in Simulation, 48(4–6):??, June 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Liu:1999:DMC

[1329] Pingzhou Liu and Xiaoying Cui. A discrete model of competition. Mathematics and Computers in Simulation, 49(1–2):1–12, July 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847549900004X.

Dumitrache:1999:GLF

[1330] Ion Dumitrache and Cătălin Buiu.

Genetic learning of fuzzy controllers. Mathematics and Computers in Simulation, 49(1–2):13–26, July 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475499000051.

Li:1999:HLM

[1331] Jichun Li and Junji Matsumoto. A heterogeneous level method for threedimensional hydrodynamics and salinity bay modelling. Mathematics and Computers in Simulation, 49(1–2): 27–39, July 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499000117.

dOvidio:1999:BAS

[1332] F. d’Ovidio, C. A. Andersen, C. N. Ernstsen, and E. Mosekilde. Bifurcation analysis of spiral growth processes in plants. Mathematics and Computers in Simulation, 49(1–2): 41–56, July 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499000105.

Kolodko:1999:SMS

[1333] A. Kolodko, K. Sabelfeld, and W. Wagner. A stochastic method for solving Smoluchowski’s coagulation equation. Mathematics and Computers in Simulation, 49(1–2): 57–79, July 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499000087.

Antoniou:1999:WSP

[1334] I. Antoniou and K. Gustafson. Wavelets and stochastic processes. Mathematics and Computers in Simulation, 49(1–2):81–104, July 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475499000099.

Schatzer:1999:VOT

[1335] C. Schätzer, W. Müller, and A. Binder.

Vector optimization of two-dimensional numerical field problems applied to the design of a wind turbine generator. Mathematics and Computers in Simulation, 49(1–2):105– 117, July 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499000129.

El-Sheikh:1999:UNS

[1336] M. G. El-Sheikh, M. E. Khalifa, and V. Gavdzinski. The unbounded normal stress of a contact problem for a cylinder. Mathematics and Computers in Simulation, 49(1–2): 119–128, July 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499000397.

Korn:1999:ISP

[1337] Granino A. Korn. Interactive solution of partial differential equations by the method-of-lines. Mathematics and Computers in Simulation, 49 (1–2):129–138, July 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000415.

Anonymous:1999:PJb

[1338] Anonymous. Pages 1–148 (July 1999).

Mathematics and Computers in Simulation, 49(1–2):??, July 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1999:Eb

[1339] Anonymous. Editorial. Mathematics and Computers in Simulation, 49(3):149, August 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000476.

Bernardin:1999:JFM

[1340] Laurent Bernardin. A Java framework for massively distributed symbolic computing. Mathematics and Computers in Simulation, 49(3):151– 160, August 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475499000488.

Blochinger:1999:OOP

[1341] Wolfgang Blochinger, Wolfgang Küchlin, Christoph Ludwig, and Andreas Weber. An object-oriented platform for distributed high-performance symbolic computation. Mathematics and Computers in Simulation, 49(3):161– 178, August 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847549900049X.

Ponitz:1999:CIG

[1342] André Pönitz. Computing invariants in graphs of small bandwidth. Mathematics and Computers in Simulation, 49(3):179–191, August 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847549900052X.

Mihailescu:1999:RDP

[1343] Preda Mihăilescu. Recent developments in primality proving. Mathematics and Computers in Simulation, 49(3):193–204, August 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000518.

Lewis:1999:SRP

[1344] Robert H. Lewis and Peter F. Stiller.

Solving the recognition problem for six lines using the Dixon resultant. Mathematics and Computers in Simulation, 49(3):205–219, August 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000506.

Monagan:1999:DAO

[1345] M. B. Monagan and J. F. Ogilvie. The diatomic anharmonic oscillator according to matrix mechanics. Mathematics and Computers in Simulation, 49 (3):221–234, August 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000464.

Sorgatz:1999:THP

[1346] Andreas Sorgatz and Stefan Wehmeier.

Towards high-performance symbolic computing: using MuPAD as a problem solving environment. Mathematics and Computers in Simulation, 49 (3):235–246, August 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000531. Spe-

cial issue on high performance symbolic

computing.

Anonymous:1999:PA

[1347] Anonymous. Pages 149–256 (August 1999). Mathematics and Computers in Simulation, 49(3):??, August 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Amberg:1999:FES

[1348] Gustav Amberg, Robert Tönhardt, and Christian Winkler. Finite element simulations using symbolic computing. Mathematics and Computers in Simulation, 49(4–5):257–274, September 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475499000543.

Vutova:1999:ETF

[1349] K. Vutova, G. Mladenov, T. Tanaka, and K. Kawabata. Evaluation of thin film surface topology shapes. Mathematics and Computers in Simulation, 49(4–5):275–283, September 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475499000440.

Paulo:1999:MMS

[1350] Eugene P. Paulo and Linda C. Malone. Mathematical model of sectoring in 3D space. Mathematics and Computers in Simulation, 49(4–5): 285–296, September 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000427.

Braselton:1999:RWT

[1351] James Braselton, John Rafter, Patricia Humphrey, and Martha Abell. Randomly walking through Wall Street: Comparing lump-sum versus dollar-cost average investment strategies. Mathematics and Computers in Simulation, 49(4–5):297– 318, September 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499000403.

Wang:1999:SDB

[1352] Guoli Wang and Heinz Unbehauen.

Symmetric dichotomy based inverse dynamic model of a high-speed flexible beam. Mathematics and Computers in Simulation, 49(4–5):319– 330, September 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499000439.

Dehghan:1999:ILO

[1353] Mehdi Dehghan. Implicit locally one-dimensional methods for twodimensional diffusion with a non-local boundary condition. Mathematics and Computers in Simulation, 49(4–5): 331–349, September 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000567.

Lara:1999:APR

[1354] Martín Lara, Antonio Elipe, and Manuel Palacios. Automatic programming of recurrent power series. Mathematics and Computers in Simulation, 49(4–5):351–362, September 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475499000877.

Potocnik:1999:EMA

[1355] Primož Potočnik and Igor Grabec. Empirical modeling of antibiotic fermentation process using neural networks and genetic algorithms. Mathematics and Computers in Simulation, 49(4– 5):363–379, September 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000452.

Fantoni:1999:SHA

[1356] Alessandro Fantoni, Manuela Vieira, and Rodrigo Martins. Simulation of hydrogenated amorphous and microcrystalline silicon optoelectronic devices. Mathematics and Computers in Simulation, 49(4–5):381– 401, September 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475499000555.

Anonymous:1999:PS

[1357] Anonymous. Pages 257–412 (September 1999). Mathematics and Computers in Simulation, 49(4–5):??, September 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Bai:1999:COS

[1358] Er-Wei Bai and Yih-Fang Huang. Convergence of optimal sequential outer bounding sets in bounded error parameter estimation. Mathematics and Computers in Simulation, 49(6): 307–317, October 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000919.

Chentsov:1999:SCC

[1359] A. G. Chentsov, S. I. Morina, and B. B. Zobnin. On some constructions of control by systems with a varying structure. Mathematics and Computers in Simulation, 49(6):319– 334, October 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499000932.

Golenko-Ginzburg:1999:DCO

[1360] Dimitri Golenko-Ginzburg, Ahazon Gonik, and Ljubisa Papic. Developing cost-optimization production control model via simulation. Mathematics and Computers in Simulation, 49 (6):335–351, October 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000907.

Mainik:1999:QPN

[1361] J. Mainik and C. P. Ortlieb. Qualitative properties of a niche-model. Mathematics and Computers in Simulation, 49(6):353–374, October 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000890.

Anonymous:1999:Ib

[1362] Anonymous. Index. Mathematics and Computers in Simulation, 49(6): 387–389, October 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847549900110X.

Anonymous:1999:PO

[1363] Anonymous. Pages 307–390 (October 1999). Mathematics and Computers in Simulation, 49(6):??, October 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1999:Ec

[1364] Anonymous. Editorial. Mathematics and Computers in Simulation, 50 (1–4):1–2, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000579.

Anonymous:1999:C

[1365] Anonymous. Conference. Mathematics and Computers in Simulation, 50 (1–4):3–10, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000749.

Axelsson:1999:ISS

[1366] Owe Axelsson. On iterative solvers in structural mechanics; separate displacement orderings and mixed variable methods. Mathematics and Computers in Simulation, 50(1–4): 11–30, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000580.

Beauwens:1999:ISB

[1367] Robert Beauwens. Implementation strategies for block recursive factorizations. Mathematics and Computers in Simulation, 50(1–4):31– 41, November 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499000592.

Eck:1999:SBE

[1368] C. Eck, O. Steinbach, and W. L.

Wendland. A symmetric boundary element method for contact problems with friction. Mathematics and Computers in Simulation, 50(1–4): 43–61, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000609.

Ewing:1999:IFV

[1369] Richard E. Ewing, Zhilin Li, Tao Lin, and Yanping Lin. The immersed finite volume element methods for the elliptic interface problems. Mathematics and Computers in Simulation, 50 (1–4):63–76, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000610.

Godlewski:1999:LSS

[1370] Edwige Godlewski and Pierre-Arnaud Raviart. The linearized stability of solutions of nonlinear hyperbolic systems of conservation laws: a general numerical approach. Mathematics and Computers in Simulation, 50 (1–4):77–95, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000622.

Lakhany:1999:SRB

[1371] A. M. Lakhany and J. R. Whiteman. Superconvergent recovery based error estimators. Mathematics and Computers in Simulation, 50(1–4): 97–114, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000634.

Bartos:1999:FES

[1372] M. Bartoš, Z. Kestřánek, Z. Kestřánek, J. Nedoma, and J. Stehlík. On the 2D and 3D finite element simulation in orthopaedy using MRI. Mathematics and Computers in Simulation, 50(1– 4):115–121, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000658.

Blaheta:1999:ACG

[1373] Radim Blaheta. Adaptive composite grid methods for problems of plasticity. Mathematics and Computers in Simulation, 50(1–4):123– 134, November 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499000646.

Bock:1999:LDV

[1374] Igor Bock. On large deflections of viscoelastic plates. Mathematics and Computers in Simulation, 50(1–4): 135–143, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847549900066X.

Brandts:1999:ICS

[1375] Jan H. Brandts. Improving the convergence of some boundary element methods. Mathematics and Computers in Simulation, 50(1–4): 145–152, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000671.

DellaCroce:1999:SCS

[1376] L. Della Croce and T. Scapolla. Solving cylindrical shell problems with a nonstandard finite element. Mathematics and Computers in Simulation, 50(1– 4):153–164, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000683.

Cerny:1999:NSS

[1377] R. Černý and P. Přikryl. Numerical solution of a Stefan-like problem in laser processing of semiconducting alloys. Mathematics and Computers in Simulation, 50(1–4):

165–173, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000695.

Georgiev:1999:MAS

[1378] A. Georgiev, S. Margenov, and M. Neytcheva. Multilevel algorithms for 3D simulation of nonlinear elasticity problems. Mathematics and Computers in Simulation, 50(1–4): 175–182, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000701.

Holubova–Tajcova:1999:MMS

[1379] Gabriela Holubová-Tajčová. Mathematical modeling of suspension bridges. Mathematics and Computers in Simulation, 50(1–4):183– 197, November 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475499000713.

Kestranek:1999:IMC

[1380] Z. Kestřánek. Iterative methods for contact problems. Mathematics and Computers in Simulation, 50(1–4): 199–204, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000750.

Kostecky:1999:MRS

[1381] Pavel Kostecký and Igor Kohút. Modelling of the rock structure stress field near the cavities and estimation of the cavity effect influence on the tidal measurements. Mathematics and Computers in Simulation, 50(1– 4):205–214, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000725.

Koukal:1999:SRF

[1382] Stanislav Koukal and Michal Křížek.

On strongly regular families of piecewise quasi-uniform triangulations of certain curved plane domains. Mathematics and Computers in Simulation, 50(1–4):215–225, November 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499000737.

Kralik:1999:ERA

[1383] Juraj Králik and Miroslav Šimonovič.

Earthquake response analysis of nuclear power plant buildings with soil– structural interaction. Mathematics and Computers in Simulation, 50(1– 4):227–236, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000762.

Krizek:1999:NEF

[1384] Michal Křížek. Numerical experience with the finite speed of gravitational interaction. Mathematics and Computers in Simulation, 50(1– 4):237–245, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000853.

Lirkov:1999:MPI

[1385] Ivan Lirkov and Svetozar Margenov. MPI parallel implementation of CBF preconditioning for 3D elasticity problems. Mathematics and Computers in Simulation, 50(1–4):247–254, November 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475499000841.

Lu:1999:SIS

[1386] Xiaowu Lu and Rudolf Schmid. Symplectic integration of Sine–Gordon type systems. Mathematics and Computers in Simulation, 50(1–4):255– 263, November 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847549900083X.

Nedoma:1999:ACS

[1387] Jiří Nedoma. Analysis of a coupled system of equations of a global geodynamic model of the earth. Mathematics and Computers in Simulation, 50(1– 4):265–283, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000865.

Nedoma:1999:NAL

[1388] J. Nedoma, M. Bartoš, H. Hornátová, Z. Kestřánek, and J. Stehlík. Numerical analysis of the loosened total hip replacements (THR). Mathematics and Computers in Simulation, 50(1– 4):285–304, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000828.

Repin:1999:UAP

[1389] Sergey I. Repin. A unified approach to a posteriori error estimation based on duality error majorants. Mathematics and Computers in Simulation, 50(1– 4):305–321, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000816.

Sanchez-Avila:1999:ALA

[1390] C. Sánchez-Ávila and J. A. García-Moreno. An adaptive LSQR algorithm for computing discontinuous solutions in deconvolution problems. Mathematics and Computers in Simulation, 50(1–4):323–329, November 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499000804.

Segeth:1999:PEE

[1391] K. Segeth. A posteriori error estimates and grid adjustment for a nonlinear parabolic equation. Mathematics and Computers in Simulation, 50(1– 4):331–338, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000798.

Vujevic:1999:SAI

[1392] Slavko Vujević and Mate Kurtović. Statistical analysis and interpretation of resistivity sounding data in earthing grid design. Mathematics and Computers in Simulation, 50(1–4): 339–349, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000786.

Zhang:1999:TDB

[1393] Ch. Zhang and A. Savaidis. Timedomain BEM for dynamic crack analysis. Mathematics and Computers in Simulation, 50(1–4):351– 362, November 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475499000774.

Anonymous:1999:PN

[1394] Anonymous. Pages 1–372 (November 1999). Mathematics and Computers in Simulation, 50(1–4):??, November 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:1999:Ed

[1395] Anonymous. Editorial. Mathematics and Computers in Simulation, 50(5–6): 373–375, December 1, 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499001433.

Lu:1999:EOW

[1396] Ya Yan Lu. Exact one-way methods for acoustic waveguides. Mathematics and Computers in Simulation, 50(5–6): 377–391, December 1, 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499001111.

Steinberg:1999:MHM

[1397] Ben Zion Steinberg and John J. Mc-Coy. A multiresolution homogenization of modal analysis with application to layered media. Mathematics and Computers in Simulation, 50(5–6): 393–417, December 1, 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000956.

Belishev:1999:REP

[1398] M. Belishev and A. Glasman. Riccati equation in projecting solenoidal fields. Mathematics and Computers in Simulation, 50(5–6):419–433, December 1, 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499000968.

He:1999:WSA

[1399] Sailing He and Vaughan H. Weston.

Wave-splitting and absorbing boundary condition for Maxwell’s equations on a curved surface. Mathematics and Computers in Simulation, 50(5–6): 435–455, December 1, 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847549900097X.

He:1999:ISF

[1400] Sailing He and Vladimir G. Romanov. Identification of small flaws in conductors using magnetostatic measurements. Mathematics and Computers in Simulation, 50(5–6):457–471, December 1, 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475499000981.

Powell:1999:ISM

[1401] Jeffrey O. Powell. An inverse scattering method for a 1-D dissipative Helmholtz equation. Mathematics and Computers in Simulation, 50(5–6):473– 488, December 1, 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000993.

Wall:1999:IPI

[1402] David J. N. Wall and Jonas Lundstedt. Inverse problems involving the one-way wave equation: medium function reconstruction. Mathematics and Computers in Simulation, 50(5–6):489– 510, December 1, 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499001007.

Cao:1999:LSP

[1403] Jiasu Cao, Sailing He, and Vaughan H. Weston. Layer-stripping and parameter reconstruction for a hyperbolic equation in a three-dimensional inhomogeneous half-space. Mathematics and Computers in Simulation, 50(5–6):

511–525, December 1, 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499001019.

Cheney:1999:RSP

[1404] Margaret Cheney, David Isaacson, Victoria I. Lytle, and Stephen F. Ackley. Recovery of surface parameters from stepped-frequency radar returns. Mathematics and Computers in Simulation, 50(5–6):527–539, December 1, 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001020.

Gustafsson:1999:WSB

[1405] Mats Gustafsson and Sailing He. A wave-splitting based optimization approach to multi-dimensional timedomain electromagnetic inverse problems. Mathematics and Computers in Simulation, 50(5–6):541–551, December 1, 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001032.

Norgren:1999:OAF

[1406] Martin Norgren. Optimization approaches to frequency-domain inverse problems for Kerr-like and non-Kerr-like nonlinear media. Mathematics and Computers in Simulation, 50(5–6): 553–575, December 1, 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499001044.

Folkow:1999:DIP

[1407] Peter D. Folkow and Kevin Kreider. Direct and inverse problems on nonlinear rods. Mathematics and Computers in Simulation, 50(5–6):577– 595, December 1, 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499001056.

Anonymous:1999:Ic

[1408] Anonymous. Index. Mathematics and Computers in Simulation, 50(5–6): 597–599, December 1, 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499001445.

Anonymous:1999:PDa

[1409] Anonymous. Pages 373–610 (1 December 1999). Mathematics and Computers in Simulation, 50(5–6):??, December 1, 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Costantini:1999:VRA

[1410] C. Costantini. Variance reduction by antithetic random numbers of Monte Carlo methods for unrestricted and reflecting diffusions. Mathematics and Computers in Simulation, 51(1–2):1– 17, December 22, 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499000944.

Liatsis:1999:TDR

[1411] P. Liatsis and H. M. Tawfik. Twodimensional road shape optimisation using genetic algorithms. Mathematics and Computers in Simulation, 51(1–2): 19–31, December 22, 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499000920.

Bella:1999:STA

[1412] G. Bella, M. Burroni, M. M. Cerimele, and F. Pistella. A simulation tool for the analysis of high speed flows. Mathematics and Computers in Simulation, 51(1–2):33–45, December 22, 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001093.

Gustafson:1999:PCF

[1413] K. Gustafson. Parallel computing forty years ago. Mathematics and Computers in Simulation, 51(1–2):47– 62, December 22, 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499001081.

Agiza:1999:MDC

[1414] Hamdy Nabih Agiza, Gian Italo Bischi, and Michael Kopel. Multistability in a dynamic Cournot game with three oligopolists. Mathematics and Computers in Simulation, 51(1–2): 63–90, December 22, 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499001068.

Hsiao:1999:SAT

[1415] Chun-Hui Hsiao and Wen-June Wang. State analysis of time-varying singular nonlinear systems via Haar wavelets. Mathematics and Computers in Simulation, 51(1–2):91–100, December 22, 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847549900107X.

Fuks:1999:PDN

[1416] Henryk Fukś and Anna T. Lawniczak. Performance of data networks with random links. Mathematics and Computers in Simulation, 51(1–2):101– 117, December 22, 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499001251.

Abonyi:1999:IFP

[1417] János Abonyi, Hans Andersen, Lajos Nagy, and Ferenc Szeifert. Inverse fuzzy-process-model based direct adaptive control. Mathematics and Computers in Simulation, 51(1–2):119– 132, December 22, 1999. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475499001421.

Anonymous:1999:PDb

[1418] Anonymous. Pages 1–142 (22 December 1999). Mathematics and Computers in Simulation, 51(1–2):??, December 22, 1999. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).